### v2

# Robust Signal Detection with Quadratically Convex Orthosymmetric Constraints

Abstract
This paper is concerned with robust signal detection in Gaussian noise under quadratically convex orthosymmetric (QCO) constraints. Specifically, the null hypothesis $H_{0}$ assumes no signal, whereas the alternative $H_{1}$ considers signal that is separated in Euclidean norm from zero and belongs to a set $K$ satisfying the QCO constraints. In addition, an adversary is allowed to inspect all the $N$ samples and replace up to $\epsilon N$ of them with arbitrary values, where $0<\epsilon<c_{0}<\frac{1}{2}$ is the corruption rate. Our main results establish the minimax rate of the separation radius $\rho_{\text{critical}}$ between $H_{0}$ and $H_{1}$ purely in terms of the geometry of $K$, the corruption rate $\epsilon$ (up to logarithmic factors in $\frac{1}{\epsilon}$) and the scale of the noise $\sigma$. We argue that the Kolmogorov widths of the constraint play a central role in determining the minimax rate. This indicates similarity with the (uncorrupted) estimation problem under QCO constraints, which was first established by Donoho et al., (1990). Moreover, the minimax lower bound reveals interesting phase transitions of the testing problem regarding $\epsilon$. Consistent with classic belief about testing and estimation, the testing problem is “easier” (even when one compares the results with recent papers studying the constrained robust estimation problem (Prasadan and Neykov,, 2025)). In addition to the main results above, where the upper bound is achieved with an intractable algorithm, inspired by Canonne et al., (2023), we develop a polynomial time algorithm which also nearly (up to logarithmic factors in $N$ and $1/\epsilon$) achieves the minimax lower bound. In contrast to Canonne et al., (2023), our algorithm works for signals of arbitrary Euclidean length, and respects the QCO constraint. Finally, all the results above are naturally extended to the $\ell_{p}$ norm testing problem for $1\leq p<2$.
1 Introduction
Data Contamination and Robust Statistics
One of the most basic tasks in statistics is to perform inference about the unknown parameters used in data generation. However, a variety of reasons can lead the observed data to deviate from the original samples of the true mechanism. These reasons could range from measurement errors in some percentage of the observations, to even malicious tampering with the data. For example, in certain biological datasets, inconsistencies in the data can be natural but beyond the level of white noise (see Martí,, 2019; Fibbi et al.,, 2023, e.g.), while in machine learning security applications, contamination can be artificial or malicious, which is also called data poisoning attacks (see Ramirez et al.,, 2022, e.g.); see also Diakonikolas and Kane, (2019) for other examples.
In this work, we take a stylized and classical statistical inference problem and study it from both constrained and robust perspectives. Specifically, we will study the Gaussian sequence model, where each coordinate $X_{ij}$ of a single sample $X_{i}$ is drawn independently from $\mathcal{N}\left(\mu_{j},\sigma^{2}\right)$, and $X_{i},i=1,2,\dots$ are i.i.d. Moreover, a constraint $K$ is imposed for the mean vector $\mu:=\left(\mu_{1},\mu_{2},\dots,\mu_{d}\right)$ s.t. $\mu\in K$, where $K$ is some appropriate set that belongs to $\mathbb{R}^{d}$ (or $K\subseteq\ell_{2}$ in the infinite-dimensional case). This extra information about the parameter encodes the prior knowledge of the statistician, and facilitates the inference meanwhile. Finally, and importantly, we also assume that there is a potential source of contamination, which is called the adversary. It is powerful in the sense that it has the knowledge of the true model and the access to the original data. For some positive constant $0\leq\epsilon<c_{0}<\frac{1}{2}$ known to both the statistician and the adversary, the adversary is allowed to inspect the original data and modify up to $\epsilon$ fraction of the samples arbitrarily. The statistician, however, does not know which samples have been modified. The goal of the statistician is to perform inference about $\mu$ based on the contaminated data. A more formal explanation of the the problem we study can be found in Section 1.1.
Performing inference under constraints and even contamination could be a formidable task. Most of the related works focus on the estimation and testing problem under several simple constraints with no contamination (Donoho and Gasko, (1992), Liu and Moitra, (2021), Liu and Moitra, (2023)). One crucial reason for this is that most of the methods and techniques heavily depend on the assumption of the true model behind the data (for example, likelihood ratio test and tests depending on the asymptotic distributions). Consider the following special case where $K=\mathbb{R}^{d}$ and $\epsilon=0$. Given $N$ samples $X_{1},X_{2},\dots,X_{N}\in\mathbb{R}^{d}$ from $\mathcal{N}(\mu,\sigma^{2}\mathbf{I}_{d})$ with unknown $\mu$ and known $\sigma\in\mathbb{R}^{+}$. In the testing problem, we are asked to distinguish between two hypotheses:
$H_{0}:\mu=0,$ (1.1)
$H_{1}:\left\lVert\mu\right\rVert_{2}\geq\rho.$
In the estimation problem, we are asked to give an estimate $\widehat{\mu}$ of $\mu$ to achieve a small error under a certain loss function (a natural one here is the $\ell_{2}$ loss). $\bar{X}=\frac{1}{N}\sum\limits_{i=1}^{N}X_{i}$ is a sufficient statistic for the model. The testing and estimation based on $\bar{X}$ is widely known to be optimal for both problems. Specifically, to control errors within the level of (small) constant, one needs $\rho\gtrsim\frac{d^{1/4}}{\sqrt{N}}\sigma$ for the testing problem (in the minimax sense) (Ingster and Suslina, (2003), Diakonikolas et al., (2022)) while $\inf\limits_{\widehat{\mu}}\sup\limits_{\mu\in\mathbb{R}^{d}}\mathbb{E}\left\lVert\widehat{\mu}-\mu\right\rVert_{2}^{2}\gtrsim\frac{d}{N}\sigma^{2}$ if one wants to optimally estimate $\mu$ in $\ell_{2}$ norm (LeCam, (1973), also see Example 15.4 in Wainwright, (2019)). However, as mentioned, the property of $\bar{X}$ is very sensitive to the authenticity of the samples. The method quickly breaks down with even modification on a single sample (the adversary may modify any single datum, thereby forcing the mean to be zero and causing considerable type-II error.). This is precisely the major reason why testing and estimation problems are very challenging when studied from a robust perspective.
Below, we provide an informal introduction to several key concepts that play a central role in our study — namely, the strong contamination model, the Kolmogorov $k$-width, and quadratically convex orthosymmetric (QCO) sets. Throughout this paper, we assume that the constraint set $K$ is a QCO set, and we operate under the strong contamination model. As will be shown later, the Kolmogorov widths of $K$ determine the critical radius of the testing problem, i.e., the minimal $\rho$ under which the problem is testable in the minimax sense.
Strong Contamination Model
Throughout the years, different types of adversary models have been proposed in the field of robust statistics. One of the earliest ones, the Huber contamination model, was introduced by Huber, (1964), where instead of directly sampling from the authentic distribution $\mathcal{P}$, it is assumed that there exists some error distribution $\mathcal{Q}$ selected by the adversary, and the data are sampled from the mixture $(1-\epsilon)\mathcal{P}+\epsilon\mathcal{Q}$. Huber’s contamination model can be regarded as a case of the larger oblivious adversary, where while the adversary has the adaptivity to some extent, the contamination process is guaranteed to be independent with the original samples. In contrast, the adversary in the strong contamination model (Diakonikolas and Kane, (2019)) is granted with strong adaptivity, specifically, with knowledge of the original data, the true model and even the algorithms. The adversary is allowed to inspect and modify the original data based on its knowledge. In this model, the fraction of the contaminated samples must be controlled to obtain useful information. Note that this quantity is not bounded in the Huber contamination model (although the expected number of corruptions is $\epsilon N$). Other types of contamination are also studied, including the perturbation of the sampling distribution with bounded TV distance (which also includes Huber’s contamination model). See Diakonikolas and Kane, (2019) for an overview.
Kolmogorov $k$-Width
Introduced by Kolmogoroff, (1936) in the study of function spaces and approximation theory, the Kolmogorov widths (also known as the Kolmogorov $k$-width) characterizes how well an element can be reached by some linear approximations of intrinsic dimension $k$, given that the element belongs to a linear space and some restrictions are applied. For a given $k$, it selects the best $k$-dimensional projection operator under the measure of minimax error. The reader can refer to Definition 2.2 for a formal presentation. Kolmogorov $k$-width is a classic concept and has seen usage in various fields. See Floater et al., (2021) for a recent work in the problem of low-rank approximation of matrices using the Kolmogorov $k$-width. Evans et al., (2009) uses this concept to study the $k$-method in isogeometric analysis. Kolmogorov $k$-width is also popular in the problems related to partial differential equations (PDEs), where the decay rate of the Kolmogorov $k$-width is intensely studied (Cohen and DeVore, (2015), Greif and Urban, (2019)). See also Papapicco et al., (2022) for an application in neural networks and PDEs, and a very recent work, Florian Arbes and Urban, (2025), for an application in optimal transportation and PDEs. Finally, see Greif and Urban, (2019) and Liang and Parks, (1996) for its power in the wave problem and wavelet representations. One can find more details and properties of the Kolmogorov $k$-width in Pinkus, (2012).
QCO Sets
Quadratically convex orthosymmetric (QCO) sets were first introduced in Donoho et al., (1990) (see also (Johnstone,, 2019, Chapter 4.8)), where the authors give many examples of such sets, such as hyperrectangles, ellipsoids (both aligned with the coordinate axes) and sets of the form $\{\theta\left\lvert\right.\sum\limits_{i\in I}a_{i}\psi(\theta_{i}^{2})\leq 1\}$, where $\psi$ is a given convex function and $a_{i}\in\mathbb{R}^{+},i=1,2,\dots$. The last example implies that all unit balls in (weighted) $\ell_{p}$ norms, for $p\geq 2$ are QCO sets. The set of QCO constraints is therefore much larger than simply the set of all ellipsoids. See Definition 2.1 for the formal definition of a QCO set. In contrast to the present work, which focuses on testing, Donoho et al., (1990) characterized the minimax estimation rates in the corresponding Gaussian sequence model under an uncorrupted setting. It turns out that the Kolmogorov widths of the set $K$ (see Definition 2.2) determine the minimax rate of estimation. Later, by taking $\epsilon=0$, we will observe that the Kolmogorov widths also determine the minimax rate of testing with QCO constraints, though the resulting rate is strictly smaller. Furthermore, we shall compare the minimax testing rate under QCO constrains with corruption to the corresponding minimax estimation rate (recently derived by Prasadan and Neykov, (2025)) and establish the same fundamental conclusion—testing is intrinsically easier than estimation.
1.1 Problem Formulation and Our Contributions
[…]
$H_{0}:\mu=0,$ (1.2)
$H_{1}:\left\lVert\mu\right\rVert_{2}\geq\rho,\mu\in K.$
Intuitively, if $\left\lVert\mu\right\rVert_{2}$ is overly close to zero or the fraction of the contaminated samples is overly large, we are not able to detect the difference in the worst case. Therefore, we are interested in characterizing the asymptotically smallest value of $\rho$ (depending on other parameters) for which the testing problem becomes feasible. In detail, for a given desired type-I and type-II errors $\alpha$, define the set of acceptable tests under $H_{0}$ as,
$A_{s}(N,d,K,\epsilon,\alpha,\sigma):=\left\{\phi:\sup\limits_{\mathcal{C}}\mathbb{P}_{0}\left(\phi(\mathcal{C}(\widetilde{\mathbf{X}}))=1\right)\leq\alpha\right\},$ (1.3)
where $\mathbb{P}_{0}$ denotes the law of $\widetilde{\mathbf{X}}$ when $\mu=\mathbf{0}$, and similar for $\mathbb{P}_{\mu}$. We write $A_{s}$ and omit the dependency hereafter when the context is clear. Let
$\rho(N,d,K,\epsilon,\alpha,\sigma):=\inf\limits_{\rho}\left\{\rho:\rho>0,\sup\limits_{\left\lVert\mu\right\rVert_{2}\geq\rho,\mu\in K}\sup\limits_{\mathcal{C}}\mathbb{P}_{\mu}\left(\phi(\mathcal{C}(\mathbf{X}))=0\right)\leq\alpha,\phi\in A_{s}\right\}$ (1.4)
be the smallest possible $\rho$ such that the testing task with both errors less than $\alpha$ is doable. We also write it as $\rho_{\text{critical}}$ for convenience when the context is clear. We are going to study the asymptotic behaviour of $\rho(N,d,K,\epsilon,\alpha,\sigma)$. See Figure 1 for an illustrative example of the problem.
XX0KK
Figure 1: Testing $H_{0}:\mu=0$ vs $H_{1}:\|\mu\|_{2}\geq\rho,\mu\in K$ for the axis aligned ellipsoid (which is a special case of a QCO set) $K$ based on a single observation $X$. Upon applying a rotation to the data, one can use the test we develop in Section 3.2 to solve the problems optimally for rotated (non-axis aligned) ellipsoids or more generally QCO sets (which are still centered around the origin). See our first remark in the discussion section for more details.
Our main result shows that the $\rho_{\text{critical}}$ is determined by three independently valid lower bounds. Moreover, the lower bounds are tight and optimal with only the exception of logarithmic terms. The first lower bound captures the geometric nature of the QCO set $K$ via the Kolmogorov $k$-width. This lower bound is independent of the adversary, and also works in the uncorrupted setting, i.e., when $\epsilon=0$. The second lower bound is obtained within the framework of differential privacy, which is introduced in Section B.3. The third lower bound is derived from a combined argument related to identical mixtures, computation of the total variation and a designed strategy of the adversary. For the upper bounds, we present two algorithms which conclude whether $H_{0}$ holds or not while ensuring type-I and type-II errors both asymptotically less than $\alpha$. The requirements on $\lVert\mu\rVert_{2}$ imposed by the first algorithm are slightly closer to the theoretical lower bounds; however, this improvement comes at the cost of exponential time complexity, rendering the method computationally intractable for large $N$ and $d$. The second algorithm, in contrast, provides a slightly looser upper bound than the first algorithm (though still nearly optimal) while achieving polynomial time complexity. Nevertheless, both upper bounds suggest that the maximum of the three lower bounds discussed previously suffices to fully characterize the behavior of $\rho_{\text{critical}}$. Finally, the results and algorithms derived in the $\ell_{2}$ norm scenario can be naturally extended to the $\ell_{p}$ case for $1\leq p<2$. Section 5 contains such relevant discussions.
1.2 Related Literature
In this section, we discuss some relevant works. We first comment on the literature on minimax testing with uncorrupted data. In their seminal body of work on nonparametric testing, Ingster, (1982); Ingster, 1993a ; Ingster, 1993b ; Ingster, 1993c ; Ingster and Suslina, (2000) showed the difference between testing and estimation in multiple problems. See also Ingster and Suslina, (2003) for a summary of the many existing results in the area. Despite various progress in goodness-of-fit testing in Gaussian models, we are not aware of a general result dedicated to QCO sets. For certain special types of ellipsoids, the problem was first studied by Ermakov, (1991), where the author derived exact minimax rates. Suslina, (1996) studied the asymptotics of the problem of signal detection in $\ell_{q}$ ellipsoids with $\ell_{p}$ ball removed. Since $\ell_{q}$ ellispoids for $q\geq 2$ are QCO sets, when $p=2$ we provide a generalization of these results. For general ellipsoids, Baraud, (2002) studied the testing problem in $\mathbb{R}^{n}$ and $\ell_{2}$ and obtained the minimax optimal critical radius. Our work can hence be seen as a generalization to part of Baraud, (2002). In an interesting recent work, Wei and Wainwright, (2020) studied local testing rates on ellipsoids. Here local means that the point of interest under the null is not necessarily the zero point, but can be an arbitrary point in the ellipsoid. Wei and Wainwright, (2020) proved upper and lower bounds based on localized Kolmogorov widths, which are not guaranteed to match in general, but do match for certain examples with Sobolev ellipsoids. In contrast, the present work focuses on more general sets, but the point of interest under the null hypothesis is always anchored at the zero point, and the critical radius is determined based on the global (instead of local) Kolmogorov widths of the set.
Subsequently, we note that the recent literature on robust testing is relatively sparse. A notable exception is the paper Canonne et al., (2023), which studies a special case of the problem we consider, where $K=\mathbb{R}^{d}$, but considers a broader class of adversaries. The author also developed a polynomial-time algorithm, the framework of which we adapt to develop a more general algorithm suitable for our settings. In addition, Diakonikolas et al., (2017) develops a statistical query (SQ) lower bound for robust estimation problem in high-dimensional setting and extends the results to the testing problem.
1.3 Organization
This paper is structured as follows. Section 2 introduces the notations and several important concepts used through the paper. Section 3 presents the lower and upper bounds for the signal detection problem and contains the main results of the paper. It also contains the discussion of the phase transitions of the problem. Section 4 presents the experimental results of the polynomial time algorithms related to Theorem 3.8 in Section 3. Section 5 contains the discussion which compares the minimax rates of testing to estimation and the extension of the main results to the $\ell_{p}$ norm separation. It also includes some open problems for future works. We attach necessary probabilistic tools and other lemmas in Appendix A. The proofs to the lower and upper bounds and other main corollaries in Section 3 and Section 5 are included in Appendix B. Appendix C presents the pseudo codes of related algorithms.
2 Preliminary
2.1 Notations and Definitions
To simplify the analysis, we assume that $K\subset\mathcal{X}=\mathbb{R}^{d}$, and therefore $\mu\in\mathbb{R}^{d},X_{i}\in\mathbb{R}^{d}$ through the paper. However, one can extend the results to the case when $K\subset\ell_{2}$ with the same techniques and algorithms with minimal efforts given the assumption $\lim\limits_{k\rightarrow\infty}D_{k}(K)=0$. See Definition 2.2 for further details. We use $\rho$ to denote the value of the $\ell_{2}$ norm of $\mu$ under $H_{1}$. $\left\lVert\cdot\right\rVert_{p}$ means the $\ell_{p}$ norm in $\mathbb{R}^{d}$ for $1\leq p\leq\infty$. When $A$ is a set, $\left|A\right|$ denotes the cardinality of $A$. We use calligraphic font $\mathcal{P},\mathcal{Q},\mathcal{R}\dots$ to represent probability distributions. $\mathcal{P}\ll\mathcal{Q}$ means that $\mathcal{P}$ is absolutely continuous with respect to $\mathcal{Q}$. $\text{TV}(\mathcal{P},\mathcal{Q})$ denotes the total variation distance and $\text{KL}(\mathcal{P},\mathcal{Q})$ for the Kullback-Leibler divergence if $\mathcal{P}\ll\mathcal{Q}$. When the underlying probability distribution is not explicitly specified, we use $\mathbb{P}$ alternatively to denote probability law in general. We use $\left\langle\cdot,\cdot\right\rangle$ and $(\cdot)^{\top}(\cdot)$ alternatively (possibly with a subscript when needed) to represent the inner product in a Hilbert space. $\text{diag}\{a\}$ represents the diagonal matrix with diagonal entries $a_{i},i=1,2,\dots$. The convenient notation $[n]$ denotes the set $\left\{1,2,\dots,n\right\}$. When $I$ is an index set, $v_{I}$ represents the restriction of $v$ on $I$ if $v$ is a vector and $\mathbf{Y}_{I}$ represents the sub-matrix of $\mathbf{Y}$ with rows selected according to $I$ if $\mathbf{Y}$ is a matrix. $\mathcal{O}$, $\varTheta$, $\varOmega$ notations are used to denote the asymptotic order of rates. The lowercase $c_{1},c_{2},\dots$ denote the universal constants and their specific values may differ from line to line.
[…]
Definition 2.1 (Quadratically convex orthosymmetric (QCO) set).
Given a set $K\subset\mathbb{R}^{d}$, we say it is a quadratically convex orthosymmetric (QCO) set if it satisfies the following conditions:
 (1). $K$ is convex;
 (2). $K$ is quadratically convex, which means that $K^{2}$ is also convex, where $K^{2}$ is defined as
$K^{2}:=\left\{\left(\theta_{1}^{2},\dots,\theta_{d}^{2}\right)^{\top}\left\lvert\right.\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K\right\};$
(3). $K$ is orthosymmetric, which means that if $\theta=\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K$, then $\theta_{\eta}:=\left(\eta_{1}\theta_{1},\dots,\eta_{d}\theta_{d}\right)^{\top}\\ \in K$, where $\eta_{i}\in\left\{-1,1\right\},1\leq i\leq d$.
Remark.
The definition of QCO set can be naturally extended to the case when $K\subset\ell_{2}$.
Definition 2.2 (Kolmogorov $k$-width).
Let $\mathcal{X}$ be a Banach space equipped with the norm $\left\lVert\cdot\right\rVert$, and $K\subset\mathcal{X}$ is a subset. The Kolmogorov $k$-width is defined as
$D_{k}(K)=\inf_{P\in\mathcal{P}_{k}}\sup_{\theta\in K}\left\lVert\theta-P\theta\right\rVert,$ (2.1)
where $\mathcal{P}_{k}$ is the set of all projection operators that project a vector onto some subspace of $\mathcal{X}$ with intrinsic dimension $k$.
Remark.
In this work, our analysis will be confined to orthogonal projections. Moreover, though $\mathcal{P}_{k}$ is the set of all orthogonal projection with intrinsic dimension $k$, we can discretize this set — only consider the projections that are aligned with the coordinate axes without loss of generality, as shown by the analysis later. For now, let us first assume that the norm leveraged in the definition is the $\ell_{2}$ norm.
Based on the definition, one can derive several basic properties of the Kolmogorov widths. It is non-increasing with respect to $k$, meaning that $D_{k_{1}}(K)\geq D_{k_{2}}(K)$ for any $0\leq k_{1}\leq k_{2}$. In addition, for a finite-dimensional space $\mathcal{X}$ with full dimension $d$, we always have $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}$ and $D_{d}(K)=0$ for any $K\subset\mathcal{X}$. For the more general case where $K\subset\ell_{2}$, we shall assume $\lim\limits_{d\rightarrow\infty}D_{d}(K)=0$. In this paper, we further define $D_{k}(K)=\infty$ when $k<0$ for the convenience of the marginal cases.
The connection between the Kolmogorov widths and our main results is established through the following definition of the first and second dimensions of the optimal projections, which are designed to identify the optimal $k$ satisfying certain marginal conditions determined by the geometric properties of the constraint set and the corruption process.
Definition 2.3 (First optimal dimension and projection).
Given the noise scale $\sigma$, the number of samples $N$ and the prior knowledge $K$, the first optimal dimension is an integer $k_{1}^{\star}(K,\sigma,N)\in\left[0,d\right]$ depending on $K$, $\sigma$, and $N$ where it satisfies that $D_{k_{1}^{\star}-1}(K)>\frac{(k_{1}^{\star})^{1/4}}{\sqrt{N}}\sigma$ but $D_{k_{1}^{\star}}(K)\leq\frac{(k_{1}^{\star}+1)^{1/4}}{\sqrt{N}}\sigma$. The projection operator corresponding to $k_{1}^{\star}(K,\sigma,N)$ is denoted as $P_{1}^{\star}(K,\sigma,N)$.
Definition 2.4 (Second optimal dimension and projection).
Given the noise scale $\sigma$, the number of samples $N$, the prior knowledge $K$ and the corruption rate $\epsilon$, the second optimal dimension is an integer $k_{2}^{\star}(K,\sigma,N,\epsilon)\in[0,d]$ depending on $K$, $\sigma$, $N$, and $\epsilon$ where it satisfies that $D_{k_{2}^{\star}-1}(K)>\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$ but $D_{k_{2}^{\star}}(K)\leq\frac{(k_{2}^{\star}+1)^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$. The projection operator corresponding to $k_{2}^{\star}(K,\sigma,N,\epsilon)$ is denoted as $P_{2}^{\star}(K,\sigma,N,\epsilon)$.
We note that from the monotone property of the Kolmogorov widths and the boundary values, such $k_{1}^{\star}(K,\sigma,N)$ and $k_{2}^{\star}(K,\sigma,N,\epsilon)$ must exist and are unique. In fact, it could be equally defined as
$k_{1}^{\star}(K,\sigma,N)$ $=\max\limits\left\{j\left\lvert\right.0\leq j\leq d,D_{j-1}(K)>\frac{j^{\frac{1}{4}}}{\sqrt{N}}\sigma\right\},$
$k_{2}^{\star}(K,\sigma,N,\epsilon)$ $=\max\limits\left\{j\left\lvert\right.0\leq j\leq d,D_{j-1}(K)>\frac{j^{\frac{1}{4}}\sqrt{\epsilon}}{N^{\frac{1}{4}}}\sigma\right\},$
or
$k_{1}^{\star}(K,\sigma,N)$ $=\min\limits\left\{j\left\lvert\right.0\leq j\leq d,D_{j}(K)\leq\frac{(j+1)^{\frac{1}{4}}}{\sqrt{N}}\sigma\right\},$
$k_{2}^{\star}(K,\sigma,N,\epsilon)$ $=\min\limits\left\{j\left\lvert\right.0\leq j\leq d,D_{j}(K)\leq\frac{(j+1)^{\frac{1}{4}}\sqrt{\epsilon}}{N^{\frac{1}{4}}}\sigma\right\}.$
[…]
Below are several concrete examples of the concepts introduced above, which serve to illustrate the definitions and provide motivation from practical applications.
Example 2.5 ($\mathbb{R}^{d}$).
The whole space $\mathbb{R}^{d}$ is trivially convex, quadratically convex and orthosymmetric. Hence, it is a QCO set. The Kolmogorov $k$-width can be easily calculated as
[…]
The Kolmogorov $k$-width is infinite until $k$ reaches $d$, which implies one cannot uniformly approximate a vector in $\mathbb{R}^{d}$ by a linear subspace well without any additional constraint. The first and second optimal dimensions are $k_{1}^{\star}=k_{2}^{\star}=d,\forall\epsilon\in[0,c_{0})$.
Example 2.6 (Hyperrectangles in $\mathbb{R}^{d}$).
The hyperrectangle in $\mathbb{R}^{d}$ is defined as
$H_{d}(a_{1},\dots,a_{d}):=\left\{x\left\lvert\right.x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d},\left|x_{i}\right|\leq a_{i},a_{i}>0,1\leq i\leq d\right\}.$
Obviously, $H_{d}(a_{1},\dots,a_{d})$ is convex and orthosymmetric. One can verify that $H_{d}^{2}(a_{1},\dots,a_{d})=\left\{x^{2}\left\lvert\right.x\in H_{d}(a_{1},\dots,a_{d})\right\}$ is also a hyperrectangle with each side being $0,a_{i}^{2}],1\leq i\leq d$. Therefore, it is a QCO set. By reindexing if necessary, we may assume that the sequence $\left\{a_{i}\right\}_{i=1}^{d}$ is arranged in decreasing order. The Kolmogorov $k$-width is
[…]
The decay rate of the Kolmogorov widths depends on the specific choice of $a_{i}$. The first optimal dimension is $k_{1}^{\star}=\max\limits\left\{j\left\lvert\right.1\leq j\leq d,\sqrt{\sum\limits_{i=j}^{d}a_{i}^{2}}>\frac{j^{1/4}}{\sqrt{N}}\sigma\right\}$ when $\sqrt{\sum\limits_{i=1}^{d}a_{i}^{2}}>\frac{1}{\sqrt{N}}\sigma$ and otherwise 0. The second optimal dimension is $k_{2}^{\star}=\max\limits\left\{j\left\lvert\right.1\leq j\leq d,\sqrt{\sum\limits_{i=j}^{d}a_{i}^{2}}>\frac{j^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma\right\}$ when $\sqrt{\sum\limits_{i=1}^{d}a_{i}^{2}}>\frac{\sqrt{\epsilon}}{N^{1/4}}\sigma$ otherwise 0.
Example 2.7 (Ellipses in $\mathbb{R}^{d}$).
An ellipse in $\mathbb{R}^{d}$ that is centered at the origin and aligns with the axes can be represented as
$E_{d}(a_{1},\dots,a_{d})=\left\{x\left\lvert\right.x\in\mathbb{R}^{d},\sum\limits_{i=1}^{d}\frac{x_{i}^{2}}{a_{i}}\leq 1\right\}.$
Any such ellipse is convex and orthosymmetric. The square of the ellipse $E_{d}(a_{1},\dots,a_{d})$ is
$E_{d}^{2}(a_{1},\dots,a_{d})=\left\{x\left\lvert\right.x\in\mathbb{R}_{+}^{d},\sum\limits_{i=1}^{d}\frac{x_{i}}{a_{i}}\leq 1\right\},$
which is part of a simplex (enclosed by several hyperplanes) in $\mathbb{R}^{d}$, and therefore convex. Figure 2 presents a depiction of $E_{d}^{2}(a_{1},\dots,a_{d})$ when $d=2$.
Figure 2: An example of an ellipse and its square-set in $\mathbb{R}^{2}$. The square-set is a triangle in $\mathbb{R}^{2}$, which is convex.
Similarly, we assume $\left\{a_{i}\right\}$ is arranged in decreasing order, by which the Kolmogorov widths are
[…]
The first optimal dimension is $k_{1}^{\star}=\max\limits\left\{j\left\lvert\right.1\leq j\leq d,a_{j}>\frac{\sqrt{j}}{N}\sigma^{2}\right\}$ when $a_{1}>\frac{1}{N}\sigma^{2}$ and 0 otherwise. The second optimal dimension $k_{2}^{\star}=\max\limits\left\{j\left\lvert\right.1\leq j\leq d,a_{j}>\frac{\sqrt{j}\epsilon}{\sqrt{N}}\sigma^{2}\right\}$ when $a_{1}>\frac{\epsilon}{\sqrt{N}}\sigma^{2}$ otherwise 0.
When $a_{i}=i^{-2s}$ for some $s>1$, the corresponding ellipse in $\mathbb{R}^{d}$ can be regarded as a special case of the more general ellipses of Sobolev types, which is defined as $S(s):=\left\{x\left\lvert\right.x\in\mathbb{R}^{\infty},\sum\limits_{i=1}^{\infty}i^{2s}x_{i}^{2}\leq 1\right\}$. This type of sets is closely related to the estimation and density approximation problem with the class of $s$-times differentiable functions in probability space. We refer readers to Wainwright, (2019) and Tsybakov, (2009) for more details.
Remark.
When a ellipse is not axes-aligned, (1), the orthosymmetric property is not guaranteed, which can be easily checked; (2), the quadratically convex property is true for $d=2$ and false for $d\geq 3$. To gain the idea of (2), consider an ellipse
$E_{d}(A)=\left\{x\left\lvert\right.x\in\mathbb{R}^{d},x^{\top}A^{-1}x\leq 1\right\},$
and its squared set denoted by $E_{d}^{2}(A)$. For any point $y$ in $E^{2}_{d}(A)$, by definition there exists a sign set $\gamma\in\left\{-1,1\right\}^{d}$ such that $\left(\gamma_{1}\sqrt{y_{1}},\dots,\gamma_{d}\sqrt{y_{d}}\right)\in E_{d}(A)$. Therefore, $E_{d}^{2}(A)$ can be expressed by
$E_{d}^{2}(A)=\left\{y\left\lvert\right.y\in\mathbb{R}_{+}^{d},\min\limits_{\gamma\in\left\{-1,1\right\}^{d}}\left\{\sum\limits_{i,j}A_{ij}\gamma_{i}\gamma_{j}\sqrt{y_{i}y_{j}}\right\}\leq 1\right\}.$
When $d=2$, by selecting $\gamma_{1}$ and $\gamma_{2}$ properly, $(y_{1},y_{2})\in\mathbb{E}_{d}^{2}(A)$ if and only if
$A_{11}y_{1}+A_{22}y_{2}-2\left|A_{12}\right|\sqrt{y_{1}y_{2}}\leq 1.$
Therefore, $E_{2}^{2}(A)$ is convex since it is a sublevel set of a convex function. However, for $d\geq 3$, it is impossible to let all coefficients of $\sqrt{y_{i}y_{j}}$ be negative, which leads to the non-convexity of the LHS.
Finally, we conclude this section with the definition of the strong $\epsilon$-contamination model studied in this paper.
Definition 2.8 (Strong $\epsilon$-contamination model).
[…]
3 Main Results
3.1 Lower Bounds
In this section, we introduce three lower bounds, i.e., Theorem 3.2, 3.3, and 3.4, which independently work on determining the lower bound of $\rho_{\text{critical}}$.
3.1.1 Based on a Geometric Property
The following lemma is an important basis for the theoretical lower bounds in Theorems 3.2 and 3.3. Exploiting the properties of QCO sets and the definition of the Kolmogorov widths, Lemma 3.1 outputs a vector with large $\ell_{2}$ norm but small $\ell_{\infty}$ norm, which is fundamental for the prior of $v$ constructed under $H_{1}$. The proof, which involves an application of the minimax theorem, is provided in Appendix B.1.
Lemma 3.1.
Let $K\subset\mathcal{X}$ be a QCO set and $c$ be a positive constant. Suppose that $D_{k-1}(K)>c\sigma$ for some $k\geq 1$. Then there exists a vector $\theta\in K$ such that $\left\lVert\theta\right\rVert_{2}=c\sigma$ while $\left\lVert\theta\right\rVert_{\infty}\leq\frac{c}{\sqrt{k}}\sigma$.
We now are ready to state the first main lower bound.
Theorem 3.2.
If $\rho\leq c(\alpha)\cdot\frac{(k_{1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, where $c$ is a constant that only depends on $\alpha$, or $k_{1}^{\star}=0$, then
$\inf_{\psi:\mathbb{P}_{0}(\psi=1)\leq\alpha}\sup_{\theta\in K,\|\theta\|\geq\rho}\mathbb{P}_{\theta}(\psi=0)\geq\alpha.$
Moreover, one possible explicit expression of $c(\alpha)$ is
$c(\alpha)=\left[2\ln\left(1+4\left(1-2\alpha\right)^{2}\right)\right]^{\frac{1}{4}}.$
Theorem 3.2 above states that for fixed $\alpha$, if $\rho\lesssim\frac{(k_{1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, it is impossible to find a test that performs well under $H_{0}$ and all possible $H_{1}$ uniformly. As a corollary, this result implies that it is also impossible to distinguish between $H_{0}$ and $H_{1}$ in the presence of an adversary $\mathcal{C}$ under the above condition, since a simple strategy for $\mathcal{C}$ is to keep all samples untouched. The proof of Theorem 3.2 relies on the selected vector in Lemma 3.1 and Le Cam’s method of “fuzzy hypothesis”.
Moreover, it is worth noting the marginal case when $k_{1}^{\star}=0$. The definition of $k_{1}^{\star}$ only relies on $K$, NN, $\sigma$, and does not depend on the original samples $\widetilde{\mathbf{X}}$ or the observations $\mathbf{X}$. Therefore, we shall assume in the following text that $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}>\frac{1}{\sqrt{N}}\sigma$, and therefore $k_{1}^{\star}\geq 1$. Otherwise, the test will not be feasible in the first place. We kindly refer the readers to Appendix B.2 for the proof of this argument and Theorem 3.2.
Recall that for uncontaminated observations without constraint, the critical value of the $\ell_{2}$ norm is $\rho_{\text{critical}}\asymp\frac{d^{1/4}}{\sqrt{N}}\sigma$ (see Section 3.3.6 in Ingster and Suslina, (2003) for more details). Theorem 3.2 generalizes this classic result in the sense that additional prior knowledge the mean is allowed and the lower bound is refined as a consequence. Indeed, Theorem 3.2 recovers the classic result when we set $\mathcal{X}=\mathbb{R}^{d}$ (where $k_{1}^{\star}=d$ according to Example 2.5). Moreover, the theorem implies that when we have non-trivial constraint $K$ on $\mu$, the essential dimensional attribute of the problem is the finite $k_{1}^{\star}$ instead of $d$. (Recall that $k_{1}^{\star}<\infty$ is ensured even if $K\subset\ell_{2}$.)
3.1.2 Based on the Framework of Differential Privacy
The concept of differential privacy was first introduced by Dwork et al., (2006) in the study of private data analysis. It was later introduced to the goodness-of-fit test (also called identity test) in statistics. See Gaboardi et al., (2016), Rogers and Kifer, (2017), Cai et al., (2017) for some earlier works of this application. Differential privacy requires a random algorithm to be robust against the disturbance of the input. Please see Section B.3 for the formal definition of differential privacy. In Theorem 3.3, we will combine Lemma 3.1 with an argument of optimal coupling and results from differential privacy to obtain another requirement for $\rho$. Please refer to Section B.3 for the techniques and proofs.
Theorem 3.3.
For the testing problem 1.2 with an adversary $\mathcal{C}$ from strong $\epsilon$-contamination model 2.8, it is impossible to derive a test that achieves type-I error and type-II error both less than 0.1 if $\rho\lesssim\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}$. Formally, if $\rho\lesssim\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}$, then we have
$\inf\limits_{\phi\in A_{s}(0.1)}\sup\limits_{\mu\in K}\sup\limits_{\mathcal{C}}\mathbb{P}_{\mu}(\phi(\mathcal{C}(\mathbf{X}^{\prime}))=0)\geq 0.1,$
where $A_{s}(0.1)$ is the set of acceptable tests under $0.1$ type-I error defined as (1.3).
We use $0.1$ for convenience, but it can be replaced by any unversal sufficiently small constant. It is worth noting that in the high-dimensional setting, where $d$, and possibly $k_{2}^{\star}$ can be substantially larger than $N$, one should be cautious that $\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}$ shall not exceeds $D_{0}(K)$, which is the maximal possible $\ell_{2}$ norm of the vector in $K$. The desired test will be impossible if this is not achieved.
3.1.3 Based on the Estimation Lower Bounds
The final lower bound, as established in Theorem 3.4 below, contrasts with Theorems 3.2 and 3.3 in that it is determined solely on the diameter of $K$ (which is $D_{0}(k)$), the fraction of contamination $\epsilon$, and the scale of the noise $\sigma$ while remaining independent of $N$.
Theorem 3.4.
For the signal detection problem (1.2), if $\epsilon\gtrsim\frac{1}{\sqrt{N}}$ and $\rho\lesssim\epsilon\sigma$, we always have
$\inf\limits_{\phi\in A_{s}}\sup\limits_{\mu\in K,\left\lVert\mu\right\rVert_{2}\geq\rho}\sup\limits_{\mathcal{C}}\mathbb{P}_{\mu}(\phi(\mathcal{C}(\widetilde{\mathbf{X}}))=0)\gtrsim 1,$ (3.1)
where $A_{s}$ is the set of acceptable tests defined as (1.3).
Remark.
The condition $\epsilon\gtrsim\frac{1}{\sqrt{N}}$ can be safely removed. See our discussions in Section 3.3.
Recall that $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}$ is the supremum of the distance between a point in $K$ and the origin. (3.1) implies that to find a test with type-I and type-II error uniformly (over the adversary $\mathcal{C}$ and the mean $\mu$) less than $\alpha$, we should have either $\rho\gtrsim\epsilon\sigma$ or $\epsilon=\mathcal{O}(1/\sqrt{N})$. Since we always have $\left\lVert\mu\right\rVert_{2}\leq D_{0}(K)$, consequently, if $\epsilon\gtrsim 1/\sqrt{N}$ and $\epsilon\sigma\gtrsim D_{0}(K)$, the desired test will be theoretically impossible.
The proof of Theorem 3.4 is motivated by a similar lower bound established for the corresponding estimation problem under star-shaped constraints by Prasadan and Neykov, (2025). It is based on a careful construction of a mixture distribution which is represented in two different ways, and a computation of the total variation between them. The full proof is provided in Section B.4. This bound aligns with intuition, as testing is generally easier than estimation. See Section 5 for further details.
3.2 Upper Bounds
3.2.1 Theoretical Algorithm
An intuitive strategy for handling the adversary is to identify subsets of $\mathbf{X}$ that exhibit a consistency property. As we will see later, such subsets behave as if they consisted entirely of authentic samples with high probability.
Definition 3.5 (Consistent subset).
A subset $S\subset\mathbf{X}=\left\{X_{1},\dots,X_{N}\right\}$ is called a consistent subset regarding the contamination fraction $\epsilon$ and a test $\phi:2^{S}\mapsto\left\{0,1\right\}$ if
 (i), $\left|S\right|\geq(1-\epsilon)N$,
 (ii), $\phi(S^{\prime})=\phi(S)$ for any $S^{\prime}\subset S$ with $\left|S^{\prime}\right|\geq(1-2\epsilon)N$.
The following theorem establishes the existence of a valid testing procedure under condition (3.2). Specifically, we demonstrate that, when (3.2) holds, there exists a desired test $\phi_{e}$ and at least one consistent subset $S$ with respect to $\phi_{e}$, upon which the entire testing task becomes feasible. To achieve this, we explicitly define an event $E_{e}$ of $\chi^{2}$-type, on which $\phi_{e}$ is determined. Furthermore, we prove that the set of original samples $\widetilde{\mathbf{X}}$, though not observed, is consistent with high probability. Finally, we show that the proposed test $\phi_{e}$ achieves uniformly small type-I and type-II errors. The detailed proof is provided in Section B.5.
Theorem 3.6.
Given a dataset $\mathbf{X}$ from the strong $\epsilon$-contamination model (see Definition 2.8) and a fixed error tolerance $\alpha$, for any $k\in[1,d]$ [^2] When $K\subset\ell_{2}$, we allow $1\leq k<\infty$., there exists a test $\phi_{e}$ for (1.2) such that the type-I and type-II errors are both uniformly less than $\alpha$ if $\rho^{2}\gtrsim E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k)$ where $E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k)$ is defined as
$E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k):=D_{k}^{2}(K)+\sigma^{2}\max\limits\left\{\frac{\sqrt{k}}{N},\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right),\sqrt{\frac{\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)k}{N}}\right\}.$ (3.2)
In other words, we have
$\rho_{\text{critical}}\lesssim\min\limits_{k}E_{\text{raw}}(\epsilon,K,N,\sigma^{2},k).$
We can further leverage the results of Theorem 3.6 by the definition of $k_{1}^{\star}$ and $k_{2}^{\star}$ and optimizing over $k$.
Corollary 3.7.
Given the same setting as Theorem 3.6, we can alternatively define $E(\epsilon,K,N,\sigma^{2})$, which is also a valid upper bound of $\rho_{\text{critical}}$, as
$E^{2}(\epsilon,K,N,\sigma^{2}):=\sigma^{2}\max\limits\left\{\frac{\sqrt{\min\limits\left\{k_{1}^{\star},k_{2}^{\star}\right\}}}{N},\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right),\sqrt{\frac{\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)\min\limits\left\{k_{1}^{\star},k_{2}^{\star}\right\}}{N}}\right\}$ (3.3)
Proof.
By the definition of $k_{1}^{\star}$, we know $D_{k_{1}^{\star}}(K)\leq\frac{\sqrt{k_{1}^{\star}+1}}{N}\sigma^{2}$. Therefore, if we set $k=k_{1}^{\star}$ in (3.2), $D_{k_{1}^{\star}}(K)$ can be absorbed into the other term involving the maximum. The logic also holds for $k_{2}^{\star}$. Finally, since (3.2) is always a valid upper bound, we select the smaller one, which leads to (3.3). ∎
Corollary 3.7 actually selects the smaller optimal dimension between $k_{1}^{\star}$ and $k_{2}^{\star}$ and consequently the corresponding optimal projection, based on which $\phi_{e}$ is performed. We will show in Section 3.3 that the upper bound (3.3) does nearly match the theoretical lower bounds in the sense that it only has a redundant $\ln\left(\frac{1}{\epsilon}\right)$ factor.
A significant improvement implied by Theorem 3.6 is that it does not require $\epsilon\rightarrow 0$, which is an essential assumption in many previous works. However, the test $\phi_{e}$ is limited in that its construction requires a full scan of all subsets of $\mathbf{X}$, which consumes exponential time with respect to $N$. This prohibitive computational cost prevents it from being practical when $N$ is large. In the next section, motivated by the filtering method given by Canonne et al., (2023), we propose an algorithm with polynomial-degree time complexity when the prior constraint $K$ and adversary $\mathcal{C}$ present.
3.2.2 Polynomial Time Algorithm
We will develop a polynomial time algorithm [^3] Here when we say “polynomial time algorithm”, we do assume that calculating the optimal projections (resulting in the optimal dimensions $k_{1}^{\star}$ and $k_{2}^{\star}$) can be done in polynomial time — as is the case for any ellipsoid or hyperrectangle. in this section to detect whether $\mu=0$ or not while requiring a slightly stronger condition compared to the bound (3.2). Specifically, we will prove the following theorem.
Theorem 3.8.
Assuming the same conditions as in Theorem 3.6, for any $1\leq k\leq d$ [^4] When $K\subset\ell_{2}$, we allow $1\leq k<\infty$., as long as the following condition is satisfied, there exists a test $\phi_{p}$ to distinguish between $H_{0}$ and $H_{1}$ with type-I and type-II errors both uniformly less than $\alpha$.
$\rho^{2}\gtrsim\underbrace{D_{k}^{2}(K)+\sigma^{2}\max\limits\left\{\frac{\epsilon\ln\left(\frac{N}{\alpha}\right)}{\sqrt{N}},\epsilon^{2}\ln\left(\frac{N}{\alpha}\right),\sqrt{\frac{\epsilon^{2}k\ln\left(\frac{N}{\alpha}\right)}{N}},\frac{\sqrt{k}\ln\left(\frac{1}{\alpha}\right)}{N}\right\}}_{:=P_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k)}.$ (3.4)
In other words, we have
$\rho_{\text{critical}}\lesssim\min\limits_{k}P_{\text{raw}}(\epsilon,K,N,\sigma^{2},k)$
In addition, the time and space complexity of the algorithm is at most polynomial in terms of the parameters $N,k,\epsilon,\frac{1}{\alpha}$.
Similar to Theorem 3.6, the following corollary characterizes the upper bound (3.4) with $k_{1}^{\star}$ and $k_{2}^{\star}$. The proof is omitted since it follows from Corollary 3.7 with minor modifications.
Corollary 3.9.
Given the same setting as Theorem 3.8, we can alternatively define $P(\epsilon,K,N,\sigma^{2})$, which is also a valid upper bound of $\rho_{\text{critical}}$, as
$P^{2}(\epsilon,K,N,\sigma^{2})=\sigma^{2}\max\limits\left\{\frac{\epsilon\ln\left(\frac{N}{\alpha}\right)}{\sqrt{N}},\epsilon^{2}\ln\left(\frac{N}{\alpha}\right),\sqrt{\frac{\epsilon^{2}\min\limits\left\{k_{1}^{\star},k_{2}^{\star}\right\}\ln\left(\frac{N}{\alpha}\right)}{N}},\frac{\sqrt{\min\limits\left\{k_{1}^{\star},k_{2}^{\star}\right\}}\ln\left(\frac{1}{\alpha}\right)}{N}\right\}.$ (3.5)
Analogous to the upper bound (3.3), (3.5) also nearly matches the theoretical lower bounds with only the exception of logarithmic terms. However, one can easily check that the upper bound (3.3) is generally tighter than (3.5), i.e., $E(\epsilon,K,N,\sigma^{2})\leq P(\epsilon,K,N,\sigma^{2})$.
We must note here that our proof of Theorem 3.8 and the algorithm are inspired by the approach undertaken by Canonne et al., (2023). However, the algorithms and results of Canonne et al., (2023) only work in the region when $\left\lVert\mu\right\rVert_{2}\lesssim 1$ and there is no constraint under the alternative, i.e., $K=\mathbb{R}^{d}$. Therefore, we crucially modified the algorithm, including techniques to allow for $\mu$ of arbitrary (Euclidean) length and the use of the projection $P^{\star}$ to allow for the prior knowledge on $\mu$. We also fixed some flaws left in the argument of Canonne et al., (2023). We kindly refer readers to Appendix B.6 for the complete proof of Theorem 3.8.
3.3 Main Results
In this section, we integrate the preceding theorems and bounds and summarize the main results. The lower bound is synthesized from Theorems 3.2, 3.3, 3.4. Although the bounds are derived through different approaches, an interesting phase transition phenomenon emerges, determined by the ranges of $\epsilon$ over which each term dominates others.
Theorem 3.10 (Main lower bound).
Given the mean testing problem (1.2) in the Gaussian sequence model, with potential contamination from a strong adversary $\mathcal{C}$ and a prior constraint on the mean $\mu$ induced by a QCO set $K$, the following condition is necessary to ensure the existence of a valid test whose type-I and type-II errors are both uniformly below a prescribed constant $\alpha$
$\rho^{2}\gtrsim\sigma^{2}\max\limits\left\{\frac{\sqrt{k_{1}^{\star}}}{N},\epsilon\sqrt{\frac{k_{2}^{\star}}{N}},\epsilon^{2}\right\},$ (3.6)
where $N$ is the sample size, $\sigma^{2}$ is the variance of the noise, $\epsilon$ is the fraction of contamination, and $k_{1}^{\star},k_{2}^{\star}$ are the optimal dimensions exclusively determined by $N$, $\sigma$, $\epsilon$ and the set $K$ as in 2.3 and 2.4.
In the proof of Corollary 3.12 below, we show that the first term, which is independent with $\epsilon$ [^5] We should recall that $k_{1}^{\star}$ is independent with $\epsilon$, dominates the other two if and only if $\epsilon\leq\frac{1}{\sqrt{N}}$. Similarly, the last term, which is independent with $N$, is the dominant if and only if $\epsilon\gtrsim\sqrt{\frac{k_{2}^{\star}}{N}}$. Therefore, the set $\left\{\epsilon\left\lvert\right.0\leq\epsilon<c_{0},\epsilon\in\left[\frac{1}{\sqrt{N}},\sqrt{\frac{k_{2}^{\star}}{N}}\right]\right\}$ [^6] We express in this form because $k_{2}^{\star}$ also depends on $\epsilon$. serves as a threshold belt. When $\epsilon$ lies below this belt, it is too small for the adversary $\mathcal{C}$ to exert its influence, and the optimal minimax rate becomes essentially independent of $\epsilon$. In other words, it is free to contaminate the dataset in this region. In contrast, if $\epsilon$ lies above the belt (for example, $\epsilon\asymp N^{-\frac{1}{4}}$), then the power of $\mathcal{C}$ can be fully utilized, to the extent that increasing the sample size $N$ does not improve the problem’s difficulty. When $\epsilon$ falls within the belt, the minimax rate is the result of the delicate balance between the adversary $\mathcal{C}$ and the testing procedures $\phi_{e},\phi_{p}$. Based on this argument, (3.6) can be equivalently written as
$\rho\gtrsim\left\{\begin{tabular}[]{lll}$\frac{\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$,&$\left(0\leq\epsilon\lesssim\frac{1}{\sqrt{N}}\right)$&(Free contamination phase),\\ $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}$,&$\left(\frac{1}{\sqrt{N}}\lesssim\epsilon\lesssim\sqrt{\frac{k_{2}^{\star}}{N}}\right)$&(Balance phase),\\ $\epsilon^{2}\sigma^{2}$,&$\left(\epsilon\gtrsim\sqrt{\frac{k_{2}^{\star}}{N}}\right)$,&(Overly contaminated phase).\end{tabular}\right$. (3.7)
Informally, a message implied by (3.7) is that after the “safe threshold” $\frac{1}{\sqrt{N}}$, every single contaminated sample in the strong $\epsilon$-contamination model invalidates the statistical testing use of $\sqrt{\frac{N}{k_{2}^{\star}}}$ uncontaminated samples on average, until the overly contaminated phase.
For the upper bound, the theorem below is a natural downstream conclusion of Theorem 3.6 and 3.8.
Theorem 3.11 (Main upper bound).
Given the same setting as in Theorem 3.10, if $\rho\gtrsim E(\epsilon,K,N,\\ \sigma^{2})$, there exists a test $\phi_{e}$ achieving type-I and type-II errors uniformly less than $\alpha$. Furthermore, if $\rho\gtrsim P(\epsilon,K,N,\sigma^{2})$ and the determination of $k_{1}^{\star}$ and $k_{2}^{\star}$ can be completed within polynomial time of $\left(N,d,\frac{1}{\epsilon}\right)$, there exists a test $\phi_{p}$ that is able to finish the testing process within polynomial time of $\left(N,d,\frac{1}{\epsilon},\frac{1}{\alpha}\right)$ while still achieving type-I and type-II errors uniformly less than $\alpha$. $E(\epsilon,K,N,\sigma^{2})$ and $P(\epsilon,K,N,\sigma^{2})$ are defined in (3.3) and (3.5) respectively.
We further establish that the main lower bound in Theorem 3.10 is minimax optimal and there is only minor difference between the conditions required by the two algorithms in Theorem 3.11. This is summarized in the following corollary.
Corollary 3.12 (Nearly matching between the lower bounds and the upper bounds).
The upper bound $E(\epsilon,K,N,\sigma^{2})$ nearly matches the theoretical lower bound (3.6) with only an exception of the logarithmic factor $\ln\left(\frac{1}{\epsilon}\right)$. The upper bound $P(\epsilon,K,N,\sigma^{2})$ of $\phi_{p}$ nearly matches $E(\epsilon,K,N,\sigma^{2})$ with only exceptions of the logarithmic factors $\ln N$, $\ln\alpha$, and $\ln\left(\frac{1}{\epsilon}\right)$.
For Corollary 3.12, it involves some refined analysis with (3.6) and $E(\epsilon,K,N,\sigma^{2})$. We kindly refer readers to Appendix B.7 for the detailed arguments.
3.4 Extension to $\ell_{p}$ Norm Separation for $1\leq p<2$
As an alternative, with the same assumptions on the samples $\widetilde{\mathbf{X}}$ and the adversary $\mathcal{C}$, we may consider the following $\ell_{p}$ robust testing problem.
$H_{0}:\mu=0,$ (3.8)
$H_{1}:\left\lVert\mu\right\rVert_{p}\geq\rho,\mu\in K.$
The classic results for the testing problem (1.2) can be naturally generalized to (3.8) when $1\leq p<2$ via the Hölder’s inequality (see Lemma A.17) by Ingster and Suslina, (2003). While Ingster’s approach cannot be directly transferred to our settings, we establish similar generalized results with the $p$-convex condition and accordingly adjusted optimal dimensions and projections. Note that in this section, we shall substitute the $\ell_{2}$ norm with the $\ell_{p}$ norm for the Kolmogorov $k$-width. Furthermore, to simplify the analysis, we consider the discretized version (see the remark after Definition 2.2).
Definition 3.13 (Discretized Kolmogorov $k$-Width).
Let $\mathcal{X}$ be a Banach space equipped with the norm $\left\lVert\cdot\right\rVert$, and $K\subset\mathcal{X}$ is a subset. The discretized Kolmogorov $k$-width is defined as
$\widetilde{D}_{k}(K)=\inf_{P\in\widetilde{\mathcal{P}}_{k}}\sup_{\theta\in K}\left\lVert\theta-P\theta\right\rVert,$ (3.9)
where $\widetilde{\mathcal{P}}_{k}$ is the set of all projection operators that align with the axes and project a vector onto some subspace of $\mathcal{X}$ with intrinsic dimension $k$.
Though discretized, it inherits most of important properties from the original definition. When the norm is $\left\lVert\cdot\right\rVert_{p}$, let us denote the corresponding Kolmogorov $k$-width for $K$ as $\widetilde{D}_{p,k}(K)$.
Definition 3.14 ($p$-convex set).
For $p\geq 1$, a set $K\subset\mathcal{X}$ is $p$-convex if
 (1). $K$ is convex;
 (2). $K^{p}$ is convex, where $K^{p}$ is defined as
$K^{p}:=\left\{\left(\left|\theta_{1}\right|^{p},\dots,\left|\theta_{d}\right|^{p}\right)^{\top}\left\lvert\right.\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K\right\}.$
Analogous to previous analysis with the $\ell_{2}$ norm, the new lower bound relies on the following revised version of Lemma 3.1. It is not difficult to check that the proof is not essentially different.
Lemma 3.15.
Let $K\subset\mathcal{X}$ be a $p$-convex orthosymmetric set, and suppose that $\widetilde{D}_{p,k-1}(K)>c\sigma$. Then there exists a vector $\theta\in K$ such that $\left\lVert\theta\right\rVert_{p}=c\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{c}{k^{1/p}}\sigma$.
We also require the following new optimal dimensions and projections for $\ell_{p}$ robust testing problem.
Definition 3.16 ($\ell_{p}$ optimal dimensions and projections).
Given the noise scale $\sigma$, the number of samples $N$, the prior knowledge $K$ and the corruption rate $\epsilon$, define the following optimal dimensions
$k_{p,1}^{\star}=\max\limits\left\{j\left\lvert\right.0\leq j\leq d,\widetilde{D}_{p,j}(K)>\frac{j^{\frac{1}{p}-\frac{1}{4}}}{\sqrt{N}}\sigma\right\},$
$k_{p,2}^{\star}=\max\limits\left\{j\left\lvert\right.0\leq j\leq d,\widetilde{D}_{p,j}(K)>\frac{\sqrt{\epsilon}j^{\frac{1}{p}-\frac{1}{4}}}{N^{\frac{1}{4}}}\sigma\right\},$
$k_{p,3}^{\star}=\max\limits\left\{j\left\lvert\right.0\leq j\leq d,\widetilde{D}_{p,j}(K)>\epsilon\sigma j^{\frac{1}{p}-\frac{1}{2}}\right\}.$
Analogous to Definition 2.3 and 2.4, the corresponding optimal projections are defined accordingly and denoted as $P_{p,1}^{\star},P_{p,2}^{\star},P_{p,3}^{\star}$.
Given the modified lemma and definitions above, we present the lower bounds for the $\ell_{p}$ robust testing problem (3.8).
Theorem 3.17 ($\ell_{p}$ main lower bound).
Given the mean testing problem (3.8) in the Gaussian sequence model, with potential contamination from a strong adversary $\mathcal{C}$ and a prior constraint on the mean $\mu$ induced by a p-convex orthosymmetric set $K$, the following condition is necessary to ensure the existence of a valid test whose type-I and type-II errors are both uniformly below a prescribed constant $\alpha$
$\left\lVert\theta\right\rVert_{p}\gtrsim\sigma\max\limits\left\{\frac{(k_{p,1}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{\sqrt{N}},\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{N^{\frac{1}{4}}},\epsilon(k_{p,3}^{\star})^{\frac{1}{p}-\frac{1}{2}}\right\}.$ (3.10)
where $N$ is the sample size, $\sigma^{2}$ is the variance of the noise, $\epsilon$ is the fraction of contamination, and $k_{p,1}^{\star},k_{p,2}^{\star},k_{p,3}^{\star}$ are the optimal dimensions exclusively determined by $N$, $\sigma$, $\epsilon$ and the set $K$ as in 3.16.
Undoubtedly, the lower bound (3.10) recovers (3.6) as expected when $p=2$. A significant difference from the $\ell_{2}$ norm separation is that we have another optimal dimension regarding the last term in (3.10). The intrinsic reason is that when covariance matrices are known and identical, the KL divergence between Gaussian distributions — an essential quantity we leverage in the proof of Theorem 3.4, depends on the $\ell_{2}$ norm between the means. When considering the $\ell_{p}$ norm separation, we require $k_{p,3}^{\star}$ to adapt to this discrepancy. We defer the proof of Theorem 3.17 to Appendix B.8.
In $\ell_{2}$ norm separation, the upper bound (3.3) matches the lower bound (3.6) except for logarithmic factors. We identify similar results in $\ell_{p}$ norm separation, which is summarized as the theorem below.
Theorem 3.18 ($\ell_{p}$ main upper bound).
If $\rho\gtrsim\sigma\max\limits\left\{\frac{(k_{p,1}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{\sqrt{N}},\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{N^{\frac{1}{4}}},\epsilon(k_{p,3}^{\star})^{\frac{1}{p}-\frac{1}{2}}\right\}$, then there exists a test to distinguish between $H_{0}$ and $H_{1}$ with uniformly small type-I and type-II errors. Moreover, the time complexity of the algorithm is polynomial degree in $\left(N,d,\frac{1}{\epsilon},\frac{1}{\alpha}\right)$. Note that in the expression above, we omit the polylogarithmic factors in $\left(N,k_{p,1}^{\star},k_{p,2}^{\star},k_{p,3}^{\star},\frac{1}{\alpha},\frac{1}{\epsilon}\right)$.
At first examination, Theorem 3.18 appears to be a direct and natural corollary of Theorem 3.11. However, a closer inspection reveals subtleties in its derivation. A crucial observation is that, though one can select the minimum $k_{p}^{\star}:=\min\limits\left\{k_{p,1}^{\star}k_{p,2}^{\star},k_{p,3}^{\star}\right\}$ and its optimal projection $P_{p}^{\star}$, and is guaranteed that there exist three vectors $\theta_{p,1},\theta_{p,2},\theta_{p,3}$ such that $\left\lVert P_{p}^{\star}\theta_{p,1}\right\rVert_{2}^{2}\gtrsim\frac{\sqrt{k^{\star}}}{N}\sigma^{2},\left\lVert P_{p}^{\star}\theta_{p,2}\right\rVert_{2}^{2}\gtrsim\epsilon\sqrt{\frac{k_{p}^{\star}}{N}}\sigma^{2},\left\lVert P_{p}^{\star}\theta_{p,3}\right\rVert_{2}^{2}\gtrsim\epsilon^{2}\sigma^{2}$, $\theta_{p,1},\theta_{p,2},\theta_{p,3}$ are not necessary to be identical. In Theorem 3.6 or Theorem 3.8, however, we require a single vector satisfying all the conditions simultaneously. In fact, similar to the $\ell_{2}$ norm case, we do need to select the minimal optimal dimension. To overcome the difficulty specified above, we again need to exploit the properties of the $p$-convex orthosymmetric set and the Kolmogorov widths. The proof of Theorem 3.18 is provided in Appendix B.9.
4 Experiments
In this section, we present the results of simulation to verify the correctness of the theorems above and algorithms in Appendix C on the application side. We will generate artificial data, design different strategies for the adversary to contaminate the data, and run the (polynomial time) algorithm to check its effectiveness. The pseudo code is attached in Appendix C.
4.1 Data Generation
We generate artificial data using the function numpy.random.multivariate in the Numpy package in Python. To fully verify the correctness of the algorithms, we consider both the classical and high-dimensional regimes separately, and explore a variety of parameter configurations. Specifically, in the classical scenario, we let the parameter tuple $\left(N,d,\epsilon\right)$ iterate over all possible combinations of $\left\{200,2000\right\}\times\left\{5,10,20,50\right\}\times\left\{0.01,0.015,0.02,0.025\right\}$; in the high-dimensional case, we let the tuple iterate over all possible combinations of $\left\{100,500\right\}\times\left\{100,500\right\}\times\left\{0.01,0.015,0.02,0.025\right\}$. We set $\sigma=1$ in all generation procedures.
4.2 Choice of Constraints
For the choice of constraints, i.e., $K$, we select the common ellipsoidal constraint:
$\mu\in K_{e}:=\left\{v\left\lvert\right.v\in\mathbb{R}^{d},\sum\limits_{i=1}^{n}\frac{v_{i}^{2}}{a_{i}}\leq 1,a_{i}=\sqrt{d}\cdot i^{-2},1\leq i\leq d\right\}.$ (4.1)
Note that our choice of the axes of the ellipse ensures that the maximal $\ell_{2}$ norm of the vectors in $K_{e}$ is greater than the theoretical lower bound (3.6). For a given positive value $c\leq\sqrt{a_{1}}$ as the $\ell_{2}$ norm of $\mu$ under $H_{1}$, the following approach uniquely determines the specific vector $\mu_{c}$ in $K_{e}$ such that $\left\lVert\mu_{c}\right\rVert_{2}=c$.
[…]
where $k^{\star}=\min\limits\left\{k_{1}^{\star},k_{2}^{\star}\right\}$. We verify in Appendix B.10 that the approach (4.2) ensures that the resulting vector $\mu_{c}$ achieves the minimal value of $\left\lVert P^{\star}\mu\right\rVert_{2}$ among the set $\left\{v\left\lvert\right.v\in\mathbb{R}^{d},v\in K_{e},\left\lVert v\right\rVert_{2}=c\right\}$, where $P^{\star}$ is the optimal projection operator corresponding to $k^{\star}$. We believe that this property places $\mu_{c}$ in a near-worst-case position for the algorithm, and thus provides an appropriate setting to test its robustness.
4.3 Configurations of Adversary
For the adversary $\mathcal{C}$ in the strong $\epsilon$-contamination model, we design the following two strategies. Note that in the strategies, $\mu_{0}$ is an unknown fake mean vector determined by $\mathcal{C}$ that can depend on the constraint $K$.
1
2import $\widetilde{\mathbf{X}},N,d,\sigma=1$
3
4Compute $\texttt{arr}\in\mathbb{R}^{N}$, $\texttt{arr}_{i}=\left\lVert\widetilde{X}_{i}\right\rVert_{2},1\leq i\leq N$
5if $H_{1}$ is true then
6 $C=\text{indices of the largest }\epsilon N\text{ entries in }\texttt{arr}$
7 $\mathbf{X}_{C}=\epsilon N\text{ i.i.d. fake data from }\mathcal{N}\left(\mathbf{0},\mathbf{I}_{d}\right)$
8 $\widetilde{\mathbf{X}}_{C}=\mathbf{X}_{C}$
9
10else
11 import $\mu_{0}$
12
13 $C=\text{indices of the smallest }\epsilon N\text{ entries in }\texttt{arr}$
14 $\mathbf{X}_{C}=\epsilon N\text{ i.i.d. fake data from }\mathcal{N}\left(\mu_{0},\mathbf{I}_{d}\right)$
15 $\widetilde{\mathbf{X}}_{C}=\mathbf{X}_{C}$
16
[…]
Algorithm 1 strategy 1 of the adversary $\mathcal{C}$
1
2import $\widetilde{\mathbf{X}},N,d,\sigma=1$
3
4if $H_{1}$ is true then
5   Compute $\texttt{dist}\in\mathbb{R}^{N}$, $\texttt{dist}_{i}=\left\lVert\widetilde{X}_{i}\right\rVert_{2},1\leq i\leq N$
6 $C=\text{indices of the largest }\epsilon N\text{ entries in }\texttt{arr}$
7 $\mathbf{X}_{C}=\epsilon N\text{ i.i.d. fake data from }\mathcal{N}\left(\mathbf{0},\mathbf{I}_{d}\right)$
8 $\widetilde{\mathbf{X}}_{C}=\mathbf{X}_{C}$
9
10else
11   Compute $\mu_{0}$ according to the constraint type
12  Compute $\texttt{dist}\in\mathbb{R}^{N}$, $\texttt{arr}_{i}=\left\lVert\widetilde{X}_{i}-\mu_{0}\right\rVert_{2}$
13 $C=\text{indices of the largest }\epsilon N\text{ entries in }\texttt{arr}$
14 $\mathbf{X}_{C}=\epsilon N\text{ i.i.d. fake data from }\mathcal{N}\left(\mu_{0},\mathbf{I}_{d}\right)$
15 $\widetilde{\mathbf{X}}_{C}=\mathbf{X}_{C}$
16
[…]
Algorithm 2 strategy 2 of the adversary $\mathcal{C}$
In strategies 1 and 2, the adversary $\mathcal{C}$ and the resulting corruption are largely determined on the arr and dist arrays, which compute the distance between the original data and the origin, and between the original data and an imagined mean $\mu_{0}$, respectively. Both strategies were implemented in our experiments, and the results indicate that the choice has minimal impact on the outcomes.
4.4 Experimental Results
The experimental results are presented in the following figures. During the experiments, we ensured that the type-I error remained near zero, while identifying the empirical $\widetilde{\rho}_{\text{critical}}$ under $H_{1}$ as the critical value of $c$ when the corresponding type-II error exactly crosses $\alpha=0.05$.
Figure 3: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and classic setting with $N=200$.
Figure 4: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and classic setting with $N=2000$.
Figure 5: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and high-dimensional setting with $N=100$.
Figure 6: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and high-dimensional setting with $N=500$.
We observed several notable phenomena from the simulation results. First, the algorithm performs reasonably well across all parameter configurations: it successfully rejects $H_{0}$ when $\left\lVert\mu\right\rVert_{2}$ is sufficiently large and fails to reject when the signal strength is too weak, in both the classical and the high-dimensional regimes. Second, focusing on $\rho_{\text{critical}}$ — the critical norm of $\mu$ that yields a type-II error of $\alpha=0.05$ — we find that $\rho_{\text{critical}}$ decreases as $N$ increases, and increases as $d$ grows. This behavior is consistent with theoretical intuition: larger sample sizes improve detectability, whereas higher dimensionality makes the testing problem more challenging. However, a more refined observation is that the influence of $d$ diminishes once the dimension becomes sufficiently large (e.g., around $d=50$ in our experiments). The reason behind this phenomenon is that, although $d$ increases to relatively large values, the optimal projection dimension $k^{\star}$ becomes increasingly insensitive to $d$ under our setting of $N=1,\sigma=1$, and the ellipsoidal constraint. A direct calculation shows that when $N=2000$, the optimal dimensions corresponding to full dimensions $[5,10,20,50]$ are $[5,9,10,12]$, respectively. This demonstrates that the gaps between the optimal dimensions can be substantially smaller than those between the actual dimensions, which leads to the phenomenon. Finally, the plots indicate that as $N$ increases, the performance of the algorithm becomes less sensitive to variations in $\epsilon$. Though not fully verified, we believe that this is attributed to the fact that for larger $N$, the eigensubspace structure of $\mathbf{X}\mathbf{X}^{\top}$ (or $\mathbf{X}^{\top}\mathbf{X}$) becomes more stable with respect to the corrupted observations.
We also remark that we computed the ratio between $\widetilde{\rho}_{\text{critical}}$ and the theoretical lower bound (3.6) in an attempt to assess the tightness of the bound. Although the preliminary results appeared encouraging, we chose not to report them or draw conclusions. The primary reason is that both (3.6) and its corresponding upper bound are derived under idealized minimax conditions that depend on the precise forms of $K$, $\mu$, and the adversary $\mathcal{C}$. In practical settings, however, these worst case quantities ($\mu$ and $\mathcal{C}$) are inherently unknown to us. Consequently, despite our simulations incorporating designed mean structures and adversarial strategies, they cannot guarantee that the conditions required for demonstrating optimality are satisfied. Identifying these conditions poses a fundamental challenge, and we leave this direction to future work.
For the hyperrectangle-type constraint, we conducted the experiments with the same settings and procedures. The results exhibited behavior similar to those obtained under the ellipsoidal constraint. Therefore, we present only the results under the ellipsoidal case.
5 Discussion
For completeness, we note that if the set $K$ is an orthogonal rotation of a QCO set — i.e., $K=UK^{\prime}$ for some orthogonal matrix $U$ — then one may transform the data by considering $U^{\top}X$ and subsequently solve the corresponding problem on the QCO set $K^{\prime}$. This follows from the fact that $U^{\top}\widetilde{X}_{i}=U^{\top}\mu+U^{\top}\xi_{i}$, where $U^{\top}\xi_{i}\sim N(0,\sigma^{2}\mathbf{I})$ and $U^{\top}\mu\in K^{\prime}$. Moreover, the computational cost of determining the optimal dimensions for $K$ and $K^{\prime}$ is essentially the same in most cases.
An alternative but equivalent way to define the optimal dimension is to simply select $k_{\text{alt}}^{\star}$ satisfying $D_{k_{\text{alt}}^{\star}-1}(K)>\left(\frac{k_{\text{alt}}^{\star}}{N}\right)^{1/4}\max\limits\left\{\sqrt{\epsilon},N^{-\frac{1}{4}}\right\}\sigma$ and $D_{k_{\text{alt}}^{\star}}(K)\leq\left(\frac{k_{\text{alt}}^{\star}+1}{N}\right)^{1/4}\max\limits\left\{\sqrt{\epsilon},N^{-\frac{1}{4}}\right\}\sigma$. It also has the equivalent definitions as the discussion following Definition 2.3 and 2.4. Once $k_{\text{alt}}^{\star}$ is determined, we simply project the data using $P^{\star}_{\text{alt}}$, the optimal projection associated with $k_{\text{alt}}^{\star}$. The remainder of the testing procedure is identical to that in Algorithms 7 and 8. By this new definition, the lower bound will be adjusted accordingly as
$\rho^{2}\gtrsim\sigma^{2}\max\limits\left\{\sqrt{\frac{k_{\text{alt}}^{\star}}{N}}\max\limits\left\{\epsilon,\frac{1}{\sqrt{N}}\right\},\epsilon^{2}\right\}.$ (5.1)
Similarly, the upper bounds $E_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ and $P_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ are obtained by substituting all $k_{1}^{\star}$ and $k_{2}^{\star}$ in $E^{2}(\epsilon,K,N,\sigma^{2})$ and $P^{2}(\epsilon,K,N,\sigma^{2})$ by $k_{\text{alt}}^{\star}$. With the proof in Appendix B.7, one can verify that $k_{\text{alt}}^{\star},P_{\text{alt}}^{\star}$ and its corresponding testing process are equivalent to that of $k_{1}^{\star}$, $k_{2}^{\star}$, $P_{1}^{\star}$, and $P_{2}^{\star}$. The new lower bound (5.1) and the upper bounds $E_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ and $P^{2}(\epsilon,K,N,\sigma^{2})$ do match the previous lower bound (3.6) and the upper bounds (3.3), (3.5) as well.
The results developed in Section 3 should be compared and contrasted with the existing results for the corresponding estimation problem. Specifically, in the estimation setting — where the same distribution assumptions and corruption model are imposed, but the goal is to estimate the mean vector $\mu$ — the minimax rates are established by Donoho et al., (1990) and Neykov, (2022). Notably, the Kolmogorov widths and the associated optimal dimensions again play a central role in these results.
Definition 5.1 (Optimal dimension and projection in estimation).
Given the same assumptions on the distribution, the constraint $K$, and the corruption model as 2.3, 2.4, and 2.8, $k_{e}^{\star}$ is defined as the dimension that satisfies $D_{k_{e}^{\star}-1}(K)>\sqrt{\frac{k_{e}^{\star}}{N}}\sigma$ but $D_{k_{e}^{\star}}\leq\sqrt{\frac{k_{e}^{\star}+1}{N}}\sigma$. It can be equivalently defined as
$k_{e}^{\star}:=\max\limits_{k}\left\{k\left\lvert\right.k\geq 0,D_{k-1}(K)\geq\sqrt{\frac{k}{N}}\sigma\right\}.$ (5.2)
Theorem 5.2 (Robust Estimation Rate, Prasadan and Neykov, (2025)).
Given the same assumptions on the distribution, the constraint $K$, and the corruption model as (1.2), the minimax rate for the estimation problem is given by
$\inf\limits_{\widehat{\nu}}\sup\limits_{\mu\in K}\mathbb{E}_{\mu}\left\lVert\widehat{\nu}(\mathbf{X})-\mu\right\rVert_{2}^{2}\asymp\sigma^{2}\max\limits\left\{\frac{k_{e}^{\star}}{N},\epsilon^{2}\right\},$ (5.3)
where $\inf$ is taken with respect to all measurable functions $\widehat{\nu}$ of the data.
When $K=\mathbb{R}^{d}$ and there is no corruption, we have $k_{e}^{\star}=d$ and $\epsilon=0$, in which case the rate above recovers the classic result that $\inf\limits_{\widehat{\nu}}\sup\limits_{\mu}\mathbb{E}_{\mu}\left\lVert\widehat{\nu}(\widetilde{\mathbf{X}})-\mu\right\rVert_{2}^{2}\asymp\frac{d}{N}\sigma^{2}$. It is also well known that, in this setting, the estimation problem is generally more difficult than the testing problem. The reason is simple and intuitive — one is able to derive a valid test with small error probabilities whenever $\left\lVert\mu\right\rVert_{2}^{2}$ is asymptotically larger than the minimax lower bound for the estimation task. This relationship, unsurprisingly, is inherited by our more general problem.
Lemma 5.3 (Testing-estimation relationship).
Given the same assumptions on the distribution, the constraint $K$, and the corruption model above, we have
$\sigma^{2}\max\limits\left\{\frac{\sqrt{k_{1}^{\star}}}{N},\epsilon\sqrt{\frac{k_{2}^{\star}}{N}},\epsilon^{2}\right\}\lesssim\sigma^{2}\max\limits\left\{\frac{k_{e}^{\star}}{N},\epsilon^{2}\right\}$ (5.4)
We refer readers to Appendix B.11 for a straightforward computational verification of Lemma 5.3. Following the discussion after Corollary 3.12, we can also analyze the phase transition of the estimation problem. It occurs around $\epsilon\asymp\sqrt{\frac{k_{e}^{\star}}{N}}$. If $\epsilon\lesssim\sqrt{\frac{k_{e}^{\star}}{N}}$, then $\epsilon$ is too small to influence the minimax rate of the estimation error. Otherwise (for example, $\epsilon\asymp N^{-\frac{1}{4}}$), the adversary is strong enough to the extent that $N$ is excluded and the minimax rate only relies on $\epsilon$ and $\sigma$.
Another interesting observation from Lemma 5.3 is that although estimation is generally more challenging than testing, the gap disappears as the corruption rate increases. In fact, both quantities stabilize at $\epsilon^{2}\sigma^{2}$ when $\epsilon$ is sufficient large. Similar phenomenon has also been observed in a recent work Kania et al., (2025). See Section 5.1 for additional details.
5.1 Open Problems and Future Work
An intriguing open direction is to investigate the localized minimax testing rates for QCO sets. A preliminary step in this direction was taken by Wei and Wainwright, (2020), who studied local testing rates in the setting of ellipsoids. It would be interesting to understand the extent to which their results can be generalized to QCO sets. Moreover, from an intuitive standpoint, the origin appears to be the most challenging point for testing under QCO constraints, as it may correspond to the largest local volume relative to other points. This raises a natural question: does the global minimax testing rate coincide with the localized rate at the origin, and thus with the rates established in this paper?
A fundamental assumption of this paper is that the covariance matrix is fully known (and taken to be $\mathbf{I}_{d}$). This, however, is rarely guaranteed in real-world applications. An important direction for future work is therefore to understand whether — and to what extent — we can still distinguish between $H_{0}$ and $H_{1}$ when $\bm{\Sigma}$ must be estimated rather than treated as prior knowledge. A recent relevant work in robust covariance matrix estimation is Li and Ye, (2020), where the authors proposed an algorithm achieving a near-optimal error rate $\mathcal{O}\left(\epsilon\ln\left(\frac{1}{\epsilon}\right)\right)$ in the Mahalanobis norm under $\epsilon$-strong contamination model while remaining computationally efficient. Nevertheless, this algorithm is designed for zero-mean distributions and unconstrained setting. It remains unknown whether reliable testing is achieveble and how the constraint $K$ can facilitate the problem when $\bm{\Sigma}$ is unknown.
In Kania et al., (2025), the authors consider a more general problem than (1.1):
$H_{0}:\left\lVert\mu\right\rVert_{2}\leq\rho_{0},$ (5.5)
$H_{1}:\left\lVert\mu\right\rVert_{2}\geq\rho_{1}.$
In other words, rather than being fixed at the origin, the mean vector under the null is allowed to vary within a neighborhood of the origin. An interesting corollary of their work is that as the neighborhood expands, the minimax testing rate converges to the minimax estimation rate — an effect that parallels a key phenomenon in our setting. Given the framework of their problem, it is natural to ask whether one can establish the minimax rate with constrained and adversarial setting of (5.5). From the perspective of the testing-estimation relationship, an even more important question is: under what conditions does the gap between testing and estimation disappear?
Another important avenue for future research will be to study $\ell_{p}$ norm separation for $p>2$. We believe there are fundamental changes compared to the problem discussed in this work. The first reason is that $\chi^{2}$-test is no longer suitable for the testing problem when $p>2$. Instead, $\chi^{p}$-test is a useful tool in this scenario, see the results by Ingster and Suslina, (2003) when there is no constraint and adversary. The second reason is that the geometric properties of the $\ell_{p}$ balls when $p>2$ are known to be dramatically different compared to those when $1\leq p\leq 2$. This affects the selection of the extreme vector as in Lemma 3.1. It is also unclear at the moment whether the Kolmogorov widths will still drive the minimax rates in that case.
Finally, a more ambitious open problem is to characterize the minimax testing rates over general convex sets that contain the origin. It would be particularly exciting to determine whether an optimal test exists for every such convex set $K$, analogous to the result of Neykov, (2022), which established the existence of a universally optimal estimation procedure for arbitrary convex sets in the corresponding estimation problem.
Acknowledgments
The authors are grateful to Siva Balakrishnan for discussions on local testing on ellipsoids and related minimax testing problems. The authors also would like to thank Clément Canonne for a helpful email exchange in which Clément suggested to us that in the unconstrained problem, $\left\lVert\mu\right\rVert_{2}\gtrsim\epsilon\sigma$ is required, which we extended to show Theorem 3.4.
Appendix A Some Basic Probabilistic Concepts and Useful Lemmas
In this section, we recall some basic probabilistic concepts and useful lemmas that are frequently mentioned and leveraged in this paper. Most of them are related to concentrations and the measures of distance between distributions.
Definition A.1 (Total variation distance).
For a $\sigma$-algebra $(\Omega,\mathscr{F})$ and the probabilities $\mathcal{P},\mathcal{Q}$ defined on $\mathscr{F}$, the total variation distance between $\mathcal{P}$ and $\mathcal{Q}$ is defined as
$\text{TV}(\mathcal{P},\mathcal{Q})=\sup\limits_{E\in\mathscr{F}}\left|\mathcal{P}(E)-\mathcal{Q}(E)\right|.$
From the definition, we immediately know that $\text{TV}(\mathcal{P},\mathcal{Q})=\text{TV}(\mathcal{Q},\mathcal{P})$ and $0\leq\text{TV}(\mathcal{P},\mathcal{Q})\leq 1$. $\text{TV}(\mathcal{P},\mathcal{Q})=0$ if and only if $\mathcal{P}=\mathcal{Q}$ with only exception of zero-measure sets. When $\mathcal{P},\mathcal{Q}$ has densities $p,q$ with respect to some measure $\nu$ and space $\Omega_{\nu}$, it can be also easily shown that
$\text{TV}(\mathcal{P},\mathcal{Q})=\displaystyle\int_{\Omega_{\nu}}\min\limits\left\{p,q\right\}d\nu=\frac{1}{2}\displaystyle\int_{\Omega_{\nu}}\left|p-q\right|d\nu.$ (A.1)
Definition A.2 (Kullback-Leibler divergence (KL divergence)).
Given the same $\sigma$-algebra and probabilities, additionally assume that $\mathcal{P}\ll\mathcal{Q}$. The KL divergence between $\mathcal{P}$ and $\mathcal{Q}$ is defined as
$\text{KL}(\mathcal{P},\mathcal{Q})=\displaystyle\int_{\Omega}\ln\frac{d\mathcal{P}}{d\mathcal{Q}}d\mathcal{P}.$
KL divergence is always non-negative, with $\text{KL}(\mathcal{P},\mathcal{Q})=0$ implying $\mathcal{P}=\mathcal{Q}$ with only exception of zero-measure sets. Contrasted with total variation distance, KL divergence is not symmetric, meaning that $\text{KL}(\mathcal{P},\mathcal{Q})\neq\text{KL}(\mathcal{Q},\mathcal{P})$ in general, even if $\frac{d\mathcal{P}}{d\mathcal{Q}}$ and $\frac{d\mathcal{Q}}{d\mathcal{P}}$ both exist.
Lemma A.3 (Pinsker’s inequality).
For the probability distributions $\mathcal{P},\mathcal{Q}$ with $\mathcal{P}\ll\mathcal{Q}$, we have
$\text{TV}(\mathcal{P},\mathcal{Q})\leq\sqrt{2\text{KL}(\mathcal{P},\mathcal{Q})}.$
When $\mathcal{P}\ll\mathcal{Q}$ and $\mathcal{Q}\ll\mathcal{P}$, it can be refined as
$\text{TV}(\mathcal{P},\mathcal{Q})\leq\sqrt{2\min\limits\left\{\text{KL}(\mathcal{P},\mathcal{Q}),\text{KL}(\mathcal{Q},\mathcal{P})\right\}}.$
There are many proofs for this lemma. One can refer Tsybakov, (2008) for a quick proof.
Definition A.4 (Coupling).
Given probability distributions $\mathcal{P},\mathcal{Q}$, a valid coupling of $\mathcal{P}$ and $\mathcal{Q}$ is any pair of random variables $(X,Y)$ where the marginal distribution of $X$ ang $Y$ is $\mathcal{P},\mathcal{Q}$ respectively.
The following classic lemma help us connect the total variation distance with the Hamming distance in the proof of Theorem 3.3.
Lemma A.5.
For any probability distributions $\mathcal{P}$ and $\mathcal{Q}$, their total variation distance equals to the infimum over the probability that $X$ is not equal to $Y$, where $(X,Y)$ is a valid coupling of $\mathcal{P}$ and $\mathcal{Q}$.
[…]
The following definitions and lemmas concern the sub-Gaussian and sub-exponential random variables, which are closely related to the concentration bounds in the proof of the main theorems.
Definition A.6 (sub-Gaussian random variables).
For a univariate random variable $X$, define its sub-Gaussian norm as $\left\lVert X\right\rVert_{\psi_{2}}:=\inf\limits\left\{t:t>0,\mathbb{E}\exp\left\{\frac{X^{2}}{t^{2}}\right\}\leq 2\right\}$. If such $\left\lVert X\right\rVert_{\psi_{2}}$ does not exist, then define $\left\lVert X\right\rVert_{\psi_{2}}=\infty$. $X$ is said to be sub-Gaussian if $\left\lVert X\right\rVert_{\psi_{2}}<\infty$.
There are many equivalent definitions for sub-Gaussian random variables. The most frequently mentioned one is $\mathbb{P}(\left|X\right|\geq t)\leq 2\exp\left\{-\frac{t^{2}}{c\left\lVert X\right\rVert_{\psi_{2}}^{2}}\right\}$.
Definition A.7 (sub-exponential random variables).
For a univariate random variable $X$, define its sub-exponential norm as $\left\lVert X\right\rVert_{\psi_{1}}:=\inf\limits\left\{t:t>0,\mathbb{E}\exp\left\{\frac{\left|X\right|}{t}\right\}\leq 2\right\}$. If such $\left\lVert X\right\rVert_{\psi_{1}}$ does not exist, then define $\left\lVert X\right\rVert_{\psi_{1}}=\infty$. $X$ is said to be sub-exponential if $\left\lVert X\right\rVert_{\psi_{1}}<\infty$.
There are also many equivalent definitions for sub-exponential random variables. A popular one is $\mathbb{P}(\left|X-\mathbb{E}X\right|\geq t)\leq 2\exp\left\{-c\min\limits\left\{\frac{t^{2}}{\left\lVert X\right\rVert_{\psi_{1}}^{2}},\frac{t}{\left\lVert X\right\rVert_{\psi_{1}}}\right\}\right\}$.
The requirement of sub-exponential random variables is weaker than that of sub-Gaussian in the sense that if $X$ is sub-Gaussian, it must be sub-exponential. Moreover, if $X$ is sub-Gaussian, then by definition $X^{2}$ is sub-exponential.
The following is the famous Bernstein’s inequality, which handles multiple independent sub-Gaussian (sub-exponential) random variables. It also generalizes the equivalent definitions mentioned above.
Lemma A.8 (Bernstein’s inequality).
Assume $X_{1},\dots,X_{n}$ are independent, zero-mean random variables. If $X_{1},\dots,X_{n}$ are sub-Gaussian, then for any $t\geq 0$, we have
$\mathbb{P}\left(\sum\limits_{i=1}^{n}X_{i}\geq t\right)\leq\exp\left\{-ct^{2}/\sum\limits_{i=1}^{n}\left\lVert X_{i}\right\rVert_{\psi_{2}}^{2}\right\}.$ (A.3)
If $X_{1},\dots,X_{n}$ are sub-exponential, then we have
$\mathbb{P}\left(\sum\limits_{i=1}^{n}X_{i}\geq t\right)\leq\exp\left\{-c\min\limits\left\{t^{2}/\sum\limits_{i=1}^{n}\left\lVert X_{i}\right\rVert_{\psi_{1}}^{2},t/\max\limits_{i}\left\{\left\lVert X_{i}\right\rVert_{\psi_{1}}\right\}\right\}\right\}.$ (A.4)
The next lemma is particularly useful when we control the $\ell_{2}$ norm of a random vector with independent sub-Gaussian coordinates.
Lemma A.9.
Assume $X=(X_{1},\dots,X_{n})\in\mathbb{R}^{d}$ where $X_{1},\dots,X_{n}$ are independent, $\mathbb{E}X_{i}=0,\text{Var}\left(X_{i}\right)=1$, and each $X_{i}$ is sub-Gaussian. Then $\left\lVert X\right\rVert_{2}-\sqrt{n}$ is sub-Gaussian with $\left\lVert\left\lVert X\right\rVert_{2}-\sqrt{n}\right\rVert_{\psi_{2}}\leq c\max\limits_{i}\left\{\left\lVert X_{i}\right\rVert_{\psi_{2}}^{2}\right\}$.
The definitions and lemmas above can be found in many modern statistics textbook. For an example, see Vershynin, (2024).
The famous Hanson-Wright inequality is a powerful tool to bound the tail probability of quadratic forms of the sub-Gaussian random variables. It can also be viewed as a generalization of the Bernstein’s inequality (A.4). See Hanson and Wright, (1971) for the initial paper and Rudelson and Vershynin, (2013) for a modern proof.
Lemma A.10 (The Hanson-Wright inequality).
Assume that $X=(X_{1},\dots,X_{n})\in\mathbb{R}^{n}$ is a random vector with independent, zero-mean and sub-Gaussian coordinates. Let $A$ be an $n\times n$ matrix and $K=\max\limits_{i}\left\lVert X_{i}\right\rVert_{\psi_{2}}$. Then for any $t\geq 0$, we have
$\mathbb{P}\left(|X^{\top}AX-\mathbb{E}X^{\top}AX|>t\right)\leq 2\exp\left\{-c\min\limits\left\{\frac{t^{2}}{K^{4}\left\lVert A\right\rVert_{F}^{2}},\frac{t}{K^{2}\left\lVert A\right\rVert_{2}}\right\}\right\},$ (A.5)
where $\left\lVert\cdot\right\rVert_{F}$ is the Frobenius norm of a matrix, and $\left\lVert\cdot\right\rVert_{2}$ is the $\ell_{2}$ operator norm of a matrix.
We now turn to the Wishart matrix, which plays a central role in covariance estimation in high-dimensional Gaussian models. Additional properties of the Wishart distribution and its applications in estimation can be found in Kubokawa and Srivastava, (2008) and Bishop et al., (2018).
Definition A.11 (The Wishart matrix).
Given a matrix $X\in\mathbb{R}^{n\times d}$ where each row is independently drawn from $\mathcal{N}(0,\bm{\Sigma})$ for some $\bm{\Sigma}\in\mathbb{R}^{d\times d}$. Then, the Wishart matrix with scale matrix $\bm{\Sigma}$ and dimension $(d,n)$, denoted as $W_{d}(\bm{\Sigma},n)$, is defined as
$W_{d}(\bm{\Sigma},n)=\mathbf{X}^{\top}\mathbf{X}.$
Specifically, when $\bm{\Sigma}=\mathbf{I}_{d}$ is the identity matrix, the notation is simplified to $W_{d}(n)$. The distribution of $W_{d}(\bm{\Sigma},n)$ is called Wishart distribution with parameter $d,n,\bm{\Sigma}$. When the meaning is clear from the context, we also use notation $W_{d}(\bm{\Sigma},n)$ (and also $W_{d}(n)$) as Wishart distribution.
For the Gaussian random matrix $X$, the following lemma can be used to bound its maximum and minimum singular values, which can be further used to establish the bound for the $\ell_{2}$ operator norm of $W_{d}(\bm{\Sigma},n)$. One can find a proof of this lemma in Wainwright, (2019) by the concentration bound for Lipschitz functions and the Sudakov-Fernique comparison inequality.
Lemma A.12.
Assume $X\in\mathbb{R}^{n\times d}$, and each row is independently drawn from $\mathcal{N}(0,\bm{\Sigma})$, then for any $\epsilon>0$, we have
$\mathbb{P}\left(\frac{\sigma_{\text{max}}(X)}{\sqrt{n}}\geq\gamma_{\text{max}}(\sqrt{\bm{\Sigma}})(1+\epsilon)+\sqrt{\frac{\text{tr}(\bm{\Sigma})}{n}}\right)\leq\exp\left\{-\frac{n\epsilon^{2}}{2}\right\}.$
When $n\geq d$, we further have the following control on minimum singular value of $X$:
$\mathbb{P}\left(\frac{\sigma_{\text{min}}(X)}{\sqrt{n}}\leq\gamma_{\text{min}}(\sqrt{\bm{\Sigma}})(1-\epsilon)-\sqrt{\frac{\text{tr}(\bm{\Sigma})}{n}}\right)\leq\exp\left\{-\frac{n\epsilon^{2}}{2}\right\}.$
Here $\sigma_{\text{max}}(X)$ (or $\sigma_{\text{min}}(X)$) is the max (or min) singular value of $X$ and $\gamma_{\text{max}}(\sqrt{\bm{\Sigma}})$ (or $\gamma_{\text{min}}(\sqrt{\bm{\Sigma}})$) is the max (or min) eigenvalue of $\sqrt{\bm{\Sigma}}$.
We can leverage Lemma A.12 to bound the $\ell_{2}$ operator norm of $W_{d}(n)$.
Corollary A.13.
Assume $W=\mathbf{X}^{\top}\mathbf{X}\in\mathbb{R}^{d\times d}$ is drawn from $W_{d}(n)$, and $n\geq d$, then for any $\delta>0$, with probability higher than $1-2\delta$, we have
$\left\lVert W-n\mathbf{I}_{d}\right\rVert_{2}\leq C\left(\sqrt{nd}+\sqrt{n\ln\left(\frac{1}{\delta}\right)}+\ln\left(\frac{1}{\delta}\right)\right)$
Proof.
Set $\bm{\Sigma}=\mathbf{I}_{d}$ in the first inequality of Lemma A.12, we have:
$\mathbb{P}\left(\frac{\sigma_{\text{max}}(X)}{\sqrt{n}}\geq(1+\epsilon)+\sqrt{\frac{d}{n}}\right)\leq\exp\left\{-\frac{n\epsilon^{2}}{2}\right\}.$
Using $\sigma_{\text{max}}(X)=\sqrt{\gamma_{\text{max}}(W_{d}(n))}$ and after some calculation, we have
$\mathbb{P}\left(\gamma_{\text{max}}(W_{d}(n))-n\geq d+2\sqrt{nd}+n\epsilon^{2}+2n\epsilon+2\sqrt{nd}\epsilon\right)\leq\exp\left\{-\frac{n\epsilon^{2}}{2}\right\}.$
Setting $\delta=\sqrt{\frac{\ln\left(\frac{1}{\epsilon}\right)}{n}}$ yields
$\mathbb{P}\left(\gamma_{\text{max}}(W_{d}(n))-n\geq d+2\sqrt{nd}+\ln\left(\frac{1}{\delta}\right)+2\sqrt{n\ln\left(\frac{1}{\delta}\right)}+2\sqrt{d\ln\left(\frac{1}{\delta}\right)}\right)\leq\delta.$
The lower bound follows from the condition that $n\geq d$. The upper bound is derived with a similar argument. The proof is completed. ∎
Under our problem setting, the following corollary is also required. It allows the mean vector $\mu$ of the normal distribution to be nonzero, but at the expense of introducing a dependence on $\left\lVert\mu\right\rVert_{2}$.
Corollary A.14.
Assume $\mathbf{X}\in\mathbb{R}^{n\times d}$ and each row is independently drawn from $\mathcal{N}(\mu,\mathbf{I}_{d})$ with $\rho:=\left\lVert\mu\right\rVert_{2}$, then for $\delta>0$, with probability higher than $1-2\delta$, we have:
(i), $n\geq d$: $\left\lVert\mathbf{X}^{\top}\mathbf{X}-n\mathbf{I}_{d}\right\rVert_{2}\leq C_{1}\left((1+\rho)\left(\sqrt{nd}+\sqrt{n\ln\left(\frac{1}{\delta}\right)}\right)+\ln\left(\frac{1}{\delta}\right)+n\rho^{2}\right)$, (ii), $n<d$: $\left\lVert\mathbf{X}\mathbf{X}^{\top}-d\mathbf{I}_{n}\right\rVert_{2}\leq C_{2}\left((1+\rho)\left(\sqrt{nd}+\sqrt{d\ln\left(\frac{1}{\delta}\right)}\right)+\ln\left(\frac{1}{\delta}\right)+n\rho^{2}\right)$.
Proof.
When $n\geq d$, let $Z_{i}=X_{i}-\mu$, we have
$\left\lVert\mathbf{X}^{\top}\mathbf{X}-n\mathbf{I}_{d}\right\rVert_{2}=\left\lVert\sum\limits_{i=1}^{n}X_{i}X_{i}^{\top}-n\mathbf{I}_{d}\right\rVert_{2}\leq\left\lVert\sum\limits_{i=1}^{n}Z_{i}Z_{i}^{\top}-n\mathbf{I}_{d}\right\rVert_{2}+2\left\lVert\sum\limits_{i=1}^{n}\mu Z_{i}^{\top}\right\rVert_{2}+n\left\lVert\mu\mu^{\top}\right\rVert_{2}.$
By Corollary A.13, we already have $\left\lVert\sum\limits_{i=1}^{n}Z_{i}Z_{i}^{\top}-n\mathbf{I}_{d}\right\rVert_{2}\leq C_{1}\left(\sqrt{nd}+\sqrt{n\ln\left(\frac{1}{\delta}\right)}+\ln\left(\frac{1}{\delta}\right)\right)$. We also have $n\left\lVert\mu\mu^{\top}\right\rVert_{2}=n\alpha^{2}$. It remains to bound $\left\lVert\sum\limits_{i=1}^{n}\mu Z_{i}^{\top}\right\rVert_{2}$.
$\left\lVert\sum\limits_{i=1}^{n}\mu Z_{i}^{\top}\right\rVert_{2}\leq\left\lVert\mu\right\rVert_{2}\cdot\left\lVert\sum\limits_{i=1}^{n}Z_{i}\right\rVert_{2}.$
Since $Z_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ for each $i$, from A.9 we know that $\left\lVert\sum\limits_{i=1}^{n}\frac{Z_{i}}{\sqrt{n}}-d\right\rVert_{2}$ is a sub-Gaussian random variable and therefore with probability higher than $1-\delta$ we have:
$\left\lVert\sum\limits_{i=1}^{n}\frac{Z_{i}}{\sqrt{n}}-d\right\rVert_{2}\leq C\sqrt{\ln\left(\frac{1}{\delta}\right)},$
which leads to $\left\lVert\sum\limits_{i=1}^{n}Y_{i}\right\rVert_{2}\leq C\left(\sqrt{nd}+\sqrt{n\ln\left(\frac{1}{\delta}\right)}\right)$. Combined with previous results on $\left\lVert\sum\limits_{i=1}^{n}Z_{i}Z_{i}^{\top}-n\mathbf{I}_{d}\right\rVert_{2}$ and $n\left\lVert\mu\mu^{\top}\right\rVert_{2}$, we complete the proof of (i).
To bound $\left\lVert\mathbf{X}\mathbf{X}^{\top}-d\mathbf{I}_{n}\right\rVert_{2}$, we notice that $\mathbf{X}\mathbf{X}^{\top}$ has the same non-trivial eigenvalues with $\mathbf{X}^{\top}\mathbf{X}=\mathbf{Z}^{\top}\mathbf{Z}+2\sum\limits_{i=1}^{n}\mu Z_{i}^{\top}+n\mu\mu^{\top}$, in this case, we only need to address the $\ell_{2}$ operator norm of $\mathbf{Z}^{\top}\mathbf{Z}$, which again equals to $\mathbf{Z}\mathbf{Z}^{\top}$. Since $\mathbf{Z}\mathbf{Z}^{\top}$ actually follows the distribution of $W_{n}(d)$, the results follow from the proof of (i). ∎
The next lemma is from Fact 4.2 in Dong et al., (2019). It is an analog of Corollary A.13, but within the field when the $\ell_{1}$ norm of a weight $\omega$ is small compared to $n$.
Lemma A.15.
Assume $\mathbf{X}\in\mathbb{R}^{n\times d}$ and each row is independently drawn from $\mathcal{N}(0,\mathbf{I}_{d})$. Then, with some constant $C_{1}>0$ and another sufficiently large constant $C_{2}$, as long as $n\geq C_{2}\cdot\frac{1}{\epsilon}$, the following holds with probability at least $1-\delta$ for any $v,\omega$ such that $\left\lVert v\right\rVert_{2}=1$, $\left\lVert\omega\right\rVert_{1}\leq C_{2}\cdot\epsilon n$:
$\left\lVert\sum\limits_{i=1}^{n}\omega_{i}X_{i}X_{i}^{\top}\right\rVert_{2}\leq C_{1}\left(\epsilon n\ln\left(\frac{1}{\epsilon}\right)+d+\ln\left(\frac{1}{\delta}\right)\right).$
Corollary A.16.
Assuming the same conditions as Lemma A.15, we further allow $X_{i}$ from $\mathcal{N}(\mu,\mathbf{I}_{d})$ with $\left\lVert\mu\right\rVert_{2}\leq\alpha$. As a consequence, the inequality will track the dependence on $\alpha$:
$\left\lVert\sum\limits_{i=1}^{n}\omega_{i}X_{i}X_{i}^{\top}\right\rVert_{2}\leq C\left[\epsilon n\left(\ln\left(\frac{1}{\epsilon}\right)+\alpha^{2}\right)+\sqrt{\epsilon n}\alpha+d+\ln\left(\frac{1}{\delta}\right)\right].$
Proof.
The proof is basically from Lemma A.15 the standard argument of concentration. For any $v$ such that $\left\lVert v\right\rVert_{2}=1$, we have:
$v^{\top}\left(\sum\limits_{i=1}^{n}\omega_{i}X_{i}X_{i}^{\top}\right)v=v^{\top}\left(\sum\limits_{i=1}^{n}(X_{i}-\mu)(X_{i}-\mu)^{\top}\right)v+\sum\limits_{i=1}^{n}\omega_{i}(v^{\top}\mu)^{2}+2(v^{\top}\mu)\sum\limits_{i=1}^{n}\omega_{i}(v^{\top}\left(X_{i}-\mu\right))$
Bound of the first term can be obtained from Lemma A.15. For the second term, by the Cauchy-Schwarz inequality we can bound it by $2\left\lVert\omega\right\rVert_{1}\cdot\alpha^{2}\leq C\cdot\epsilon n\alpha^{2}$. For the third term, since $\sum\limits_{i=1}^{n}\omega_{i}v^{\top}(X_{i}-\mu)$ follows from $\mathcal{N}(0,\left\lVert\omega\right\rVert_{2}^{2})$, we know we can bound it by $C\cdot\alpha\left\lVert\omega\right\rVert_{2}\leq C\cdot\alpha\sqrt{\epsilon n}$. Combining them together yields the result. ∎
Hölder’s inequality is a fundamentally important tool in probability, mathematical analysis and functional analysis. We briefly review it here.
Lemma A.17 (Hölder’s inequality).
Let $(\Omega,\mathscr{F})$ be a $\sigma$-algebra equipped with a measure $\mu$. For a pair of conjugate numbers $p,q$: $\frac{1}{p}+\frac{1}{q},1\leq p,q\leq\infty$, and $f\in L_{p}(\Omega,\mu)$, $g\in L_{q}(\Omega,\mu)$, we have
$\left\lVert fg\right\rVert_{1}=\displaystyle\int_{\Omega}\left|fg\right|d\mu\leq\left(\displaystyle\int_{\Omega}\left|f\right|^{p}d\mu\right)^{\frac{1}{p}}\left(\displaystyle\int_{\Omega}\left|g\right|^{q}d\mu\right)^{\frac{1}{q}}=\left\lVert f\right\rVert_{p}\left\lVert g\right\rVert_{q}.$
The equality holds if and only if $a\left|f\right|^{p}=b\left|g\right|^{q}$ a.s. for some constants $a,b$.
The next lemma is a famous result for the combination number.
Lemma A.18.
For any $0\leq k\leq n$, we have $\sum\limits_{i=0}^{k}\binom{n}{i}\leq\left(\frac{en}{k}\right)^{k}$.
Proof.
By the Taylor’s expansion of $e^{k}$, we know $e^{k}=\sum\limits_{i=0}^{\infty}\frac{k^{i}}{i!}\geq\sum\limits_{i=0}^{k}\frac{k^{i}}{i!}$. Therefore, it suffices to show that $\frac{k^{k}}{n^{k}}\binom{n}{i}\leq\frac{k^{i}}{i!}$ for any $0\leq i\leq k$, and this is equivalent to showing that $\frac{k^{k}}{n^{k}}\prod\limits_{j=0}^{i-1}(n-j)\leq k^{i}$. However, the last statement can be directly obtained from the fact that $\frac{k^{k-i}}{n^{k-i}}\leq 1\leq\frac{n^{i}}{\prod\limits_{j=0}^{i-1}(n-j)}$. ∎
Finally, we introduce an intuitive and useful result in convexity and combinatorial mathematics. We note that Canonne et al., (2023) first introduced this result in their paper but their proof contains a flaw which we fix below.
Lemma A.19.
Denote $0,1]^{N}$ to be the set $\left\{(x_{1},x_{2},\dots,x_{N}):0\leq x_{i}\leq 1,\forall i\in[N]\right\}$. Given any nonnegative integer $k\leq N$ and any vector $x\in[0,1]^{N}$ with condition $\left\lVert x\right\rVert_{1}\leq k$, there exists a positive integer $M$, sets $T_{1},T_{2},\dots,T_{M}\subset[N]$, and positive numbers $a_{1},a_{2},\dots,a_{M}$ such that the following statements hold:
(i), $\left|T_{i}\right|\leq k,\forall i\in[M]$; (ii), $0\leq a_{i}\leq 1,\forall i\in[M],\sum\limits_{i=1}^{M}a_{i}\leq 1$; (iii), $x=\sum\limits_{i=1}^{M}a_{i}\mathbf{1}_{T_{i}}$,
where same as before $\mathbf{1}_{T_{i}}$ is the indicator vector of the set $T_{i}$.
Proof.
We first show the following result: for any vector $(x_{1},\ldots,x_{N})\in[0,1]^{n}$ and an given integer $k\leq N$ such that $\left\lVert x\right\rVert_{1}=k$, and $0\leq x_{i}\leq 1$ there exist sets $T_{i}$ of cardinality $|T_{i}|=k$ and weights $a_{i}\geq 0$ with $\sum\limits_{i=1}^{M}a_{i}=1$ such that $x=\sum\limits_{i=1}^{M}a_{i}\mathbf{1}_{T_{i}}$.
The proof of this statement uses Birkhoff-von Neumann’s theorem. Construct the following $N\times N$ matrix
$W:=\begin{pmatrix}\frac{x_{1}}{k}&\frac{x_{2}}{k}&\dots&\frac{x_{N}}{k}\\ \frac{x_{1}}{k}&\frac{x_{2}}{k}&\dots&\frac{x_{N}}{k}\\ \vdots&\vdots&\dots&\vdots\\ \frac{1-x_{1}}{N-k}&\frac{1-x_{2}}{N-k}&\dots&\frac{1-x_{N}}{N-k}\\ \vdots&\vdots&\dots&\vdots\\ \frac{1-x_{1}}{N-k}&\frac{1-x_{2}}{N-k}&\dots&\frac{1-X_{N}}{N-k}\end{pmatrix}$,
where the $W_{ij}$ is defined as
[…]
It is easy to check that this is a doubly stochastic matrix. By Birkhoff-von Neumann’s theorem, there exists a set of permutation matrices $P_{1},\dots,P_{M}$ such that
$W=\sum\limits_{i=1}^{M}\omega_{i}P_{i}.$
Since the sum of the first $k$ rows of $W$ is $x$, now summing all the first $k$ rows of the matrices $P_{i}$ shows the desired result.
For the general case, suppose that $k-1<\left\lVert x\right\rVert_{1}<k$. Consider the augmented vector $\widetilde{x}=(x_{1},\ldots,x_{N},k-\sum\limits_{i=1}^{N}x_{i})$, and apply the logic above to that vector. This gives us a representation $\widetilde{x}=\sum\limits_{i=1}^{\widetilde{M}}w_{i}v_{i}$, where $v_{i}$ consists of $k$-ones and (N+1k)(N+1-k)-zeros. Erasing the last entry of the $v_{i}$’s, we obtain a representation of $x$ with desired vectors and weights. The proof is completed. ∎
Appendix B Deferred Proofs of the Lemmas and Main Theorems
B.1 Proof of Lemma 3.1
The proof of Lemma 3.1 leverages key properties of QCO sets and applies the minimax theorem to a suitably truncated vector. A similar argument appears in Section 3.3.2 of Neykov, (2022).
Proof of Lemma 3.1.
By the conditions and the definition of the Kolmogorov $k$-width, the following inequality holds:
$\inf_{P\in\mathcal{P}_{k-1}}\sup_{\theta\in K}\left\lVert\theta-P\theta\right\rVert_{2}^{2}>c^{2}\sigma^{2}.$
If we only consider the projections aligned with the coordinate axes, there are $n_{k}:=\binom{d}{k}$ such projections (and countably many when $\mathcal{X}\subset\ell_{2}$). Denote the index set of the whole space $\mathcal{X}$ by $I$ — i.e., $I=[d]$ when $\mathcal{X}=\mathbb{R}^{d}$ and $\mathbb{N}$ when $\mathcal{X}\subset\ell_{2}$ — and the index sets of such projections by $J$. We have
$\inf_{P\in\mathcal{P}_{k-1}}\sup_{\theta\in K}=\left\lVert\theta-P\theta\right\rVert_{2}^{2}\leq\inf_{S\in J}\sup_{\theta\in K}\sum_{i\in I}\theta^{2}_{i}-\sum_{i\in S}\theta_{i}^{2},$
where the minimum over $S$ is taken with respect to all elements in $J$. Since $K$ is quadratically convex, the RHS above can be written as
$\inf_{S\in J}\sup_{\theta\in K}\sum_{i\in I}\theta^{2}_{i}-\sum_{i\in S}\theta_{i}^{2}\leq\inf_{S\in J}\sup_{t\in K^{2}}\sum_{i\in I}t_{i}-\sum_{i\in S}t_{i}=\inf_{w\in S_{k}}\sup_{t\in K^{2}}\left\lVert t\right\rVert_{1}-\langle w,t\rangle,$
where $w$ ranges in the set $S_{k}:=\left\{e\left\lvert\right.e\in\mathbb{R}^{d}\text{ (or $\mathbb{R}^{\infty}$)},e_{i}\in\left\{0,1\right\},\left\lVert e\right\rVert_{1}=k-1\right\}$. It follows that for each small enough $\varepsilon>0$ and each $w_{i}\in S_{k}$ there exists a $t_{i}$ such that $\left\lVert t_{i}\right\rVert_{1}-\langle w_{i},t_{i}\rangle\geq c^{2}\sigma^{2}-\varepsilon$. Since $K$ is convex and orthosymmetric, we may assume without loss of generality that $t_{i}\geq 0$ for all $i\in I$, and $t_{i}=0$ on the support of $w_{i}$. In this way, after appropriately scaling down the coordinates of $t_{i}$ (which is permissible since $K$ is convex), we obtain
$\left\lVert t_{i}\right\rVert_{1}=\left\lVert t_{i}\right\rVert_{1}-\left\langle w_{i},t_{i}\right\rangle=c^{2}\sigma^{2}-\epsilon,$
where the second equality follows from the fact that $\mathbf{0}\in K$. We now argue that there exists a convex combination $t_{\alpha}:=\sum\limits_{i\in J}\alpha_{i}t_{i}$ such that $\left\lVert t_{\alpha}\right\rVert_{\infty}\leq\frac{c^{2}\sigma^{2}}{k}$. First, we observe that since all $t_{i}$ have nonnegative entries, we have $\left\lVert t_{\alpha}\right\rVert_{\infty}=\sup\limits_{r:r\geq 0,\left\lVert r\right\rVert_{1}=1}\langle r,t_{\alpha}\rangle$. Hence, it suffices to show that
$\inf_{\alpha:\alpha\geq 0,\left\lVert\alpha\right\rVert_{1}=1}\sup_{r:r\geq 0,\|r\|_{1}=1}\langle r,t_{\alpha}\rangle\leq\frac{c^{2}\sigma^{2}}{k}.$
Since both sets over which the optimization is performed are convex, and the function $\langle r,t_{\alpha}\rangle=\sum\limits_{i\in J}\alpha_{i}\langle r,t_{i}\rangle$ is convex-concave (in fact, it is linear in both arguments), by the minimax theorem, we have
$\inf_{\alpha:\alpha\geq 0,\left\lVert\alpha\right\rVert_{1}=1}\sup_{r:r\geq 0,\left\lVert r\right\rVert_{1}=1}\langle r,t_{\alpha}\rangle=\sup_{r:r\geq 0,\left\lVert r\right\rVert_{1}=1}\inf_{\alpha:\alpha\geq 0,\left\lVert\alpha\right\rVert_{1}=1}\langle r,t_{\alpha}\rangle=\sup_{r:r\geq 0,\left\lVert r\right\rVert_{1}=1}\min_{i\in J}\langle r,t_{i}\rangle.$
Let $U$ denote the index set of the largest k1k-1 entries of $r$. Observe that $\min\limits_{i\in J}\langle r,t_{i}\rangle\leq\langle r,t_{r}\rangle$, where $t_{r}$ is chosen such that $t_{r,U}=\mathbf{0}$. Consequently,
$\langle r,t_{r}\rangle=\sum\limits_{i\not\in U}r_{i}t_{r,i}\leq\left\lVert r_{\backslash U}\right\rVert_{\infty}\left\lVert t_{r}\right\rVert_{1}=\left\lVert r_{\backslash U}\right\rVert_{\infty}(c^{2}\sigma^{2}-\varepsilon),$
[…]
$\left\lVert t^{\star}\right\rVert_{\infty}\leq\frac{c^{2}\sigma^{2}-\epsilon}{k}<\frac{c^{2}\sigma^{2}}{k}.$
In addition, since $t^{*}$ is a convex combination of vectors $t_{i}$, we must have $\left\lVert t^{\star}\right\rVert_{1}=c^{2}\sigma^{2}-\varepsilon$. Taking $\varepsilon\rightarrow 0$ and noticing $\left\lVert t^{\star}\right\rVert_{\infty}=\left\lVert\theta^{\star}\right\rVert_{\infty}^{2},\left\lVert t^{\star}\right\rVert_{1}=\left\lVert\theta^{\star}\right\rVert_{2}^{2}$ complete the proof. ∎
B.2 Proof of Theorem 3.2
Proof of Theorem 3.2.
We first assume $k^{\star}\geq 1$, and start by a classical treatment of minimax testing lower bounds which we take from Baraud, (2002). Let $\nu$ denote an arbitrary prior on the set $\left\{\mu\left\lvert\right.\mu\in K,\left\lVert\mu\right\rVert_{2}\geq\rho\right\}$, and $\mathbb{P}_{\nu}$ denote the marginal joint distribution of $X_{1},X_{2},\dots,X_{N}$ given the prior $\nu$. We have that
$\inf_{\psi:\mathbb{P}_{0}(\psi=1)\leq\alpha}\sup_{\mu\in K,\left\lVert\mu\right\rVert_{2}\geq\rho}\mathbb{P}_{\mu}(\psi=0)$ $\geq\inf_{\psi:\mathbb{P}_{0}(\psi=1)\leq\alpha}\mathbb{P}_{\nu}(\psi=0)$
$\geq\inf_{\psi:\mathbb{P}_{0}(\psi=1)\leq\alpha}1-\mathbb{P}_{\nu}(\psi=1)$
$\geq\inf_{\psi:\mathbb{P}_{0}(\psi=1)\leq\alpha}1-\mathbb{P}_{0}(\psi=1)+\left(\mathbb{P}_{0}(\psi=1)-\mathbb{P}_{\nu}(\psi=1)\right)$
$\geq 1-\alpha-\text{TV}\left(\mathbb{P}_{0},\mathbb{P}_{\nu}\right)$
$\overset{\text{(i)}}{=}1-\alpha-\frac{1}{2}\left\lVert\mathbb{P}_{0}-\mathbb{P}_{\nu}\right\rVert_{1},$
where (i) uses (A.1). Now assuming $\mathbb{P}_{\nu}\ll\mathbb{P}_{0}$, and let $L_{\nu}=\frac{d\mathbb{P}_{\nu}}{d\mathbb{P}_{0}}$ be the Radon-Nikodym derivative of $\mathbb{P}_{\nu}$ to $\mathbb{P}_{0}$. By Cauchy-Schwartz inequality, we have
$\left\lVert\mathbb{P}_{0}-\mathbb{P}_{\nu}\right\rVert_{1}=\mathbb{E}_{0}\left|L_{\nu}-1\right|\leq\sqrt{\mathbb{E}_{0}\left(L_{\nu}-1\right)^{2}}=\sqrt{\mathbb{E}_{0}\left[L_{\nu}^{2}\right]-1}.$
Hence we conclude
$\inf_{\psi:\mathbb{P}_{0}(\psi=1)\leq\alpha}\sup_{\mu\in K,\left\lVert\mu\right\rVert_{2}\geq\rho}\mathbb{P}_{\mu}(\psi=0)\geq 1-\alpha-\frac{1}{2}\sqrt{\mathbb{E}_{0}\left[L_{\nu}^{2}\right]-1}.$
Let $\theta$ be the vector from Lemma 3.1 when $k=k_{1}^{\star}$, which means $\left\lVert\theta\right\rVert_{2}=\frac{\left(k_{1}^{\star}\right)^{1/4}}{\sqrt{N}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{1}{\sqrt{N}\left(k_{1}^{\star}\right)^{1/4}}\sigma$. We will use the prior $\left(\kappa\gamma_{i}\theta_{i}\right)_{i\in I}\in K$, where $\gamma_{i}$ are i.i.d. Rademacher random variables for $i\in I$ [^7] I=[d]I=[d] when $K\subset\mathbb{R}^{d}$ and $I=\mathbb{N}$ when $K\subset\ell_{2}$. and $0<\kappa<1$ is some small constant. Observe that all these vectors have squared Euclidean norm equal to $\rho^{2}=\frac{\kappa^{2}\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$.
By the definition of $\mathbb{P}_{\nu}$, calculating the $\chi^{2}$-divergence of this mixture yields:
$\mathbb{E}_{0}(L_{\nu})^{2}-1$ $=\mathbb{E}_{X_{1},X_{2},\dots,X_{N}\sim\text{i.i.d.}\mathcal{N}(0,\sigma^{2}\mathbf{I}_{d})}\left(\frac{d\mathbb{P}_{\nu}}{d\mathbb{P}_{0}}\right)^{2}-1$
$=\mathbb{E}_{\mu,\mu^{\prime}}\mathbb{E}_{X_{1},X_{2},\dots,X_{N}\sim\text{i.i.d.}\mathcal{N}(0,\sigma^{2}\mathbf{I}_{d})}\frac{\prod\limits_{i=1}^{N}\exp\left\{\frac{-\left\lVert X_{i}-\mu\right\rVert_{2}^{2}}{2\sigma^{2}}\right\}\prod\limits_{i=1}^{N}\exp\left\{-\frac{\left\lVert X_{i}-\mu^{\prime}\right\rVert_{2}^{2}}{2\sigma^{2}}\right\}}{\prod\limits_{i=1}^{N}\exp\left\{-\frac{\left\lVert X_{i}\right\rVert_{2}^{2}}{\sigma^{2}}\right\}}-1$
$=\mathbb{E}_{\mu,\mu^{\prime}}\exp\left\{\frac{N\mu^{\top}\mu^{\prime}}{\sigma^{2}}\right\}-1,$
where by $\mathbb{E}_{\mu,\mu^{\prime}}$ we mean expectation over two i.i.d. draws from the prior distribution $\nu$ that we specified above. Using the symmetry of Rademacher random variable, we have
$\mathbb{E}_{\mu,\mu^{\prime}}\exp\left\{\frac{N\mu^{\top}\mu^{\prime}}{\sigma^{2}}\right\}-1$ $=\mathbb{E}_{\mu,\mu^{\prime}}\exp\left\{N\sum\limits_{i\in I}\frac{\mu_{i}\mu_{i}^{\prime}}{\sigma^{2}}\right\}-1$
$=\prod_{i\in I}\cosh\left(\frac{N\kappa^{2}\theta_{i}^{2}}{\sigma^{2}}\right)-1$
$\leq\exp\left\{\sum_{i\in I}\frac{N^{2}\theta_{i}^{4}\kappa^{4}}{2\sigma^{4}}\right\}-1$
$\leq\exp\left\{\frac{N^{2}\left\lVert\theta\right\rVert_{\infty}^{2}\left\lVert\theta\right\rVert_{2}^{2}\kappa^{4}}{2\sigma^{4}}\right\}-1$
$\leq\exp\left\{\frac{\kappa^{4}}{2}\right\}-1,$
where we used the well known inequality $\cosh(x)\leq\exp\left\{\frac{x^{2}}{2}\right\}$, which can be checked simply by comparing the Taylor series, and the properties of $\theta$ from Lemma 3.1. It follows that the minimax risk for testing is lower bounded by
$1-\alpha-\frac{1}{2}\sqrt{\exp\left\{\frac{\kappa^{4}}{2}\right\}-1}.$
Recall that the $\ell_{2}$ norm of the selected $\mu$ is $\rho=\left\lVert\mu\right\rVert_{2}=\frac{(k_{1}^{\star})^{1/4}\kappa}{\sqrt{N}}\sigma$. If we let the expression above equal to $\alpha$ and solve for $\kappa$, we obtain $\kappa=c(\alpha)$. Consequently, we conclude that if $\kappa\leq c(\alpha)$, and therefore $\rho\leq c(\alpha)\cdot\frac{(k_{1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, testing with uniform power $\alpha$ is not possible.
The remaining corner case is when $k_{1}^{\star}=0$, or equivalently $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}\leq\frac{1}{\sqrt{N}}\sigma$. In this case, any point $\theta\in K$ satisfies $\left\lVert\theta\right\rVert_{2}\leq\frac{1}{\sqrt{N}}\sigma$. Taking a point $\theta_{0}$ that (nearly) achieves the equality, and considering the mixture formed by $\gamma\theta_{0}$, where $\gamma$ is a Rademacher random variable, one obtains that the resulting $\chi^{2}$-divergence is bounded by $\cosh\left(\frac{N\left\lVert\theta\right\rVert_{2}^{2}}{\sigma^{2}}\right)-1\leq\exp\left\{\frac{N^{2}\left\lVert\theta\right\rVert_{2}^{4}}{2\sigma^{4}}\right\}-1\leq\sqrt{e}-1$ and consequently
$\inf_{\psi:\mathbb{P}_{0}(\psi=1)\leq\alpha}\sup_{\mu\in K}\mathbb{P}_{\mu}(\psi=0)\geq 1-\alpha-\frac{1}{2}\sqrt{\sqrt{e}-1},$
implying that it is impossible to test the problem. Hence, $k_{1}^{\star}$ must be at least positive for sufficiently small $\alpha$. ∎
B.3 Proof of Theorem 3.3
In this section, we prove Theorem 3.3. Our proof relies on Lemma B.1 and a carefully designed optimal coupling. To establish Lemma B.1, we first recall the notion of differential privacy in B.2 and present two auxiliary lemmas, Lemma B.3 and Lemma B.4. Based on these tools, we then construct a symmetric set using the QCO properties and compute the Hamming distance as an application of Lemma A.5. The computation proceeds by constructing a sequence of couplings that reduces the evaluation of the Hamming distance into that of the total variation distance.
Lemma B.1 (main lemma).
Suppose $\mathbf{X}=(X_{1},X_{2},\dots,X_{N})^{\top}$ and $\mathbf{X}^{\prime}=(X^{\prime}_{1},X_{2}^{\prime},\dots,X_{N}^{\prime})^{\top}$ are random matrices with rows $X_{i},X^{\prime}_{i}\in\mathbb{R}^{d}$. Assume the laws of $\mathbf{X}$ and $\mathbf{X}^{\prime}$ are $\mathcal{U}$ and $\mathcal{V}$, respectively. Additionally, assume that there exists a coupling between $\mathcal{U}$ and $\mathcal{V}$ under which the expectation of the Hamming distance between $\mathbf{X}$ and $\mathbf{X}^{\prime}$ is $\mathcal{O}(\epsilon N)$, i.e.,
$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\lesssim\epsilon N.$ (B.1)
Then, there exists no robust test capable of distinguishing between $\mathcal{U}$ and $\mathcal{V}$ under contamination level of $\varOmega(\epsilon)$ while achieving type-I and type-II error probabilities below $0.1$ simultaneously and uniformly.
To prove Lemma B.1, we start with the notion of $(\epsilon,\delta)$-differential privacy.
Definition B.2 ($(\epsilon,\delta)$-differential privacy).
For a pair of non-negative real numbers $(\epsilon,\delta)$, a randomized algorithm $\mathcal{A}$ that is defined on the sample space $\Omega$ is a $(\epsilon,\delta)$-differentially private algorithm if for any $S\subset\text{range}(\mathcal{A})$ and any $X,X^{\prime}\in\Omega$ with $d_{H}(X,X^{\prime})\leq 1$, we have
$\mathbb{P}\left(\mathcal{A}(X)\in S\right)\leq e^{\epsilon}\mathbb{P}\left(\mathcal{A}(X^{\prime})\in S\right)+\delta.$ (B.2)
When $\delta=0$, the algorithm is called pure differential privacy; when $\epsilon=0$, the algorithm is called $\delta$-fully approximate differential privacy.
With this concept and motivated by Theorem 11 in Acharya et al., (2018), we are able to prove a lemma that is very close to Lemma B.1.
Lemma B.3.
Assume that there is a coupling between $\mathcal{U}$ and $\mathcal{V}$ under which that the expectation of the Hamming distance between $\mathbf{X}$ and $\mathbf{X}^{\prime}$ is smaller or equal than $D$, i.e.,
$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\leq D.$ (B.3)
Then, if a robust test $\mathcal{A}:\Omega\mapsto\left\{0,1\right\}$ between $\mathcal{U}$ and $\mathcal{V}$ is a $\delta$-fully approximate differential private algorithm, with type-I and type-II error probabilities both below $0.1$, then we must have $D=\varOmega(\frac{1}{\delta})$.
Proof.
By the assumption of the lemma, suppose that $(\mathbf{X},\mathbf{X}^{\prime})$ is the random variable from the coupling. We have $\mathbb{P}\left(\mathcal{A}(\mathbf{X})=0\right)\geq 0.9$ and $\mathbb{P}\left(\mathcal{A}(\mathbf{X}^{\prime})=1\right)\geq 0.9$. By Markov’s inequality, we have $\mathbb{P}\left(d_{H}(\mathbf{X},\mathbf{X}^{\prime})>10D\right)\leq\frac{\mathbb{E}d_{H}(\mathbf{X},\mathbf{X}^{\prime})}{10D}\leq 0.1$. From the union bound, we obtain that
$\mathbb{P}\left(\mathcal{A}(\mathbf{X})=0\cap\mathcal{A}(\mathbf{X}^{\prime})=1\cap d_{H}(\mathbf{X},\mathbf{X}^{\prime})<10D\right)\geq 0.7.$ (B.4)
By the definition of differential privacy, we can construct a chain from $\mathbf{X}$ to $\mathbf{X}^{\prime}$ such that each neighbour points differ on at most one coordinate, by which we have
$|\mathbb{P}\left(\mathcal{A}(\mathbf{X})=0\right)-\mathbb{P}\left(\mathcal{A}(\mathbf{X}^{\prime})=0\right)|\leq\delta\cdot d_{H}(\mathbf{X},\mathbf{X}^{\prime}).$ (B.5)
It is worth noting that the probability in (B.4) includes two parts of randomness: the randomness of $\mathbf{X}$ and $\mathbf{X}^{\prime}$ from $\Omega$, and the randomness of $\mathcal{A}$. Now take a realization of $\mathbf{X}$ and $\mathbf{X}^{\prime}$, denoted by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, that satisfy (B.4), we have
$1-0.7\geq\mathbb{P}(\mathcal{A}(\mathbf{x}^{\prime})=0)\geq\mathbb{P}(\mathcal{A}(\mathbf{x})=0)-\delta\cdot d_{H}(\mathbf{x},\mathbf{x}^{\prime})\geq 0.7-\delta\cdot 10D,$ (B.6)
from which we obtain that $10D\geq 0.4\frac{1}{\delta}$. The proof is completed. ∎
The next lemma shows how to obtain a robust and differentially private test from a given robust test between $\mathcal{U}$ and $\mathcal{V}$ using a black-box robustness-to-privacy transformation. The techniques are from Canonne et al., (2023) and Hopkins et al., (2023).
Lemma B.4.
Suppose $\mathcal{A}$ is a robust test for the mean testing problem (1.2) achieving type-I and type-II errors at most $0.1$ under a contamination level at most $\epsilon$. Then, for $\delta=\frac{1}{\epsilon N}$, there exists a $\delta$-fully approximate differential privacy $\widetilde{\mathcal{A}}$ whose error probabilities are also at most $0.1$ for any dataset with size $N$.
Proof.
For any algorithm $\mathcal{A}$ that attempts to distinguish between $H_{0}$ and $H_{1}$ based on $\mathbf{X}$, define $\mathcal{S}(X,\mathcal{A})$ as the smallest number such that there exists another dataset $\mathbf{X}^{\prime}$ with $\mathcal{A}(\mathbf{X}^{\prime})=1$ and $d_{H}(\mathbf{X},\mathbf{X}^{\prime})=\mathcal{S}(\mathbf{X},\mathcal{A})$. Using this quantity, we construct a new algorithm $\widetilde{\mathcal{A}}$ that outputs $1$ with probability $\max\left(0,1-\delta\cdot\mathcal{S}(X,\mathcal{A})\right)$. We now prove that $\widetilde{\mathcal{A}}$ is the desired algorithm.
First, we notice that for any two datasets $\mathbf{X}^{\prime}$, $\mathbf{X}^{\prime\prime}$ that satisfy $d_{H}(\mathbf{X}^{\prime},\mathbf{X}^{\prime\prime})\leq 1$, we always have $|\mathcal{S}(X^{\prime},\mathcal{A})-\mathcal{S}(X^{\prime\prime},\mathcal{A})|\leq 1$ for any algorithm $\mathcal{A}$. This is foreseen since if we denote $\mathbf{X}^{{\dagger}}$ as the specific dataset that satisfies $\mathcal{A}(\mathbf{X}^{{\dagger}})=1$ and $d_{H}(\mathbf{X}^{\prime},\mathbf{X}^{{\dagger}})=\mathcal{S}(\mathbf{X}^{\prime},\mathcal{A})$, then by the triangular inequality we have $d_{H}(\mathbf{X}^{\prime\prime},\mathbf{X}^{{\dagger}})\leq\mathcal{S}(\mathbf{X}^{\prime},\mathcal{A})+1$, and consequently $\mathcal{S}(\mathbf{X}^{\prime\prime},\mathcal{A})\leq\mathcal{S}(\mathbf{X}^{\prime},\mathcal{A})+1$. In the same way we also have $\mathcal{S}(\mathbf{X}^{\prime\prime},\mathcal{A})\geq\mathcal{S}(\mathbf{X}^{\prime},\mathcal{A})-1$. The statement is proved.
By the statement above we know that the constructed algorithm $\widetilde{\mathcal{A}}$ is a $(0,\delta)$-differentially private algorithm since for any two dataset $\mathbf{X}$ and $\mathbf{X}^{\prime}$ with $d_{H}(\mathbf{X},\mathbf{X}^{\prime})\leq 1$, we have
$|\mathbb{P}\left(\widetilde{\mathcal{A}}(\mathbf{X})=1\right)-\mathbb{P}\left(\widetilde{\mathcal{A}}(\mathbf{X}^{\prime})=1\right)|$ $=|\max\left\{0,1-\delta\cdot\mathcal{S}(\mathbf{X},\mathcal{A})\right\}-\max\left\{0,1-\delta\cdot\mathcal{S}(\mathbf{X}^{\prime},\mathcal{A})\right\}|$
$\leq|\delta\cdot\mathcal{S}(\mathbf{X},\mathcal{A})-\delta\cdot\mathcal{S}(\mathbf{X}^{\prime},\mathcal{A})|$
$\leq\delta.$
Now assume $\widetilde{\mathbf{X}}$ is from $H_{0}$. By the definition of robust testing, with probability at least $0.9$, we can arbitrarily modify $\widetilde{\mathbf{X}}$ and obtain the same result as long as the number of contamination does not exceed $\epsilon N$. This equivalently says that $\mathcal{S}(\widetilde{\mathbf{X}},\mathcal{A})\geq\epsilon N$, which, by the construction of $\widetilde{\mathcal{A}}$, implies that $\widetilde{\mathcal{A}}$ will output $1$ with probability 0 in this case. On the other hand, if $\widetilde{\mathbf{X}}$ is from $H_{1}$, then with probability at least $0.9$, we will have $\mathcal{A}(X)=1$, in which case $\mathcal{S}(X,\mathcal{A})=0$ and consequently $\widetilde{\mathcal{A}}$ will output $1$ with probability $1$. This shows that $\widetilde{\mathcal{A}}$ is a $\delta$-fully approximate differential privacy with error probabilities at most $0.1$ under both hypotheses. ∎
We are now prepared to prove Lemma B.1.
Proof of Lemma B.1.
Suppose we already have a robust test $\mathcal{A}$ for the mean testing problem 1.2 with the fraction of contamination at most $\epsilon$. Then by Lemma B.4, we can contruct a $\delta$-fully approximate differential privacy $\widetilde{\mathcal{A}}$ with errors at most $0.1$. Therefore, the constructed algorithm $\widetilde{\mathcal{A}}$ satifies the condition of Lemma B.3. Take $D=\mathbb{E}d_{H}(\mathbf{X},\mathbf{X}^{\prime})$ in Lemma B.3, and by the conclusion of Lemma B.3 we should have $\mathbb{E}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\gtrsim\frac{1}{1/\epsilon N}=\epsilon N$. This necessary condition finishes the proof. ∎
Proof of Theorem 3.3.
Take $\mathcal{U}$ to be the law of $N$ i.i.d. samples from $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$ and $\mathcal{V}$ to be the law of $N$ i.i.d. samples from $\mathcal{N}\left(v,\sigma^{2}\mathbf{I}\right)$ with $v$ drawn from the uniform distribution on the set
$V:=\left\{v\left\lvert\right.v_{i}=\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},1\leq i\in I\right\},$
where $\theta$ is the vector from Lemma 3.1 setting $c=\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$, i.e., $\left\lVert\theta\right\rVert_{2}=\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{\sqrt{\epsilon}}{(k_{2}^{\star})^{1/4}N^{1/4}}\sigma$. We are going to compute $d_{H}(\mathbf{X},\mathbf{X}^{\prime})$ using Lemma A.5. Specifically, we compare (in terms of total variation) a distribution of $k$ exchangeable variables generated from $\mathcal{V}$ to a joint distribution of k1k-1 exchangeable variables generated from $\mathcal{N}\left(v,\sigma^{2}\mathbf{I}\right)$ and the last one from $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$ for $1\leq k\leq N$. We approach this by decomposing both $\mathbf{X}$ and $\mathbf{X}^{\prime}$ by their sequence of conditional probabilities and constructing the optimal coupling between $X_{k}$ and $X^{\prime}_{k}$ conditional on $X_{1},\dots,X_{k-1}$ using Lemma A.5 such that
$\mathbb{P}\left(X_{k}\neq X^{\prime}_{k}|X_{1},\dots,X_{k-1}\right)=\text{TV}(p(x_{k}|x_{1},\dots,x_{k-1}),q(x_{k})),$
where $p(x_{k}|x_{1},\dots,x_{k-1})$ is the density of conditional distribution of the first $k$ coordinates of $X^{\prime}$ and $q(x_{k})$ is the density of standard normal distribution. By the definition of total variation distance, it is not hard to see under this optimal coupling that
$\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim(\mathcal{U},\mathcal{V})}\mathbf{1}_{\left\{X_{k}\neq X^{\prime}_{k}\right\}}$ $=\mathbb{E}\text{TV}(p(x_{k}|x_{1},\dots,x_{k-1}),q(x_{k}))$
$=\text{TV}(p(x_{1},\dots,x_{k}),q(x_{1},\dots,x_{k})),$
where $p(x_{1},\dots,x_{k})$ is the joint distribution of the first $k$ coordinates of $X^{\prime}$ and $q(x_{1},\dots,x_{k}):=p(x_{1},\dots,x_{k-1})\cdot q(x_{k})$. Since
$\mathbb{E}_{(X,X^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(X,X^{\prime})=\mathbb{E}_{X,X^{\prime}\sim(\mathcal{U},\mathcal{V})}\sum\limits_{i=1}^{N}\mathbf{1}_{\left\{X_{i}\neq X^{\prime}_{i}\right\}},$
to evaluate $\mathbb{E}_{(X,X^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(X,X^{\prime})$, it suffices to evaluate
$\sum\limits_{k=1}^{N}\text{TV}(p(x_{1},\dots,x_{k}),q(x_{1},\dots,x_{k})).$
We calculate each individual term. By Pinsker’s inequality A.3, we have
$2\text{TV}^{2}(p(x_{1},\dots,x_{k}),q(x_{1},\dots,x_{k}))$ $\leq\text{KL}(p(x_{1},\dots,x_{k})|q(x_{1},\dots,x_{k}))$
$=\mathbb{E}\ln\left(\frac{\frac{1}{2^{|V|}}\sum\limits_{v\in V}\prod\limits_{i=1}^{k}\exp\left\{-\frac{\left\lVert X_{i}-v\right\rVert_{2}^{2}}{2\sigma^{2}}\right\}}{\frac{1}{2^{|V|}}\sum\limits_{v\in V}\prod\limits_{i=1}^{k-1}\exp\left\{-\frac{\left\lVert X_{i}-v\right\rVert_{2}^{2}}{2\sigma^{2}}\right\}\cdot\exp\left\{-\frac{\left\lVert X_{k}\right\rVert_{2}^{2}}{2\sigma^{2}}\right\}}\right)$
$\overset{\text{(i)}}{=}\mathbb{E}\ln\left(\frac{\frac{1}{2^{|V|}}\sum\limits_{v\in V}\exp\left\{\left(\sum\limits_{i=1}^{k}X_{i}\right)^{\top}v/\sigma^{2}\right\}}{\frac{1}{2^{|V|}}\sum\limits_{v\in V}\exp\left\{\left(\sum\limits_{i=1}^{k-1}X_{i}\right)^{\top}v/\sigma^{2}\right\}}\right)-\frac{\left\lVert v\right\rVert_{2}^{2}}{2\sigma^{2}}$
$\overset{\text{(ii)}}{=}\mathbb{E}\ln\left(\frac{\prod\limits_{j=1}^{d}\cosh\left\{\left(\sum\limits_{i=1}^{k}X_{i}\right)_{j}v_{j}/\sigma^{2}\right\}}{\prod\limits_{j=1}^{d}\cosh\left\{\left(\sum\limits_{i=1}^{k-1}X_{i}\right)_{j}v_{j}/\sigma^{2}\right\}}\right)-\frac{\left\lVert v\right\rVert_{2}^{2}}{2\sigma^{2}},$
where $(\cdot)_{j}$ denotes the $j$-th coordinate, (i) is from that $\left\lVert v\right\rVert_{2}^{2}$ does not depend on the specfic choice of $v\in V$ by the construction of $V$ and (ii) is from an important observation that
$\frac{1}{2^{|V|}}\sum\limits_{v\in V}\exp\left\{\left(\sum\limits_{i=1}^{k}X_{i}\right)^{\top}v\right\}=\prod\limits_{j=1}^{d}\cosh\left\{\left(\sum\limits_{i=1}^{k}X_{i}\right)_{j}v_{j}\right\}.$
Therefore, we have
$\sum\limits_{k=1}^{N}\text{KL}(p(x_{1},\dots,x_{k})|q(x_{1},\dots,x_{k}))$ $=\sum\limits_{k=1}^{N}\mathbb{E}\left[\ln\left(\frac{\prod\limits_{j=1}^{d}\cosh\left\{\left(\sum\limits_{i=1}^{k}X_{i}\right)_{j}v_{j}/\sigma^{2}\right\}}{\prod\limits_{j=1}^{d}\cosh\left\{\left(\sum\limits_{i=1}^{k-1}X_{i}\right)_{j}v_{j}/\sigma^{2}\right\}}\right)-\frac{\left\lVert v\right\rVert_{2}^{2}}{2\sigma^{2}}\right]$
$\overset{\text{(iii)}}{=}\sum\limits_{j=1}^{d}\mathbb{E}\ln\left(\cosh\left\{\left(\sum\limits_{i=1}^{N}X_{i}\right)_{j}v_{j}/\sigma^{2}\right\}\right)-N\cdot\frac{\left\lVert v\right\rVert_{2}^{2}}{2\sigma^{2}},$
where (iii) is from telescoping and the exchange of summation and expectation. Using the inequality that $\ln\left(\cosh x\right)\leq\frac{x^{2}}{2},\forall x\in\mathbb{R}$, which can be easily proved by inspecting the Taylor’s expansion of $\ln\left(\cosh x\right)$, and the fact that $\left(\sum\limits_{i=1}^{N}X_{i}\right)_{j}\sim\mathcal{N}\left(Nv_{j},N\sigma^{2}\right)$ for any $j$, we can bound the expression above by
$\sum\limits_{k=1}^{N}\text{KL}(p(x_{1},\dots,x_{k})|q(x_{1},\dots,x_{k}))$ $\leq\frac{1}{2}\sum\limits_{j=1}^{d}\mathbb{E}\left[\left(\sum\limits_{i=1}^{N}X_{i}\right)_{j}v_{j}/\sigma^{2}\right]^{2}-N\cdot\frac{\left\lVert v\right\rVert_{2}^{2}}{2\sigma^{2}}$
$\leq\frac{1}{2}\sum\limits_{j=1}^{d}\frac{v_{j}^{2}}{\sigma^{4}}\left(N^{2}v_{j}^{2}+N\sigma^{2}\right)-N\cdot\frac{\left\lVert v\right\rVert_{2}^{2}}{2\sigma^{2}}$
$=\frac{N^{2}\left\lVert v\right\rVert_{4}^{4}}{2\sigma^{4}}.$
Combining all of the expressions and inequalities above together, we arrive at the following
$\frac{1}{N}\sum\limits_{k=1}^{N}\text{TV}(p(x_{1},\dots,x_{k})|q(x_{1},\dots,x_{k}))$ $\lesssim\frac{1}{N}\sum\limits_{k=1}^{N}\sqrt{\text{KL}(p(x_{1},\dots,x_{k})|q(x_{1},\dots,x_{k}))}$
$\overset{\text{(i)}}{\leq}\sqrt{\frac{1}{N}\cdot\sum\limits_{k=1}^{N}\text{KL}(p(x_{1},\dots,x_{k})|q(x_{1},\dots,x_{k}))}$
$\lesssim\frac{\sqrt{N}\left\lVert v\right\rVert_{4}^{2}}{\sigma^{2}},$
where (i) follows from the Jensen’s inequality. Recall that by Lemma 3.1 we have $\left\lVert\theta\right\rVert_{2}=\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{\sqrt{\epsilon}}{(k_{2}^{\star})^{1/4}N^{1/4}}\sigma$. Now that $\theta_{i}^{4}\leq\theta_{i}^{2}\cdot\left\lVert\theta\right\rVert_{\infty}^{2}$ for all $i\in\left\{1,2,\dots,d\right\}$, we know $\left\lVert\theta\right\rVert_{4}^{2}\leq\left\lVert\theta\right\rVert_{2}\left\lVert\theta\right\rVert_{\infty}\leq\frac{\epsilon}{\sqrt{N}}\sigma^{2}$. By the definition of $V$, we know that $\left\lVert v\right\rVert_{4}^{2}=\left\lVert\theta\right\rVert_{4}^{2}$. Therefore, we have $\frac{\sqrt{N}\left\lVert v\right\rVert_{4}^{2}}{\sigma^{2}}\leq\frac{\sqrt{N}}{\sigma^{2}}\cdot\frac{\epsilon}{\sqrt{N}}\sigma^{2}\leq\epsilon$. By Lemma B.1, we have
$\inf\limits_{\phi\in A_{s}(0.1)}\sup\limits_{\mu\in K}\sup\limits_{\mathcal{C}}\mathbb{P}_{\mu}(\phi(\mathcal{C}(\widetilde{\mathbf{X}}))=0)\geq 0.1.$
The proof is completed. ∎
B.4 Proof of Theorem 3.4
In this section, we prove Theorem 3.4. We reach the result by constructing two mixture distributions which are, in fact, identical. We then evaluate the total variation distance between two truncated binomial distributions, which enables us to construct an adaptive adversary, leading high type-II error. Similar techniques also appear in Prasadan and Neykov, (2025) and Chen et al., (2017).
Proof of Theorem 3.4.
We prove the theorem by several steps.
Step I: control the TV distance. Set $\epsilon^{\prime}=\epsilon-1/\sqrt{2N}$. Since $\epsilon\gtrsim 1/\sqrt{N}$, we have $\epsilon\asymp\epsilon^{\prime}$. Now fix a $\mu_{0}\in K$ such that $\left\lVert\mu_{0}\right\rVert_{2}^{2}\leq\frac{4\sigma^{2}\epsilon^{\prime 2}}{(1-\epsilon^{\prime})^{2}}$. Consider two Gaussian measures $\mathbb{P}_{0}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I}_{d})$ and $\mathbb{P}_{\mu_{0}}\sim\mathcal{N}(\mu_{0},\sigma^{2}\mathbf{I}_{d})$. By the Pinsker’s inequality we have $\text{TV}\left(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}}\right)\leq\sqrt{\frac{1}{2}D_{\text{KL}}\left(\mathbb{P}_{0}\left\lVert\right.\mathbb{P}_{\mu_{0}}\right)}=\frac{\epsilon^{\prime}}{1-\epsilon^{\prime}}$. Since $r(x):=\frac{x}{1-x}$ is an increasing function for $x\in(0,1)$, there exists some $\epsilon^{\prime\prime}\in(0,\epsilon^{\prime})$ such that $\text{TV}\left(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}}\right)=\frac{\epsilon^{\prime\prime}}{1-\epsilon^{\prime\prime}}$.
Step II: construct a pair of identical mixtures. Let $p_{0}$ and $p_{\mu_{0}}$ be the probability density functions of $\mathbb{P}_{0}$ and $\mathbb{P}_{\mu_{0}}$, construct two distributions $Q_{1},Q_{2}$ with density
$q_{1}=\frac{(p_{\mu_{0}}-p_{0})\mathbf{1}_{\left\{p_{\mu_{0}}\geq p_{0}\right\}}}{\text{TV}\left(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}}\right)},\ q_{2}=\frac{(p_{0}-p_{\mu_{0}})\mathbf{1}_{\left\{p_{0}>p_{\mu_{0}}\right\}}}{\text{TV}(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}})}.$ (B.7)
Consider the mixture measures
$R_{1}:=(1-\epsilon^{\prime\prime})\mathbb{P}_{0}+\epsilon^{\prime\prime}Q_{1},\ R_{2}:=(1-\epsilon^{\prime\prime})\mathbb{P}_{\mu_{0}}+\epsilon^{\prime\prime}Q_{2}.$ (B.8)
It is not hard to verify that actually $R_{1}=R_{2}$, i.e., they represent the same distribution. Consequently $\text{TV}\left(\prod\limits_{i=1}^{N}R_{1}^{i},\prod\limits_{i=1}^{N}R_{2}^{i}\right)=0$, where $\prod\limits_{i=1}^{N}R_{j}^{i},j\in\left\{1,2\right\}$ is the $N$-fold product of $R_{j}$.
Step III: expand the TV distance. Define
$\binom{[N]}{s}:=\left\{(i_{1},\dots,i_{s}):i_{1},\dots,i_{s}\in\left\{1,2,\dots,N\right\},1\leq i_{1}<i_{2}<\dots<i_{s}\leq N\right\}.$
Further define the quantities
$S_{1}^{L}=\sum\limits_{s=0}^{N\left(\epsilon^{\prime\prime}+\frac{1}{\sqrt{2N}}\right)}\sum\limits_{I\in\binom{[N]}{s}}(1-\epsilon^{\prime\prime})^{N-s}\epsilon^{\prime\prime s}\prod\limits_{i\in[N]\backslash I}p_{0}(x_{i})\prod\limits_{i\in I}q_{1}(x_{i}),$
$S_{2}^{L}=\sum\limits_{s=0}^{N\left(\epsilon^{\prime\prime}+\frac{1}{\sqrt{2N}}\right)}\sum\limits_{I\in\binom{[N]}{s}}(1-\epsilon^{\prime\prime})^{N-s}\epsilon^{\prime\prime s}\prod\limits_{i\in[N]\backslash I}p_{\mu_{0}}(x_{i})\prod\limits_{i\in I}q_{2}(x_{i}),$
$S_{1}^{U}=\sum\limits_{s=N\left(\epsilon^{\prime\prime}+\frac{1}{\sqrt{2N}}\right)+1}^{N}\sum\limits_{I\in\binom{[N]}{s}}(1-\epsilon^{\prime\prime})^{N-s}\epsilon^{\prime\prime s}\prod\limits_{i\in[N]\backslash I}p_{0}(x_{i})\prod\limits_{i\in I}q_{1}(x_{i}),$
$S_{2}^{U}=\sum\limits_{s=N\left(\epsilon^{\prime\prime}+\frac{1}{\sqrt{2N}}\right)+1}^{N}\sum\limits_{I\in\binom{[N]}{s}}(1-\epsilon^{\prime\prime})^{N-s}\epsilon^{\prime\prime s}\prod\limits_{i\in[N]\backslash I}p_{\mu_{0}}(x_{i})\prod\limits_{i\in I}q_{2}(x_{i}),$
where without loss of generality we assume that all bounds of summation are integers. For $S_{1}^{U}$ and $S_{2}^{U}$, we have
$\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{U}-S_{2}^{U}\right|d\mathbf{x}\leq\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}(S_{1}^{U}+S_{2}^{U})d\mathbf{x}\leq\mathbb{P}\left(B>N\left(\epsilon^{\prime\prime}+\frac{1}{\sqrt{2N}}\right)\right)\overset{\text{(i)}}{\leq}e^{-1}.$
Here $B$ is a random variable that follows a binomial distribution with parameters $N$ and $\epsilon^{\prime\prime}$, and (i) follows from the Hoeffding’s inequality.
By expanding the TV distance between the $N$-fold products of $R_{1}$ and $R_{2}$ and after some calculation, we have
$0=\text{TV}\left(\prod\limits_{i=1}^{N}R_{1}^{i},\prod\limits_{i=1}^{N}R_{2}^{i}\right)\overset{\text{(ii)}}{\geq}\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{L}-S_{2}^{L}\right|d\mathbf{x}-\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{U}-S_{2}^{U}\right|d\mathbf{x},$
where (ii) is from the definition of TV distance and the triangle inequality. Combining with previous argument, we obtain
$\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{L}-S_{2}^{L}\right|d\mathbf{x}\leq e^{-1}.$
Step IV: conditional argument. Consider the adversary $\mathcal{C}_{B,\epsilon^{\prime\prime}}$ with following strategy. Upon observing the original data $\widetilde{\mathbf{X}}$ and its underlying law (which is either $\mathbb{P}_{0}$ or $\mathbb{P}_{\mu_{0}}$ in our case), the adversary first samples a number $n_{c}$ from the conditional law of $B|B\leq N(\epsilon^{\prime\prime}+1/\sqrt{2N})$, then sample $n_{c}$ replacements from the mixture distribution corresponding to the true underlying law. Specifically, if $\widetilde{\mathbf{X}}$ is from $\mathbb{P}_{0}$, then replacements will be from $Q_{1}$, otherwise the replacements will be from $Q_{2}$. In this way, if we denote $\widetilde{\mathbb{P}}_{0}$ to be the law of the contaminated samples from $\mathbb{P}_{0}$ and $Q_{1}$, and $\widetilde{\mathbb{P}}_{\mu_{0}}$ to be the law of the contaminated samples from $\mathbb{P}_{\mu_{0}}$ and $Q_{2}$, then with minor calculation we have
$\text{TV}(\widetilde{\mathbb{P}}_{0},\widetilde{\mathbb{P}}_{\mu_{0}})=\frac{\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{L}-S_{2}^{L}\right|d\mathbf{x}}{\mathbb{P}\left(B\leq N\left(\epsilon^{\prime\prime}+\frac{1}{\sqrt{2N}}\right)\right)}\overset{\text{(iii)}}{\leq}\frac{e^{-1}}{1-e^{-1}}<1,$
where (iii) is again from the Hoeffding’s inequality and the argument from the step III.
Step V: finish the proof. Now for any $\widetilde{\mu}$ that $\left\lVert\widetilde{\mu}\right\rVert_{2}\leq\left\lVert\mu_{0}\right\rVert_{2}=\frac{2\sigma\epsilon^{\prime}}{1-\epsilon^{\prime}}$. Following a similar argument in the proof of Theorem 3.2, for any $\phi\in A_{s}$, we have
$\sup\limits_{\mu\in K,\left\lVert\mu\right\rVert_{2}\geq\rho}\sup\limits_{\mathcal{C}}\mathbb{P}_{\mu}\left(\phi(\mathcal{C}(\mathbf{X}^{\prime}))=0\right)$ $\geq\mathbb{P}_{\widetilde{\mu}}\left(\phi(\mathcal{C}_{B,\epsilon^{\prime\prime}}(\mathbf{X}^{\prime}))=0\right)$
$=1-\mathbb{P}_{\widetilde{\mu}}\left(\phi(\mathcal{C}_{B,\epsilon^{\prime\prime}}(\mathbf{X}^{\prime}))=1\right)+\mathbb{P}_{0}\left(\phi(\mathcal{C}_{B,\epsilon^{\prime\prime}}(\mathbf{X}^{\prime}))=1\right)$
$-\mathbb{P}_{0}\left(\phi(\mathcal{C}_{B,\epsilon^{\prime\prime}}(\mathbf{X}^{\prime}))=1\right)$
$\geq 1-\alpha-\left(\mathbb{P}_{\widetilde{\mu}}\left(\phi(\mathcal{C}_{B,\epsilon^{\prime\prime}}(\mathbf{X}^{\prime}))=1\right)-\mathbb{P}_{0}\left(\phi(\mathcal{C}_{B,\epsilon^{\prime\prime}}(\mathbf{X}^{\prime}))=1\right)\right)$
$=1-\alpha-\left(\widetilde{\mathbb{P}}_{\widetilde{\mu}}\left(\phi=1\right)-\widetilde{\mathbb{P}}_{0}(\phi=1)\right)$
[…]
$\geq 1-\alpha-\text{TV}\left(\widetilde{\mathbb{P}}_{0},\widetilde{\mathbb{P}}_{\mu_{0}}\right)\asymp 1.$
Taking infimum over $\phi\in A_{s}$ on both sides completes the proof. ∎
B.5 Proof of Theorem 3.6
We start with a recall of the definition of the consistent subset, which plays a central role in the proof.
Definition B.5 (Consistent subset).
A subset $S\subset\mathbf{X}=\left\{X_{1},\dots,X_{N}\right\}$ is called a consistent subset regarding the contamination fraction $\epsilon$ and a test $\phi:2^{S}\mapsto\left\{0,1\right\}$ if
 (i), $\left|S\right|\geq(1-\epsilon)N$,
 (ii), $\phi(S^{\prime})=\phi(S)$ for any $S^{\prime}\subset S$ with $\left|S^{\prime}\right|\geq(1-2\epsilon)N$.
[…]
Proof of Theorem 3.6.
We will first prove the fact that for a properly fixed error tolerance $\alpha$, the consistent subset exists with probability at least $1-\alpha$. Specifically, consider the following $\chi^{2}$-type event
$E_{e}:=\left\{\left\lVert P^{\star}\sum\limits_{i\in T}X_{i}\right\rVert_{2}^{2}-k^{\star}\left|T\right|\sigma^{2}\geq c\left|T\right|^{2}E^{2}(\epsilon,K,N,\sigma^{2})\right\},$ (B.9)
where $T$ is any subset of $\mathbf{X}$, and $c\in\mathbb{R}$ is some universal, properly chosen constant. On the subset $T$, the test $\phi_{e}$ rejects $H_{0}$ when $E_{e}$ holds and fails to reject otherwise.
For the rest of the proof, let us first focus on $\widetilde{\mathbf{X}}$ and any its subset $\widetilde{\mathbf{X}}_{S^{\prime}}$ with $\left|S^{\prime}\right|\geq(1-2\epsilon)N$. We will prove that with $\widetilde{\mathbf{X}}$ is consistent with all such $S^{\prime}$ if (3.2) is satisfied. The key to the proof is carefully analyzing the $\ell_{2}$ norm of the sum of the projected original samples
$\left\lVert P^{\star}\sum\limits_{i\in S^{\prime}}\widetilde{X}_{i}\right\rVert_{2}^{2}=\underbrace{\left\lVert P^{\star}\sum\limits_{i\in S^{\prime}}(\widetilde{X}_{i}-\mu)\right\rVert_{2}^{2}}_{:=\text{I}}+\underbrace{\left|S^{\prime}\right|^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}}_{:=\text{II}}+2\underbrace{\left|S^{\prime}\right|\sum\limits_{i\in S^{\prime}}\left[P^{\star}(\widetilde{X}_{i}-\mu)\right]^{\top}\mu}_{:=\text{III}}.$ (B.10)
We attempt to show that the parts I and III, which are $\ell_{2}$ norm of the centralized projected vectors and the crossing term, are both asymptotically less than the part II under the condition (3.2), which vanishes under the null.
Part I. By representing $S^{\prime}=[N]\backslash\left([N]\backslash S^{\prime}\right)$, we have
$\text{I}=\underbrace{\left\lVert P^{\star}\sum\limits_{i\in[N]}(\widetilde{X}_{i}-\mu)\right\rVert_{2}^{2}}_{:=\text{I}_{1}}+\underbrace{\left\lVert P^{\star}\sum\limits_{i\in[N]\backslash S^{\prime}}(\widetilde{X}_{i}-\mu)\right\rVert_{2}^{2}}_{:=\text{I}_{2,S^{\prime}}}-2\underbrace{\sum\limits_{i\in[N]}(\widetilde{X}_{i}-\mu)^{\top}P^{\star}\sum\limits_{i\in[N]\backslash S^{\prime}}(\widetilde{X}_{i}-\mu)}_{:=\text{I}_{3,S^{\prime}}}.$
For $\text{I}_{1}$, from the property of the Gaussian distribution, we know $\text{I}_{1}\sim N\sigma^{2}\cdot\chi_{k^{\star}}^{2}$. Therefore $\mathbb{E}\text{I}_{1}=k^{\star}N\sigma^{2},\text{Var}\left(\text{I}_{1}\right)=2k^{\star}N^{2}\sigma^{4}$. By the Chebyshev’s inequality we know that
$\left|\text{I}_{1}-k^{\star}N\sigma^{2}\right|\leq\sqrt{\frac{2k^{\star}}{\alpha}}N\sigma^{2}:=t_{\text{I}_{1}}.$
with probability greater than $1-\alpha$.
For $\text{I}_{2,S^{\prime}}$, since we allow $S^{\prime}$ to be any subset of $S$, refined analysis is needed. Using the Hanson-Wright inequality A.10 and observing that $\left\lVert P^{\star}\right\rVert_{F}=\sqrt{k},\left\lVert P^{\star}\right\rVert_{2}\leq 1$, for each fixed $S^{\prime}\subset S$, we have
$\mathbb{P}\left(\left|\text{I}_{2}-k^{\star}\left|[N]\backslash S^{\prime}\right|\sigma^{2}\right|\geq t\right)\leq 2\exp\left\{-c\min\limits\left\{\frac{t^{2}}{k^{\star}\left|[N]\backslash S^{\prime}\right|^{2}\sigma^{4}},\frac{t}{\left|[N]\backslash S^{\prime}\right|\sigma^{2}}\right\}\right\}.$
[…]
If we select $t_{\text{I}_{2}}$ to be
$t_{\text{I}_{2}}=\max\limits\left\{\sqrt{\frac{4\left[2\epsilon N\ln\left(\frac{e}{2\epsilon}\right)+\ln\left(\frac{2}{\alpha}\right)\right]k^{\star}\epsilon^{2}N^{2}\sigma^{4}}{c}},\frac{2\left[2\epsilon N\ln\left(\frac{e}{\epsilon}\right)+\ln\left(\frac{2}{\alpha}\right)\right]\epsilon N\sigma^{2}}{c}\right\}$
then it can be verified that $\left|\text{I}_{2,S^{\prime}}-k^{\star}\left|[N]\backslash S^{\prime}\right|\sigma^{2}\right|\leq t_{\text{I}_{2}}$ holds for all $S^{\prime}$ with probability greater than $1-\alpha$.
For $\text{I}_{3,S^{\prime}}$, we use the following argument of conditional probability to effectively bound its tail. First, observe that the joint distribution of $\sum\limits_{i\in[N]\backslash S^{\prime}}(\widetilde{X}_{i}-\mu)$ and $\sum\limits_{i\in[N]}(\widetilde{X}_{i}-\mu)$ is a multivariate Gaussian
$\left(\sum\limits_{i\in[N]\backslash S^{\prime}}(\widetilde{X}_{i}-\mu)^{\top},\sum\limits_{i\in[N]}(\widetilde{X}_{i}-\mu)^{\top}\right)\sim\mathcal{N}\left(0,\begin{pmatrix}\left|[N]\backslash S^{\prime}\right|\mathbf{I}_{d}&\left|[N]\backslash S^{\prime}\right|\mathbf{I}_{d}\\ \left|[N]\backslash S^{\prime}\right|\mathbf{I}_{d}&N\mathbf{I}_{d}\end{pmatrix}\sigma^{2}\right).$
By the property of multivariate Gaussian distribution, conditional on the event $\left\{\sum\limits_{i\in[N]}(\widetilde{X}_{i}-\mu)=x\right\}$, the conditional distribution of $\sum\limits_{i\in[N]\backslash S^{\prime}}(\widetilde{X}_{i}-\mu)$ is
$\left[\sum\limits_{i\in[N]\backslash S^{\prime}}(\widetilde{X}_{i}-\mu)\left\lvert\right.\sum\limits_{i\in[N]}(\widetilde{X}_{i}-\mu)=x\right]\sim\mathcal{N}\left(\frac{\left|[N]\backslash S^{\prime}\right|}{N}x,\sigma^{2}\left(\left|[N]\backslash S^{\prime}\right|-\frac{\left|[N]\backslash S^{\prime}\right|^{2}}{N}\right)\mathbf{I}_{d}\right).$
By the maximal inequality for (sub-)Gaussian random variable that $\mathbb{P}\left(\max\limits_{1\leq i\leq N}\widetilde{X}_{i}\geq\sqrt{2\sigma^{2}(\ln N+t)}\right)\leq e^{-t}$, and conditional on the event $\left\{\sum\limits_{i\in[N]}(\widetilde{X}_{i}-\mu)=x\right\}$, we have the probability of the following event
$\left\{\text{I}_{3,S^{\prime}}\geq\frac{\left|[N]\backslash S^{\prime}\right|}{N}x^{\top}P^{\star}x+\sqrt{2x^{\top}P^{\star}x\left(\left|[N]\backslash S^{\prime}\right|-\frac{\left|[N]\backslash S^{\prime}\right|^{2}}{N}\right)\sigma^{2}\cdot\left(2\epsilon N\ln\left(\frac{e}{2\epsilon}\right)+t\right)}\text{ for some $S^{\prime}\subset S$}\right\}$
less than $e^{-t}$. We also need the following basic inequality of conditional probability for any random variables $X$ and $Y$
$\mathbb{P}\left(Y\geq a\right)$ $=\mathbb{E}\mathbf{1}_{\left\{Y\geq a\right\}}$
$=\mathbb{E}\left[\mathbf{1}_{\left\{Y\geq a\right\}}\cdot\left(\mathbf{1}_{\left\{f(X)\geq a\right\}}+\mathbf{1}_{\left\{f(X)<a\right\}}\right)\right]$
$=\mathbb{E}\left[\mathbb{E}\left[\mathbf{1}_{\left\{Y\geq a\right\}}\cdot\mathbf{1}_{\left\{f(X)\geq a\right\}}\left|\right.X\right]+\mathbb{E}\left[\mathbf{1}_{\left\{Y\geq a\right\}}\cdot\mathbf{1}_{\left\{f(X)<a\right\}}\left|\right.X\right]\right]$
$\leq\mathbb{E}\left[\mathbb{E}\left[\mathbf{1}_{\left\{f(X)\geq a\right\}}\left|\right.X\right]+\mathbb{E}\left[\mathbf{1}_{\left\{Y\geq f(X)\right\}}\cdot\mathbf{1}_{\left\{f(X)<a\right\}}\left|\right.X\right]\right]$
$\leq\mathbb{P}\left(f(X)\geq a\right)+\mathbb{E}\left[\mathbb{P}\left(Y\geq f(X)\left|\right.X\right)\right].$
Now observe that $\left|[N]\backslash S^{\prime}\right|\leq 2\epsilon N$ for all $S^{\prime}$, and take $Y=\max\limits_{S^{\prime}}\left\{\text{I}_{3,S^{\prime}}\right\}$, $X=\sum\limits_{i\in[N]}(\widetilde{X}_{i}-\mu)$, $f(x)=g(x^{\top}P^{\star}x)$ where $g(x)=\frac{\left|[N]\backslash S^{\prime}\right|}{N}x+\sqrt{4x\cdot\epsilon N\left(\epsilon N\ln\left(\frac{e}{\epsilon}\right)+t\right)}$ and $a=g(Nk^{\star}\sigma^{2}+\sqrt{\frac{4k^{\star}}{\alpha}}N\sigma^{2})$. By the argument above and the fact that $g(\cdot)$ is an increasing function on the positive real line, we know
$\mathbb{P}\left(f(X)\geq a\right)=\mathbb{P}\left(x^{\top}P^{\star}x\geq Nk^{\star}\sigma^{2}+\sqrt{\frac{4k^{\star}}{\alpha}}N\sigma^{2}\right)\leq\frac{\alpha}{2}.$
From the maximal inequality above, we also have $\mathbb{E}\left[\mathbb{P}\left(Y\geq f(X)\left|\right.X\right)\right]\leq\mathbb{E}e^{-t}=e^{-t}$. Taking $t=\ln\left(\frac{2}{\alpha}\right)$, and combining the results above, we conclude that with probability greater than $1-\alpha$ the following event holds for all $S^{\prime}$.
$\left\{\text{I}_{3,S^{\prime}}\leq\left|[N]\backslash S^{\prime}\right|k^{\star}\sigma^{2}+\left|[N]\backslash S^{\prime}\right|\sqrt{\frac{4k^{\star}}{\alpha}}\sigma^{2}+\sqrt{8\left(Nk^{\star}\sigma^{2}+\sqrt{\frac{4k^{\star}}{\alpha}}N\sigma^{2}\right)\epsilon N\left(2\epsilon N\ln\left(\frac{e}{2\epsilon}\right)+\ln\left(\frac{2}{\alpha}\right)\right)\sigma^{2}}\right\}.$
By the symmetrical nature of $\text{I}_{3,S^{\prime}}$, same technique and argument can also be applied to the probability of the lower tail of $\text{III}_{3,S^{\prime}}$.
Part III. We rewrite the part III as
[…]
Dealing with $\text{III}_{1,S^{\prime}}$ is relatively easier. Observe that $\text{III}_{1,S^{\prime}}$ actually follows a Gaussian distribution with $\mathbb{E}\text{III}_{1,S^{\prime}}=0$ and $\text{Var}\left(\text{III}_{1,S^{\prime}}\right)=\left|S^{\prime}\right|^{2}N\sigma^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\leq N^{3}\sigma^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}$. By the Chebyshev’s inequality, we have
$\mathbb{P}\left(\left|\text{III}_{1,S^{\prime}}\right|\leq\sqrt{\frac{N^{3}\sigma^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}}{\alpha}}\right)\geq 1-\alpha.$
[…]
$\mathbb{P}\left(\left|\text{III}_{2,S^{\prime}}\right|\leq\sqrt{2\epsilon N^{3}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\sigma^{2}}\cdot\sqrt{2\epsilon N\ln\left(\frac{e}{2\epsilon}\right)+\ln\left(\frac{1}{\alpha}\right)},\forall S^{\prime}\subset S\right)\geq 1-\alpha.$
In summary, each tail behaviour of the term above consists of two parts: the mean of the term and the variation around the mean. To clearly display the results above, the following table can be helpful.
Terms Mean Variation Tail
$\text{I}_{1}$ $k^{\star}N\sigma^{2}$ $\sqrt{\frac{2k^{\star}}{\alpha}}N\sigma^{2}$ $\alpha$
$\text{I}_{2,S^{\prime}}$ $k^{\star}\left|[N]\backslash S^{\prime}\right|\sigma^{2}$ $\max\limits\left\{\sqrt{\frac{4\left[2\epsilon N\ln\left(\frac{e}{2\epsilon}\right)+\ln\left(\frac{2}{\alpha}\right)\right]k^{\star}\epsilon^{2}N^{2}\sigma^{4}}{c}},\frac{2\left[2\epsilon N\ln\left(\frac{e}{\epsilon}\right)+\ln\left(\frac{2}{\alpha}\right)\right]\epsilon N\sigma^{2}}{c}\right\}$ $\alpha$
$\text{I}_{3,S^{\prime}}$ $k^{\star}\left|[N]\backslash S^{\prime}\right|\sigma^{2}$ $left|[N]\backslash S^{\prime}\right|\sqrt{\frac{4k^{\star}}{\alpha}}+\sqrt{8\left(Nk^{\star}+\sqrt{\frac{4k^{\star}}{\alpha}}N\right)\epsilon N\left(2\epsilon N\ln\left(\frac{e}{2\epsilon}\right)+\ln\left(\frac{2}{\alpha}\right)\right)}\right]\sigma^{2}$ $2\alpha$
$\text{III}_{1,S^{\prime}}$ 0 $\sqrt{\frac{N^{3}\sigma^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}}{\alpha}}$ $\alpha$
$\text{III}_{2,S^{\prime}}$ 0 $\sqrt{2\epsilon N^{3}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\sigma^{2}}\cdot\sqrt{2\epsilon N\ln\left(\frac{e}{2\epsilon}\right)}$ $\alpha$
Table 1: terms from (B.10) and corresponding means and variations.
To complete the proof, we finally need to compare the variation of each term above with the part II.
$\text{I}_{1}$. Require: $\left|S^{\prime}\right|^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\sqrt{\frac{2k^{\star}}{\alpha}}N\sigma^{2}$. This is equivalent to $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\frac{\sqrt{k^{\star}}}{N}\sigma^{2}$.
$\text{I}_{2,S^{\prime}}$. Require: $\left|S^{\prime}\right|^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\max\limits\left\{\sqrt{\frac{4\left[2\epsilon N\ln\left(\frac{e}{\epsilon}\right)+\ln\left(\frac{2}{\alpha}\right)\right]k^{\star}\epsilon^{2}N^{2}\sigma^{4}}{c}},\frac{2\left[2\epsilon N\ln\left(\frac{e}{\epsilon}\right)+\ln\left(\frac{2}{\alpha}\right)\right]\epsilon N\sigma^{2}}{c}\right\}$. This is equivalent to $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\sqrt{\frac{\epsilon^{3}\ln\left(\frac{1}{\epsilon}\right)k^{\star}\sigma^{4}}{N}}$ and $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)\sigma^{2}$.
$\text{I}_{3,S^{\prime}}$. Require: $\left|S^{\prime}\right|^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\left|[N]\backslash S^{\prime}\right|\sqrt{\frac{4k^{\star}}{\alpha}}\sigma^{2}\vee\sqrt{8\left(Nk^{\star}+\sqrt{\frac{4k^{\star}}{\alpha}}N\right)\epsilon N\left(2\epsilon N\ln\left(\frac{1}{2\epsilon}\right)+\ln\left(\frac{2}{\alpha}\right)\right)}\sigma^{2}$. This is equivalent to $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\frac{\epsilon\sqrt{k^{\star}}}{N}\sigma^{2}$ and $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\sqrt{\frac{\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)k^{\star}\sigma^{4}}{N}}$.
$\text{III}_{1,S^{\prime}}$. Require: $\left|S^{\prime}\right|^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\sqrt{\frac{N^{3}\sigma^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}}{\alpha}}$. This is equivalent to $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\frac{\sigma^{2}}{N}$.
$\text{III}_{2,S^{\prime}}$. Require: $\left|S^{\prime}\right|^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\sqrt{2\epsilon N^{3}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\sigma^{2}}\cdot\sqrt{2\epsilon N\ln\left(\frac{e}{2\epsilon}\right)+\ln\left(\frac{1}{\alpha}\right)}$. This is equivalent to $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\max\limits\left\{\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)\sigma^{2},\frac{\epsilon\sigma^{2}}{N}\right\}$.
In the argument above, we have used the assumption that $(1-2\epsilon)N\leq\left|S^{\prime}\right|\leq N$ and therefore $0\leq\left|[N]\backslash S^{\prime}\right|\leq 2\epsilon N$. Summarizing the table and the argument above, it is required that
$\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\sigma^{2}\max\limits\left\{\frac{\sqrt{k^{\star}}}{N},\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right),\sqrt{\frac{\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)k^{\star}}{N}}\right\}.$
Finally, by the Pythagorean theorem, we have $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}=\rho^{2}-\left\lVert\mu-P^{\star}\mu\right\rVert_{2}^{2}\geq\rho^{2}-D_{k^{\star}}^{2}(K)$. Therefore, in order that $\left\lVert P_{k}^{\star}\mu\right\rVert_{2}^{2}$ satifies the condition above, it suffices that
$\rho^{2}\gtrsim D_{k^{\star}}^{2}(K)+\sigma^{2}\max\limits\left\{\frac{\sqrt{k^{\star}}}{N},\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right),\sqrt{\frac{\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)k^{\star}}{N}}\right\}=E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2}).$
[…]
B.6 Proof of Theorem 3.8
Motivated by Canonne et al., (2023), we demonstrate the rationale of Algorithm 8 in this section. Same as the proof of Theorem 3.6, $k^{\star}$ is any fixed dimension and $P^{\star}$ is its corresponding projection operator. For the conciseness of the notation, we define $\widetilde{Y}_{i}:=P^{\star}\widetilde{X}_{i}$, $Y_{i}:=P^{\star}X_{i}$, and $\nu:=P^{\star}\mu$. We assume without loss of generality that $\widetilde{Y}_{i},Y_{i}\in\mathbb{R}^{k^{\star}}$ and $\text{Var}\left(\widetilde{Y}_{i}\right)=\mathbf{I}_{k^{\star}}$ with the fact in mind that any projected multivariate Gaussian random variable can be regarded as embedded into a lower dimension subspace with covariance matrix $\text{diag}\{(1,...,1,0,\dots,0)\}$ by rotating the axes. We also assume $\sigma=1$ in this section (one can scale the results by $\sigma$ anytime if needed).
To clarify the structure of this section and highlight how the definitions, lemmas, and algorithms fit together, we provide the following roadmap. We depart from the definition of regularity, and arrive at the final complete testing procedure, Algorithm 8. The pseudo code is provided in Appendix C.
Figure 7: roadmap of this section.
We start with the definition of regularity and $\omega$-regularity, which controls the behaviour of the $\ell_{2}$ norm of the individual and sum, and the inner product.
Definition B.6 (regularity).
$\mathbf{Y}=\left\{Y_{1},\dots,Y_{N}\right\}$ is said to be $(\epsilon,\beta_{1},\beta_{2})$-regular if for all subsets $S\subset[N]$ with $\left|S\right|\leq\epsilon N$, we have the following properties:
(i), $\left|\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\beta_{1}$, (ii), $\left|\left\lVert\sum\limits_{i\in S}Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\beta_{2}$, and (iii), $\left|\left\langle\sum\limits_{i\in S}Y_{i},\sum\limits_{j\in[N]}Y_{j}\right\rangle-\left|S\right|k^{\star}\right|\leq c\sqrt{N}\beta_{1}.$
$\omega$-regularity is a natural generalization of the definition above, which involves the interaction of a weight vector $\omega$ with $\mathbf{Y}$.
Definition B.7 ($\omega$-regularity).
Given a weight vector $\omega=(\omega_{1},\dots,\omega_{N})^{\top}$, $\mathbf{Y}$ is said to be $(\epsilon,\beta_{1},\beta_{2})$-regular if for all subsets $S\subset[N]$ with $\left|S\right|\leq\epsilon N$, we have the following properties:
(i), $\left|\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\beta_{1}$, (ii), $\left|\left\lVert\sum\limits_{i\in S}\sqrt{\omega}_{i}Y_{i}\right\rVert_{2}^{2}-\left\lVert\omega_{S}\right\rVert_{1}k^{\star}\right|\leq c\beta_{2}$, and (iii), $\left|\left\langle\sum\limits_{i\in S}\sqrt{\omega}_{i}Y_{i},\sum\limits_{j\in[N]}\sqrt{\omega}_{j}Y_{j}\right\rangle-\left\lVert\omega_{S}\right\rVert_{1}k^{\star}\right|\leq c\sqrt{N}\beta_{1}.$
Recall that $\omega_{S}\in\mathbb{R}^{N}$ means the restriction of $\omega$ on the set $S$.
B.7 immediately recovers B.6 if $\omega=\mathbf{1}_{S}$.
The next lemma states that any uncontaminated $\mathbf{Y}$ from $\mathcal{N}\left(\nu,\mathbf{I}_{k^{\star}}\right)$ is in fact $(\epsilon,\beta_{1},\beta_{2})$-regular with high probability and specified $\beta_{1}$ and $\beta_{2}$ below depending on $\epsilon,\left\lVert\nu\right\rVert_{2},N,k^{\star}$.
Lemma B.8.
Assume $Y_{1},\dots,Y_{N}$ are i.i.d. samples from $\mathcal{N}\left(\nu,\mathbf{I}_{k^{\star}}\right)$, then with probability greater than $1-\alpha$, $\mathbf{Y}=\left\{Y_{1},\dots,Y_{N}\right\}$ is $(\epsilon,\beta_{1},\beta_{2})$-regular with
$\beta_{1}=\epsilon N\left(\sqrt{k^{\star}}+\sqrt{N}\left\lVert\nu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\sqrt{N}\left\lVert\nu\right\rVert_{2}^{2},$ (B.11)
$\beta_{2}=\epsilon N\sqrt{\epsilon Nk^{\star}\ln\left(\frac{1}{\epsilon}\right)}+(\epsilon N)^{2}\ln\left(\frac{1}{\epsilon}\right)+\left\lVert\nu\right\rVert_{2}(\epsilon N)^{2}\sqrt{\ln\left(\frac{1}{\epsilon}\right)}+\left\lVert\nu\right\rVert_{2}^{2}(\epsilon N)^{2},$
given the condition that $N\geq\frac{\ln(1/\alpha)}{\epsilon\ln(1/\epsilon)}$.
Proof.
The last condition is of no importance since it only depends on the predetermined constants $\alpha$, $\epsilon$, and the fact that $\epsilon\geq\frac{1}{N}$. We prove this lemma by direct check of the definition.
Check of (i). By adding and subtracting $\nu$, for a single sample $Y$, we have
$\left\lVert Y\right\rVert_{2}^{2}-k^{\star}=\left\lVert Y-\nu\right\rVert_{2}^{2}-k^{\star}+\left\lVert\nu\right\rVert_{2}^{2}+2\nu^{\top}(Y-\nu).$
By assumption of the distribution of $Y$, $\left\lVert Y-\mu\right\rVert_{2}^{2}$ follows a chi-square distribution with degree of freedom $k^{\star}$. By standard concentration bound, we have:
$\mathbb{P}\left(\left|\left\lVert Y-\nu\right\rVert_{2}^{2}-k^{\star}\right|\geq t\right)\leq 2\exp\left\{-C\min\limits\left\{\frac{t^{2}}{k^{\star}},t\right\}\right\}.$
Now set the probability to be $\frac{\alpha}{6N}$, we have $t=c\max\limits\left\{\sqrt{k^{\star}\ln\left(\frac{N}{\alpha}\right)},\ln\left(\frac{N}{\alpha}\right)\right\}\leq c\left(\sqrt{k^{\star}\ln\left(\frac{N}{\alpha}\right)}+\ln\left(\frac{N}{\alpha}\right)\right)$ for some appropriate $c$.
$\nu^{\top}(Y-\nu)$ follows a Gaussian distribution with mean zero and variance $\left\lVert\nu\right\rVert_{2}^{2}$. By the concentration bound for the Gaussian, we have
$\mathbb{P}\left(\left|2\nu^{\top}(Y-\nu)\right|\geq t\right)\leq 2\exp\left\{-\frac{t^{2}}{8\left\lVert\nu\right\rVert_{2}^{2}}\right\}.$
Setting the probability to be $\frac{\alpha}{6N}$ as the same, we find $t=c\sqrt{\left\lVert\nu\right\rVert_{2}^{2}\ln\left(\frac{N}{\alpha}\right)}$. Combining the results above, we have
$\left|\left\lVert Y_{i}\right\rVert_{2}^{2}-k^{\star}\right|\leq c\left(\left(\sqrt{k^{\star}}+\left\lVert\nu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\ln\left(\frac{N}{\alpha}\right)+\left\lVert\nu\right\rVert_{2}^{2}\right)$
with probability greater than $1-\frac{\alpha}{3N}$. Consequently, this holds for all $Y_{i},1\leq i\leq N$ with probability greater than $1-\frac{\alpha}{3}$ by the union bound. Therefore, we have
$\left|\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\left[\epsilon N\sqrt{(k^{\star}+\left\lVert\nu\right\rVert_{2}^{2})\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\left\lVert\nu\right\rVert_{2}^{2}\right]$ (B.12)
Check of (ii). By the same trick, we have $\left\lVert\sum\limits_{i\in S}Y_{i}\right\rVert_{2}^{2}=\left\lVert\sum\limits_{i\in S}(Y_{i}-\nu)\right\rVert_{2}^{2}+\left\lVert\nu\right\rVert_{2}^{2}\left|S\right|^{2}+2\left|S\right|\nu^{\top}\sum\limits_{i\in S}(Y_{i}-\nu)$. Similarly, we have $\left\lVert\sum\limits_{i\in S}(Y_{i}-\nu)\right\rVert_{2}^{2}=\left|S\right|\left\lVert\frac{1}{\sqrt{\left|S\right|}}\sum\limits_{i\in S}(Y_{i}-\nu)\right\rVert_{2}^{2}$, and
$\mathbb{P}\left(\left|\left\lVert\frac{1}{\sqrt{\left|S\right|}}\sum\limits_{i\in S}(Y_{i}-\nu)\right\rVert_{2}^{2}-k^{\star}\right|\geq t\right)\leq 2\exp\left\{-c\min\limits\left\{\frac{t^{2}}{k^{\star}},t\right\}\right\}.$
By Lemma A.18, we have at most $\left(\frac{e}{\epsilon}\right)^{\epsilon N}$ different possible $S$ in total. Setting the probability to be $\frac{\alpha}{6}\left(\frac{\epsilon}{e}\right)^{\epsilon N}$ and solving for $t$, we obtain
$t$ $=c\max\limits\left\{\sqrt{k^{\star}\left[\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right]},\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right\}$
$\leq c\left[\sqrt{k^{\star}\left[\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right]}+\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right]$
$\leq c\left[\sqrt{k^{\star}\epsilon N\ln\left(\frac{1}{\epsilon}\right)}+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right],$
where the last step use the condition that $N\geq\frac{\ln(1/\alpha)}{\epsilon\ln(1/\epsilon)}$.
For $\nu^{\top}\sum\limits_{i\in S}(Y_{i}-\nu)$, by the same concentration bound and technique, we will obtain
$\mathbb{P}\left(\left|\nu^{\top}\sum\limits_{i\in S}(Y_{i}-\nu)\right|\geq c\left\lVert\nu\right\rVert_{2}\sqrt{\epsilon N}\cdot\sqrt{\epsilon N\ln\left(\frac{1}{\epsilon}\right)}\right)\leq\frac{\alpha}{6}\left(\frac{\epsilon}{e}\right)^{\epsilon N}.$
Combining the results together, and using the union bound over all possible $T$, we have
$\left|\left\lVert\sum\limits_{i\in S}Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\left[\epsilon N\sqrt{\epsilon Nk^{\star}\ln\left(\frac{1}{\epsilon}\right)}+(\epsilon N)^{2}\ln\left(\frac{1}{\epsilon}\right)+\left\lVert\nu\right\rVert_{2}(\epsilon N)^{2}\sqrt{\ln\left(\frac{1}{\epsilon}\right)}+\left\lVert\nu\right\rVert_{2}^{2}(\epsilon N)^{2}\right]$
with probability greater than $1-\frac{\alpha}{3}$. The RHS is exactly $\beta_{2}$ in (B.11).
Check of (iii). Finally, we are going to evaluate the term $\left\langle\sum\limits_{i\in S}Y_{i},\sum\limits_{j\in S}Y_{j}\right\rangle$. Since this is equal to $\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}+\sum\limits_{i\in S}\left\langle Y_{i},\sum\limits_{j\in S\backslash i}Y_{j}\right\rangle$. We have obtained the bound for $\left\lVert Y_{i}\right\rVert_{2}^{2}$ in (i). For the term of inner product, we have
$\left\langle Y_{i},\sum\limits_{j\in S\backslash i}Y_{j}\right\rangle=\left\langle Y_{i}-\nu,\sum\limits_{j\in S\backslash i}(Y_{j}-\nu)\right\rangle+\left\langle Y_{i}-\nu,(N-1)\nu\right\rangle+\left\langle\nu,\sum\limits_{j\in S\backslash i}(Y_{j}-\nu)\right\rangle+(N-1)\left\lVert\nu\right\rVert_{2}^{2}.$
Applying the sub-exponential concentration bound for the first term and sub-Gaussian concentration bound for the second as the previous handling, this is bounded by
$c\left[\sqrt{Nk^{\star}\ln\left(\frac{N}{\alpha}\right)}+\ln\left(\frac{N}{\alpha}\right)+(N-1)\left\lVert\nu\right\rVert_{2}\sqrt{\ln\left(\frac{N}{\alpha}\right)}+(N-1)\left\lVert\nu\right\rVert_{2}^{2}\right].$
With the bound of $\left\lVert Y_{i}\right\rVert_{2}^{2}$ in (i), the whole thing is bounded by
$c_{1}\left[\epsilon N\left(\sqrt{k^{\star}}+\left\lVert\nu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\left\lVert\nu\right\rVert_{2}^{2}\right]$
$+$ $c_{2}\left[\epsilon N\sqrt{Nk^{\star}\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N^{2}\left\lVert\nu\right\rVert_{2}\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right]$
$\leq$ $c\left[\epsilon N\left(\sqrt{Nk^{\star}}+N\left\lVert\nu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right]$
Comparing with (B.12) and Definition B.6, we find that the required $\beta_{1}$ is
$\epsilon N\left(\sqrt{k^{\star}}+\sqrt{N}\left\lVert\nu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\sqrt{N}\left\lVert\nu\right\rVert_{2}^{2},$
which is exactly $\beta_{1}$ in (B.11). The proof is completed. ∎
By Lemma B.8, we should be able to control the inner product between the sum of samples on two sets $S$ and $S^{\prime}$ by the polarization identity. The next corollary exactly states this benefit.
Corollary B.9.
Assume $\mathbf{Y}$ is $(2\epsilon,\beta_{1},\beta_{2})$-regular. Then for any subsets $S,S^{\prime}\subset[N]$ with $\left|S\right|\leq\epsilon N,\left|S^{\prime}\right|\leq\epsilon N$, we have with probability greater than $1-\alpha$:
$\left|\left\langle\sum\limits_{i\in S}Y_{i},\sum\limits_{j\in S^{\prime}}Y_{j}\right\rangle-\left|S\cap S^{\prime}\right|k^{\star}\right|\leq c\beta_{2}.$
Proof.
By the polarization identity, we know
$\left\langle\sum\limits_{i\in S}Y_{i},\sum\limits_{j\in S^{\prime}}Y_{j}\right\rangle=\frac{1}{2}\left[\left\lVert\sum\limits_{i\in S\cup S^{\prime}}Y_{i}\right\rVert_{2}^{2}+\left\lVert\sum\limits_{i\in S\cap S^{\prime}}Y_{i}\right\rVert_{2}^{2}-\left\lVert\sum\limits_{i\in S\backslash S^{\prime}}Y_{i}\right\rVert_{2}^{2}-\left\lVert\sum\limits_{i\in S^{\prime}\backslash S}Y_{i}\right\rVert_{2}^{2}\right].$
Since each of $\S,S\SS\cup S^{\prime},S\cap S^{\prime},S\backslash$ $^{\prime},S^{\prime}\backslash$ S has cardinality no more than $2\epsilon N$ by assumption, by Definition B.6 and Lemma B.8, we know
$\left|\left\langle\sum\limits_{i\in S}Y_{i},\sum\limits_{j\in S^{\prime}}Y_{j}\right\rangle-\frac{1}{2}\left(\left|S\cup S^{\prime}\right|+\left|S\cap S^{\prime}\right|-\left|S\backslash S^{\prime}\right|-\left|S^{\prime}\backslash S\right|\right)k^{\star}\right|\leq C\beta_{2}.$
[…]
With Lemma A.19, it is not hard to extend the result in Corollary B.9 to the case dealing with the convex combination of the indicator vectors.
Corollary B.10.
Assume $\mathbf{Y}$ is $(2\epsilon,\beta_{1},\beta_{2})$-regular, then for any vectors $a,b\in[0,1]^{N}$ with $\left\lVert a\right\rVert_{1}\leq\epsilon N,\left\lVert b\right\rVert_{1}\leq\epsilon N$, we have with probability greater than $1-\alpha$:
(i), $\left|\left\langle\sum\limits_{i=1}^{N}a_{i}Y_{i},\sum\limits_{j=1}^{N}b_{j}Y_{j}\right\rangle-\left\langle a,b\right\rangle k^{\star}\right|\leq c_{1}\beta_{2}$; (ii), $\left|\left\langle\sum\limits_{i=1}^{N}a_{i}Y_{i},\sum\limits_{j=1}^{N}Y_{j}\right\rangle-\left\lVert a\right\rVert_{1}k^{\star}\right|\leq c_{2}\sqrt{N}\beta_{1}$
In Corollary B.9, there is no corresponding (ii) since that is exactly the third definition of regularity.
Proof.
We proof by direct calculation. By Lemma A.19, the vectors $a$ and $b$ can be represent as the convex combination of some indicator vectors, respectively. Let $a=\sum\limits_{k=1}^{M_{a}}\xi_{k}\mathbf{1}_{S_{k}},b=\sum\limits_{l=1}^{M_{b}}\eta_{l}\mathbf{1}_{S_{l}}$, where $S_{k},S_{l}\subset[N],1\leq k\leq N,1\leq l\leq N$. We have:
$\left\langle\sum\limits_{i=1}^{N}a_{i}Y_{i},\sum\limits_{j=1}^{N}b_{j}Y_{j}\right\rangle$ $=\left\langle\sum\limits_{i=1}^{N}\sum\limits_{k=1}^{M_{a}}\xi_{k}\mathbf{1}_{S_{k},i}Y_{i},\sum\limits_{j=1}^{N}\sum\limits_{l=1}^{M_{b}}\eta_{l}\mathbf{1}_{S_{l},j}Y_{j}\right\rangle$
$=\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}\left\langle\sum\limits_{i\in S_{k}}Y_{i},\sum\limits_{j\in S_{l}}Y_{j}\right\rangle$
$\overset{(1)}{=}\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}\left(\left|S_{k}\cap S_{l}\right|k^{\star}+R_{k,l}\right)$
$=k^{\star}\cdot\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}\sum\limits_{i=1}^{N}\mathbf{1}_{S_{k},i}\cdot\mathbf{1}_{S_{l},i}+\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}R_{k,l}$
$=k^{\star}\cdot\sum\limits_{i=1}^{N}\left(\sum\limits_{k=1}^{M_{a}}\xi_{k}\mathbf{1}_{S_{k},i}\right)\left(\sum\limits_{l=1}^{M_{b}}\eta_{l}\mathbf{1}_{S_{l},i}\right)+\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}R_{k,l}$
$=k^{\star}\cdot\sum\limits_{i=1}^{N}a_{i}b_{i}+\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}R_{k,l}.$
Here in the equation, (1) is from Corollary B.9, and $R_{k,l}\leq C\beta_{2}$ for some constant $C$. Now that by Lemma A.19, $0\leq\sum\limits_{k=1}^{M_{a}}\xi_{k}\leq 1,0\leq\sum\limits_{l=1}^{M_{b}}\eta_{l}\leq 1$, we know $\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}R_{k,l}\leq C\beta_{2}$. Therefore, (i) is checked. With very similar argument, we can also check (ii). The proof is completed. ∎
At this stage, we have all tools needed for the proof. The goal of the following Algorithms 3, 4, 5 is to filter the weight and samples such that the resulting weight and samples meet the requirements of $\omega$-regularity. We start with the first requirement, which is the simplest one.
1Set $\gamma_{1}=c\left[\sqrt{k^{\star}\ln\left(\frac{N}{\alpha}\right)}+\ln\left(\frac{N}{\alpha}\right)\right]$, $\texttt{count}=0,\texttt{i}=0$
2while $\texttt{i}<N$ do
3 if $\left|\left\lVert Y_{i}\right\rVert_{2}^{2}-k^{\star}\right|>\gamma_{1}$ then
4
5    $\texttt{count}=\texttt{count}+1$
6    if $\texttt{count}>\epsilon N$ then
7       return None
8     Delete $Y_{i}$ from $\mathbf{Y}$
9 $\texttt{i}=\texttt{i}+1$
10
return $\mathbf{Y}$
Algorithm 3 Prefiltering.
Lemma B.11.
Upon finishing the algorithm 3, all of samples satisfy
$\left|\left\lVert Y_{i}\right\rVert_{2}^{2}-k^{\star}\right|\leq\gamma_{1}.$
Moreover, under the null hypothesis, the algorithm will terminate without rejecting with probability at least $1-\alpha$.
Proof.
The satisfaction of the inequality comes directly from the definition of Algorithm 3. According to Lemma B.8 where setting $\left\lVert\nu\right\rVert_{2}=0$, all uncontaminated samples will satisfy the inequality with probability at least $1-\alpha$ and therefore will not be deleted. The remaining follows. ∎
Remark.
Rigorously speaking, after the prefiltering, the number of samples left should be adjusted to $N-\texttt{count}$, where count is possible to be positive. However, if the prefiltering finishes without rejection, count is at most $\epsilon N$, upon which, and for the conciseness of the notations, we continue to use $N$ in the proofs later. One can verify that this does not affect the correctness of the following proofs.
The next two algorithms both deal with the second requirement of $\omega$-regularity. However, different techniques are required for the classic scenario $N>k^{\star}$ and the high-dimensional scenario $N\leq k^{\star}$. Assume first $N>k^{\star}$.
Let
$\gamma_{2}:=c\left[\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right].$ (B.13)
Denote $D(\omega)=k^{\star}\cdot\text{diag}\{\omega\}$, and $\lambda(\omega)=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$. The major part of the algorithm is a “while” iteration, whose control condition is $\lambda(\omega^{(t)})\geq\gamma_{2}$, where $\omega(t)$ represents the value of $\omega$ in the $t$-th iteration. Inside each iteration, let $v^{(t)}$ denote the singular vector associated with $\lambda(\omega^{(t)})$. We first compute $\tau_{i}^{(t)}=\left\langle v,Y_{i}\right\rangle^{2}\mathbf{1}_{\left\{\omega_{i}^{(t)}>0\right\}}$, then sort the samples according to the decreasing order of $\tau_{i}^{(t)}$. (Without any loss of generality, we assume $\tau_{i}^{(t)}$ are already sorted.) Next, let $I$ be the smallest index such that $\sum\limits_{i=1}^{I}\omega_{i}^{(t)}\geq 2\epsilon N$ and update $\omega_{i}^{(t+1)}$ as following:
[…]
Finally, we check the condition whether $\left\lVert\omega\right\rVert_{1}\leq N(1-2\epsilon)$ and reject the null hypothesis directly if this holds and pass to the following step if not.
1Set $\gamma_{2}$ as (B.13), and $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
2while $\lambda\geq\gamma_{2}$ do
3   Set $v$ to be the unit singular vector associated with $\lambda$
4  Compute $\tau_{i}=\left\langle v,Y_{i}\right\rangle^{2}\mathbf{1}_{\left\{\omega_{i}>0\right\}}$ for $1\leq i\leq N$
5  Set II be the smallest index such that $\sum\limits_{i=1}^{I}\omega_{i}\geq 2\epsilon N$
6  Update $w$ according to (B.14)
7 if $\left\lVert\omega\right\rVert_{1}<N(1-2\epsilon)$ then
8
9    return None
10
11  Set $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$
12
return $\omega$
Algorithm 4 Sample filtering when $N>k^{\star}$.
We are going to prove that if the algorithm finishes without rejecting the null, the second condition of $\omega$-regularity is guaranteed with high probability. This is formalized as the following lemma.
Lemma B.12.
Given the algorithm 4, we have the following facts:
 (i), the algorithm will terminate within finite rounds,
 (ii), under $H_{0}$, the algorithm will not output rejection with high probability;
 (iii), once finished without rejection, for any $S\subset[N]$ with $\left|S\right|\leq\epsilon N$, we have $\left\lVert\sum\limits_{i\in S}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}\leq 2\gamma_{2}\epsilon N$ with probability higher than $1-\alpha$.
Proof.
Proof of (i) is relatively easy. By the update (B.14), the $\omega_{i}$ corresponding to the maximal $\tau_{i}$ will be set to zero in each iteration. Consequently, $\left\lVert\omega\right\rVert_{1}$ will be less than $N(1-2\epsilon)$ in at most $2\epsilon N$ rounds, followed by the rejection.
[…]
Now it remains to prove the statement. Suppose $\omega^{(t)}\in\Lambda_{1}$ and we are still in the iteration. By the definition of $\Lambda_{1}$ and the update (B.14), it suffices to prove $\sum\limits_{i\in[I]\cap C}\tau_{i}^{(t)}\omega_{i}^{(t)}\geq\sum\limits_{i\in[I]\cap([N]\backslash C)}\tau_{i}^{(t)}\omega^{(t)}$, i.e.,
$v^{\top}\left(\sum\limits_{i\in[I]\cap C}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}\right)v\geq v^{\top}\left(\sum\limits_{i\in[I]\cap([N]\backslash C)}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}\right)v.$
By the control condition, we have
$v^{\top}\left(\sum\limits_{i=1}^{N}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}-N\mathbf{I}_{k^{\star}}\right)v\geq\gamma_{2}.$
By Corollary A.14 (set $\left\lVert\nu\right\rVert_{2}=0$) and the definition of $\gamma_{2}$ (B.13), with high probability, we have
[…]
$\geq v^{\top}\left(\sum\limits_{i=1}^{N}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}-N\mathbf{I}_{k^{\star}}\right)v-v^{\top}\left(\sum\limits_{i=1}^{N}\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}-N\mathbf{I}_{k^{\star}}\right)v$
$\geq(1-c)\gamma_{2}.$
By the definition of $I$ we know $2\epsilon N\leq\sum\limits_{i\in[I]}\omega_{i}<2\epsilon N+1$. Because $\omega^{(t)}\in\Lambda_{1}$, we know $\left\lVert\mathbf{1}-\omega^{(t)}\right\rVert_{1}\leq 2\left\lVert\mathbf{1}_{C}-\omega_{C}^{(t)}\right\rVert_{1}\leq 2\epsilon N$. Hence $I-(2\epsilon N+1)\leq\sum\limits_{i\in[I]}(1-\omega_{i}^{(t)})\leq 2\epsilon N$, which leads to $I\leq 4\epsilon N+1$. Therefore, according to Lemma A.15, with high probability, we have
$v^{\top}\left(\sum\limits_{i\in[I]\cap([N]\backslash C)}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}\right)v\leq c\gamma_{2}.$ (B.15)
Now that we sorted $\tau_{i}$ by decreasing order, $\sum\limits_{i\in[I]}\omega_{i}\geq 2\epsilon N$, and $\left|C\right|\leq\epsilon N$, we have
$\sum\limits_{i\in C}\omega_{i}^{(t)}\tau_{i}^{(t)}-\sum\limits_{i\in[I]}\omega_{i}^{(t)}\tau_{i}^{(t)}\leq\sum\limits_{i\in C\backslash[I]}\omega_{i}^{(t)}\tau_{i}^{(t)}\leq\frac{1}{2}\sum\limits_{i\in[I]}\omega_{i}^{(t)}\tau_{i}^{(t)}.$
Therefore $\sum\limits_{i\in[I]}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{2}{3}\sum\limits_{i\in C}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{2(1-c)}{3}\gamma_{2}$. Note that by selecting the constant in (B.13) sufficiently large, we can let $c$ sufficiently small. Hence, we can assume $c\leq\frac{1}{4}$, by which we have $\sum\limits_{i\in[I]}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{1}{2}\gamma_{2}$. Combine with (B.15), we find
$\sum\limits_{i\in[I]\cap C}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{1}{4}\gamma_{2}\geq\sum\limits_{i\in[I]\cap([N]\backslash C)}\omega_{i}^{(t)}\tau_{i}^{(t)}.$
Hence, (ii) is proved.
To prove (iii), we need the following lemma, which exploits the property of the norm of the mean from the algorithm.
Lemma B.13.
If $\left\lVert\nu\right\rVert_{2}\geq c$ for some sufficient large but universal constant $c$, then Algorithm 4 will reject $H_{0}$ with high probability.
Proof.
First, we fix the vector $v=\frac{\nu}{\left\lVert\nu\right\rVert_{2}}$. If the algorithm exits without rejection, we have
$\left\lVert\sum\limits_{i=1}^{N}\omega_{i}Y_{i}Y_{i}^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}<\gamma_{2}.$
This suggests that $\sum\limits_{i=1}^{N}\omega_{i}(v^{\top}Y_{i})^{2}-N<\gamma_{2}$, based on which we have
$\frac{\sum\limits_{i\in[N]\backslash C}\omega_{i}(v^{\top}Y_{i})^{2}}{N}-1\leq\frac{\sum\limits_{i=1}^{N}\omega_{i}(v^{\top}Y_{i})^{2}}{N}-1<\frac{\gamma_{2}}{N}\leq c,$
where the last step is from the definition of $\gamma_{2}$ (B.13). If we expand $v^{\top}Y_{i}$, this merely says
$\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]^{2}+\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}+2\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]\left\lVert\nu\right\rVert_{2}\leq cN.$ (B.16)
By the fact that $ab\geq-\frac{1}{4}a^{2}-b^{2}$ for $a,b\in\mathbb{R}$, we know
$2\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]\left\lVert\nu\right\rVert_{2}\geq-\frac{1}{2}\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}-2\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]^{2}.$
Adding both sides with $\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}$ and plugging (B.16) in, we have
$\frac{1}{2}\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}\leq cN+\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]^{2}\leq cN+\sum\limits_{i\in[N]\backslash C}\left[v^{\top}(Y_{i}-\nu)\right]^{2}.$
Now since $\left\lVert v\right\rVert_{2}=1$, we know that $\left[v^{\top}(Y_{i}-\nu)\right]^{2}$ is subject to a chi-square distribution with degree of freedom 1. By the standard argument for the sub-exponential random variables (see A.7), we know that $\sum\limits_{i\in[N]\backslash C}\left[v^{\top}(Y_{i}-\nu)\right]^{2}\leq N+c\sqrt{N}$ with (high) constant probability. On the other hand, by the definition of the algorithm we have $\left\lVert\omega_{[N]\backslash C}\right\rVert_{1}\geq\left\lVert\omega\right\rVert_{1}-\epsilon N\geq(1-3\epsilon)N$. Therefore, we have with (high) constant probability,
$\frac{1}{2}(1-3\epsilon)N\left\lVert\nu\right\rVert_{2}^{2}\leq c_{1}\cdot N+c_{2}\cdot\sqrt{N}.$
Now deviding both sides by $\frac{1}{2}(1-3\epsilon)N$, we see that Lemma B.13 holds if we take $c=\frac{2(c_{1}+c_{2})}{1-3\epsilon}$
We are going to prove (iii) by contradiction. Suppose that at the end of the algorithm without rejection, we have $\left\lVert\sum\limits_{i\in S}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}>2\gamma_{2}\epsilon N$ for some $S\subset[N]$ with $\left|S\right|\leq\epsilon N$. This indicates that there exists a vector $v\in\mathbb{R}^{k^{\star}}$ with $\left\lVert v\right\rVert_{2}=1$ such that $\sum\limits_{i\in S}\sqrt{\omega_{i}}v^{\top}Y_{i}>\sqrt{2\gamma_{2}\epsilon N}$. Therefore, by the Cauchy-Schwarz inequality, we have $\sum\limits_{i\in S\cup C}\omega_{i}(v^{\top}Y_{i})^{2}\geq 2\gamma_{2}$. This implies that
$\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}Y_{i}Y_{i}^{\top}+\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}\right\rVert_{2}\geq\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}Y_{i}Y_{i}^{\top}\right\rVert_{2}\geq 2\gamma_{2}.$ (B.17)
Now consider the quantity $\left\lVert\sum\limits_{i\in[N]\backslash(S\cup C)}\omega_{i}Y_{i}Y_{i}^{\top}-\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert$, by the triangular inequality, we have
$\left\lVert\sum\limits_{i\in[N]\backslash(S\cup C)}\omega_{i}Y_{i}Y_{i}^{\top}-\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert$ $\leq\left\lVert\sum\limits_{i=1}^{N}\omega_{i}\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}-N\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$
$+\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}-\left|S\cup C\right|\nu\nu^{\top}\right\rVert_{2}$
$\leq\underbrace{\left\lVert\sum\limits_{i=1}^{N}\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}-N\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}}_{\text{I}}+\underbrace{\left\lVert\sum\limits_{i=1}^{N}(1-\omega_{i})\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}\right\rVert_{2}}_{\text{II}}$
$+\underbrace{\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}\widetilde{Y_{i}}\widetilde{Y_{i}}^{\top}\right\rVert}_{\text{III}}+\left|S\cup C\right|\left\lVert\nu\right\rVert_{2}^{2}.$
By Lemma B.13, without loss of generality we can assume $\left\lVert\nu\right\rVert_{2}\leq c_{0}$ for some constant $c_{0}$. For the last term, since $\left|S\cup C\right|\leq 2\epsilon N$, obviously it can be bounded by $c\gamma_{2}$. We are going to deal with the other three parts respectively.
Bound on I. We have:
$\text{I}\leq\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)(\widetilde{Y}_{i}-\nu)^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}+2\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)^{\top}\nu\right\rVert_{2}.$
Set $\left\lVert\nu\right\rVert_{2}=0$ in Corollary A.14, we have
$\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)(\widetilde{Y}_{i}-\nu)^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\leq C_{1}\left(\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)\right).$ (B.18)
By normalization and standard concentration tools (see A.6), we have
$\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)^{\top}\nu\right\rVert_{2}=\left\lVert\nu\right\rVert_{2}\cdot\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)\right\rVert_{2}=C_{0}\cdot\sqrt{Nk^{\star}}\cdot\left\lVert\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\frac{\widetilde{Y}_{i}-\nu}{\sqrt{k^{\star}}}\right\rVert_{2}\leq c_{2}\sqrt{Nk^{\star}}$ (B.19)
with (high) constant probability. Now combine (B.18) and (B.19) to get
$\text{I}\leq(c_{1}+2c_{2})\left(\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)\right)\leq c\gamma_{2}.$
Bound on II. Since upon the finish the algorithm, we have $\left\lVert\mathbf{1}-\omega\right\rVert_{1}\leq 2\epsilon N$, by Corollary A.16, we know
$\text{II}\leq c_{3}\left[\epsilon N\ln\left(\frac{1}{\epsilon}\right)+k^{\star}+\ln\left(\frac{1}{\alpha}\right)\right]\leq c\gamma_{2}.$
Bound on III. Since $\left|S\cup C\right|\leq 2\epsilon N$, by Corollary A.16 and an argument similar to the one in II, we also have
$\text{III}\leq c\gamma_{2}.$
Therefore, we conclude that
$\left\lVert\sum\limits_{i\in[N]\backslash(S\cup C)}\omega_{i}Y_{i}^{\top}Y_{i}-\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\leq 4c\gamma_{2}.$ (B.20)
Now combine (B.17) and (B.20) and use the triangular inequality, we find that
$\left\lVert\sum\limits_{i=1}^{N}\omega_{i}Y_{i}^{\top}Y_{i}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\geq(2-4c)\gamma_{2}.$
Again, by choosing the constant in the definition of $\gamma_{2}$ (B.13) sufficiently large, we can make $c$ above sufficiently small. Therefore, we can assume $c<\frac{1}{4}$, which leads to $\left\lVert\sum\limits_{i=1}^{N}\omega_{i}Y_{i}^{\top}Y_{i}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}>\gamma_{2}$, a contradiction of the control condition of Algorithm 4. The proof is completed. ∎
We now turn to the high-dimensional scenario where $N\leq k^{\star}$. In this case, we let
$\gamma_{3}=c\left(\sqrt{Nk^{\star}}+\sqrt{k^{\star}\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right)$ (B.21)
for some universal constant $c$. The control condition of the iteration, in this case, is
$\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}\geq\gamma_{3},$
where $D(\omega)=k^{\star}\cdot\text{diag}\{\omega\}$. Inside each iteration, similar as before, we find $v$, which is the eigenvector associated with $\lambda$, and compute $\tau_{i}=\frac{v_{i}^{2}}{\omega_{i}}\mathbf{1}_{\left\{\omega_{i}>0\right\}},1\leq i\leq N$. Next, we update $\omega$ as following
$\omega_{i}=\left(1-\frac{\tau_{i}}{\max\limits_{i}\tau_{i}}\right)\omega_{i},1\leq i\leq N.$ (B.22)
Finally, we check the condition whether $\left\lVert\omega\right\rVert_{1}\leq N(1-6\epsilon)$ and reject the null directly if this holds and pass to the following step if not.
1Set $\gamma_{3}$ as (B.21), and $\lambda=\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
2while $\lambda\geq\gamma_{3}$ do
3   Set $v$ to be the unit singular vector associated with $\lambda$
4  Compute $\tau_{i}=\frac{v_{i}^{2}}{\omega_{i}}\mathbf{1}_{\left\{w_{i}>0\right\}}$
5  Update $w$ according to (B.22)
6 if $\left\lVert\omega\right\rVert_{1}<N(1-6\epsilon)$ then
7
8    return None
9
10  Set $\lambda=\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}$
11
return $\omega$
Algorithm 5 Sample filtering when $N\leq k^{\star}$.
Similar to the algorithm 4, we will have the following results to guarantee the satisfaction of the second condition of $\omega$-regularity.
Lemma B.14.
Given the algorithm 5, the following statements hold:
 (i), the algorithm will terminate within finite rounds;
 (ii), under $H_{0}$, the algorithm will not output rejection with high probability;
 (iii), upon the finish of the algorithm, for any $S\subset[N]$ with $\left|S\right|\leq\epsilon N$, we have
$\left|\left\lVert\sum\limits_{i\in S}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}-\left\lVert\omega_{S}\right\rVert_{1}k^{\star}\right|\leq c\epsilon N\gamma_{3}$
with probability higher than $1-\alpha$.
Proof.
Proof of (i) is same as the one in the proof of Lemma B.12.
[…]
Now assume $\omega^{(t)}\in\Lambda_{2}$ and we are still in the iteration, we will prove $\omega^{(t+1)}\in\Lambda_{2}$. By the definition, we only need to prove $5\sum\limits_{i\in C}v_{i}^{2}\geq\sum\limits_{i\in[N]\backslash C}v_{i}^{2}$. Suppose this does not hold, then $\left\lVert v_{C}\right\rVert_{2}^{2}<\frac{1}{6}$. By Corollary A.14 (set $\left\lVert\nu\right\rVert_{2}=0$), we have $\left\lVert\sqrt{D(\omega_{[N]\backslash C}^{(t)})}Y_{[N]\backslash C}Y_{[N]\backslash C}^{{}^{\top}}\sqrt{D(\omega_{[N]\backslash C}^{(t)})}-D(\omega_{[N]\backslash C}^{(t)})\right\rVert_{2}\leq c\gamma_{3}\leq c\lambda^{(t)}$ with high probability. However, since $v$ is the eigenvector corresponding to the maximal eigenvalue, we have
$\lambda^{(t)}$ $=\left|v^{\top}\left(\sqrt{D(\omega^{(t)})}Y^{\top}Y\sqrt{D(\omega^{(t)})}-D(\omega^{(t)})\right)v\right|$
$\leq c\lambda^{(t)}\left\lVert v_{[N]\backslash C}\right\rVert_{2}^{2}+2\lambda^{(t)}\left\lVert v_{[N]\backslash C}\right\rVert_{2}\left\lVert v_{C}\right\rVert_{2}+\lambda^{(t)}\left\lVert v_{C}\right\rVert_{2}^{2}.$
Noting that $\left\lVert v_{C}\right\rVert_{2}^{2}+\left\lVert v_{[N]\backslash C}\right\rVert_{2}^{2}=1$, define $f_{c}(x)=c+(1-c)x+2\sqrt{x(1-x)},0\leq x<\frac{1}{6}$. Note that by selecting the constant in (B.21) sufficiently large, we can let $c\rightarrow 0$. It can be proved that in this case $0\leq f_{0}(x)<1$, which is a contradiction. (ii) is proved.
Upon the termination of the algorithm without rejection, we have $\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}<\gamma_{3}$ and $\left\lVert\omega\right\rVert_{1}\geq(1-6\epsilon)N$. Therefore, we have
$\left|\left\lVert\sum\limits_{i\in S}\sqrt{\omega}_{i}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|$ $=\left|\sum\limits_{i\in S}\sum\limits_{j\in S}\sqrt{\omega_{i}\omega_{j}}Y_{i}^{\top}Y_{j}-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|$
$=\left|\mathbf{1}_{S}^{\top}\left[\sqrt{D(\omega)}YY^{\top}\sqrt{D(\omega)}-D(\omega)\right]\mathbf{1}_{S}\right|\overset{\text{(i)}}{\leq}c\epsilon N\gamma_{3}.$
Here (i) uses the condition that $\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}\leq\gamma_{3}$. The proof is completed. ∎
After Algorithms 4 or 5, we have filtered the samples to ensure the second condition of $\omega$-regularity. Now we turn to the check of the final condition.
Actually, the weight $\omega$ output by Algorithms 4 or 5 already provides us with some useful insight.
Lemma B.15.
Set $\beta_{1}$, $\beta_{2}$ according to (B.11) and $\gamma_{i},i\in\left\{2,3\right\}$ according to (B.13) or (B.21) depending on whether $N\geq k^{\star}$ or not. For the weight $\omega$ output by the Algorithms 4 or 5 and any S[$\CS\subset[N]\backslash$ C with $\left|S\right|\leq\epsilon N$, we have
$\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|\leq c(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma)$
with high probability.
Proof.
By Lemma B.8, with high probability the original samples $\widetilde{\mathbf{Y}}$ are $(2\epsilon,\beta_{1},\beta_{2})$-regular. By spliting $N]=([N]\backslash C)\cup C$, we have
$\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|$ $=\underbrace{\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},\widetilde{Y}_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|}_{\text{I}}$
$+\underbrace{\left|\sum\limits_{i\in S}\sum\limits_{j\in C}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},\widetilde{Y}_{j}\right\rangle\right|}_{\text{II}}+\underbrace{\left|\sum\limits_{i\in S}\sum\limits_{j\in C}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle\right|}_{\text{III}}$
For the term I, we further have
$\text{I}\leq\underbrace{\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}}\left\langle Y_{i},\widetilde{Y}_{j}\right\rangle-k^{\star}\left\lVert\sqrt{\omega_{S}}\right\rVert_{1}\right|}_{\text{I}_{1}}+\underbrace{\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}}(1-\sqrt{\omega_{j}})\left\langle Y_{i},\widetilde{Y}_{j}\right\rangle-k^{\star}\left\lVert\sqrt{\omega_{S}}(1-\sqrt{\omega_{S}})\right\rVert_{1}\right|}_{\text{I}_{2}}$
We now bound the terms $\text{I}_{1},\text{I}_{2},\text{II},\text{III}$ respectively.
Bound on $\text{I}_{1}$. Because $\left|S\right|\leq\epsilon N$, from $\omega$-regularity and Corollary B.10, we know $\text{I}_{1}\leq c_{1}\sqrt{N}\beta_{1}$.
Bound on $\text{I}_{2}$. Because $\omega$ is output by the Algorithms 4 or 5, we have $\left\lVert 1-\sqrt{\omega}\right\rVert_{1}\leq\left\lVert 1-\omega\right\rVert_{1}\leq\epsilon N$. By Corollary B.10 again, we have $\text{I}_{2}\leq c_{2}\beta_{2}$.
Bound on II. From Corollary B.10 and the fact $S\cap C=\varnothing$, we know $\text{II}\leq c_{3}\beta_{2}$.
Bound on III. From Lemma B.12, B.14, and the polarization equality, we have $\text{III}\leq c_{4}\epsilon N\gamma$.
Combining them together, we come to the conclusion of the lemma. ∎
We are now prepared for the final algorithm, which is designed to ensure the third condition of $\omega$-regularity. Given samples $S$ and weight $\omega$ by the algorithms 4 or 5, we compute $\tau_{i}=\left|\left\langle\sqrt{\omega_{i}}Y_{i},\sum\limits_{j=1}^{N}\sqrt{\omega_{j}}Y_{j}\right\rangle-\omega_{i}k^{\star}\right|\mathbf{1}_{\left\{\omega_{i}>0\right\}}$. Then, we set $\omega_{i}$ to zero for the indices $i$ corresponding to the $\epsilon N$ maximal $\tau_{i}$.
1Compute $\tau_{i}=\left|\left\langle\sqrt{\omega_{i}}Y_{i},\sum\limits_{j=1}^{N}\sqrt{\omega_{j}}Y_{j}\right\rangle-\omega_{i}k^{\star}\right|\mathbf{1}_{\left\{\omega_{i}>0\right\}},1\leq i\leq N$
2Sort $\tau_{i}$ by desreasing order and find the indices $\left\{i_{1},i_{2},\dots,i_{\epsilon N}\right\}$ corresponding to the first $\epsilon N$ maximal $\tau_{i}$
3Set $\omega_{i_{1}},\omega_{i_{2}},\dots,\omega_{i_{\epsilon N}}$ to zero
return $\omega$
Algorithm 6 Weight filtering.
Lemma B.16.
For the weight output by Algorithm 6 and any $S\subset[N]$ with $\left|S\right|\leq\epsilon N$, we have
$\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|\leq c(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma).$
This lemma is very similar to Lemma B.15 but stronger in the sense that we allow any $S\subset[N]$ rather than S[$\CS\subset[N]\backslash$ C. This advantage is exactly granted by the sorting and pruning processes in Algorithm 6.
Proof.
The proof of this lemma is not hard but depends on the following fact. Because we have proved in Lemma B.15 that for any S[$\CS\subset[N]\backslash$ C and $\left|S\right|\leq\epsilon N$, we have $\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|\leq c(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma)$, we know that this bound also holds for any $S\subset[N]\backslash\left\{1,2,\dots,\epsilon N\right\}$. With the proning on $\omega$ to remove the first $\epsilon N$ coordinates, we can assume without loss of generality that $S\subset[N]\backslash\left\{1,2,\dots\epsilon N\right\}$. We have
$\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{T}\right\rVert_{1}\right|=\left|\sum\limits_{i\in S}\sum\limits_{j=\epsilon N+1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{T}\right\rVert_{1}\right|\overset{\text{(1)}}{\leq}c(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma).$
Here (1) can be obtained by considering the subtraction of summation on the original weight with indices between $1$ and $\epsilon N$ from the summation on the whole original weight, and using the fact that $S\cap\left\{1,2,\dots,\epsilon N\right\}=\varnothing$. ∎
We have finished the proofs of Lemma B.11, B.12, B.14, and B.16, which means we have verified the correctness of $\omega$-regularity for our samples with high probability. The final step is to exploit the property of $\omega$-regularity to distinguish between $H_{0}$ and $H_{1}$.
Lemma B.17.
Let $\omega$ be the weights after running through the algorithm 3, 4 or 5, and 6, and $\beta_{1},\beta_{2}$ seleced as in Lemma B.8. Then, with high probability we have
$\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}-N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right|\leq c\left(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma+\underbrace{\left(N^{\frac{3}{2}}\left\lVert\nu\right\rVert_{2}+\sqrt{Nk^{\star}}\right)\sqrt{\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)}_{:=\text{R}}\right).$ (B.23)
Proof.
The proof of this lemma is from the definition of $\omega$-regularity and straightforward calculations. To evaluate $\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}$, we write it as:
$\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}$ $=\left\lVert\sum\limits_{i\in[N]\backslash C}\sqrt{\omega_{i}}Y_{i}+\sum\limits_{i\in C}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}$
$=\underbrace{\left\lVert\sum\limits_{i\in[N]\backslash C}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}}_{\text{I}}-\underbrace{\left\lVert\sum\limits_{i\in C}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}}_{\text{II}}+\underbrace{2\sum\limits_{i=1}^{N}\sum\limits_{j\in C}\sqrt{\omega_{i}\omega_{j}}Y_{i}^{\top}Y_{j}}_{\text{III}}$
We are going to bound the terms individually.
Bound on I. Define the vector $r\in\mathbb{R}^{N}$ where $r_{i}=1$ for $i\in C$ and $1-\sqrt{\omega_{i}}$ for i[$\Ci\in[N]\backslash$ C. By Algorithm 4 or 5, we have $\left\lVert r\right\rVert_{1}\leq\epsilon N+6\epsilon N=7\epsilon N$. For the term I, we have
$\left\lVert\sum\limits_{i\in[N]\backslash C}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}=\underbrace{\left\lVert\sum\limits_{i=1}^{N}\widetilde{Y}_{i}\right\rVert_{2}^{2}}_{\text{I}_{1}}+\underbrace{\left\lVert\sum\limits_{i=1}^{N}r_{i}\widetilde{Y}_{i}\right\rVert_{2}^{2}}_{\text{I}_{2}}-\underbrace{2\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{N}r_{j}\widetilde{Y}_{i}^{\top}\widetilde{Y}_{j}}_{\text{I}_{3}}$
For $\text{I}_{1}$, we have
$\text{I}_{1}=\sum\limits_{i=1}^{N}\left\lVert\widetilde{Y}_{i}-\nu\right\rVert_{2}^{2}+N^{2}\left\lVert\nu\right\rVert_{2}^{2}+2N\nu^{\top}\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu).$
By the standard concentration bounds (see A.6 and A.7), with high probability we have
$\left|\text{I}_{1}-Nk^{\star}-N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right|\leq c_{1}\left(N^{\frac{3}{2}}\left\lVert\nu\right\rVert_{2}\sqrt{\ln\left(\frac{1}{\alpha}\right)}+\sqrt{Nk^{\star}\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right).\right)$
For $\text{I}_{2}$, by the argument above, Lemma B.8, and Corollary B.10 with appropriate choice of $\epsilon$, with high probability we have
$\left|\text{I}_{2}-k^{\star}\left\lVert r\right\rVert_{2}^{2}\right|\leq c_{2}\beta_{2}.$
Again by Lemma B.8 and Corollary B.10, with high probability we have
$\left|\text{I}_{3}-2k^{\star}\left\lVert r\right\rVert_{1}\right|\leq c_{3}\sqrt{N}\beta_{1}.$
Combining the results above and using the fact that $Nk^{\star}+k^{\star}\left\lVert r\right\rVert_{2}^{2}+2k^{\star}\left\lVert r\right\rVert_{1}=k^{\star}\left\lVert\mathbf{1}-r\right\rVert_{2}^{2}=k^{\star}\left\lVert\omega_{[N]\backslash C}\right\rVert_{1}$, we have
$\left|\text{I}-k^{\star}\left\lVert\omega_{[N]\backslash C}\right\rVert_{1}-N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right|\leq c\left[\sqrt{N}\beta_{1}+\beta_{2}+\left(N^{\frac{3}{2}}\left\lVert\nu\right\rVert_{2}+\sqrt{Nk^{\star}}\right)\sqrt{\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)\right].$
Bound on II. By Lemma B.12 (recall the definition of $\gamma_{2}$ (B.13) and the fact that $N>k^{\star}$ in this case) or B.14, with high probability we have
$\left|\text{II}-k^{\star}\left\lVert\omega_{C}\right\rVert_{1}\right|\leq c\epsilon N\gamma.$
Bound on III. By Lemma B.16, with high probability we have
$\left|\text{III}-2k^{\star}\left\lVert\omega_{C}\right\rVert_{1}\right|\leq c\left(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma\right).$
Now combining the bounds of I, II, and III together, we have
$\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}-N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right|\leq c\left(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma+\left(N^{\frac{3}{2}}\left\lVert\nu\right\rVert_{2}+\sqrt{Nk^{\star}}\right)\sqrt{\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)\right),$
which is the desired results. ∎
Proof of Theorem 3.8.
Substituting $\nu$ with $P^{\star}\mu$, the key is to require each term on the right-hand-side in (B.23) to be asymptotically smaller than the term $N^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}$. For convenience, we rewrite the definition of $\beta_{1}$ and $\beta_{2}$ here.
$\beta_{1}=\epsilon N\left(\sqrt{k^{\star}}+\sqrt{N}\left\lVert P^{\star}\mu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\sqrt{N}\left\lVert P^{\star}\mu\right\rVert_{2}^{2},$
$\beta_{2}=\epsilon N\sqrt{\epsilon Nk^{\star}\ln\left(\frac{1}{\epsilon}\right)}+(\epsilon N)^{2}\ln\left(\frac{1}{\epsilon}\right)+\left\lVert P^{\star}\mu\right\rVert_{2}(\epsilon N)^{2}\sqrt{\ln\left(\frac{1}{\epsilon}\right)}+\left\lVert P^{\star}\mu\right\rVert_{2}^{2}(\epsilon N)^{2}.$
The table below presents the terms and corresponding required conditions for $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}$. Although there are many terms, only a few of them are useful indeed.
Sources No. Terms Required Conditions for $\left\lVert P^{\star}\mu\right\rVert_{2}$
$\sqrt{N}\beta_{1}$ 1 $\epsilon N\sqrt{Nk^{\star}}\sqrt{\ln\left(\frac{N}{\alpha}\right)}$ $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\frac{\epsilon\sqrt{k^{\star}\ln\left(\frac{N}{\alpha}\right)}}{\sqrt{N}}$
2 $\epsilon N^{2}\left\lVert P^{\star}\mu\right\rVert_{2}\sqrt{\ln\left(\frac{N}{\alpha}\right)}$ $\left\lVert P^{\star}\mu\right\rVert_{2}\gtrsim\epsilon\sqrt{\ln\left(\frac{N}{\alpha}\right)}$
3 $\epsilon N\sqrt{N}\ln\left(\frac{N}{\alpha}\right)$ $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\frac{\epsilon\ln\left(\frac{N}{\alpha}\right)}{\sqrt{N}}$
4 $\epsilon N^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}$ $\epsilon\lesssim 1$
$\beta_{2}$ 5 $\epsilon N\sqrt{\epsilon Nk^{\star}\ln\left(\frac{1}{\epsilon}\right)}$ $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\frac{\epsilon\sqrt{\epsilon k^{\star}\ln\left(\frac{1}{\epsilon}\right)}}{\sqrt{N}}$
6 $\epsilon^{2}N^{2}\ln\left(\frac{1}{\epsilon}\right)$ $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)$
7 $\left\lVert P^{\star}\mu\right\rVert_{2}\epsilon^{2}N^{2}\sqrt{\ln\left(\frac{1}{\epsilon}\right)}$ $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\epsilon^{2}\sqrt{\ln\left(\frac{1}{\epsilon}\right)}$
8 $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\epsilon^{2}N^{2}$ $\epsilon\lesssim 1$
$\epsilon N\gamma$ 9 $\epsilon N\sqrt{Nk^{\star}}$ $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\frac{\epsilon\sqrt{k^{\star}}}{\sqrt{N}}$
10 $\epsilon N\sqrt{N\ln\left(\frac{1}{\alpha}\right)}$ $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\frac{\epsilon\sqrt{\ln\left(\frac{1}{\alpha}\right)}}{\sqrt{N}}$
11 $\epsilon N\ln\left(\frac{1}{\alpha}\right)$ Dominated by the term 3
12 $\epsilon^{2}N^{2}\ln\left(\frac{1}{\epsilon}\right)$ $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)$
R 13 $N^{3/2}\left\lVert P^{\star}\mu\right\rVert_{2}\sqrt{\ln\frac{1}{\alpha}}$ $\left\lVert P^{\star}\mu\right\rVert_{2}\gtrsim\frac{\sqrt{\ln\left(\frac{1}{\alpha}\right)}}{\sqrt{N}}$
14 $N\sqrt{k^{\star}}\ln\left(\frac{1}{\alpha}\right)$ $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\frac{\sqrt{k^{\star}}\ln\left(\frac{1}{\alpha}\right)}{N}$
15 $\ln\left(\frac{1}{\alpha}\right)$ Dominated by the term 3
Table 2: terms from (B.23) and required conditions on $\left\lVert P^{\star}\mu\right\rVert_{2}$.
Inspecting the column “Required Conditions” and computing the maximal, we conclude the final requirement for $\left\lVert P^{\star}\mu\right\rVert_{2}$:
$\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\max\limits\left\{\frac{\epsilon\ln\left(\frac{N}{\alpha}\right)}{\sqrt{N}},\epsilon^{2}\ln\left(\frac{N}{\alpha}\right),\sqrt{\frac{\epsilon^{2}k^{\star}\ln\left(\frac{N}{\alpha}\right)}{N}},\frac{\sqrt{k^{\star}}\ln\left(\frac{1}{\alpha}\right)}{N}\right\}.$ (B.24)
Once this requirement is satisfied, by Lemma B.17 and the inequality (B.23), we know that $\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}P^{\star}X_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}\right|$ will be less than $c_{1}N^{2}P(\epsilon,k^{\star},N,1)$ under $H_{0}$ and greater than $c_{2}N^{2}P(\epsilon,k^{\star},N,1)$ under $H_{1}$ for some $0<c_{1}<c_{2}<1$, and both with high probability. $c_{1}$ and $c_{2}$ are universal and their values can be determined following proper choices of $\epsilon,\alpha,N,k^{\star}$. From the argument at the end of the proof of Theorem 3.6, we know (B.24) leads to (3.4). The proof is completed. ∎
B.7 Proof of Corollary 3.12
Proof.
By Definition 2.3 and 2.4, we can rewrite $k_{1}^{\star}$ and $k_{2}^{\star}$ as
$k_{1}^{\star}=\max\limits_{k}\left\{k\left\lvert\right.k\geq 0,\frac{D_{k-1}(K)}{k^{1/4}}\geq\frac{\sigma}{\sqrt{N}}\right\},$
$k_{2}^{\star}=\max\limits_{k}\left\{k\left\lvert\right.k\geq 0,\frac{D_{k-1}(K)}{k^{1/4}}\geq\frac{\sqrt{\epsilon}}{N^{\frac{1}{4}}}\sigma\right\}.$
For a fixed $K$, define function $g(k)=\frac{D_{k-1}(K)}{k^{1/4}}$. By the property of Kolmogorov widths, we know $g$ is a decreasing function. Therefore, in the alternative definition above, if $\frac{\sigma}{\sqrt{N}}\geq\frac{\sqrt{\epsilon}}{N^{1/4}}\sigma$, i.e., $\epsilon\leq\frac{1}{\sqrt{N}}$, we have $k_{1}^{\star}\leq k_{2}^{\star}$; similarly, if $\epsilon\geq\frac{1}{\sqrt{N}}$, we have $k_{1}^{\star}\geq k_{2}^{\star}$. We discuss the two possibilities in the following.
(i): $\epsilon\leq\frac{1}{\sqrt{N}}$. In this case, the dominant term in $E^{2}(\epsilon,K,N,\sigma^{2})$ is $\frac{\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$. If $k_{1}^{\star}=k_{2}^{\star}$. Obviously, we know $\frac{\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$ in (3.6) will also be the dominant. Otherwise, we have $k_{1}^{\star}+1\leq k_{2}^{\star}$. By the definition and the monotone property of Kolmogorov widths, we have
$\frac{\left(k_{2}^{\star}\right)^{\frac{1}{4}}\sqrt{\epsilon}}{N^{\frac{1}{4}}}\sigma\leq D_{k_{2}^{\star}-1}\leq D_{k_{1}^{\star}}(K)\leq\frac{\left(k_{1}^{\star}+1\right)^{\frac{1}{4}}}{\sqrt{N}}\sigma.$
This leads to
$\epsilon\leq\frac{1}{\sqrt{N}}\sqrt{\frac{k_{1}^{\star}+1}{k_{2}^{\star}}}\lesssim\frac{1}{\sqrt{N}}\sqrt{\frac{k_{1}^{\star}}{k_{2}^{\star}}}.$
With a little calculation, one can verify that this is equivalent to saying $\frac{\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$ is the dominant in (3.6) again. Hence, (3.6) matches $E(\epsilon,K,N,\sigma^{2})$ in this possibility.
(ii): $\epsilon\geq\frac{1}{\sqrt{N}}$. In this case, the dominant term in $E^{2}(\epsilon,K,N,\sigma^{2})$ is either $\epsilon\sqrt{\ln\left(\frac{1}{\epsilon}\right)\frac{k_{2}^{\star}}{N}}\sigma^{2}$ or $\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)\sigma^{2}$. If $k_{1}^{\star}=k_{2}^{\star}$, then $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}$ will dominate over $\frac{\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$ and consequently the dominant term in (3.6) is either $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}$ or $\epsilon^{2}\sigma^{2}$, which nearly matches $E^{2}(\epsilon,K,N,\sigma^{2})$ with only an exception of a $\ln\left(\frac{1}{\epsilon}\right)$ factor. Otherwise, we have $k_{1}^{\star}\geq k_{2}^{\star}+1$. Again, by the definition and the monotone property of Kolmogorov widths, we have
$\frac{\left(k_{1}^{\star}\right)^{\frac{1}{4}}}{\sqrt{N}}\sigma\leq D_{k_{1}^{\star}-1}(K)\leq D_{k_{2}^{\star}}(K)\leq\frac{(k_{2}^{\star}+1)^{\frac{1}{4}}\sqrt{\epsilon}}{N^{\frac{1}{4}}}\sigma,$
which leads to
$\epsilon\geq\frac{1}{\sqrt{N}}\sqrt{\frac{k_{1}^{\star}}{k_{2}^{\star}+1}}.$
If $k_{2}^{\star}=0$, then the dominant terms in (3.6) and $E^{2}(\epsilon,K,N,\sigma^{2})$ are $\epsilon^{2}\sigma^{2}$ and $\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)\sigma^{2}$, respectively, which match except for a $\ln\left(\frac{1}{\epsilon}\right)$ term. Otherwise, we will have
$\epsilon\gtrsim\frac{1}{\sqrt{N}}\sqrt{\frac{k_{1}^{\star}}{k_{2}^{\star}}},$
which again can be used to verify that $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}$ dominates over $\frac{\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$ in (3.6). In both cases, (3.6) and $E^{2}(\epsilon,K,N,\sigma^{2})$ match with only an exception of a $\ln\left(\frac{1}{\epsilon}\right)$ factor. The other part of the corollary is directly from the expression of $E^{2}(\epsilon,K,N,\sigma^{2})$ and $P^{2}(\epsilon,K,N,\sigma^{2})$. The proof is completed. ∎
B.8 Proof of Theorem 3.17
Proof.
The techniques are parallel to the proof of the lower bounds above. By Lemma 3.15, for $k=k_{p,1}^{\star},c=\frac{(k_{p,1}^{\star})^{1/p-1/4}}{\sqrt{N}}\sigma$, there exists a $\theta\in K$ such that $\left\lVert\theta\right\rVert_{p}=\frac{(k_{p,1}^{\star})^{1/p-1/4}}{\sqrt{N}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{1}{(k_{p,1}^{\star})^{1/4}\sqrt{N}}\sigma$. Again, we impose a uniform prior on the set $\left\{\mu\left\lvert\right.\mu\in K,\mu_{i}=\kappa\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},i\in I\right\}$ for some positive constant $\kappa$. By the same arguments of Appendix B.2, if $\mathbb{E}_{\mu,\mu^{\prime}}\exp\left\{\frac{N\mu^{\top}\mu^{\prime}}{\sigma^{2}}\right\}-1$, or equivalently $\frac{N^{2}\left\lVert\theta\right\rVert_{4}^{4}\kappa^{4}}{2\sigma^{4}}$ is sufficiently small, then there is no test with uniformly small type-I and type-II errors. We have
$\frac{N^{2}\left\lVert\theta\right\rVert_{4}^{4}\kappa^{4}}{2\sigma^{4}}\leq\frac{N^{2}\sum\limits_{i=1}^{d}\left\lVert\theta\right\rVert_{\infty}^{4-p}\left|\theta_{i}\right|^{p}\kappa^{4}}{2\sigma^{4}}=\frac{N^{2}\left\lVert\theta\right\rVert_{\infty}^{4-p}\left\lVert\theta\right\rVert_{p}^{p}\kappa^{4}}{2\sigma^{4}}\leq\kappa^{4}.$
This implies that it is necessary to ensure $\kappa\gtrsim 1$, from which we establish the first lower bound that $\rho_{\text{critical}}\gtrsim\frac{(k_{p,1}^{\star})^{1/p-1/4}}{\sqrt{N}}\sigma$.
Similarly, it suffices to bound $\frac{\sqrt{N}\left\lVert v\right\rVert_{4}^{2}}{\sigma^{2}}$ for the second term, where $v$ is uniform from the set
$V:=\left\{v\left\lvert\right.v_{i}=\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},i\in I\right\},$
and $\theta\in K$ from Lemma 3.15 satisfies $\left\lVert\theta\right\rVert_{p}=\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{1/p-1/4}}{N^{1/4}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{\sqrt{\epsilon}}{(Nk_{p,2}^{\star})^{1/4}}\sigma$. Consequently, it is not difficult to verify that $\frac{\sqrt{N}\left\lVert v\right\rVert_{4}^{2}}{\sigma^{2}}\leq\epsilon$. By Lemma B.1, we establish the second lower bound that $\rho_{\text{critical}}\gtrsim\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{1/p-1/4}}{N^{1/4}}\sigma$.
Suppose now that $\theta$ is the vector from Lemma 3.15 when $k=k_{p,3}^{\star}$. Subsequently, we know $\left\lVert\theta\right\rVert_{2}^{2}\leq\left\lVert\theta\right\rVert_{p}^{p}\cdot\left\lVert\theta\right\rVert_{\infty}^{2-p}\leq\epsilon^{2}\sigma^{2}$. Following the same reasoning in Appendix B.4, we know there always exists a strategy for $\mathcal{C}$ to discourage the testing. Therefore, we establish the third lower bound. The proof is completed. ∎
B.9 Proof of Theorem 3.18
Before we prove Theorem 3.18, we need the following observation of the p-convex orthosymmetric set.
Lemma B.18 (Truncation of a $p$-convex orthosymmetric set).
Assume $K\subset\mathbb{R}^{d}$ is a $p$-convex orthosymmetric set. For any $1\leq j\leq d$ and $J:=\left\{i_{1},\dots,i_{j}\right\}\subset\left\{1,\dots,d\right\}$, define the Truncation of $K$ on $J$ as
$K_{J}:=\left\{\theta_{J}\left\lvert\right.\theta_{J}=\left(\theta_{i_{1}},\dots,\theta_{i_{j}}\right)\in\mathbb{R}^{j},\exists\beta\in K,\beta_{J}=\theta_{J}\right\},$
where $\beta_{J}=(\beta_{i_{1}},\dots,\beta_{i_{j}})$ is the truncated $\beta$ on the index set $J$. Then $K_{J}\subset\mathbb{R}^{j}$ is also $p$-convex orthosymmetric.
Proof.
Lemma B.18 is straightforward from the definition. For any $\theta_{J,1},\theta_{J,2}\in K_{J}$, suppose $\theta_{J,1}=\beta_{1,J},\theta_{J,2}=\beta_{2,J},\beta_{1},\beta_{2}\in K$, then we know for any $\lambda\in[0,1]$, we have $\lambda\theta_{J,1}^{p}+(1-\lambda)\theta_{J,2}^{p}=\lambda\beta_{1,J}^{p}+(1-\lambda)\beta_{2,J}^{p}=(\lambda\beta_{1}^{p}+(1-\lambda)\beta_{2}^{p})_{J}\overset{\text{(i)}}{=}\widetilde{\beta}_{J}^{p}\in K_{J}^{p}$, where $\widetilde{\beta}\in K^{p}$ and (i) is from that $K$ is p-convex. Consequently, $K_{J}$ is p-convex.
Suppose $\bm{\xi}=(\xi_{1},\dots,\xi_{d}),\xi_{i}\in\left\{-1,1\right\}$, then for any $\theta_{J}\in K_{J}$, we have $\bm{\xi}_{J}\odot\theta_{J}=\bm{\xi}_{J}\odot\beta_{J}=(\bm{\xi}\odot\beta)_{J}\overset{(\text{ii})}{=}\bar{\beta}_{J}\in K_{J}$, where $\bar{\beta}\in K$, $\odot$ is element-wise multiplication and (ii) is from that $K$ is orthosymmetric. Consequently, $K_{J}$ is also orthosymmetric. ∎
We are now prepared to prove the original theorem.
Proof.
Suppose first $\epsilon\leq\frac{1}{\sqrt{N}}$. In this case, we have $\frac{j^{1/p-1/4}}{\sqrt{N}}\sigma\geq\frac{\sqrt{\epsilon}j^{1/p-1/4}}{N^{1/4}}\sigma$ for any $0\leq j\leq d$. By the monotonicity of Kolmogorov widths and the definitions of $k_{p,1}^{\star}$ and $k_{p,2}^{\star}$, we know $k_{p,1}^{\star}\leq k_{p,2}^{\star}$. Later calculations show that in this case we also have $k_{p,2}^{\star}\leq k_{p,3}^{\star}$. This implies $k_{p,1}^{\star}\leq k_{p,2}^{\star}\leq k_{p,3}^{\star}$. For $k_{p,1}^{\star}$ and its corresponding $P_{p,1}^{\star}$, without loss of generality we assume $P_{p,1}^{\star}\theta=(\theta_{1},\dots,\theta_{k_{p,1}^{\star}},0,\dots,0)$. Regard $P_{p,1}^{\star}\theta$ as a vector in $\mathbb{R}^{k_{p,1}^{\star}}$, by the Hölder’s inequality A.17, we know $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$ for any $\theta\in K$. Further define $k_{p,1,1}^{\star}:=\max\limits\left\{j\left\lvert\right.0\leq j\leq k_{p,1}^{\star},D_{j-1}(K)>\frac{j^{1/4}}{\sqrt{N}}\sigma\right\}$. Since $\epsilon\leq\frac{1}{\sqrt{N}}$, by the proof of Corollary 3.12 and Lemma B.18, we know that if $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, then there exists a desired test. However, this is already guaranteed since $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$ and $k_{p,1}^{\star}\geq k_{p,1,1}^{\star}$.
Suppose now $\epsilon\geq\frac{1}{\sqrt{N}}$. Analogous to the discussion above, we know that $k_{p,2}^{\star}\leq k_{p,1}^{\star}$. If $k_{p,2}^{\star}=k_{p,3}^{\star}$, then we can assume $P_{p,2}^{\star}=P_{p,3}^{\star}$, and we have
$\left\lVert P_{p,2}^{\star}\theta\right\rVert_{2}\gtrsim\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{\frac{1}{4}}}{N^{\frac{1}{4}}}\sigma,$
$\left\lVert P_{p,2}^{\star}\theta\right\rVert_{2}\gtrsim\epsilon\sigma.$
Since $\epsilon\geq\frac{1}{\sqrt{N}}$, we also have
$\left\lVert P_{p,2}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,2}^{\star})^{1/4}}{\sqrt{N}}\sigma.$
Similarly, we can regard $P_{p,2}^{\star}\theta$ as a vector in $\mathbb{R}^{k_{p,2}^{\star}}$ and define $k_{p,2,1}^{\star}$ and $k_{p,2,2}^{\star}$ accordingly as we did for $k_{1}^{\star}$ and $k_{2}^{\star}$. Since $k_{p,2}^{\star}\geq\max\limits\left\{k_{p,2,1}^{\star},k_{p,2,2}^{\star}\right\}$, we are again guaranteed to have a desired test by the inequalities above and the results in the $\ell_{2}$ norm case.
Assume now $k_{p,2}^{\star}\neq k_{p,3}^{\star}$. Suppose we have $k_{p,2}^{\star}\geq k_{p,3}^{\star}+1$. By the monotonicity of Kolmogorov widths and the definitions of $k_{p,2}^{\star}$ and $k_{p,3}^{\star}$, we have
$\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{N^{\frac{1}{4}}}\sigma\leq\widetilde{D}_{p,k_{p,2}^{\star}-1}(K)\leq\widetilde{D}_{p,k_{p,3}^{\star}}(K)<\epsilon(k_{p,3}^{\star})^{\frac{1}{p}-\frac{1}{2}}\sigma,$
from which we obtain
$\epsilon\gtrsim\frac{(k_{p,2}^{\star})^{\frac{2}{p}-\frac{1}{2}}}{(k_{p,3}^{\star})^{\frac{2}{p}-1}}\cdot\frac{1}{\sqrt{N}}\gtrsim\sqrt{\frac{k_{p,3}^{\star}}{N}}.$
For $k_{p,3}^{\star}$, we already have $\left\lVert P_{p,3}^{\star}\theta\right\rVert_{2}\gtrsim\epsilon\sigma$. By the inequalities above, it is not hard to verify that $\left\lVert P_{p,3}^{\star}\theta\right\rVert_{2}\gtrsim\frac{\sqrt{\epsilon}(k_{p,3}^{\star})^{1/4}}{N^{1/4}}\sigma$ and $\left\lVert P_{p,3}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,3}^{\star})^{1/4}}{\sqrt{N}}\sigma$. Therefore, again by the discussion above and the results in the $\ell_{2}$ norm case, we are guaranteed to have a desired test in $\mathbb{R}^{k_{p,3}^{\star}}$.
The last case is $k_{p,2}^{\star}\leq k_{p,3}^{\star}-1$. The analysis is essentially the same as the case above. We have
$\epsilon(k_{p,3}^{\star})^{\frac{1}{p}-\frac{1}{2}}\sigma\leq\widetilde{D}_{p,k_{p,3}^{\star}-1}(K)\leq\widetilde{D}_{p,k_{p,2}^{\star}}(K)\leq\frac{\sqrt{\epsilon}(k_{p,2}^{\star}+1)^{\frac{1}{p}-\frac{1}{4}}}{N^{\frac{1}{4}}}\sigma,$
which leads to
$\epsilon\lesssim\frac{(k_{p,2}^{\star})^{\frac{2}{p}-\frac{1}{2}}}{\sqrt{N}(k_{p,3}^{\star})^{\frac{2}{p}}-1}\lesssim\sqrt{\frac{k_{p,2}^{\star}}{N}}.$
Since we have $\left\lVert P_{p,2}^{\star}\theta\right\rVert_{2}\gtrsim\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{\frac{1}{4}}}{N^{\frac{1}{4}}}\sigma$, we also have $\left\lVert P_{p,2}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,2}^{\star})^{1/4}}{\sqrt{N}}\sigma$ and $\left\lVert P_{p,2}^{\star}\theta\right\rVert_{2}\gtrsim\epsilon\sigma$, which are sufficient for us to have a desired test in the space $\mathbb{R}^{k_{p,2}^{\star}}$. ∎
B.10 Verification of the Minimality in Section 4.2
Proof.
Without any loss of generality we assume $a_{i}$’s are sorted in decreasing order. If $k^{\star}=d$, the assertion is obvious since $\left\lVert P^{\star}\mu\right\rVert_{2}=\left\lVert\mu\right\rVert_{2}=c$. Suppose $k^{\star}<d$ in the following. If $c^{2}\leq a_{k^{\star}+1}$, then the vector
$\mu=(\underbrace{0,\dots,0}_{k^{\star}\text{ zeros}},c,0,\dots,0)$
satisfies $P^{\star}\mu=\mathbf{0}$. Therefore, the assertion holds again. Let us now assume that $c^{2}>a_{k^{\star}+1}$. Since $\left\lVert\mu\right\rVert_{2}=c$, we have $\sum\limits_{i=k^{\star}+1}^{d}\mu_{i}^{2}=c^{2}-\sum\limits_{i=1}^{k^{\star}}\mu_{i}^{2}$. On the other hand, since $\sum\limits_{i=1}^{d}\frac{\mu_{i}^{2}}{a_{i}}\leq 1$ and $a_{i}$’s are decreasing, we have
$\frac{c^{2}-\sum\limits_{i=1}^{k^{\star}+1}\mu_{i}^{2}}{a_{k^{\star}+1}}\leq\sum\limits_{i=k^{\star}+1}^{d}\frac{\mu_{i}^{2}}{a_{i}}\leq 1-\sum\limits_{i=1}^{k^{\star}}\frac{\mu_{i}^{2}}{a_{i}}.$
Rearranging gives
$\frac{c^{2}}{a_{k^{\star}+1}}-1\leq\sum\limits_{i=1}^{k^{\star}}\mu_{i}^{2}\left(\frac{1}{a_{k^{\star}+1}}-\frac{1}{a_{i}}\right)\leq\sum\limits_{i=1}^{k^{\star}}\mu_{i}^{2}\left(\frac{1}{a_{k^{\star}+1}}-\frac{1}{a_{1}}\right).$
Hence, we conclude that $\left\lVert P^{\star}\mu\right\rVert_{2}^{2}=\sum\limits_{i=1}^{k^{\star}}\mu_{i}^{2}\geq\frac{c^{2}/a_{k^{\star}+1}-1}{1/a_{k^{\star}+1}-1/a_{1}}$. It is clear that the vector given in Section 4.2 achieves this bound. The only remaining thing is to see that the vector indeed has $\ell_{2}$ norm $c$ and it belongs to $K$, which can be verified by simple algebra. The proof is completed. ∎
B.11 Proof of Lemma 5.3
Proof.
By the definition of $k_{1}^{\star}$ and $k_{e}^{\star}$ and the decreasing property of Kolmogorov widths, since $\frac{k^{1/4}}{\sqrt{N}}\leq\sqrt{\frac{k}{N}}$ for any $k,N$, we have $k_{1}^{\star}\geq k_{e}^{\star}$. On the other side, for any $j+1\geq(k_{e}^{\star}+1)^{2}$, we have $D_{j}(K)\leq D_{k_{e}^{\star}}(K)\leq\sqrt{k_{e}^{\star}+1}\sigma\leq(j+1)^{\frac{1}{4}}\sigma$. Therefore, we know $k_{1}^{\star}\lesssim(k_{e}^{\star})^{2}$. To finish the proof, we discuss different cases of $\epsilon$.
(i): $0\leq\epsilon\leq\frac{1}{\sqrt{N}}$. In this case, the dominant terms in (3.6) and (5.3) are $\frac{\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$ and $\frac{k_{e}^{\star}}{N}\sigma^{2}$, respectively. Since $k_{1}^{\star}\lesssim(k_{e}^{\star})^{2}$, the claim follows.
(ii): $\epsilon\geq\frac{1}{\sqrt{N}}$. In this case, we need to further discuss the relationship between $k_{2}^{\star}$ and $k_{e}^{\star}$.
(1): $k_{2}^{\star}=k_{e}^{\star}$. When $\frac{1}{\sqrt{N}}\leq\epsilon\leq\sqrt{\frac{k_{2}^{\star}}{N}}$, the dominant terms in (3.6) and (5.3) are $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}$ and $\frac{k_{e}^{\star}}{N}\sigma^{2}$, respectively. However, since $\epsilon\leq\frac{k_{2}^{\star}}{N}$, we know $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}\leq\frac{k_{e}^{\star}}{N}\sigma^{2}$. Otherwise, the dominant terms will be both $\epsilon^{2}\sigma^{2}$. Therefore, the claim holds for this case.
(2): $k_{2}^{\star}+1\leq k_{e}^{\star}$. We have
$\sqrt{\frac{k_{e}^{\star}}{N}}\sigma\leq D_{k_{e}^{\star}-1}\leq D_{k_{2}^{\star}}\leq\left(\frac{k_{2}^{\star}+1}{N}\right)^{\frac{1}{4}}\sqrt{\epsilon}\sigma,$
which leads to
$\epsilon\geq\frac{1}{\sqrt{N}}\cdot\frac{k_{e}^{\star}}{\sqrt{k_{2}^{\star}+1}}\geq\sqrt{\frac{k_{e}^{\star}}{N}}.$
Once this holds, the dominant terms will be both $\epsilon^{2}\sigma^{2}$. Therefore, the claim holds for this case.
(3): $k_{e}^{\star}+1\leq k_{2}^{\star}$. We have
$\left(\frac{k_{2}^{\star}}{N}\right)^{\frac{1}{4}}\sqrt{\epsilon}\sigma\leq D_{k_{2}^{\star}-1}\leq D_{k_{e}^{\star}}\leq\sqrt{\frac{k_{e}^{\star}+1}{N}}\sigma,$
which leads to
$\frac{1}{\sqrt{N}}\leq\epsilon\leq\frac{1}{\sqrt{N}}\cdot\frac{k_{e}^{\star}+1}{\sqrt{k_{2}^{\star}}}\lesssim\sqrt{\frac{k_{e}^{\star}}{N}}.$
Similarly, this implies that the dominant terms in (3.6) and (5.3) are $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}$ and $\frac{k_{e}^{\star}}{N}\sigma^{2}$, respectively, and $\epsilon^{2}\sigma^{2}$ will never dominate. Since $\epsilon\lesssim\frac{1}{\sqrt{N}}\cdot\frac{k_{e}^{\star}}{\sqrt{k_{2}^{\star}}}$ by the inequalities above, we know $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}\lesssim\frac{k_{e}^{\star}}{N}\sigma^{2}$.
The discussion above shows that the minimax rate of estimation is greater than the minimax rate of testing in all possible cases. The proof is completed. ∎
Appendix C Theoretical and Polynomial Time Algorithms
C.1 Theoretical Algorithm
According to Theorem 3.6 and Section B.5, we first attempt to search for a consistent set $T$ with respect to $\phi_{e}$. $\phi_{e}(\mathbf{X}_{T})$ then determines whether we are to accept or reject $H_{0}$.
1
2import $\mathbf{X},N,K,\sigma$
3
4Compute $k_{1}^{\star},P_{1}^{\star},k_{2}^{\star},P_{2}^{\star}$ from $N,K,\sigma,\epsilon$ as 2.3 and 2.4
5if $k_{1}^{\star}\leq k_{2}^{\star}$ then
6 $k^{\star}=k_{1}^{\star},P^{\star}=P_{1}^{\star}$
7
8else
9 $k^{\star}=k_{2}^{\star},P^{\star}=P_{2}^{\star}$
10
[…]
12for $T$ in $\left\{\text{all subsets with cardinality greater than }(1-\epsilon)N\right\}$ do
13
14 consistent_flag == True
15 if $\left\lVert\sum\limits_{i\in T}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left|T\right|\sigma^{2}\geq c\left|T\right|^{2}E^{2}(\epsilon,K,N,\sigma)$ then
16    subset_test == True
17 else
18    subset_test == False
[…]
20
21    if $\left\lVert\sum\limits_{i\in T^{\prime}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left|T^{\prime}\right|\sigma^{2}\geq c\left|T^{\prime}\right|^{2}E^{2}(\epsilon,K,N,\sigma)$ then
22       subsubset_test == True
23    else
24       subsubset_test == False
25    if subset_test XOR subsubset_test then
26       consistent_flag == False
27       break
28
29 if consistent_flag then
30    if subset_test then
31         Reject $H_{0}$
32       return
33    else
34         Accept $H_{0}$
35       return
36
37
38Accept $H_{0}$
return
Algorithm 7 exponential degree algorithm when $\rho\gtrsim E(\epsilon,K,N,\sigma^{2})$
C.2 Polynomial Time Algorithm
The polynomial-degree algorithm combines the three sub-algorithms in Section B.6 together to filter the observations $\mathbf{X}$ and through the weight $\omega$. The theoretical guarantee is provided in Section B.6.
1
2import $\mathbf{X},N,K,\sigma=1$
3import Algorithm 3, 4, 5, 6
4
[…]
6Compute $k_{1}^{\star},P_{1}^{\star},k_{2}^{\star},P_{2}^{\star}$ from $N,K,\sigma=1,\epsilon$ as 2.3 and 2.4
7if $k_{1}^{\star}\leq k_{2}^{\star}$ then
8 $k^{\star}=k_{1}^{\star},P^{\star}=P_{1}^{\star}$
9
10else
11 $k^{\star}=k_{2}^{\star},P^{\star}=P_{2}^{\star}$
12
[…]
14Execute Algorithm 3, and record the returning value $\texttt{R}_{1}$
15if $\texttt{R}_{1}$ is None then
16   Reject $H_{0}$
17 return
18
19else
20 $\mathbf{Y}=\texttt{R}_{1}$
21
22Set $\omega=\mathbf{1}_{[k^{\star}]}$
23if $N>k^{\star}$ then
24   Execute Algorithm 4, and record the returning value $\texttt{R}_{2}$
25else
26   Execute Algorithm 5, and record the returning value $\texttt{R}_{2}$
27if $\texttt{R}_{2}$ is None then
28   Reject $H_{0}$
29 return
30
31else
32 $\omega=\texttt{R}_{2}$
33
34Execute Algorithm 6, and record the returning value $\texttt{R}_{3}$
[…]
36if $\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}\right|\geq c_{2}N^{2}P(\epsilon,k^{\star},N,1)$ then
37   Reject $H_{0}$
38
39else
40   Accept $H_{0}$
41
return
Algorithm 8 polynomial degree algorithm when $\rho\gtrsim P(\epsilon,k^{\star},N,1)$.

### v3

@@
 Abstract
-This paper is concerned with robust signal detection in Gaussian noise under quadratically convex orthosymmetric (QCO) constraints.
-Specifically, the null hypothesis $H_{0}$ assumes no signal, whereas the alternative $H_{1}$ considers signal that is separated in Euclidean norm from zero and belongs to a set $K$ satisfying the QCO constraints.
-In addition, an adversary is allowed to inspect all the $N$ samples and replace up to $\epsilon N$ of them with arbitrary values, where $0<\epsilon<c_{0}<\frac{1}{2}$ is the corruption rate.
-Our main results establish the minimax rate of the separation radius $\rho_{\text{critical}}$ between $H_{0}$ and $H_{1}$ purely in terms of the geometry of $K$, the corruption rate $\epsilon$ (up to logarithmic factors in $\frac{1}{\epsilon}$) and the scale of the noise $\sigma$.
-We argue that the Kolmogorov widths of the constraint play a central role in determining the minimax rate.
-This indicates similarity with the (uncorrupted) estimation problem under QCO constraints, which was first established by Donoho et al., (1990).
-Moreover, the minimax lower bound reveals interesting phase transitions of the testing problem regarding $\epsilon$.
-Consistent with classic belief about testing and estimation, the testing problem is “easier” (even when one compares the results with recent papers studying the constrained robust estimation problem (Prasadan and Neykov,, 2025)).
-In addition to the main results above, where the upper bound is achieved with an intractable algorithm, inspired by Canonne et al., (2023), we develop a polynomial time algorithm which also nearly (up to logarithmic factors in $N$ and $1/\epsilon$) achieves the minimax lower bound.
-In contrast to Canonne et al., (2023), our algorithm works for signals of arbitrary Euclidean length, and respects the QCO constraint.
-Finally, all the results above are naturally extended to the $\ell_{p}$ norm testing problem for $1\leq p<2$.
+This paper studies the problem of robust signal detection in Gaussian noise under quadratically convex orthosymmetric (QCO) constraints.
+We consider a minimax testing framework where the signal belongs to a QCO set and is separated from zero in Euclidean norm, while an adversary is allowed to arbitrarily corrupt a fraction $\epsilon$ of the samples.
+We establish the minimax separation radius between the null and alternative purely in terms of the constraint geometry, sample size, corruption rate, and noise scale.
+Our analysis argues that the Kolmogorov widths of the constraint set play a central role in determining the detection limits, paralleling to classic results in estimation problem.
+The derived lower bounds exhibit phase transitions with respect to the corruption rate and confirm that robust testing is statistically easier than robust estimation.
+While the information-theoretic upper bound is achieved by a computationally intractable test, we develop a polynomial-time algorithm that achieves the minimax lower bound up to logarithmic factors.
+Unlike prior work, our algorithm handles signals of arbitrary Euclidean length while respecting the QCO constraints.
+Finally, we extend these results to the robust $\ell_{p}$ norm testing for $1\leq p<2$.
 1 Introduction
@@
 Performing inference under constraints and even contamination could be a formidable task.
-Most of the related works focus on the estimation and testing problem under several simple constraints with no contamination (Donoho and Gasko, (1992), Liu and Moitra, (2021), Liu and Moitra, (2023)).
+Most of the related works focus on the estimation and testing problem under several simple constraints with no contamination (Donoho and Gasko, (1992); Liu and Moitra, (2021, 2023)).
 One crucial reason for this is that most of the methods and techniques heavily depend on the assumption of the true model behind the data (for example, likelihood ratio test and tests depending on the asymptotic distributions).
@@
 The testing and estimation based on $\bar{X}$ is widely known to be optimal for both problems.
-Specifically, to control errors within the level of (small) constant, one needs $\rho\gtrsim\frac{d^{1/4}}{\sqrt{N}}\sigma$ for the testing problem (in the minimax sense) (Ingster and Suslina, (2003), Diakonikolas et al., (2022)) while $\inf\limits_{\widehat{\mu}}\sup\limits_{\mu\in\mathbb{R}^{d}}\mathbb{E}\left\lVert\widehat{\mu}-\mu\right\rVert_{2}^{2}\gtrsim\frac{d}{N}\sigma^{2}$ if one wants to optimally estimate $\mu$ in $\ell_{2}$ norm (LeCam, (1973), also see Example 15.4 in Wainwright, (2019)).
+Specifically, to control errors within the level of (small) constant, one needs $\rho\gtrsim\frac{d^{1/4}}{\sqrt{N}}\sigma$ for the testing problem (in the minimax sense) (Ingster and Suslina, (2003); Diakonikolas et al., (2022)) while $\inf\limits_{\widehat{\mu}}\sup\limits_{\mu\in\mathbb{R}^{d}}\mathbb{E}\left\lVert\widehat{\mu}-\mu\right\rVert_{2}^{2}\gtrsim\frac{d}{N}\sigma^{2}$ if one wants to optimally estimate $\mu$ in $\ell_{2}$ norm (LeCam, (1973), also see Example 15.4 in Wainwright, (2019)).
 However, as mentioned, the property of $\bar{X}$ is very sensitive to the authenticity of the samples.
@@
 Kolmogorov $k$-Width
-Introduced by Kolmogoroff, (1936) in the study of function spaces and approximation theory, the Kolmogorov widths (also known as the Kolmogorov $k$-width) characterizes how well an element can be reached by some linear approximations of intrinsic dimension $k$, given that the element belongs to a linear space and some restrictions are applied.
+Introduced by Kolmogoroff, (1936) in the study of function spaces and approximation theory, the Kolmogorov $k$-width (also known as the Kolmogorov $N$-width) characterizes how well an element can be reached by some linear approximations of intrinsic dimension $k$, given that the element belongs to a linear space and some restrictions are applied.
 For a given $k$, it selects the best $k$-dimensional projection operator under the measure of minimax error.
-The reader can refer to Definition 2.2 for a formal presentation.
+Readers can refer to Definition 2.2 for a formal presentation.
 Kolmogorov $k$-width is a classic concept and has seen usage in various fields.
@@
 Evans et al., (2009) uses this concept to study the $k$-method in isogeometric analysis.
-Kolmogorov $k$-width is also popular in the problems related to partial differential equations (PDEs), where the decay rate of the Kolmogorov $k$-width is intensely studied (Cohen and DeVore, (2015), Greif and Urban, (2019)).
+Kolmogorov $k$-width is also popular in the problems related to partial differential equations (PDEs), where the decay rate of the Kolmogorov $k$-width is intensely studied (Cohen and DeVore, (2015); Greif and Urban, (2019)).
 See also Papapicco et al., (2022) for an application in neural networks and PDEs, and a very recent work, Florian Arbes and Urban, (2025), for an application in optimal transportation and PDEs.
@@
 See Figure 1 for an illustrative example of the problem.
-XX0KK
-Figure 1: Testing $H_{0}:\mu=0$ vs $H_{1}:\|\mu\|_{2}\geq\rho,\mu\in K$ for the axis aligned ellipsoid (which is a special case of a QCO set) $K$ based on a single observation $X$.
+Figure 1: Testing $H_{0}:\mu=0$ vs $H_{1}:\|\mu\|_{2}\geq\rho,\mu\in K$ for the axis aligned ellipsoid (which is a special case of the QCO set) $K$ based on a single observation $X$.
 Upon applying a rotation to the data, one can use the test we develop in Section 3.2 to solve the problems optimally for rotated (non-axis aligned) ellipsoids or more generally QCO sets (which are still centered around the origin).
 See our first remark in the discussion section for more details.
-Our main result shows that the $\rho_{\text{critical}}$ is determined by three independently valid lower bounds.
+Our main result shows that $\rho_{\text{critical}}$ is determined by three independently valid lower bounds.
 Moreover, the lower bounds are tight and optimal with only the exception of logarithmic terms.
 The first lower bound captures the geometric nature of the QCO set $K$ via the Kolmogorov $k$-width.
-This lower bound is independent of the adversary, and also works in the uncorrupted setting, i.e., when $\epsilon=0$.
-The second lower bound is obtained within the framework of differential privacy, which is introduced in Section B.3.
-The third lower bound is derived from a combined argument related to identical mixtures, computation of the total variation and a designed strategy of the adversary.
-For the upper bounds, we present two algorithms which conclude whether $H_{0}$ holds or not while ensuring type-I and type-II errors both asymptotically less than $\alpha$.
-The requirements on $\lVert\mu\rVert_{2}$ imposed by the first algorithm are slightly closer to the theoretical lower bounds; however, this improvement comes at the cost of exponential time complexity, rendering the method computationally intractable for large $N$ and $d$.
-The second algorithm, in contrast, provides a slightly looser upper bound than the first algorithm (though still nearly optimal) while achieving polynomial time complexity.
-Nevertheless, both upper bounds suggest that the maximum of the three lower bounds discussed previously suffices to fully characterize the behavior of $\rho_{\text{critical}}$.
+This lower bound is independent of the adversary, therefore also works in the uncorrupted setting, i.e., when $\epsilon=0$.
+The second lower bound is obtained within the framework of differential privacy, which is introduced in Section 3.1.2.
+The third lower bound is derived from a combined argument related to the identical mixture distributions, computation of the total variation and a designed strategy of the adversary.
+For the upper bounds, we present two algorithms which conclude whether $H_{0}$ holds or not while ensuring type-I and type-II errors both less than $\alpha$.
+The requirements on $\lVert\mu\rVert_{2}$ imposed by the theoretical algorithm are slightly closer to the lower bounds.
+However, this improvement comes at the cost of exponential time complexity, rendering the method computationally intractable for large $N$ and $d$.
+The polynomial-time algorithm, in contrast, provides a slightly weaker upper bound than the theoretical algorithm (though still nearly optimal) while achieving polynomial-time complexity.
+Nevertheless, both upper bounds suggest that the three lower bounds previously discussed suffice to fully characterize the behavior of $\rho_{\text{critical}}$.
+Moreover, such minimax optimal rates also imply a phase transition phenomenon regarding $\epsilon$ for the contamination process.
 Finally, the results and algorithms derived in the $\ell_{2}$ norm scenario can be naturally extended to the $\ell_{p}$ case for $1\leq p<2$.
@@
 Suslina, (1996) studied the asymptotics of the problem of signal detection in $\ell_{q}$ ellipsoids with $\ell_{p}$ ball removed.
-Since $\ell_{q}$ ellispoids for $q\geq 2$ are QCO sets, when $p=2$ we provide a generalization of these results.
+Since $\ell_{q}$ ellipsoid for $q\geq 2$ are QCO sets, when $p=2$ we provide a generalization of these results.
 For general ellipsoids, Baraud, (2002) studied the testing problem in $\mathbb{R}^{n}$ and $\ell_{2}$ and obtained the minimax optimal critical radius.
@@
 It also contains the discussion of the phase transitions of the problem.
-Section 4 presents the experimental results of the polynomial time algorithms related to Theorem 3.8 in Section 3.
-Section 5 contains the discussion which compares the minimax rates of testing to estimation and the extension of the main results to the $\ell_{p}$ norm separation.
+Section 4 is the extension of the main results in Section 3 to the robust $\ell_{p}$ norm testing problem.
+Section 5 presents the experiment procedures, including the data generation, constraint and adversary setups, the experimental results of the polynomial-time algorithms related to Theorem 3.11 and relevant discussion.
+Section 6 contains the discussion which compares the minimax rates of robust testing to robust estimation.
 It also includes some open problems for future works.
-We attach necessary probabilistic tools and other lemmas in Appendix A.
-The proofs to the lower and upper bounds and other main corollaries in Section 3 and Section 5 are included in Appendix B.
-Appendix C presents the pseudo codes of related algorithms.
+The pseudo code of the relevant algorithms discussed in the paper and the additional experimental results of Section 5 are provided in Appendix A.
+We attach necessary probabilistic tools and auxiliary lemmas in Appendix B.
+The proofs of the lower and upper bounds and other main theorems, lemmas, and corollaries in Section 3, Section 4 and Section 6 are included in Appendix C.
 2 Preliminary
@@
 $\text{TV}(\mathcal{P},\mathcal{Q})$ denotes the total variation distance and $\text{KL}(\mathcal{P},\mathcal{Q})$ for the Kullback-Leibler divergence if $\mathcal{P}\ll\mathcal{Q}$.
-When the underlying probability distribution is not explicitly specified, we use $\mathbb{P}$ alternatively to denote probability law in general.
+When the underlying probability distribution is not explicitly specified, we use $\mathbb{P}$ alternatively to denote its probability law in general.
 We use $\left\langle\cdot,\cdot\right\rangle$ and $(\cdot)^{\top}(\cdot)$ alternatively (possibly with a subscript when needed) to represent the inner product in a Hilbert space.
@@
 Definition 2.1 (Quadratically convex orthosymmetric (QCO) set).
-Given a set $K\subset\mathbb{R}^{d}$, we say it is a quadratically convex orthosymmetric (QCO) set if it satisfies the following conditions:
- (1). $K$ is convex;
- (2). $K$ is quadratically convex, which means that $K^{2}$ is also convex, where $K^{2}$ is defined as
+Given a set $K\subset\mathbb{R}^{d}$, we say $K$ is a quadratically convex orthosymmetric (QCO) set if it satisfies the following conditions:
+ (1), $K$ is convex;
+ (2), $K$ is quadratically convex, which means that $K^{2}$ is also convex, where $K^{2}$ is defined as
 $K^{2}:=\left\{\left(\theta_{1}^{2},\dots,\theta_{d}^{2}\right)^{\top}\left\lvert\right.\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K\right\};$
-(3).
-$K$ is orthosymmetric, which means that if $\theta=\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K$, then $\theta_{\eta}:=\left(\eta_{1}\theta_{1},\dots,\eta_{d}\theta_{d}\right)^{\top}\\ \in K$, where $\eta_{i}\in\left\{-1,1\right\},1\leq i\leq d$.
+(3), $K$ is orthosymmetric, which means that if $\theta=\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K$, then $\theta_{\eta}:=\left(\eta_{1}\theta_{1},\dots,\eta_{d}\theta_{d}\right)^{\top}\in K$, where $\eta_{i}\in\left\{-1,1\right\},1\leq i\leq d$.
 Remark.
-The definition of QCO set can be naturally extended to the case when $K\subset\ell_{2}$.
-Definition 2.2 (Kolmogorov $k$-width).
+The definition of QCO set can be naturally extended when $K\subset\ell_{2}$, as we only need to replace $[d]$ by the positive integers $\mathbb{N}^{+}$.
+Definition 2.2 (Kolmogorov $k$-width, also known as Kolmogorov $N$-width).
 Let $\mathcal{X}$ be a Banach space equipped with the norm $\left\lVert\cdot\right\rVert$, and $K\subset\mathcal{X}$ is a subset. The Kolmogorov $k$-width is defined as
@@
 In this work, our analysis will be confined to orthogonal projections.
-Moreover, though $\mathcal{P}_{k}$ is the set of all orthogonal projection with intrinsic dimension $k$, we can discretize this set — only consider the projections that are aligned with the coordinate axes without loss of generality, as shown by the analysis later.
+Moreover, though $\mathcal{P}_{k}$ is the set of all orthogonal projection with intrinsic dimension $k$, we discretize this set — only consider the projections that are aligned with the coordinate axes without loss of generality, as shown by the analysis later.
 For now, let us first assume that the norm leveraged in the definition is the $\ell_{2}$ norm.
 Based on the definition, one can derive several basic properties of the Kolmogorov widths.
-It is non-increasing with respect to $k$, meaning that $D_{k_{1}}(K)\geq D_{k_{2}}(K)$ for any $0\leq k_{1}\leq k_{2}$.
+They are non-increasing with respect to $k$, meaning that $D_{k_{1}}(K)\geq D_{k_{2}}(K)$ for any $0\leq k_{1}\leq k_{2}$.
 In addition, for a finite-dimensional space $\mathcal{X}$ with full dimension $d$, we always have $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}$ and $D_{d}(K)=0$ for any $K\subset\mathcal{X}$.
 For the more general case where $K\subset\ell_{2}$, we shall assume $\lim\limits_{d\rightarrow\infty}D_{d}(K)=0$.
-In this paper, we further define $D_{k}(K)=\infty$ when $k<0$ for the convenience of the marginal cases.
-The connection between the Kolmogorov widths and our main results is established through the following definition of the first and second dimensions of the optimal projections, which are designed to identify the optimal $k$ satisfying certain marginal conditions determined by the geometric properties of the constraint set and the corruption process.
+In this paper, we further define $D_{k}(K)=\infty$ when $k<0$ and $D_{k}(K)=0$ for $k>d$ for the convenience of the marginal cases.
+The connection between the Kolmogorov widths and our main results is established through the following definition of the first and second dimensions of the optimal projections, which are designed to identify the optimal $k$ satisfying certain marginal conditions determined by the geometric properties of the constraint set $K$ and the corruption process.
 Definition 2.3 (First optimal dimension and projection).
@@
 The hyperrectangle in $\mathbb{R}^{d}$ is defined as
-$H_{d}(a_{1},\dots,a_{d}):=\left\{x\left\lvert\right.x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d},\left|x_{i}\right|\leq a_{i},a_{i}>0,1\leq i\leq d\right\}.$
+$H_{d}(a_{1},\dots,a_{d}):=\left\{x\left\lvert\right.x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d},\left|x_{i}\right|\leq a_{i},a_{i}\geq 0,1\leq i\leq d\right\}.$
 Obviously, $H_{d}(a_{1},\dots,a_{d})$ is convex and orthosymmetric.
@@
 By reindexing if necessary, we may assume that the sequence $\left\{a_{i}\right\}_{i=1}^{d}$ is arranged in decreasing order.
-The Kolmogorov $k$-width is
+The Kolmogorov widths are
 […]
@@
 Figure 2: An example of an ellipse and its square-set in $\mathbb{R}^{2}$. The square-set is a triangle in $\mathbb{R}^{2}$, which is convex.
-Similarly, we assume $\left\{a_{i}\right\}$ is arranged in decreasing order, by which the Kolmogorov widths are
+Again, we assume $\left\{a_{i}\right\}_{i=1}^{d}$ is arranged in decreasing order, by which the Kolmogorov widths are
 […]
 The first optimal dimension is $k_{1}^{\star}=\max\limits\left\{j\left\lvert\right.1\leq j\leq d,a_{j}>\frac{\sqrt{j}}{N}\sigma^{2}\right\}$ when $a_{1}>\frac{1}{N}\sigma^{2}$ and 0 otherwise.
-The second optimal dimension $k_{2}^{\star}=\max\limits\left\{j\left\lvert\right.1\leq j\leq d,a_{j}>\frac{\sqrt{j}\epsilon}{\sqrt{N}}\sigma^{2}\right\}$ when $a_{1}>\frac{\epsilon}{\sqrt{N}}\sigma^{2}$ otherwise 0.
+The second optimal dimension $k_{2}^{\star}=\max\limits\left\{j\left\lvert\right.1\leq j\leq d,a_{j}>\frac{\sqrt{j}\epsilon}{\sqrt{N}}\sigma^{2}\right\}$ when $a_{1}>\frac{\epsilon}{\sqrt{N}}\sigma^{2}$ and 0 otherwise.
 When $a_{i}=i^{-2s}$ for some $s>1$, the corresponding ellipse in $\mathbb{R}^{d}$ can be regarded as a special case of the more general ellipses of Sobolev types, which is defined as $S(s):=\left\{x\left\lvert\right.x\in\mathbb{R}^{\infty},\sum\limits_{i=1}^{\infty}i^{2s}x_{i}^{2}\leq 1\right\}$.
@@
 Remark.
-When a ellipse is not axes-aligned, (1), the orthosymmetric property is not guaranteed, which can be easily checked; (2), the quadratically convex property is true for $d=2$ and false for $d\geq 3$.
+When an ellipse is not axes-aligned: (1), the orthosymmetric property is not guaranteed, which can be easily checked; (2), the quadratically convex property is true for $d=2$ and false for $d\geq 3$.
 To gain the idea of (2), consider an ellipse
@@
 3.1 Lower Bounds
-In this section, we introduce three lower bounds, i.e., Theorem 3.2, 3.3, and 3.4, which independently work on determining the lower bound of $\rho_{\text{critical}}$.
+In this section, we introduce three independent lower bounds, i.e., Theorem 3.2, 3.5, and 3.6, which jointly and fully determine the lower bound of $\rho_{\text{critical}}$.
 3.1.1 Based on a Geometric Property
-The following lemma is an important basis for the theoretical lower bounds in Theorems 3.2 and 3.3.
-Exploiting the properties of QCO sets and the definition of the Kolmogorov widths, Lemma 3.1 outputs a vector with large $\ell_{2}$ norm but small $\ell_{\infty}$ norm, which is fundamental for the prior of $v$ constructed under $H_{1}$.
-The proof, which involves an application of the minimax theorem, is provided in Appendix B.1.
+The following lemma serves as an important basis for the theoretical lower bounds in Theorems 3.2 and 3.5.
+The underlying intuition is that the orthosymmetry and convexity of the prior constraint $K$ imply the existence of vectors with small individual coordinates that nonetheless yield a significant aggregate norm.
+Such “spread-out” vectors, and the maximal norm they can attain, are effectively characterized by the Kolmogorov widths of $K$.
+Specifically, Lemma 3.1 guarantees the existence of a vector with a large $\ell_{2}$ norm but a small $\ell_{\infty}$ norm by leveraging the properties of QCO sets and the definition of the Kolmogorov widths.
 Lemma 3.1.
@@
 Then there exists a vector $\theta\in K$ such that $\left\lVert\theta\right\rVert_{2}=c\sigma$ while $\left\lVert\theta\right\rVert_{\infty}\leq\frac{c}{\sqrt{k}}\sigma$.
-We now are ready to state the first main lower bound.
-Theorem 3.2.
+The proof of Lemma 3.1 is derived from the key properties of the QCO set $K$, combined with an application of the minimax theorem to a suitably truncated vector in $K^{2}$.
+The formal proof is provided in Appendix C.1.
+A similar argument also appears in Section 3.3.2 of Neykov, (2022).
+Such existence in Lemma 3.1 is fundamental for the construction of the least favorable prior, where we define a uniform mixture over the coordinate-wise sign flips of this vector to induce a small total variation distance between the mixture and the null distribution $H_{0}$, thereby establishing one of the minimax lower bounds.
+This standard technique (see, e.g., LeCam, (1973)) can be adapted to our current problem with the prior constraint $K$ via Lemma 3.1.
+Theorem 3.2 (First lower bound).
 If $\rho\leq c(\alpha)\cdot\frac{(k_{1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, where $c$ is a constant that only depends on $\alpha$, or $k_{1}^{\star}=0$, then
 $\inf_{\psi:\mathbb{P}_{0}(\psi=1)\leq\alpha}\sup_{\theta\in K,\|\theta\|\geq\rho}\mathbb{P}_{\theta}(\psi=0)\geq\alpha.$
-Moreover, one possible explicit expression of $c(\alpha)$ is
+Moreover, one possible explicit expressions of $c(\alpha)$ is
 $c(\alpha)=\left[2\ln\left(1+4\left(1-2\alpha\right)^{2}\right)\right]^{\frac{1}{4}}.$
-Theorem 3.2 above states that for fixed $\alpha$, if $\rho\lesssim\frac{(k_{1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, it is impossible to find a test that performs well under $H_{0}$ and all possible $H_{1}$ uniformly.
-As a corollary, this result implies that it is also impossible to distinguish between $H_{0}$ and $H_{1}$ in the presence of an adversary $\mathcal{C}$ under the above condition, since a simple strategy for $\mathcal{C}$ is to keep all samples untouched.
-The proof of Theorem 3.2 relies on the selected vector in Lemma 3.1 and Le Cam’s method of “fuzzy hypothesis”.
+For a fixed level $\alpha$, Theorem 3.2 establishes that if $\rho\lesssim\frac{(k_{1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, no test can uniformly distinguish $H_{1}$ from $H_{0}$ with non-trivial power.
+As an immediate consequence, distinguishing $H_{0}$ and $H_{1}$ remains impossible in the presence of an adversary $\mathcal{C}$ under the same condition, since a valid strategy for $\mathcal{C}$ is simply applying the identity map (i.e., leaving the samples unperturbed).
 Moreover, it is worth noting the marginal case when $k_{1}^{\star}=0$.
-The definition of $k_{1}^{\star}$ only relies on $K$, NN, $\sigma$, and does not depend on the original samples $\widetilde{\mathbf{X}}$ or the observations $\mathbf{X}$.
-Therefore, we shall assume in the following text that $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}>\frac{1}{\sqrt{N}}\sigma$, and therefore $k_{1}^{\star}\geq 1$.
+Informally speaking, $k_{1}^{\star}=0$ means that the constraint set $K$ is insufficiently rich to support a signal that is statistically distinguishable from $H_{0}$.
+Since the definition of $k_{1}^{\star}$ only relies on $K$, NN, $\sigma$, and does not depend on the original samples $\widetilde{\mathbf{X}}$ or the observations $\mathbf{X}$, we shall assume in the following text that $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}>\frac{1}{\sqrt{N}}\sigma$, and therefore $k_{1}^{\star}\geq 1$.
 Otherwise, the test will not be feasible in the first place.
-We kindly refer the readers to Appendix B.2 for the proof of this argument and Theorem 3.2.
-Recall that for uncontaminated observations without constraint, the critical value of the $\ell_{2}$ norm is $\rho_{\text{critical}}\asymp\frac{d^{1/4}}{\sqrt{N}}\sigma$ (see Section 3.3.6 in Ingster and Suslina, (2003) for more details).
-Theorem 3.2 generalizes this classic result in the sense that additional prior knowledge the mean is allowed and the lower bound is refined as a consequence.
+We kindly refer readers to Appendix C.2 for the proof of this argument and Theorem 3.2, which relies on the selected vector in Lemma 3.1 and Le Cam’s method of “fuzzy hypothesis”.
+Recall that for uncontaminated observations without constraint, the critical value for $\left\lVert\mu\right\rVert_{2}$ is $\rho_{\text{critical}}\asymp\frac{d^{1/4}}{\sqrt{N}}\sigma$ (see Section 3.3.6 in Ingster and Suslina, (2003)).
+Theorem 3.2 generalizes this classic result in the sense that additional prior knowledge of the mean is allowed and the lower bound is refined as a consequence.
 Indeed, Theorem 3.2 recovers the classic result when we set $\mathcal{X}=\mathbb{R}^{d}$ (where $k_{1}^{\star}=d$ according to Example 2.5).
-Moreover, the theorem implies that when we have non-trivial constraint $K$ on $\mu$, the essential dimensional attribute of the problem is the finite $k_{1}^{\star}$ instead of $d$.
+Moreover, the theorem implies that when we have non-trivial constraint $K$ on $\mu$, one of the essential dimensional attributes of the problem is the finite $k_{1}^{\star}$ instead of $d$.
 (Recall that $k_{1}^{\star}<\infty$ is ensured even if $K\subset\ell_{2}$.)
 3.1.2 Based on the Framework of Differential Privacy
-The concept of differential privacy was first introduced by Dwork et al., (2006) in the study of private data analysis.
-It was later introduced to the goodness-of-fit test (also called identity test) in statistics.
-See Gaboardi et al., (2016), Rogers and Kifer, (2017), Cai et al., (2017) for some earlier works of this application.
-Differential privacy requires a random algorithm to be robust against the disturbance of the input.
-Please see Section B.3 for the formal definition of differential privacy.
-In Theorem 3.3, we will combine Lemma 3.1 with an argument of optimal coupling and results from differential privacy to obtain another requirement for $\rho$.
-Please refer to Section B.3 for the techniques and proofs.
-Theorem 3.3.
-For the testing problem 1.2 with an adversary $\mathcal{C}$ from strong $\epsilon$-contamination model 2.8, it is impossible to derive a test that achieves type-I error and type-II error both less than 0.1 if $\rho\lesssim\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}$.
-Formally, if $\rho\lesssim\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}$, then we have
+Differential privacy was first introduced by Dwork et al., (2006) in the context of private data analysis.
+Formally, it requires the output distribution of a randomized algorithm to be insensitive to small perturbations in the input dataset.
+A key insight in this field is the strong connection between robust statistics and differential privacy; specifically, under mild conditions, robust estimators can be converted into differentially private algorithms using the propose-test-release framework.
+Subsequently, differential privacy was applied to the goodness-of-fit testing (also called identity testing) in statistics.
+Pioneering works in this direction include Gaboardi et al., (2016); Rogers and Kifer, (2017); Cai et al., (2017).
+Definition 3.3 ($(\epsilon,\delta)$-differential privacy, Dwork et al., (2006)).
+For a pair of non-negative real numbers $(\epsilon,\delta)$, a randomized algorithm $\mathcal{A}$ defined on the sample space $\Omega$ is a $(\epsilon,\delta)$-differentially private algorithm if for any $S\subset\text{range}(\mathcal{A})$ and any $X,X^{\prime}\in\Omega$ with the Hamming distance $d_{H}(X,X^{\prime})\leq 1$, we have
+$\mathbb{P}\left(\mathcal{A}(X)\in S\right)\leq e^{\epsilon}\mathbb{P}\left(\mathcal{A}(X^{\prime})\in S\right)+\delta.$ (3.1)
+When $\delta=0$, the notion is called pure differential privacy; when $\delta>0$, it is referred to as approximate differential privacy.
+Our second lower bound relies on a fundamental coupling lemma: if two datasets $\mathbf{X}$ and $\mathbf{X}^{\prime}$, drawn from $\mathcal{U}$ and $\mathcal{V}$ respectively, can be coupled to overlap on a significant fraction of indices (i.e., $X_{i}=X^{\prime}_{i}$ for most $i$), then distinguishing between $\mathcal{U}$ and $\mathcal{V}$ is statistically hard.
+Lemma 3.4 (Core lemma).
+Suppose $\mathbf{X}=(X_{1},X_{2},\dots,X_{N})^{\top}$ and $\mathbf{X}^{\prime}=(X^{\prime}_{1},X_{2}^{\prime},\dots,X_{N}^{\prime})^{\top}$ are random matrices with rows $X_{i},X^{\prime}_{i}\in\mathbb{R}^{d}$.
+Assume the laws of $\mathbf{X}$ and $\mathbf{X}^{\prime}$ are $\mathcal{U}$ and $\mathcal{V}$, respectively.
+Additionally, assume that there exists a coupling between $\mathcal{U}$ and $\mathcal{V}$ under which the expectation of the Hamming distance between $\mathbf{X}$ and $\mathbf{X}^{\prime}$ is $\mathcal{O}(\epsilon N)$, i.e.,
+$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\lesssim\epsilon N.$ (3.2)
+Then, there exists no robust test capable of distinguishing between $\mathcal{U}$ and $\mathcal{V}$ under contamination level of $\varOmega(\epsilon)$ while achieving type-I and type-II error probabilities below $0.1$ simultaneously and uniformly.
+The proof of Lemma 3.4 relies on a result established by Acharya et al., (2018), which states that if (3.2) holds for a $(0,\delta)$-differentially private algorithm, then necessarily $\epsilon N=\Omega(1/\delta)$.
+Furthermore, for any robust test attempting to distinguish between $\mathcal{U}$ and $\mathcal{V}$, one can construct a $(0,\delta)$-differentially private algorithm via the black-box robustness-to-privacy transformation proposed in Canonne et al., (2023) and Hopkins et al., (2023).
+Combining these two arguments yields a necessary condition for the expected Hamming distance between $\mathbf{X}$ and $\mathbf{X}^{\prime}$.
+The following theorem, which establishes the second lower bound for the testing problem (1.2), builds upon Lemma 3.4 and the selected vector $\theta\in K$ associated with the second optimal dimension $k_{2}^{\star}$ (see Lemma 3.1).
+Specifically, let $\mathcal{U}=\mathbb{P}_{0}^{\otimes N}$ and $\mathcal{V}=\mathbb{P}_{1}^{\otimes N}$, where $\mathbb{P}_{0}$ denotes $\mathcal{N}(0,\sigma^{2})$ and $\mathbb{P}_{1}$ is the mixture of $\mathcal{N}(v,\sigma^{2})$ with $v$ drawn uniformly from the set
+$V:=\left\{v\left\lvert\right.v_{i}=\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},i\geq 1,i\in I\right\}.$
+We construct a chain of distributions interpolating between $\mathcal{U}$ and $\mathcal{V}$ based on the optimal coupling (Lemma B.5).
+Consequently, the computation of the expected Hamming distance is reduced to the sum of the total variation distances between adjacent distributions in the chain.
+We kindly refer readers to Appendix C.3 for the formal proofs of Lemma 3.4 and Theorem 3.5.
+Theorem 3.5 (Second lower bound).
+Consider the testing problem (1.2) under the strong $\epsilon$-contamination model with an adversary $\mathcal{C}$.
+If the signal strength satisfies $\rho\lesssim\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}$, then no test can simultaneously achieve type-I and type-II errors bounded by $0.1$.
+Formally, under this condition, we have
 $\inf\limits_{\phi\in A_{s}(0.1)}\sup\limits_{\mu\in K}\sup\limits_{\mathcal{C}}\mathbb{P}_{\mu}(\phi(\mathcal{C}(\mathbf{X}^{\prime}))=0)\geq 0.1,$
-where $A_{s}(0.1)$ is the set of acceptable tests under $0.1$ type-I error defined as (1.3).
-We use $0.1$ for convenience, but it can be replaced by any unversal sufficiently small constant.
-It is worth noting that in the high-dimensional setting, where $d$, and possibly $k_{2}^{\star}$ can be substantially larger than $N$, one should be cautious that $\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}$ shall not exceeds $D_{0}(K)$, which is the maximal possible $\ell_{2}$ norm of the vector in $K$.
-The desired test will be impossible if this is not achieved.
+where $A_{s}(0.1)$ denotes the set of tests with type-I error at most $0.1$, as defined in (1.3).
+In Lemma 3.4 and Theorem 3.5, we fix the constant at $0.1$ for simplicity, though it can be replaced by any sufficiently small universal constant.
+In the high-dimensional setting where $d$ (and possibly $k_{2}^{\star}$) may substantially exceed $N$, one must ensure that
+$\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}\sigma}{N^{1/4}}\leq D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}.$
+If this condition is violated, the testing problem becomes statistically impossible.
 3.1.3 Based on the Estimation Lower Bounds
-The final lower bound, as established in Theorem 3.4 below, contrasts with Theorems 3.2 and 3.3 in that it is determined solely on the diameter of $K$ (which is $D_{0}(k)$), the fraction of contamination $\epsilon$, and the scale of the noise $\sigma$ while remaining independent of $N$.
-Theorem 3.4.
-For the signal detection problem (1.2), if $\epsilon\gtrsim\frac{1}{\sqrt{N}}$ and $\rho\lesssim\epsilon\sigma$, we always have
-$\inf\limits_{\phi\in A_{s}}\sup\limits_{\mu\in K,\left\lVert\mu\right\rVert_{2}\geq\rho}\sup\limits_{\mathcal{C}}\mathbb{P}_{\mu}(\phi(\mathcal{C}(\widetilde{\mathbf{X}}))=0)\gtrsim 1,$ (3.1)
+The third lower bound, as established in Theorem 3.6 below, contrasts with Theorems 3.2 and 3.5 in that it is determined solely by the diameter of $K$ (denoted as $D_{0}(K)$), the fraction of contamination $\epsilon$, and the noise scale $\sigma$, while remaining independent of $N$.
+Theorem 3.6 (Third lower bound).
+For the signal detection problem (1.2) under the strong $\epsilon$-contamination model with an adversary $\mathcal{C}$, if $\epsilon\gtrsim\frac{1}{\sqrt{N}}$ and $\rho\lesssim\epsilon\sigma$, we always have
+$\inf\limits_{\phi\in A_{s}}\sup\limits_{\mu\in K,\left\lVert\mu\right\rVert_{2}\geq\rho}\sup\limits_{\mathcal{C}}\mathbb{P}_{\mu}(\phi(\mathcal{C}(\widetilde{\mathbf{X}}))=0)=\varTheta(1),$ (3.3)
 where $A_{s}$ is the set of acceptable tests defined as (1.3).
 Remark.
-The condition $\epsilon\gtrsim\frac{1}{\sqrt{N}}$ can be safely removed. See our discussions in Section 3.3.
-Recall that $D_{0}(K)=\sup\limits_{\theta\in K}\left\lVert\theta\right\rVert_{2}$ is the supremum of the distance between a point in $K$ and the origin.
-(3.1) implies that to find a test with type-I and type-II error uniformly (over the adversary $\mathcal{C}$ and the mean $\mu$) less than $\alpha$, we should have either $\rho\gtrsim\epsilon\sigma$ or $\epsilon=\mathcal{O}(1/\sqrt{N})$.
-Since we always have $\left\lVert\mu\right\rVert_{2}\leq D_{0}(K)$, consequently, if $\epsilon\gtrsim 1/\sqrt{N}$ and $\epsilon\sigma\gtrsim D_{0}(K)$, the desired test will be theoretically impossible.
-The proof of Theorem 3.4 is motivated by a similar lower bound established for the corresponding estimation problem under star-shaped constraints by Prasadan and Neykov, (2025).
-It is based on a careful construction of a mixture distribution which is represented in two different ways, and a computation of the total variation between them.
-The full proof is provided in Section B.4.
-This bound aligns with intuition, as testing is generally easier than estimation.
-See Section 5 for further details.
+(1).
+This lower bound is notable as it coincides with the minimax rate for the estimation problem under the strong $\epsilon$-contamination model (see, e.g., Prasadan and Neykov, (2025)).
+This finding stands in contrast to the conventional wisdom that testing is typically statistically easier than estimation (see Ingster, 1993a ; Baraud, (2002)).
+However, our subsequent analysis confirms that in this setting, the minimax rates for estimation and testing indeed match (or are of the same order) when certain conditions are met.
+(2). The condition $\epsilon\gtrsim\frac{1}{\sqrt{N}}$ can be safely removed. See our discussions in Section 3.3.
+Recall that $D_{0}(K)=\sup_{\theta\in K}\lVert\theta\rVert_{2}$ represents the radius of $K$ (i.e., the maximal distance from the origin).
+Theorem 3.6 implies that a necessary condition for the existence of a test with type-I and type-II errors uniformly bounded by $\alpha$ is that either $\rho\gtrsim\epsilon\sigma$ or $\epsilon=\mathcal{O}(1/\sqrt{N})$.
+Noting that the signal strength is inherently bounded by $\lVert\mu\rVert_{2}\leq D_{0}(K)$, it follows that if $\epsilon\gtrsim 1/\sqrt{N}$ and $\epsilon\sigma\gtrsim D_{0}(K)$, consistent testing is statistically impossible.
+The proof of Theorem 3.6 is inspired by a similar lower bound established for the estimation problem under star-shaped constraints in Prasadan and Neykov, (2025).
+It relies on the construction of a specific distribution that admits two distinct mixture representations, with components corresponding to $H_{0}$ and $H_{1}$ respectively.
+Crucially, the adversary $\mathcal{C}$ can exploit this structural ambiguity to design a contamination strategy that renders the resulting distributions statistically indistinguishable when the lower bound condition is violated.
+The full proof is detailed in Appendix C.4.
 3.2 Upper Bounds
+The upper bounds are established constructively.
+Specifically, we design explicit testing procedures and identify the required signal strength $\rho=\lVert\mu\rVert_{2}$ that guarantees test validity (i.e., uniformly bounded type-I and type-II errors).
 3.2.1 Theoretical Algorithm
-An intuitive strategy for handling the adversary is to identify subsets of $\mathbf{X}$ that exhibit a consistency property.
-As we will see later, such subsets behave as if they consisted entirely of authentic samples with high probability.
+Since the adversary $\mathcal{C}$ is allowed to modify only a limited number of samples, a natural strategy to mitigate the contamination is to identify subsets of $\mathbf{X}$ that exhibit a consistency property.
 Definition 3.7 (Consistent subset).
@@
  (ii), $\phi(S^{\prime})=\phi(S)$ for any $S^{\prime}\subset S$ with $\left|S^{\prime}\right|\geq(1-2\epsilon)N$.
-The following theorem establishes the existence of a valid testing procedure under condition (3.2).
-Specifically, we demonstrate that, when (3.2) holds, there exists a desired test $\phi_{e}$ and at least one consistent subset $S$ with respect to $\phi_{e}$, upon which the entire testing task becomes feasible.
-To achieve this, we explicitly define an event $E_{e}$ of $\chi^{2}$-type, on which $\phi_{e}$ is determined.
-Furthermore, we prove that the set of original samples $\widetilde{\mathbf{X}}$, though not observed, is consistent with high probability.
-Finally, we show that the proposed test $\phi_{e}$ achieves uniformly small type-I and type-II errors.
-The detailed proof is provided in Section B.5.
-Theorem 3.6.
-Given a dataset $\mathbf{X}$ from the strong $\epsilon$-contamination model (see Definition 2.8) and a fixed error tolerance $\alpha$, for any $k\in[1,d]$ [^2] When $K\subset\ell_{2}$, we allow $1\leq k<\infty$., there exists a test $\phi_{e}$ for (1.2) such that the type-I and type-II errors are both uniformly less than $\alpha$ if $\rho^{2}\gtrsim E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k)$ where $E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k)$ is defined as
+[…]
+[…]
+The following theorem formally establishes the existence of a valid testing procedure under the condition (3.4), and consequently provides an upper bound for the testing problem (1.2).
+Specifically, we demonstrate that, when (3.4) holds, there exists a desired test $\phi_{e}$ and at least one consistent subset $\mathbf{X}_{S}$ with respect to $\phi_{e}$, upon which the entire testing task becomes feasible.
+To achieve this, we explicitly define an event $E_{e}$ of $\chi^{2}$-type, on which $\phi_{e}$ is determined, and prove that the original samples $\widetilde{\mathbf{X}}$ is consistent with high probability.
+The detailed proof is provided in Appendix C.5.
+Theorem 3.8 (Theoretical upper bound).
+Given a dataset $\mathbf{X}$ from the strong $\epsilon$-contamination model and a fixed error tolerance $\alpha$, for any $k\in[1,d]$ [^2] When $K\subset\ell_{2}$, we allow $1\leq k<\infty$., there exists a test $\phi_{e}$ for (1.2) such that the type-I and type-II errors are both uniformly less than $\alpha$ if $\rho^{2}\gtrsim E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k)$ where $E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k)$ is defined as
 $E_{\text{raw}}^{2}(\epsilon,K,N,\sigma^{2},k):=D_{k}^{2}(K)+\sigma^{2}\max\limits\left\{\frac{\sqrt{k}}{N},\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right),\sqrt{\frac{\epsilon^{2}\ln\left(\frac{1}{\epsilon}\right)k}{N}}\right\}.$ (3.4)
@@
 $\rho_{\text{critical}}\lesssim\min\limits_{k}E_{\text{raw}}(\epsilon,K,N,\sigma^{2},k).$
-We can further leverage the results of Theorem 3.6 by the definition of $k_{1}^{\star}$ and $k_{2}^{\star}$ and optimizing over $k$.
+We can further leverage the results of Theorem 3.8 by optimizing over $k$ and applying the definition of $k_{1}^{\star}$ and $k_{2}^{\star}$.
 Corollary 3.9.
@@
 ∎
-Corollary 3.7 actually selects the smaller optimal dimension between $k_{1}^{\star}$ and $k_{2}^{\star}$ and consequently the corresponding optimal projection, based on which $\phi_{e}$ is performed.
-We will show in Section 3.3 that the upper bound (3.3) does nearly match the theoretical lower bounds in the sense that it only has a redundant $\ln\left(\frac{1}{\epsilon}\right)$ factor.
-A significant improvement implied by Theorem 3.6 is that it does not require $\epsilon\rightarrow 0$, which is an essential assumption in many previous works.
-However, the test $\phi_{e}$ is limited in that its construction requires a full scan of all subsets of $\mathbf{X}$, which consumes exponential time with respect to $N$.
-This prohibitive computational cost prevents it from being practical when $N$ is large.
-In the next section, motivated by the filtering method given by Canonne et al., (2023), we propose an algorithm with polynomial-degree time complexity when the prior constraint $K$ and adversary $\mathcal{C}$ present.
-3.2.2 Polynomial Time Algorithm
-We will develop a polynomial time algorithm [^3] Here when we say “polynomial time algorithm”, we do assume that calculating the optimal projections (resulting in the optimal dimensions $k_{1}^{\star}$ and $k_{2}^{\star}$) can be done in polynomial time — as is the case for any ellipsoid or hyperrectangle.
-in this section to detect whether $\mu=0$ or not while requiring a slightly stronger condition compared to the bound (3.2).
-Specifically, we will prove the following theorem.
-Theorem 3.8.
+Corollary 3.9 effectively selects the smaller of the optimal dimensions $k_{1}^{\star}$ and $k_{2}^{\star}$, and consequently determines the optimal projection on which $\phi_{e}$ is based.
+We show in Section 3.3 that the upper bound (3.5) matches the theoretical lower bounds up to a logarithmic factor of $\ln(1/\epsilon)$.
+A significant improvement implied by Theorem 3.8 is that it does not require the asymptotic condition $\epsilon\rightarrow 0$, which is an essential assumption in many previous works.
+However, the test $\phi_{e}$ is limited by its reliance on an exhaustive search over all subsets of $\mathbf{X}$, resulting in exponential computational complexity with respect to $N$.
+This prohibitive cost renders the test computationally infeasible for large sample sizes.
+In the next section, motivated by the filtering method in Canonne et al., (2023), we propose a polynomial-time algorithm designed to handle the prior constraint $K$ in the presence of the adversary $\mathcal{C}$.
+3.2.2 Polynomial-Time Algorithm
+In this section, we develop a polynomial-time algorithm [^3] By polynomial-time”, we assume that computing the optimal projections (determining the optimal dimensions $k_{1}^{\star}$ and $k_{2}^{\star}$) can be performed in polynomial time, as is the case for any ellipsoid or hyperrectangle.
+to detect whether $\mu=\mathbf{0}$ under a slightly stronger condition than the bound (3.4).
+In contrast to the theoretical algorithm which requires an exhaustive search, the proposed algorithm leverages weights to adaptively filter potentially corrupted samples.
+Specifically, this filtering strategy and the algorithm are built and developed upon the framework of $\omega$-regularity, established in prior works (e.g., Canonne et al., (2023); Dong et al., (2019)).
+Definition 3.10 ($\omega$-regularity).
+Given a weight vector $\omega=(\omega_{1},\dots,\omega_{N})^{\top}$, $\mathbf{Y}$ is said to be $(\epsilon,\beta_{1},\beta_{2})$-regular if for all subsets $S\subset[N]$ with $\left|S\right|\leq\epsilon N$, we have the following properties:
+(i), $\left|\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\beta_{1}$, (ii), $\left|\left\lVert\sum\limits_{i\in S}\sqrt{\omega}_{i}Y_{i}\right\rVert_{2}^{2}-\left\lVert\omega_{S}\right\rVert_{1}k^{\star}\right|\leq c\beta_{2}$, and (iii), $\left|\left\langle\sum\limits_{i\in S}\sqrt{\omega}_{i}Y_{i},\sum\limits_{j\in[N]}\sqrt{\omega}_{j}Y_{j}\right\rangle-\left\lVert\omega_{S}\right\rVert_{1}k^{\star}\right|\leq c\sqrt{N}\beta_{1}$,
+where we recall that $\omega_{S}\in\mathbb{R}^{N}$ means the restriction of $\omega$ on the set $S$.
+The intuition behind the polynomial-time algorithm parallels that of the theoretical construction, although establishing this connection relies on more advanced random matrix theory concerning (sub-)Gaussian random variables (see Appendix B for technical preliminaries).
+We first establish that the authentic samples satisfy the $\omega$-regularity condition for specified parameters $\beta_{1},\beta_{2}$ with high probability (setting the initial weights $\omega=\mathbf{1}_{[N]}$).
+Our analysis demonstrates that, under sufficient conditions on the signal strength $\rho=\lVert\mu\rVert_{2}$, the re-weighted samples effectively mimic the behavior of the unobserved authentic data $\widetilde{\mathbf{X}}$.
+Consequently, we are able to recover the test result corresponding to $\widetilde{\mathbf{X}}$.
+Such conditions are formalized in the following theorem.
+Theorem 3.11 (Polynomial upper bound).
 Assuming the same conditions as in Theorem 3.8, for any $1\leq k\leq d$ [^4] When $K\subset\ell_{2}$, we allow $1\leq k<\infty$., as long as the following condition is satisfied, there exists a test $\phi_{p}$ to distinguish between $H_{0}$ and $H_{1}$ with type-I and type-II errors both uniformly less than $\alpha$.
@@
 $\rho_{\text{critical}}\lesssim\min\limits_{k}P_{\text{raw}}(\epsilon,K,N,\sigma^{2},k)$
-In addition, the time and space complexity of the algorithm is at most polynomial in terms of the parameters $N,k,\epsilon,\frac{1}{\alpha}$.
+In addition, the time and space complexity of the algorithm is at most polynomial in terms of the parameters $N,k,\frac{1}{\epsilon},\frac{1}{\alpha}$.
 Similar to Theorem 3.8, the following corollary characterizes the upper bound (3.6) with $k_{1}^{\star}$ and $k_{2}^{\star}$.
@@
 However, one can easily check that the upper bound (3.5) is generally tighter than (3.7), i.e., $E(\epsilon,K,N,\sigma^{2})\leq P(\epsilon,K,N,\sigma^{2})$.
-We must note here that our proof of Theorem 3.8 and the algorithm are inspired by the approach undertaken by Canonne et al., (2023).
-However, the algorithms and results of Canonne et al., (2023) only work in the region when $\left\lVert\mu\right\rVert_{2}\lesssim 1$ and there is no constraint under the alternative, i.e., $K=\mathbb{R}^{d}$.
-Therefore, we crucially modified the algorithm, including techniques to allow for $\mu$ of arbitrary (Euclidean) length and the use of the projection $P^{\star}$ to allow for the prior knowledge on $\mu$.
-We also fixed some flaws left in the argument of Canonne et al., (2023).
-We kindly refer readers to Appendix B.6 for the complete proof of Theorem 3.8.
+We acknowledge that the proof of Theorem 3.11 and the associated algorithm build upon the framework established in Canonne et al., (2023).
+However, the results in Canonne et al., (2023) are restricted to the regime where $\lVert\mu\rVert_{2}\lesssim 1$ in the unconstrained setting (i.e., $K=\mathbb{R}^{d}$).
+Consequently, we have significantly extended the algorithm to accommodate arbitrary signal magnitudes and to incorporate the prior constraint $K$ via the optimal projection $P^{\star}$.
+Furthermore, our analysis refines the arguments of Canonne et al., (2023) to address certain technical issues.
+We kindly refer readers to Appendix C.6 for the complete proof of the properties regarding the $\omega$-regularity and Theorem 3.11 and Appendix A for the pseudo code of the polynomial-time algorithm.
 3.3 Main Results
 In this section, we integrate the preceding theorems and bounds and summarize the main results.
-The lower bound is synthesized from Theorems 3.2, 3.3, 3.4.
+The lower bound is synthesized from Theorems 3.2, 3.5, 3.6.
 Although the bounds are derived through different approaches, an interesting phase transition phenomenon emerges, determined by the ranges of $\epsilon$ over which each term dominates others.
@@
 $\rho^{2}\gtrsim\sigma^{2}\max\limits\left\{\frac{\sqrt{k_{1}^{\star}}}{N},\epsilon\sqrt{\frac{k_{2}^{\star}}{N}},\epsilon^{2}\right\},$ (3.8)
-where $N$ is the sample size, $\sigma^{2}$ is the variance of the noise, $\epsilon$ is the fraction of contamination, and $k_{1}^{\star},k_{2}^{\star}$ are the optimal dimensions exclusively determined by $N$, $\sigma$, $\epsilon$ and the set $K$ as in 2.3 and 2.4.
+where $N$ is the sample size, $\sigma^{2}$ is the variance of the noise, $\epsilon$ is the fraction of contamination, and $k_{1}^{\star},k_{2}^{\star}$ are the optimal dimensions exclusively determined by $N$, $\sigma$, $\epsilon$ and the set $K$ as in Definition 2.3, 2.4.
+Roughly speaking, Theorem 3.13 demonstrates that the minimax lower bound of the testing problem (1.2) is jointly determined by three parts — the geometric properties of the set $K$, which is represented by the factors $k_{1}^{\star},k_{2}^{\star}$; the properties of the original data, which is implied by $N$ and $\sigma$; and the corruption process, which is implied by $\epsilon$.
+The corruption fraction $\epsilon$ connects the three components in the lower bound.
 In the proof of Corollary 3.15 below, we show that the first term, which is independent with $\epsilon$ [^5] We should recall that $k_{1}^{\star}$ is independent with $\epsilon$, dominates the other two if and only if $\epsilon\leq\frac{1}{\sqrt{N}}$.
@@
 Based on this argument, (3.8) can be equivalently written as
-$\rho\gtrsim\left\{\begin{tabular}[]{lll}$\frac{\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$,&$\left(0\leq\epsilon\lesssim\frac{1}{\sqrt{N}}\right)$&(Free contamination phase),\\ $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}$,&$\left(\frac{1}{\sqrt{N}}\lesssim\epsilon\lesssim\sqrt{\frac{k_{2}^{\star}}{N}}\right)$&(Balance phase),\\ $\epsilon^{2}\sigma^{2}$,&$\left(\epsilon\gtrsim\sqrt{\frac{k_{2}^{\star}}{N}}\right)$,&(Overly contaminated phase).\end{tabular}\right$.
+$\rho^{2}\gtrsim\left\{\begin{tabular}[]{lll}$\frac{\sqrt{k_{1}^{\star}}}{N}\sigma^{2}$,&$\left(0\leq\epsilon\lesssim\frac{1}{\sqrt{N}}\right)$&(Free contamination phase),\\ $\epsilon\sqrt{\frac{k_{2}^{\star}}{N}}\sigma^{2}$,&$\left(\frac{1}{\sqrt{N}}\lesssim\epsilon\lesssim\sqrt{\frac{k_{2}^{\star}}{N}}\right)$&(Balance phase),\\ $\epsilon^{2}\sigma^{2}$,&$\left(\epsilon\gtrsim\sqrt{\frac{k_{2}^{\star}}{N}}\right)$,&(Overly contaminated phase).\end{tabular}\right$.
 (3.9)
 Informally, a message implied by (3.9) is that after the “safe threshold” $\frac{1}{\sqrt{N}}$, every single contaminated sample in the strong $\epsilon$-contamination model invalidates the statistical testing use of $\sqrt{\frac{N}{k_{2}^{\star}}}$ uncontaminated samples on average, until the overly contaminated phase.
-For the upper bound, the theorem below is a natural downstream conclusion of Theorem 3.6 and 3.8.
+For the upper bound, Theorem 3.14 below is a natural downstream conclusion of Theorem 3.8 and 3.11.
 Theorem 3.14 (Main upper bound).
-Given the same setting as in Theorem 3.10, if $\rho\gtrsim E(\epsilon,K,N,\\ \sigma^{2})$, there exists a test $\phi_{e}$ achieving type-I and type-II errors uniformly less than $\alpha$.
-Furthermore, if $\rho\gtrsim P(\epsilon,K,N,\sigma^{2})$ and the determination of $k_{1}^{\star}$ and $k_{2}^{\star}$ can be completed within polynomial time of $\left(N,d,\frac{1}{\epsilon}\right)$, there exists a test $\phi_{p}$ that is able to finish the testing process within polynomial time of $\left(N,d,\frac{1}{\epsilon},\frac{1}{\alpha}\right)$ while still achieving type-I and type-II errors uniformly less than $\alpha$.
+Given the same setting as in Theorem 3.13, if $\rho\gtrsim E(\epsilon,K,N,\sigma^{2})$, there exists a test $\phi_{e}$ achieving type-I and type-II errors uniformly less than $\alpha$.
+Furthermore, if $\rho\gtrsim P(\epsilon,K,N,\sigma^{2})$ and the determination of $k_{1}^{\star}$ and $k_{2}^{\star}$ can be completed within polynomial-time of $\left(N,d,\frac{1}{\epsilon}\right)$, there exists a test $\phi_{p}$ that is able to finish the testing process within polynomial-time of $\left(N,k_{1}^{\star},k_{2}^{\star},\frac{1}{\epsilon},\frac{1}{\alpha}\right)$ while still achieving type-I and type-II errors uniformly less than $\alpha$.
 $E(\epsilon,K,N,\sigma^{2})$ and $P(\epsilon,K,N,\sigma^{2})$ are defined in (3.5) and (3.7) respectively.
@@
 The upper bound $P(\epsilon,K,N,\sigma^{2})$ of $\phi_{p}$ nearly matches $E(\epsilon,K,N,\sigma^{2})$ with only exceptions of the logarithmic factors $\ln N$, $\ln\alpha$, and $\ln\left(\frac{1}{\epsilon}\right)$.
-For Corollary 3.12, it involves some refined analysis with (3.6) and $E(\epsilon,K,N,\sigma^{2})$. We kindly refer readers to Appendix B.7 for the detailed arguments.
-3.4 Extension to $\ell_{p}$ Norm Separation for $1\leq p<2$
+For Corollary 3.15, it involves some refined analysis with the main lower bound (3.8) and $E(\epsilon,K,N,\sigma^{2})$. We kindly refer readers to Appendix C.7 for the detailed arguments.
+4 Extension to $\ell_{p}$ Norm Separation for $1\leq p<2$
 As an alternative, with the same assumptions on the samples $\widetilde{\mathbf{X}}$ and the adversary $\mathcal{C}$, we may consider the following $\ell_{p}$ robust testing problem.
@@
 $H_{1}:\left\lVert\mu\right\rVert_{p}\geq\rho,\mu\in K.$
-The classic results for the testing problem (1.2) can be naturally generalized to (3.8) when $1\leq p<2$ via the Hölder’s inequality (see Lemma A.17) by Ingster and Suslina, (2003).
+The classic results for the testing problem (1.2) can be naturally generalized to (4.1) when $1\leq p<2$ via the Hölder’s inequality by Ingster and Suslina, (2003).
 While Ingster’s approach cannot be directly transferred to our settings, we establish similar generalized results with the $p$-convex condition and accordingly adjusted optimal dimensions and projections.
 Note that in this section, we shall substitute the $\ell_{2}$ norm with the $\ell_{p}$ norm for the Kolmogorov $k$-width.
-Furthermore, to simplify the analysis, we consider the discretized version (see the remark after Definition 2.2).
+Furthermore, to simplify the analysis, we only consider the discretized version (see the remark after Definition 2.2).
 Definition 4.1 (Discretized Kolmogorov $k$-Width).
@@
 When the norm is $\left\lVert\cdot\right\rVert_{p}$, let us denote the corresponding Kolmogorov $k$-width for $K$ as $\widetilde{D}_{p,k}(K)$.
-Definition 3.14 ($p$-convex set).
-For $p\geq 1$, a set $K\subset\mathcal{X}$ is $p$-convex if
+Definition 4.2 ($p$-convex orthosymmetric (PCO) set).
+For $p\geq 1$, a set $K\subset\mathcal{X}$ is a $p$-convex orthosymmetric set if
  (1). $K$ is convex;
  (2). $K^{p}$ is convex, where $K^{p}$ is defined as
-$K^{p}:=\left\{\left(\left|\theta_{1}\right|^{p},\dots,\left|\theta_{d}\right|^{p}\right)^{\top}\left\lvert\right.\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K\right\}.$
+$K^{p}:=\left\{\left(\left|\theta_{1}\right|^{p},\dots,\left|\theta_{d}\right|^{p}\right)^{\top}\left\lvert\right.\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K\right\};$
+(3), $K$ is orthosymmetric, which means that if $\theta=\left(\theta_{1},\dots,\theta_{d}\right)^{\top}\in K$, then $\theta_{\eta}:=\left(\eta_{1}\theta_{1},\dots,\eta_{d}\theta_{d}\right)^{\top}\in K$, where $\eta_{i}\in\left\{-1,1\right\},1\leq i\leq d$.
+It is not difficult to verify that the whole space, hyperrectangles and ellipsoids are still all PCO sets for $1\leq p<2$.
 Analogous to previous analysis with the $\ell_{2}$ norm, the new lower bound relies on the following revised version of Lemma 3.1.
@@
 $\left\lVert\theta\right\rVert_{p}\gtrsim\sigma\max\limits\left\{\frac{(k_{p,1}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{\sqrt{N}},\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{N^{\frac{1}{4}}},\epsilon(k_{p,3}^{\star})^{\frac{1}{p}-\frac{1}{2}}\right\}.$ (4.3)
-where $N$ is the sample size, $\sigma^{2}$ is the variance of the noise, $\epsilon$ is the fraction of contamination, and $k_{p,1}^{\star},k_{p,2}^{\star},k_{p,3}^{\star}$ are the optimal dimensions exclusively determined by $N$, $\sigma$, $\epsilon$ and the set $K$ as in 3.16.
+where $N$ is the sample size, $\sigma^{2}$ is the variance of the noise, $\epsilon$ is the fraction of contamination, and $k_{p,1}^{\star},k_{p,2}^{\star},k_{p,3}^{\star}$ are the optimal dimensions exclusively determined by $N$, $\sigma$, $\epsilon$ and the set $K$ as in Definition 4.4.
 Undoubtedly, the lower bound (4.3) recovers (3.8) as expected when $p=2$.
@@
 When considering the $\ell_{p}$ norm separation, we require $k_{p,3}^{\star}$ to adapt to this discrepancy.
-We defer the proof of Theorem 3.17 to Appendix B.8.
+We defer the proof of Theorem 4.5 to Appendix C.8.
 In $\ell_{2}$ norm separation, the upper bound (3.5) matches the lower bound (3.8) except for logarithmic factors.
-We identify similar results in $\ell_{p}$ norm separation, which is summarized as the theorem below.
+We identify similar results in $\ell_{p}$ norm separation, which is summarized as Theorem 4.6 below.
 Theorem 4.6 ($\ell_{p}$ main upper bound).
 If $\rho\gtrsim\sigma\max\limits\left\{\frac{(k_{p,1}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{\sqrt{N}},\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{\frac{1}{p}-\frac{1}{4}}}{N^{\frac{1}{4}}},\epsilon(k_{p,3}^{\star})^{\frac{1}{p}-\frac{1}{2}}\right\}$, then there exists a test to distinguish between $H_{0}$ and $H_{1}$ with uniformly small type-I and type-II errors.
-Moreover, the time complexity of the algorithm is polynomial degree in $\left(N,d,\frac{1}{\epsilon},\frac{1}{\alpha}\right)$.
+Moreover, the time complexity of the algorithm is polynomial-time in $\left(N,d,\frac{1}{\epsilon},\frac{1}{\alpha}\right)$.
 Note that in the expression above, we omit the polylogarithmic factors in $\left(N,k_{p,1}^{\star},k_{p,2}^{\star},k_{p,3}^{\star},\frac{1}{\alpha},\frac{1}{\epsilon}\right)$.
@@
 To overcome the difficulty specified above, we again need to exploit the properties of the $p$-convex orthosymmetric set and the Kolmogorov widths.
-The proof of Theorem 3.18 is provided in Appendix B.9.
+The proof of Theorem 4.6 is provided in Appendix C.9.
 5 Experiments
-In this section, we present the results of simulation to verify the correctness of the theorems above and algorithms in Appendix C on the application side.
-We will generate artificial data, design different strategies for the adversary to contaminate the data, and run the (polynomial time) algorithm to check its effectiveness.
-The pseudo code is attached in Appendix C.
+In this section, we present the results of simulation to verify the correctness of the theorems above and algorithms in Appendix A on the application side.
+We generate artificial data, design different strategies for the adversary to contaminate the data, and run the (polynomial-time) algorithm to check its effectiveness.
+The pseudo code is attached in Appendix A.
 5.1 Data Generation
@@
 where $k^{\star}=\min\limits\left\{k_{1}^{\star},k_{2}^{\star}\right\}$.
-We verify in Appendix B.10 that the approach (4.2) ensures that the resulting vector $\mu_{c}$ achieves the minimal value of $\left\lVert P^{\star}\mu\right\rVert_{2}$ among the set $\left\{v\left\lvert\right.v\in\mathbb{R}^{d},v\in K_{e},\left\lVert v\right\rVert_{2}=c\right\}$, where $P^{\star}$ is the optimal projection operator corresponding to $k^{\star}$.
+We verify in Appendix C.10 that the approach (5.2) ensures that the resulting vector $\mu_{c}$ achieves the minimal value of $\left\lVert P^{\star}\mu\right\rVert_{2}$ among the set $\left\{v\left\lvert\right.v\in\mathbb{R}^{d},v\in K_{e},\left\lVert v\right\rVert_{2}=c\right\}$, where $P^{\star}$ is the optimal projection operator corresponding to $k^{\star}$.
 We believe that this property places $\mu_{c}$ in a near-worst-case position for the algorithm, and thus provides an appropriate setting to test its robustness.
@@
 5   Compute $\texttt{dist}\in\mathbb{R}^{N}$, $\texttt{dist}_{i}=\left\lVert\widetilde{X}_{i}\right\rVert_{2},1\leq i\leq N$
-6 $C=\text{indices of the largest }\epsilon N\text{ entries in }\texttt{arr}$
+6 $C=\text{indices of the largest }\epsilon N\text{ entries in }\texttt{dist}$
 7 $\mathbf{X}_{C}=\epsilon N\text{ i.i.d. fake data from }\mathcal{N}\left(\mathbf{0},\mathbf{I}_{d}\right)$
@@
 11   Compute $\mu_{0}$ according to the constraint type
-12  Compute $\texttt{dist}\in\mathbb{R}^{N}$, $\texttt{arr}_{i}=\left\lVert\widetilde{X}_{i}-\mu_{0}\right\rVert_{2}$
-13 $C=\text{indices of the largest }\epsilon N\text{ entries in }\texttt{arr}$
+12  Compute $\texttt{dist}\in\mathbb{R}^{N}$, $\texttt{dist}_{i}=\left\lVert\widetilde{X}_{i}-\mu_{0}\right\rVert_{2}$
+13 $C=\text{indices of the largest }\epsilon N\text{ entries in }\texttt{dist}$
 14 $\mathbf{X}_{C}=\epsilon N\text{ i.i.d. fake data from }\mathcal{N}\left(\mu_{0},\mathbf{I}_{d}\right)$
@@
 5.4 Experimental Results
-The experimental results are presented in the following figures.
-During the experiments, we ensured that the type-I error remained near zero, while identifying the empirical $\widetilde{\rho}_{\text{critical}}$ under $H_{1}$ as the critical value of $c$ when the corresponding type-II error exactly crosses $\alpha=0.05$.
-Figure 3: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and classic setting with $N=200$.
-Figure 4: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and classic setting with $N=2000$.
-Figure 5: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and high-dimensional setting with $N=100$.
-Figure 6: plot of rate of correct reject versus $\ell_{2}$ norm of the mean under $H_{1}$ and high-dimensional setting with $N=500$.
+Experimental results for $N=200$ in the classical case and $N=100$ in the high-dimensional case are presented in the following figures.
+Results for other parameter configurations are similar and are deferred to Appendix A.
+In our experiments, we maintained the type-I error near zero while identifying the empirical $\widetilde{\rho}_{\text{critical}}$ under $H_{1}$ as the critical value of $\rho$ when the corresponding type-II error exactly crosses $\beta=0.05$.
+Figure 3: Top: Rejection rate (or Empirical power) versus the $\ell_{2}$ norm of the mean under $H_{1}$ in the classic setting ($N=200$).
+Bottom: Rejection rate versus the $\ell_{2}$ norm of the mean in the high-dimensional setting ($N=100$).
 We observed several notable phenomena from the simulation results.
-First, the algorithm performs reasonably well across all parameter configurations: it successfully rejects $H_{0}$ when $\left\lVert\mu\right\rVert_{2}$ is sufficiently large and fails to reject when the signal strength is too weak, in both the classical and the high-dimensional regimes.
+First, the algorithm performs reasonably well across all parameter configurations, even for the moderate sample sizes: it successfully rejects $H_{0}$ when $\left\lVert\mu\right\rVert_{2}$ is sufficiently large and fails to reject when the signal strength is too weak, in both the classical and the high-dimensional regimes.
 Second, focusing on $\rho_{\text{critical}}$ — the critical norm of $\mu$ that yields a type-II error of $\alpha=0.05$ — we find that $\rho_{\text{critical}}$ decreases as $N$ increases, and increases as $d$ grows.
@@
 6 Discussion
+6.1 Rotated QCO Sets
 For completeness, we note that if the set $K$ is an orthogonal rotation of a QCO set — i.e., $K=UK^{\prime}$ for some orthogonal matrix $U$ — then one may transform the data by considering $U^{\top}X$ and subsequently solve the corresponding problem on the QCO set $K^{\prime}$.
@@
 Moreover, the computational cost of determining the optimal dimensions for $K$ and $K^{\prime}$ is essentially the same in most cases.
+6.2 Equivalence in Optimal Dimensions
 An alternative but equivalent way to define the optimal dimension is to simply select $k_{\text{alt}}^{\star}$ satisfying $D_{k_{\text{alt}}^{\star}-1}(K)>\left(\frac{k_{\text{alt}}^{\star}}{N}\right)^{1/4}\max\limits\left\{\sqrt{\epsilon},N^{-\frac{1}{4}}\right\}\sigma$ and $D_{k_{\text{alt}}^{\star}}(K)\leq\left(\frac{k_{\text{alt}}^{\star}+1}{N}\right)^{1/4}\max\limits\left\{\sqrt{\epsilon},N^{-\frac{1}{4}}\right\}\sigma$.
@@
 Similarly, the upper bounds $E_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ and $P_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ are obtained by substituting all $k_{1}^{\star}$ and $k_{2}^{\star}$ in $E^{2}(\epsilon,K,N,\sigma^{2})$ and $P^{2}(\epsilon,K,N,\sigma^{2})$ by $k_{\text{alt}}^{\star}$.
-With the proof in Appendix B.7, one can verify that $k_{\text{alt}}^{\star},P_{\text{alt}}^{\star}$ and its corresponding testing process are equivalent to that of $k_{1}^{\star}$, $k_{2}^{\star}$, $P_{1}^{\star}$, and $P_{2}^{\star}$.
+With the proof in Appendix C.7, one can verify that $k_{\text{alt}}^{\star},P_{\text{alt}}^{\star}$ and its corresponding testing process are equivalent to that of $k_{1}^{\star}$, $k_{2}^{\star}$, $P_{1}^{\star}$, and $P_{2}^{\star}$.
 The new lower bound (6.1) and the upper bounds $E_{\text{alt}}^{2}(\epsilon,K,N,\sigma^{2})$ and $P^{2}(\epsilon,K,N,\sigma^{2})$ do match the previous lower bound (3.8) and the upper bounds (3.5), (3.7) as well.
+6.3 Testing-Estimation Relationship
 The results developed in Section 3 should be compared and contrasted with the existing results for the corresponding estimation problem.
 Specifically, in the estimation setting — where the same distribution assumptions and corruption model are imposed, but the goal is to estimate the mean vector $\mu$ — the minimax rates are established by Donoho et al., (1990) and Neykov, (2022).
+Although the authors do not explicitly employ the concepts of Kolmogorov widths and optimal dimensions, their results can nonetheless be interpreted within our framework.
 Notably, the Kolmogorov widths and the associated optimal dimensions again play a central role in these results.
@@
 $\sigma^{2}\max\limits\left\{\frac{\sqrt{k_{1}^{\star}}}{N},\epsilon\sqrt{\frac{k_{2}^{\star}}{N}},\epsilon^{2}\right\}\lesssim\sigma^{2}\max\limits\left\{\frac{k_{e}^{\star}}{N},\epsilon^{2}\right\}$ (6.4)
-We refer readers to Appendix B.11 for a straightforward computational verification of Lemma 5.3.
+We refer readers to Appendix C.11 for a straightforward computational verification of Lemma 6.3.
 Following the discussion after Corollary 3.15, we can also analyze the phase transition of the estimation problem.
@@
 An interesting corollary of their work is that as the neighborhood expands, the minimax testing rate converges to the minimax estimation rate — an effect that parallels a key phenomenon in our setting.
+Such match in estimation and testing rate is also observed by other work (e.g., Ndaoud, (2019)).
 Given the framework of their problem, it is natural to ask whether one can establish the minimax rate with constrained and adversarial setting of (6.5).
@@
 The authors also would like to thank Clément Canonne for a helpful email exchange in which Clément suggested to us that in the unconstrained problem, $\left\lVert\mu\right\rVert_{2}\gtrsim\epsilon\sigma$ is required, which we extended to show Theorem 3.6.
-Appendix A Some Basic Probabilistic Concepts and Useful Lemmas
+Appendix A Theoretical and Polynomial-Time Algorithms
+A.1 Theoretical Algorithm
+According to Theorem 3.8 and Section C.5, we first attempt to search for a consistent set $T$ with respect to $\phi_{e}$.
+$\phi_{e}(\mathbf{X}_{T})$ then determines whether we are to accept or reject $H_{0}$.
+1
+2import $\mathbf{X},N,K,\sigma$
+3
+4Compute $k_{1}^{\star},P_{1}^{\star},k_{2}^{\star},P_{2}^{\star}$ from $N,K,\sigma,\epsilon$ as Definition 2.3 and 2.4
+5if $k_{1}^{\star}\leq k_{2}^{\star}$ then
+6 $k^{\star}=k_{1}^{\star},P^{\star}=P_{1}^{\star}$
+7
+8else
+9 $k^{\star}=k_{2}^{\star},P^{\star}=P_{2}^{\star}$
+10
+[…]
+12for $T$ in $\left\{\text{all subsets with cardinality greater than }(1-\epsilon)N\right\}$ do
+13
+14 consistent_flag == True
+15 if $\left\lVert\sum\limits_{i\in T}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left|T\right|\sigma^{2}\geq c\left|T\right|^{2}E^{2}(\epsilon,K,N,\sigma)$ then
+16    subset_test == True
+17 else
+18    subset_test == False
+[…]
+20
+21    if $\left\lVert\sum\limits_{i\in T^{\prime}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left|T^{\prime}\right|\sigma^{2}\geq c\left|T^{\prime}\right|^{2}E^{2}(\epsilon,K,N,\sigma)$ then
+22       subsubset_test == True
+23    else
+24       subsubset_test == False
+25    if subset_test XOR subsubset_test then
+26       consistent_flag == False
+27       break
+28
+29 if consistent_flag then
+30    if subset_test then
+31         Reject $H_{0}$
+32       return
+33    else
+34         Accept $H_{0}$
+35       return
+36
+37
+38Accept $H_{0}$
+return
+Algorithm 3 theoretical algorithm when $\rho\gtrsim E(\epsilon,K,N,\sigma^{2})$
+A.2 Polynomial-Time Algorithm
+The polynomial-time algorithm combines the three sub-algorithms in Section C.6 together to filter the observations $\mathbf{X}$ and through the weight $\omega$.
+The theoretical guarantee is provided in Section C.6.
+1Set $\gamma_{1}=c\left[\sqrt{k^{\star}\ln\left(\frac{N}{\alpha}\right)}+\ln\left(\frac{N}{\alpha}\right)\right]$, $\texttt{count}=0,\texttt{i}=0$
+2while $\texttt{i}<N$ do
+3 if $\left|\left\lVert Y_{i}\right\rVert_{2}^{2}-k^{\star}\right|>\gamma_{1}$ then
+4
+5    $\texttt{count}=\texttt{count}+1$
+6    if $\texttt{count}>\epsilon N$ then
+7       return None
+8     Delete $Y_{i}$ from $\mathbf{Y}$
+9 $\texttt{i}=\texttt{i}+1$
+10
+return $\mathbf{Y}$
+Algorithm 4 Prefiltering.
+1Set $\gamma_{2}$ as (C.13), and $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
+2while $\lambda\geq\gamma_{2}$ do
+3   Set $v$ to be the unit singular vector associated with $\lambda$
+4  Compute $\tau_{i}=\left\langle v,Y_{i}\right\rangle^{2}\mathbf{1}_{\left\{\omega_{i}>0\right\}}$ for $1\leq i\leq N$
+5  Set II be the smallest index such that $\sum\limits_{i=1}^{I}\omega_{i}\geq 2\epsilon N$
+6  Update $w$ according to (C.14)
+7 if $\left\lVert\omega\right\rVert_{1}<N(1-2\epsilon)$ then
+8
+9    return None
+10
+11  Set $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$
+12
+return $\omega$
+Algorithm 5 Sample filtering when $N>k^{\star}$.
+1Set $\gamma_{3}$ as (C.21), and $\lambda=\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
+2while $\lambda\geq\gamma_{3}$ do
+3   Set $v$ to be the unit singular vector associated with $\lambda$
+4  Compute $\tau_{i}=\frac{v_{i}^{2}}{\omega_{i}}\mathbf{1}_{\left\{w_{i}>0\right\}}$
+5  Update $w$ according to (C.22)
+6 if $\left\lVert\omega\right\rVert_{1}<N(1-6\epsilon)$ then
+7
+8    return None
+9
+10  Set $\lambda=\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}$
+11
+return $\omega$
+Algorithm 6 Sample filtering when $N\leq k^{\star}$.
+1Compute $\tau_{i}=\left|\left\langle\sqrt{\omega_{i}}Y_{i},\sum\limits_{j=1}^{N}\sqrt{\omega_{j}}Y_{j}\right\rangle-\omega_{i}k^{\star}\right|\mathbf{1}_{\left\{\omega_{i}>0\right\}},1\leq i\leq N$
+2Sort $\tau_{i}$ by desreasing order and find the indices $\left\{i_{1},i_{2},\dots,i_{\epsilon N}\right\}$ corresponding to the first $\epsilon N$ maximal $\tau_{i}$
+3Set $\omega_{i_{1}},\omega_{i_{2}},\dots,\omega_{i_{\epsilon N}}$ to zero
+return $\omega$
+Algorithm 7 Weight filtering.
+1
+2import $\mathbf{X},N,K,\sigma=1$
+3import Algorithm 4, 5, 6, 7
+4
+[…]
+6Compute $k_{1}^{\star},P_{1}^{\star},k_{2}^{\star},P_{2}^{\star}$ from $N,K,\sigma=1,\epsilon$ as Definition 2.3 and 2.4
+7if $k_{1}^{\star}\leq k_{2}^{\star}$ then
+8 $k^{\star}=k_{1}^{\star},P^{\star}=P_{1}^{\star}$
+9
+10else
+11 $k^{\star}=k_{2}^{\star},P^{\star}=P_{2}^{\star}$
+12
+[…]
+14Execute Algorithm 4, and record the returning value $\texttt{R}_{1}$
+15if $\texttt{R}_{1}$ is None then
+16   Reject $H_{0}$
+17 return
+18
+19else
+20 $\mathbf{Y}=\texttt{R}_{1}$
+21
+22Set $\omega=\mathbf{1}_{[k^{\star}]}$
+23if $N>k^{\star}$ then
+24   Execute Algorithm 5, and record the returning value $\texttt{R}_{2}$
+25else
+26   Execute Algorithm 6, and record the returning value $\texttt{R}_{2}$
+27if $\texttt{R}_{2}$ is None then
+28   Reject $H_{0}$
+29 return
+30
+31else
+32 $\omega=\texttt{R}_{2}$
+33
+34Execute Algorithm 7, and record the returning value $\texttt{R}_{3}$
+[…]
+36if $\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}\right|\geq c_{2}N^{2}P(\epsilon,k^{\star},N,1)$ then
+37   Reject $H_{0}$
+38
+39else
+40   Accept $H_{0}$
+41
+return
+Algorithm 8 polynomial-time algorithm when $\rho\gtrsim P(\epsilon,k^{\star},N,1)$.
+A.3 Deferred Experiment Results
+The following figures display additional results deferred from Section 5.4.
+Figure 4: Top: Rejection rate (or Empirical power) versus the $\ell_{2}$ norm of the mean under $H_{1}$ in the classic setting ($N=2000$).
+Bottom: Rejection rate versus the $\ell_{2}$ norm of the mean in the high-dimensional setting ($N=500$).
+Appendix B Some Basic Probabilistic Concepts and Useful Lemmas
 In this section, we recall some basic probabilistic concepts and useful lemmas that are frequently mentioned and leveraged in this paper.
 Most of them are related to concentrations and the measures of distance between distributions.
-Definition A.1 (Total variation distance).
+Definition B.1 (Total variation distance).
 For a $\sigma$-algebra $(\Omega,\mathscr{F})$ and the probabilities $\mathcal{P},\mathcal{Q}$ defined on $\mathscr{F}$, the total variation distance between $\mathcal{P}$ and $\mathcal{Q}$ is defined as
@@
 When $\mathcal{P},\mathcal{Q}$ has densities $p,q$ with respect to some measure $\nu$ and space $\Omega_{\nu}$, it can be also easily shown that
-$\text{TV}(\mathcal{P},\mathcal{Q})=\displaystyle\int_{\Omega_{\nu}}\min\limits\left\{p,q\right\}d\nu=\frac{1}{2}\displaystyle\int_{\Omega_{\nu}}\left|p-q\right|d\nu.$ (A.1)
-Definition A.2 (Kullback-Leibler divergence (KL divergence)).
+$\text{TV}(\mathcal{P},\mathcal{Q})=\displaystyle\int_{\Omega_{\nu}}\min\limits\left\{p,q\right\}d\nu=\frac{1}{2}\displaystyle\int_{\Omega_{\nu}}\left|p-q\right|d\nu.$ (B.1)
+Definition B.2 (the Kullback-Leibler divergence (KL divergence)).
 Given the same $\sigma$-algebra and probabilities, additionally assume that $\mathcal{P}\ll\mathcal{Q}$. The KL divergence between $\mathcal{P}$ and $\mathcal{Q}$ is defined as
@@
 Contrasted with total variation distance, KL divergence is not symmetric, meaning that $\text{KL}(\mathcal{P},\mathcal{Q})\neq\text{KL}(\mathcal{Q},\mathcal{P})$ in general, even if $\frac{d\mathcal{P}}{d\mathcal{Q}}$ and $\frac{d\mathcal{Q}}{d\mathcal{P}}$ both exist.
-Lemma A.3 (Pinsker’s inequality).
+Lemma B.3 (Pinsker’s inequality).
 For the probability distributions $\mathcal{P},\mathcal{Q}$ with $\mathcal{P}\ll\mathcal{Q}$, we have
-$\text{TV}(\mathcal{P},\mathcal{Q})\leq\sqrt{2\text{KL}(\mathcal{P},\mathcal{Q})}.$
+$\text{TV}(\mathcal{P},\mathcal{Q})\leq\sqrt{\frac{1}{2}\text{KL}(\mathcal{P},\mathcal{Q})}.$
 When $\mathcal{P}\ll\mathcal{Q}$ and $\mathcal{Q}\ll\mathcal{P}$, it can be refined as
-$\text{TV}(\mathcal{P},\mathcal{Q})\leq\sqrt{2\min\limits\left\{\text{KL}(\mathcal{P},\mathcal{Q}),\text{KL}(\mathcal{Q},\mathcal{P})\right\}}.$
+$\text{TV}(\mathcal{P},\mathcal{Q})\leq\sqrt{\frac{1}{2}\min\limits\left\{\text{KL}(\mathcal{P},\mathcal{Q}),\text{KL}(\mathcal{Q},\mathcal{P})\right\}}.$
 There are many proofs for this lemma. One can refer Tsybakov, (2009) for a quick proof.
-Definition A.4 (Coupling).
+Definition B.4 (Coupling).
 Given probability distributions $\mathcal{P},\mathcal{Q}$, a valid coupling of $\mathcal{P}$ and $\mathcal{Q}$ is any pair of random variables $(X,Y)$ where the marginal distribution of $X$ ang $Y$ is $\mathcal{P},\mathcal{Q}$ respectively.
 The following classic lemma help us connect the total variation distance with the Hamming distance in the proof of Theorem 3.5.
-Lemma A.5.
+Lemma B.5.
 For any probability distributions $\mathcal{P}$ and $\mathcal{Q}$, their total variation distance equals to the infimum over the probability that $X$ is not equal to $Y$, where $(X,Y)$ is a valid coupling of $\mathcal{P}$ and $\mathcal{Q}$.
 […]
+The coupling that achieves the equality above is referred as the optimal coupling between $\mathcal{P}$ and $\mathcal{Q}$.
 The following definitions and lemmas concern the sub-Gaussian and sub-exponential random variables, which are closely related to the concentration bounds in the proof of the main theorems.
-Definition A.6 (sub-Gaussian random variables).
+Definition B.6 (Sub-Gaussian random variables).
 For a univariate random variable $X$, define its sub-Gaussian norm as $\left\lVert X\right\rVert_{\psi_{2}}:=\inf\limits\left\{t:t>0,\mathbb{E}\exp\left\{\frac{X^{2}}{t^{2}}\right\}\leq 2\right\}$.
@@
 The most frequently mentioned one is $\mathbb{P}(\left|X\right|\geq t)\leq 2\exp\left\{-\frac{t^{2}}{c\left\lVert X\right\rVert_{\psi_{2}}^{2}}\right\}$.
-Definition A.7 (sub-exponential random variables).
+Definition B.7 (Sub-exponential random variables).
 For a univariate random variable $X$, define its sub-exponential norm as $\left\lVert X\right\rVert_{\psi_{1}}:=\inf\limits\left\{t:t>0,\mathbb{E}\exp\left\{\frac{\left|X\right|}{t}\right\}\leq 2\right\}$.
@@
 The following is the famous Bernstein’s inequality, which handles multiple independent sub-Gaussian (sub-exponential) random variables. It also generalizes the equivalent definitions mentioned above.
-Lemma A.8 (Bernstein’s inequality).
+Lemma B.8 (Bernstein’s inequality).
 Assume $X_{1},\dots,X_{n}$ are independent, zero-mean random variables. If $X_{1},\dots,X_{n}$ are sub-Gaussian, then for any $t\geq 0$, we have
-$\mathbb{P}\left(\sum\limits_{i=1}^{n}X_{i}\geq t\right)\leq\exp\left\{-ct^{2}/\sum\limits_{i=1}^{n}\left\lVert X_{i}\right\rVert_{\psi_{2}}^{2}\right\}.$ (A.3)
+$\mathbb{P}\left(\sum\limits_{i=1}^{n}X_{i}\geq t\right)\leq\exp\left\{-ct^{2}/\sum\limits_{i=1}^{n}\left\lVert X_{i}\right\rVert_{\psi_{2}}^{2}\right\}.$ (B.3)
 If $X_{1},\dots,X_{n}$ are sub-exponential, then we have
-$\mathbb{P}\left(\sum\limits_{i=1}^{n}X_{i}\geq t\right)\leq\exp\left\{-c\min\limits\left\{t^{2}/\sum\limits_{i=1}^{n}\left\lVert X_{i}\right\rVert_{\psi_{1}}^{2},t/\max\limits_{i}\left\{\left\lVert X_{i}\right\rVert_{\psi_{1}}\right\}\right\}\right\}.$ (A.4)
+$\mathbb{P}\left(\sum\limits_{i=1}^{n}X_{i}\geq t\right)\leq\exp\left\{-c\min\limits\left\{t^{2}/\sum\limits_{i=1}^{n}\left\lVert X_{i}\right\rVert_{\psi_{1}}^{2},t/\max\limits_{i}\left\{\left\lVert X_{i}\right\rVert_{\psi_{1}}\right\}\right\}\right\}.$ (B.4)
 The next lemma is particularly useful when we control the $\ell_{2}$ norm of a random vector with independent sub-Gaussian coordinates.
-Lemma A.9.
+Lemma B.9.
 Assume $X=(X_{1},\dots,X_{n})\in\mathbb{R}^{d}$ where $X_{1},\dots,X_{n}$ are independent, $\mathbb{E}X_{i}=0,\text{Var}\left(X_{i}\right)=1$, and each $X_{i}$ is sub-Gaussian.
@@
 The definitions and lemmas above can be found in many modern statistics textbook. For an example, see Vershynin, (2024).
-The famous Hanson-Wright inequality is a powerful tool to bound the tail probability of quadratic forms of the sub-Gaussian random variables.
-It can also be viewed as a generalization of the Bernstein’s inequality (A.4).
+The famous Hanson–Wright inequality is a powerful tool to bound the tail probability of quadratic forms of the sub-Gaussian random variables.
+It can also be viewed as a generalization of the Bernstein’s inequality (B.4).
 See Hanson and Wright, (1971) for the initial paper and Rudelson and Vershynin, (2013) for a modern proof.
-Lemma A.10 (The Hanson-Wright inequality).
+Lemma B.10 (Hanson–Wright inequality).
 Assume that $X=(X_{1},\dots,X_{n})\in\mathbb{R}^{n}$ is a random vector with independent, zero-mean and sub-Gaussian coordinates.
@@
 Then for any $t\geq 0$, we have
-$\mathbb{P}\left(|X^{\top}AX-\mathbb{E}X^{\top}AX|>t\right)\leq 2\exp\left\{-c\min\limits\left\{\frac{t^{2}}{K^{4}\left\lVert A\right\rVert_{F}^{2}},\frac{t}{K^{2}\left\lVert A\right\rVert_{2}}\right\}\right\},$ (A.5)
+$\mathbb{P}\left(|X^{\top}AX-\mathbb{E}X^{\top}AX|>t\right)\leq 2\exp\left\{-c\min\limits\left\{\frac{t^{2}}{K^{4}\left\lVert A\right\rVert_{F}^{2}},\frac{t}{K^{2}\left\lVert A\right\rVert_{2}}\right\}\right\},$ (B.5)
 where $\left\lVert\cdot\right\rVert_{F}$ is the Frobenius norm of a matrix, and $\left\lVert\cdot\right\rVert_{2}$ is the $\ell_{2}$ operator norm of a matrix.
@@
 Additional properties of the Wishart distribution and its applications in estimation can be found in Kubokawa and Srivastava, (2008) and Bishop et al., (2018).
-Definition A.11 (The Wishart matrix).
+Definition B.11 (Wishart matrix).
 Given a matrix $X\in\mathbb{R}^{n\times d}$ where each row is independently drawn from $\mathcal{N}(0,\bm{\Sigma})$ for some $\bm{\Sigma}\in\mathbb{R}^{d\times d}$.
@@
 One can find a proof of this lemma in Wainwright, (2019) by the concentration bound for Lipschitz functions and the Sudakov-Fernique comparison inequality.
-Lemma A.12.
+Lemma B.12.
 Assume $X\in\mathbb{R}^{n\times d}$, and each row is independently drawn from $\mathcal{N}(0,\bm{\Sigma})$, then for any $\epsilon>0$, we have
@@
 Here $\sigma_{\text{max}}(X)$ (or $\sigma_{\text{min}}(X)$) is the max (or min) singular value of $X$ and $\gamma_{\text{max}}(\sqrt{\bm{\Sigma}})$ (or $\gamma_{\text{min}}(\sqrt{\bm{\Sigma}})$) is the max (or min) eigenvalue of $\sqrt{\bm{\Sigma}}$.
-We can leverage Lemma A.12 to bound the $\ell_{2}$ operator norm of $W_{d}(n)$.
-Corollary A.13.
+We can leverage Lemma B.12 to bound the $\ell_{2}$ operator norm of $W_{d}(n)$.
+Corollary B.13.
 Assume $W=\mathbf{X}^{\top}\mathbf{X}\in\mathbb{R}^{d\times d}$ is drawn from $W_{d}(n)$, and $n\geq d$, then for any $\delta>0$, with probability higher than $1-2\delta$, we have
@@
 Proof.
-Set $\bm{\Sigma}=\mathbf{I}_{d}$ in the first inequality of Lemma A.12, we have:
+Set $\bm{\Sigma}=\mathbf{I}_{d}$ in the first inequality of Lemma B.12, we have:
 $\mathbb{P}\left(\frac{\sigma_{\text{max}}(X)}{\sqrt{n}}\geq(1+\epsilon)+\sqrt{\frac{d}{n}}\right)\leq\exp\left\{-\frac{n\epsilon^{2}}{2}\right\}.$
@@
 It allows the mean vector $\mu$ of the normal distribution to be nonzero, but at the expense of introducing a dependence on $\left\lVert\mu\right\rVert_{2}$.
-Corollary A.14.
+Corollary B.14.
 Assume $\mathbf{X}\in\mathbb{R}^{n\times d}$ and each row is independently drawn from $\mathcal{N}(\mu,\mathbf{I}_{d})$ with $\rho:=\left\lVert\mu\right\rVert_{2}$, then for $\delta>0$, with probability higher than $1-2\delta$, we have:
@@
 $\left\lVert\mathbf{X}^{\top}\mathbf{X}-n\mathbf{I}_{d}\right\rVert_{2}=\left\lVert\sum\limits_{i=1}^{n}X_{i}X_{i}^{\top}-n\mathbf{I}_{d}\right\rVert_{2}\leq\left\lVert\sum\limits_{i=1}^{n}Z_{i}Z_{i}^{\top}-n\mathbf{I}_{d}\right\rVert_{2}+2\left\lVert\sum\limits_{i=1}^{n}\mu Z_{i}^{\top}\right\rVert_{2}+n\left\lVert\mu\mu^{\top}\right\rVert_{2}.$
-By Corollary A.13, we already have $\left\lVert\sum\limits_{i=1}^{n}Z_{i}Z_{i}^{\top}-n\mathbf{I}_{d}\right\rVert_{2}\leq C_{1}\left(\sqrt{nd}+\sqrt{n\ln\left(\frac{1}{\delta}\right)}+\ln\left(\frac{1}{\delta}\right)\right)$.
+By Corollary B.13, we already have $\left\lVert\sum\limits_{i=1}^{n}Z_{i}Z_{i}^{\top}-n\mathbf{I}_{d}\right\rVert_{2}\leq C_{1}\left(\sqrt{nd}+\sqrt{n\ln\left(\frac{1}{\delta}\right)}+\ln\left(\frac{1}{\delta}\right)\right)$.
 We also have $n\left\lVert\mu\mu^{\top}\right\rVert_{2}=n\alpha^{2}$.
@@
 $\left\lVert\sum\limits_{i=1}^{n}\mu Z_{i}^{\top}\right\rVert_{2}\leq\left\lVert\mu\right\rVert_{2}\cdot\left\lVert\sum\limits_{i=1}^{n}Z_{i}\right\rVert_{2}.$
-Since $Z_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ for each $i$, from A.9 we know that $\left\lVert\sum\limits_{i=1}^{n}\frac{Z_{i}}{\sqrt{n}}-d\right\rVert_{2}$ is a sub-Gaussian random variable and therefore with probability higher than $1-\delta$ we have:
+Since $Z_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ for each $i$, from B.9 we know that $\left\lVert\sum\limits_{i=1}^{n}\frac{Z_{i}}{\sqrt{n}}-d\right\rVert_{2}$ is a sub-Gaussian random variable and therefore with probability higher than $1-\delta$ we have:
 $\left\lVert\sum\limits_{i=1}^{n}\frac{Z_{i}}{\sqrt{n}}-d\right\rVert_{2}\leq C\sqrt{\ln\left(\frac{1}{\delta}\right)},$
@@
 ∎
-The next lemma is from Fact 4.2 in Dong et al., (2019). It is an analog of Corollary A.13, but within the field when the $\ell_{1}$ norm of a weight $\omega$ is small compared to $n$.
-Lemma A.15.
+The next lemma is from Fact 4.2 in Dong et al., (2019). It is an analog of Corollary B.13, but within the field when the $\ell_{1}$ norm of a weight $\omega$ is small compared to $n$.
+Lemma B.15.
 Assume $\mathbf{X}\in\mathbb{R}^{n\times d}$ and each row is independently drawn from $\mathcal{N}(0,\mathbf{I}_{d})$.
@@
 $\left\lVert\sum\limits_{i=1}^{n}\omega_{i}X_{i}X_{i}^{\top}\right\rVert_{2}\leq C_{1}\left(\epsilon n\ln\left(\frac{1}{\epsilon}\right)+d+\ln\left(\frac{1}{\delta}\right)\right).$
-Corollary A.16.
-Assuming the same conditions as Lemma A.15, we further allow $X_{i}$ from $\mathcal{N}(\mu,\mathbf{I}_{d})$ with $\left\lVert\mu\right\rVert_{2}\leq\alpha$.
+Corollary B.16.
+Assuming the same conditions as Lemma B.15, we further allow $X_{i}$ from $\mathcal{N}(\mu,\mathbf{I}_{d})$ with $\left\lVert\mu\right\rVert_{2}\leq\alpha$.
 As a consequence, the inequality will track the dependence on $\alpha$:
@@
 Proof.
-The proof is basically from Lemma A.15 the standard argument of concentration. For any $v$ such that $\left\lVert v\right\rVert_{2}=1$, we have:
+The proof is basically from Lemma B.15 the standard argument of concentration. For any $v$ such that $\left\lVert v\right\rVert_{2}=1$, we have:
 $v^{\top}\left(\sum\limits_{i=1}^{n}\omega_{i}X_{i}X_{i}^{\top}\right)v=v^{\top}\left(\sum\limits_{i=1}^{n}(X_{i}-\mu)(X_{i}-\mu)^{\top}\right)v+\sum\limits_{i=1}^{n}\omega_{i}(v^{\top}\mu)^{2}+2(v^{\top}\mu)\sum\limits_{i=1}^{n}\omega_{i}(v^{\top}\left(X_{i}-\mu\right))$
-Bound of the first term can be obtained from Lemma A.15.
-For the second term, by the Cauchy-Schwarz inequality we can bound it by $2\left\lVert\omega\right\rVert_{1}\cdot\alpha^{2}\leq C\cdot\epsilon n\alpha^{2}$.
+Bound of the first term can be obtained from Lemma B.15.
+For the second term, by the Cauchy–Schwarz inequality we can bound it by $2\left\lVert\omega\right\rVert_{1}\cdot\alpha^{2}\leq C\cdot\epsilon n\alpha^{2}$.
 For the third term, since $\sum\limits_{i=1}^{n}\omega_{i}v^{\top}(X_{i}-\mu)$ follows from $\mathcal{N}(0,\left\lVert\omega\right\rVert_{2}^{2})$, we know we can bound it by $C\cdot\alpha\left\lVert\omega\right\rVert_{2}\leq C\cdot\alpha\sqrt{\epsilon n}$.
@@
 Hölder’s inequality is a fundamentally important tool in probability, mathematical analysis and functional analysis. We briefly review it here.
-Lemma A.17 (Hölder’s inequality).
+Lemma B.17 (Hölder’s inequality).
 Let $(\Omega,\mathscr{F})$ be a $\sigma$-algebra equipped with a measure $\mu$.
@@
 The next lemma is a famous result for the combination number.
-Lemma A.18.
+Lemma B.18.
 For any $0\leq k\leq n$, we have $\sum\limits_{i=0}^{k}\binom{n}{i}\leq\left(\frac{en}{k}\right)^{k}$.
@@
 ∎
-Finally, we introduce an intuitive and useful result in convexity and combinatorial mathematics.
-We note that Canonne et al., (2023) first introduced this result in their paper but their proof contains a flaw which we fix below.
-Lemma A.19.
+Finally, we present an intuitive and useful result at the intersection of convexity and combinatorics.
+While this result was originally introduced in Canonne et al., (2023), their proof contains a technical gap which we address below.
+Lemma B.19.
 Denote $0,1]^{N}$ to be the set $\left\{(x_{1},x_{2},\dots,x_{N}):0\leq x_{i}\leq 1,\forall i\in[N]\right\}$.
@@
 ∎
-Appendix B Deferred Proofs of the Lemmas and Main Theorems
-B.1 Proof of Lemma 3.1
-The proof of Lemma 3.1 leverages key properties of QCO sets and applies the minimax theorem to a suitably truncated vector. A similar argument appears in Section 3.3.2 of Neykov, (2022).
+Appendix C Deferred Proofs of the Lemmas and Main Theorems
+C.1 Proof of Lemma 3.1
 Proof of Lemma 3.1.
@@
 ∎
-B.2 Proof of Theorem 3.2
+C.2 Proof of Theorem 3.2
 Proof of Theorem 3.2.
@@
 $\overset{\text{(i)}}{=}1-\alpha-\frac{1}{2}\left\lVert\mathbb{P}_{0}-\mathbb{P}_{\nu}\right\rVert_{1},$
-where (i) uses (A.1).
-Now assuming $\mathbb{P}_{\nu}\ll\mathbb{P}_{0}$, and let $L_{\nu}=\frac{d\mathbb{P}_{\nu}}{d\mathbb{P}_{0}}$ be the Radon-Nikodym derivative of $\mathbb{P}_{\nu}$ to $\mathbb{P}_{0}$.
-By Cauchy-Schwartz inequality, we have
+where (i) uses (B.1).
+Now assuming $\mathbb{P}_{\nu}\ll\mathbb{P}_{0}$, and let $L_{\nu}=\frac{d\mathbb{P}_{\nu}}{d\mathbb{P}_{0}}$ be the Radon–Nikodym derivative of $\mathbb{P}_{\nu}$ to $\mathbb{P}_{0}$.
+By Cauchy–Schwartz inequality, we have
 $\left\lVert\mathbb{P}_{0}-\mathbb{P}_{\nu}\right\rVert_{1}=\mathbb{E}_{0}\left|L_{\nu}-1\right|\leq\sqrt{\mathbb{E}_{0}\left(L_{\nu}-1\right)^{2}}=\sqrt{\mathbb{E}_{0}\left[L_{\nu}^{2}\right]-1}.$
@@
 implying that it is impossible to test the problem. Hence, $k_{1}^{\star}$ must be at least positive for sufficiently small $\alpha$. ∎
-B.3 Proof of Theorem 3.3
-In this section, we prove Theorem 3.3.
-Our proof relies on Lemma B.1 and a carefully designed optimal coupling.
-To establish Lemma B.1, we first recall the notion of differential privacy in B.2 and present two auxiliary lemmas, Lemma B.3 and Lemma B.4.
-Based on these tools, we then construct a symmetric set using the QCO properties and compute the Hamming distance as an application of Lemma A.5.
-The computation proceeds by constructing a sequence of couplings that reduces the evaluation of the Hamming distance into that of the total variation distance.
-Lemma B.1 (main lemma).
+C.3 Proof of Theorem 3.5
+In this section, we formally prove Theorem 3.5.
+We first recall the core lemma.
+To establish Lemma C.1, two auxiliary lemmas, Lemma C.3 and Lemma C.4, are presented.
+Based on these tools, we then construct a symmetric set served as the prior for the mean $\mu$ via the QCO properties and compute the Hamming distance as an application of Lemma B.5 via a sequential optimal coupling.
+Finally, the lower bound is yielded by the Pinsker’s inequality and a concrete calculation of the KL divergence.
+Lemma C.1 (Core lemma).
 Suppose $\mathbf{X}=(X_{1},X_{2},\dots,X_{N})^{\top}$ and $\mathbf{X}^{\prime}=(X^{\prime}_{1},X_{2}^{\prime},\dots,X_{N}^{\prime})^{\top}$ are random matrices with rows $X_{i},X^{\prime}_{i}\in\mathbb{R}^{d}$.
@@
 Additionally, assume that there exists a coupling between $\mathcal{U}$ and $\mathcal{V}$ under which the expectation of the Hamming distance between $\mathbf{X}$ and $\mathbf{X}^{\prime}$ is $\mathcal{O}(\epsilon N)$, i.e.,
-$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\lesssim\epsilon N.$ (B.1)
+$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\lesssim\epsilon N.$ (C.1)
 Then, there exists no robust test capable of distinguishing between $\mathcal{U}$ and $\mathcal{V}$ under contamination level of $\varOmega(\epsilon)$ while achieving type-I and type-II error probabilities below $0.1$ simultaneously and uniformly.
-To prove Lemma B.1, we start with the notion of $(\epsilon,\delta)$-differential privacy.
-Definition B.2 ($(\epsilon,\delta)$-differential privacy).
-For a pair of non-negative real numbers $(\epsilon,\delta)$, a randomized algorithm $\mathcal{A}$ that is defined on the sample space $\Omega$ is a $(\epsilon,\delta)$-differentially private algorithm if for any $S\subset\text{range}(\mathcal{A})$ and any $X,X^{\prime}\in\Omega$ with $d_{H}(X,X^{\prime})\leq 1$, we have
-$\mathbb{P}\left(\mathcal{A}(X)\in S\right)\leq e^{\epsilon}\mathbb{P}\left(\mathcal{A}(X^{\prime})\in S\right)+\delta.$ (B.2)
-When $\delta=0$, the algorithm is called pure differential privacy; when $\epsilon=0$, the algorithm is called $\delta$-fully approximate differential privacy.
-With this concept and motivated by Theorem 11 in Acharya et al., (2018), we are able to prove a lemma that is very close to Lemma B.1.
-Lemma B.3.
+We start with a review of the notion of $(\epsilon,\delta)$-differential privacy for convenience.
+Definition C.2 ($(\epsilon,\delta)$-differential privacy, Dwork et al., (2006)).
+For a pair of non-negative real numbers $(\epsilon,\delta)$, a randomized algorithm $\mathcal{A}$ defined on the sample space $\Omega$ is an $(\epsilon,\delta)$-differentially private algorithm if for any $S\subset\text{range}(\mathcal{A})$ and any $X,X^{\prime}\in\Omega$ with the Hamming distance $d_{H}(X,X^{\prime})\leq 1$, we have
+$\mathbb{P}\left(\mathcal{A}(X)\in S\right)\leq e^{\epsilon}\mathbb{P}\left(\mathcal{A}(X^{\prime})\in S\right)+\delta.$ (C.2)
+When $\delta=0$, the notion is called pure differential privacy; when $\delta>0$, it is referred to as approximate differential privacy.
+Lemma C.3 is motivated by Theorem 11 in Acharya et al., (2018), and shares a nearly identical conclusion with the core lemma.
+Lemma C.3.
 Assume that there is a coupling between $\mathcal{U}$ and $\mathcal{V}$ under which that the expectation of the Hamming distance between $\mathbf{X}$ and $\mathbf{X}^{\prime}$ is smaller or equal than $D$, i.e.,
-$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\leq D.$ (B.3)
-Then, if a robust test $\mathcal{A}:\Omega\mapsto\left\{0,1\right\}$ between $\mathcal{U}$ and $\mathcal{V}$ is a $\delta$-fully approximate differential private algorithm, with type-I and type-II error probabilities both below $0.1$, then we must have $D=\varOmega(\frac{1}{\delta})$.
+$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\leq D.$ (C.3)
+Then, if a test $\mathcal{A}$ to distinguish between $\mathcal{U}$ and $\mathcal{V}$: $\Omega\mapsto\left\{0,1\right\}$ is a $(0,\delta)$-differentially private algorithm, with type-I and type-II error probabilities both below $0.1$, then we must have $D=\varOmega(\frac{1}{\delta})$.
 Proof.
@@
 From the union bound, we obtain that
-$\mathbb{P}\left(\mathcal{A}(\mathbf{X})=0\cap\mathcal{A}(\mathbf{X}^{\prime})=1\cap d_{H}(\mathbf{X},\mathbf{X}^{\prime})<10D\right)\geq 0.7.$ (B.4)
+$\mathbb{P}\left(\mathcal{A}(\mathbf{X})=0\cap\mathcal{A}(\mathbf{X}^{\prime})=1\cap d_{H}(\mathbf{X},\mathbf{X}^{\prime})<10D\right)\geq 0.7.$ (C.4)
 By the definition of differential privacy, we can construct a chain from $\mathbf{X}$ to $\mathbf{X}^{\prime}$ such that each neighbour points differ on at most one coordinate, by which we have
-$|\mathbb{P}\left(\mathcal{A}(\mathbf{X})=0\right)-\mathbb{P}\left(\mathcal{A}(\mathbf{X}^{\prime})=0\right)|\leq\delta\cdot d_{H}(\mathbf{X},\mathbf{X}^{\prime}).$ (B.5)
-It is worth noting that the probability in (B.4) includes two parts of randomness: the randomness of $\mathbf{X}$ and $\mathbf{X}^{\prime}$ from $\Omega$, and the randomness of $\mathcal{A}$.
-Now take a realization of $\mathbf{X}$ and $\mathbf{X}^{\prime}$, denoted by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, that satisfy (B.4), we have
-$1-0.7\geq\mathbb{P}(\mathcal{A}(\mathbf{x}^{\prime})=0)\geq\mathbb{P}(\mathcal{A}(\mathbf{x})=0)-\delta\cdot d_{H}(\mathbf{x},\mathbf{x}^{\prime})\geq 0.7-\delta\cdot 10D,$ (B.6)
+$|\mathbb{P}\left(\mathcal{A}(\mathbf{X})=0\right)-\mathbb{P}\left(\mathcal{A}(\mathbf{X}^{\prime})=0\right)|\leq\delta\cdot d_{H}(\mathbf{X},\mathbf{X}^{\prime}).$ (C.5)
+It is worth noting that the probability in (C.4) includes two parts of randomness: the randomness of $\mathbf{X}$ and $\mathbf{X}^{\prime}$ from $\Omega$, and the randomness of $\mathcal{A}$.
+Now take a realization of $\mathbf{X}$ and $\mathbf{X}^{\prime}$, denoted by $\mathbf{x}$ and $\mathbf{x}^{\prime}$, that satisfy (C.4), we have
+$1-0.7\geq\mathbb{P}(\mathcal{A}(\mathbf{x}^{\prime})=0)\geq\mathbb{P}(\mathcal{A}(\mathbf{x})=0)-\delta\cdot d_{H}(\mathbf{x},\mathbf{x}^{\prime})\geq 0.7-\delta\cdot 10D,$ (C.6)
 from which we obtain that $10D\geq 0.4\frac{1}{\delta}$. The proof is completed. ∎
-The next lemma shows how to obtain a robust and differentially private test from a given robust test between $\mathcal{U}$ and $\mathcal{V}$ using a black-box robustness-to-privacy transformation.
+Lemma C.4 shows how to obtain a $(0,\delta)$-differentially private test from a given robust test between $\mathcal{U}$ and $\mathcal{V}$ using the black-box robustness-to-privacy transformation.
 The techniques are from Canonne et al., (2023) and Hopkins et al., (2023).
-Lemma B.4.
+Lemma C.4.
 Suppose $\mathcal{A}$ is a robust test for the mean testing problem (1.2) achieving type-I and type-II errors at most $0.1$ under a contamination level at most $\epsilon$.
-Then, for $\delta=\frac{1}{\epsilon N}$, there exists a $\delta$-fully approximate differential privacy $\widetilde{\mathcal{A}}$ whose error probabilities are also at most $0.1$ for any dataset with size $N$.
+Then, for $\delta=\frac{1}{\epsilon N}$, there exists a $(0,\delta)$-differentially private algorithm $\widetilde{\mathcal{A}}$ whose error probabilities are also at most $0.1$ for any dataset with size $N$.
 Proof.
@@
 Now assume $\widetilde{\mathbf{X}}$ is from $H_{0}$.
-By the definition of robust testing, with probability at least $0.9$, we can arbitrarily modify $\widetilde{\mathbf{X}}$ and obtain the same result as long as the number of contamination does not exceed $\epsilon N$.
+By the definition of robust testing, with probability at least $0.9$, we can arbitrarily modify $\widetilde{\mathbf{X}}$ and obtain the same result as long as the number of modifications does not exceed $\epsilon N$.
 This equivalently says that $\mathcal{S}(\widetilde{\mathbf{X}},\mathcal{A})\geq\epsilon N$, which, by the construction of $\widetilde{\mathcal{A}}$, implies that $\widetilde{\mathcal{A}}$ will output $1$ with probability 0 in this case.
 On the other hand, if $\widetilde{\mathbf{X}}$ is from $H_{1}$, then with probability at least $0.9$, we will have $\mathcal{A}(X)=1$, in which case $\mathcal{S}(X,\mathcal{A})=0$ and consequently $\widetilde{\mathcal{A}}$ will output $1$ with probability $1$.
-This shows that $\widetilde{\mathcal{A}}$ is a $\delta$-fully approximate differential privacy with error probabilities at most $0.1$ under both hypotheses.
+This shows that $\widetilde{\mathcal{A}}$ is a $(0,\delta)$-differential privacy with error probabilities at most $0.1$ under both hypotheses.
 ∎
-We are now prepared to prove Lemma B.1.
-Proof of Lemma B.1.
-Suppose we already have a robust test $\mathcal{A}$ for the mean testing problem 1.2 with the fraction of contamination at most $\epsilon$.
-Then by Lemma B.4, we can contruct a $\delta$-fully approximate differential privacy $\widetilde{\mathcal{A}}$ with errors at most $0.1$.
-Therefore, the constructed algorithm $\widetilde{\mathcal{A}}$ satifies the condition of Lemma B.3.
-Take $D=\mathbb{E}d_{H}(\mathbf{X},\mathbf{X}^{\prime})$ in Lemma B.3, and by the conclusion of Lemma B.3 we should have $\mathbb{E}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\gtrsim\frac{1}{1/\epsilon N}=\epsilon N$.
+We are now prepared to prove the core lemma and Theorem 3.5.
+Proof of Lemma C.1.
+Suppose we already have a robust test $\mathcal{A}$ for the mean testing problem (1.2) with the fraction of contamination at most $\epsilon$.
+Then by Lemma C.4, we can contruct a $(0,\delta)$-differential privacy $\widetilde{\mathcal{A}}$ with errors at most $0.1$.
+Therefore, the constructed algorithm $\widetilde{\mathcal{A}}$ satifies the condition of Lemma C.3.
+Take $D=\mathbb{E}d_{H}(\mathbf{X},\mathbf{X}^{\prime})$ in Lemma C.3, and by the conclusion of Lemma C.3 we should have $\mathbb{E}d_{H}(\mathbf{X},\mathbf{X}^{\prime})\gtrsim\frac{1}{1/\epsilon N}=\epsilon N$.
 This necessary condition finishes the proof.
@@
 samples from $\mathcal{N}\left(v,\sigma^{2}\mathbf{I}\right)$ with $v$ drawn from the uniform distribution on the set
-$V:=\left\{v\left\lvert\right.v_{i}=\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},1\leq i\in I\right\},$
+$V:=\left\{v\left\lvert\right.v_{i}=\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},i\geq 1,i\in I\right\},$
 where $\theta$ is the vector from Lemma 3.1 setting $c=\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$, i.e., $\left\lVert\theta\right\rVert_{2}=\frac{(k_{2}^{\star})^{1/4}\sqrt{\epsilon}}{N^{1/4}}\sigma$ and $\left\lVert\theta\right\rVert_{\infty}\leq\frac{\sqrt{\epsilon}}{(k_{2}^{\star})^{1/4}N^{1/4}}\sigma$.
-We are going to compute $d_{H}(\mathbf{X},\mathbf{X}^{\prime})$ using Lemma A.5.
+We are going to compute $d_{H}(\mathbf{X},\mathbf{X}^{\prime})$ using Lemma B.5.
 Specifically, we compare (in terms of total variation) a distribution of $k$ exchangeable variables generated from $\mathcal{V}$ to a joint distribution of k1k-1 exchangeable variables generated from $\mathcal{N}\left(v,\sigma^{2}\mathbf{I}\right)$ and the last one from $\mathcal{N}\left(0,\sigma^{2}\mathbf{I}\right)$ for $1\leq k\leq N$.
-We approach this by decomposing both $\mathbf{X}$ and $\mathbf{X}^{\prime}$ by their sequence of conditional probabilities and constructing the optimal coupling between $X_{k}$ and $X^{\prime}_{k}$ conditional on $X_{1},\dots,X_{k-1}$ using Lemma A.5 such that
+We approach this by decomposing both $\mathbf{X}$ and $\mathbf{X}^{\prime}$ by their sequence of conditional probabilities and constructing the optimal coupling between $X_{k}$ and $X^{\prime}_{k}$ conditional on $X_{1},\dots,X_{k-1}$ using Lemma B.5 such that
 $\mathbb{P}\left(X_{k}\neq X^{\prime}_{k}|X_{1},\dots,X_{k-1}\right)=\text{TV}(p(x_{k}|x_{1},\dots,x_{k-1}),q(x_{k})),$
 where $p(x_{k}|x_{1},\dots,x_{k-1})$ is the density of conditional distribution of the first $k$ coordinates of $X^{\prime}$ and $q(x_{k})$ is the density of standard normal distribution.
-By the definition of total variation distance, it is not hard to see under this optimal coupling that
-$\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim(\mathcal{U},\mathcal{V})}\mathbf{1}_{\left\{X_{k}\neq X^{\prime}_{k}\right\}}$ $=\mathbb{E}\text{TV}(p(x_{k}|x_{1},\dots,x_{k-1}),q(x_{k}))$
+If we denote the optimal coupling above by $\mathcal{P}_{k}=\mathcal{P}_{k}(X_{k},X^{\prime}_{k}|X_{1},...,X_{k-1})$, it is not hard to verify that
+$\widetilde{\mathcal{P}}(\mathbf{X},\mathbf{X}^{\prime}):=\mathcal{P}_{1}(X_{1},X^{\prime}_{1})\cdot\prod\limits_{i=2}^{N}\mathcal{P}_{i}(X_{k},X^{\prime}_{k}|X_{1},\dots,X_{k-1})$
+is a valid coupling between $\mathcal{U}$ and $\mathcal{V}$. By the property of the optimal coupling and the definition of total variation distance, we have under $\widetilde{\mathcal{P}}$ that
+$\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim\widetilde{\mathcal{P}}}\mathbf{1}_{\left\{X_{k}\neq X^{\prime}_{k}\right\}}$ $=\mathbb{E}\text{TV}(p(x_{k}|x_{1},\dots,x_{k-1}),q(x_{k}))$
 $=\text{TV}(p(x_{1},\dots,x_{k}),q(x_{1},\dots,x_{k})),$
 where $p(x_{1},\dots,x_{k})$ is the joint distribution of the first $k$ coordinates of $X^{\prime}$ and $q(x_{1},\dots,x_{k}):=p(x_{1},\dots,x_{k-1})\cdot q(x_{k})$. Since
-$\mathbb{E}_{(X,X^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(X,X^{\prime})=\mathbb{E}_{X,X^{\prime}\sim(\mathcal{U},\mathcal{V})}\sum\limits_{i=1}^{N}\mathbf{1}_{\left\{X_{i}\neq X^{\prime}_{i}\right\}},$
-to evaluate $\mathbb{E}_{(X,X^{\prime})\sim(\mathcal{U},\mathcal{V})}d_{H}(X,X^{\prime})$, it suffices to evaluate
+$\mathbb{E}_{(\mathbf{X},\mathbf{X}^{\prime})\sim\widetilde{\mathcal{P}}}d_{H}(\mathbf{X},\mathbf{X}^{\prime})=\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim\widetilde{\mathcal{P}}}\sum\limits_{i=1}^{N}\mathbf{1}_{\left\{X_{i}\neq X^{\prime}_{i}\right\}},$
+to evaluate $\mathbb{E}_{(\mathbf{X},\mathbb{X}^{\prime})\sim\widetilde{\mathcal{P}}}d_{H}(\mathbf{X},\mathbf{X}^{\prime})$, it suffices to evaluate
 $\sum\limits_{k=1}^{N}\text{TV}(p(x_{1},\dots,x_{k}),q(x_{1},\dots,x_{k})).$
-We calculate each individual term. By Pinsker’s inequality A.3, we have
+We bound each individual term by the Pinsker’s inequality, where we have
 $2\text{TV}^{2}(p(x_{1},\dots,x_{k}),q(x_{1},\dots,x_{k}))$ $\leq\text{KL}(p(x_{1},\dots,x_{k})|q(x_{1},\dots,x_{k}))$
-$=\mathbb{E}\ln\left(\frac{\frac{1}{2^{|V|}}\sum\limits_{v\in V}\prod\limits_{i=1}^{k}\exp\left\{-\frac{\left\lVert X_{i}-v\right\rVert_{2}^{2}}{2\sigma^{2}}\right\}}{\frac{1}{2^{|V|}}\sum\limits_{v\in V}\prod\limits_{i=1}^{k-1}\exp\left\{-\frac{\left\lVert X_{i}-v\right\rVert_{2}^{2}}{2\sigma^{2}}\right\}\cdot\exp\left\{-\frac{\left\lVert X_{k}\right\rVert_{2}^{2}}{2\sigma^{2}}\right\}}\right)$
 $\overset{\text{(i)}}{=}\mathbb{E}\ln\left(\frac{\frac{1}{2^{|V|}}\sum\limits_{v\in V}\exp\left\{\left(\sum\limits_{i=1}^{k}X_{i}\right)^{\top}v/\sigma^{2}\right\}}{\frac{1}{2^{|V|}}\sum\limits_{v\in V}\exp\left\{\left(\sum\limits_{i=1}^{k-1}X_{i}\right)^{\top}v/\sigma^{2}\right\}}\right)-\frac{\left\lVert v\right\rVert_{2}^{2}}{2\sigma^{2}}$
 $\overset{\text{(ii)}}{=}\mathbb{E}\ln\left(\frac{\prod\limits_{j=1}^{d}\cosh\left\{\left(\sum\limits_{i=1}^{k}X_{i}\right)_{j}v_{j}/\sigma^{2}\right\}}{\prod\limits_{j=1}^{d}\cosh\left\{\left(\sum\limits_{i=1}^{k-1}X_{i}\right)_{j}v_{j}/\sigma^{2}\right\}}\right)-\frac{\left\lVert v\right\rVert_{2}^{2}}{2\sigma^{2}},$
-where $(\cdot)_{j}$ denotes the $j$-th coordinate, (i) is from that $\left\lVert v\right\rVert_{2}^{2}$ does not depend on the specfic choice of $v\in V$ by the construction of $V$ and (ii) is from an important observation that
+where $(\cdot)_{j}$ denotes the $j$-th coordinate, (i) is from that $\left\lVert v\right\rVert_{2}^{2}$ does not depend on the specific choice of $v\in V$ by the construction of $V$ and (ii) is from an important observation that
 $\frac{1}{2^{|V|}}\sum\limits_{v\in V}\exp\left\{\left(\sum\limits_{i=1}^{k}X_{i}\right)^{\top}v\right\}=\prod\limits_{j=1}^{d}\cosh\left\{\left(\sum\limits_{i=1}^{k}X_{i}\right)_{j}v_{j}\right\}.$
@@
 Therefore, we have $\frac{\sqrt{N}\left\lVert v\right\rVert_{4}^{2}}{\sigma^{2}}\leq\frac{\sqrt{N}}{\sigma^{2}}\cdot\frac{\epsilon}{\sqrt{N}}\sigma^{2}\leq\epsilon$.
-By Lemma B.1, we have
+By the core lemma, we have
 $\inf\limits_{\phi\in A_{s}(0.1)}\sup\limits_{\mu\in K}\sup\limits_{\mathcal{C}}\mathbb{P}_{\mu}(\phi(\mathcal{C}(\widetilde{\mathbf{X}}))=0)\geq 0.1.$
 The proof is completed. ∎
-B.4 Proof of Theorem 3.4
+C.4 Proof of Theorem 3.6
 In this section, we prove Theorem 3.6.
-We reach the result by constructing two mixture distributions which are, in fact, identical.
+We reach the result by constructing two mixture distributions based on $H_{0}$ and $H_{1}$ which are, in fact, identical.
 We then evaluate the total variation distance between two truncated binomial distributions, which enables us to construct an adaptive adversary, leading high type-II error.
@@
 Let $p_{0}$ and $p_{\mu_{0}}$ be the probability density functions of $\mathbb{P}_{0}$ and $\mathbb{P}_{\mu_{0}}$, construct two distributions $Q_{1},Q_{2}$ with density
-$q_{1}=\frac{(p_{\mu_{0}}-p_{0})\mathbf{1}_{\left\{p_{\mu_{0}}\geq p_{0}\right\}}}{\text{TV}\left(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}}\right)},\ q_{2}=\frac{(p_{0}-p_{\mu_{0}})\mathbf{1}_{\left\{p_{0}>p_{\mu_{0}}\right\}}}{\text{TV}(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}})}.$ (B.7)
+$q_{1}=\frac{(p_{\mu_{0}}-p_{0})\mathbf{1}_{\left\{p_{\mu_{0}}\geq p_{0}\right\}}}{\text{TV}\left(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}}\right)},\ q_{2}=\frac{(p_{0}-p_{\mu_{0}})\mathbf{1}_{\left\{p_{0}>p_{\mu_{0}}\right\}}}{\text{TV}(\mathbb{P}_{0},\mathbb{P}_{\mu_{0}})}.$ (C.7)
 Consider the mixture measures
-$R_{1}:=(1-\epsilon^{\prime\prime})\mathbb{P}_{0}+\epsilon^{\prime\prime}Q_{1},\ R_{2}:=(1-\epsilon^{\prime\prime})\mathbb{P}_{\mu_{0}}+\epsilon^{\prime\prime}Q_{2}.$ (B.8)
+$R_{1}:=(1-\epsilon^{\prime\prime})\mathbb{P}_{0}+\epsilon^{\prime\prime}Q_{1},\ R_{2}:=(1-\epsilon^{\prime\prime})\mathbb{P}_{\mu_{0}}+\epsilon^{\prime\prime}Q_{2}.$ (C.8)
 It is not hard to verify that actually $R_{1}=R_{2}$, i.e., they represent the same distribution.
-Consequently $\text{TV}\left(\prod\limits_{i=1}^{N}R_{1}^{i},\prod\limits_{i=1}^{N}R_{2}^{i}\right)=0$, where $\prod\limits_{i=1}^{N}R_{j}^{i},j\in\left\{1,2\right\}$ is the $N$-fold product of $R_{j}$.
+Consequently $\text{TV}\left(R_{1}^{\otimes N},R_{2}^{\otimes N}\right)=0$, where $R_{j}^{\otimes N},j\in\left\{1,2\right\}$ is the $N$-fold product of $R_{j}$.
 Step III: expand the TV distance. Define
@@
 $\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{U}-S_{2}^{U}\right|d\mathbf{x}\leq\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}(S_{1}^{U}+S_{2}^{U})d\mathbf{x}\leq\mathbb{P}\left(B>N\left(\epsilon^{\prime\prime}+\frac{1}{\sqrt{2N}}\right)\right)\overset{\text{(i)}}{\leq}e^{-1}.$
-Here $B$ is a random variable that follows a binomial distribution with parameters $N$ and $\epsilon^{\prime\prime}$, and (i) follows from the Hoeffding’s inequality.
+Here $B$ is a binomial random variable with parameters $N$ and $\epsilon^{\prime\prime}$, and (i) follows from the Hoeffding’s inequality.
 By expanding the TV distance between the $N$-fold products of $R_{1}$ and $R_{2}$ and after some calculation, we have
-$0=\text{TV}\left(\prod\limits_{i=1}^{N}R_{1}^{i},\prod\limits_{i=1}^{N}R_{2}^{i}\right)\overset{\text{(ii)}}{\geq}\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{L}-S_{2}^{L}\right|d\mathbf{x}-\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{U}-S_{2}^{U}\right|d\mathbf{x},$
+$0=\text{TV}\left(R_{1}^{\otimes N},R_{2}^{\otimes N}\right)\overset{\text{(ii)}}{\geq}\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{L}-S_{2}^{L}\right|d\mathbf{x}-\frac{1}{2}\displaystyle\int_{\mathbb{R}^{d}}\left|S_{1}^{U}-S_{2}^{U}\right|d\mathbf{x},$
 where (ii) is from the definition of TV distance and the triangle inequality. Combining with previous argument, we obtain
@@
 Step IV: conditional argument.
-Consider the adversary $\mathcal{C}_{B,\epsilon^{\prime\prime}}$ with following strategy.
+Consider a adversary $\mathcal{C}_{B,\epsilon^{\prime\prime}}$ with following strategy.
 Upon observing the original data $\widetilde{\mathbf{X}}$ and its underlying law (which is either $\mathbb{P}_{0}$ or $\mathbb{P}_{\mu_{0}}$ in our case), the adversary first samples a number $n_{c}$ from the conditional law of $B|B\leq N(\epsilon^{\prime\prime}+1/\sqrt{2N})$, then sample $n_{c}$ replacements from the mixture distribution corresponding to the true underlying law.
@@
 […]
-$\geq 1-\alpha-\text{TV}\left(\widetilde{\mathbb{P}}_{0},\widetilde{\mathbb{P}}_{\mu_{0}}\right)\asymp 1.$
+$\geq 1-\alpha-\text{TV}\left(\widetilde{\mathbb{P}}_{0},\widetilde{\mathbb{P}}_{\mu_{0}}\right)=\varTheta(1).$
 Taking infimum over $\phi\in A_{s}$ on both sides completes the proof. ∎
-B.5 Proof of Theorem 3.6
+C.5 Proof of Theorem 3.8
 We start with a recall of the definition of the consistent subset, which plays a central role in the proof.
-Definition B.5 (Consistent subset).
+Definition C.5 (Consistent subset).
 A subset $S\subset\mathbf{X}=\left\{X_{1},\dots,X_{N}\right\}$ is called a consistent subset regarding the contamination fraction $\epsilon$ and a test $\phi:2^{S}\mapsto\left\{0,1\right\}$ if
@@
 Specifically, consider the following $\chi^{2}$-type event
-$E_{e}:=\left\{\left\lVert P^{\star}\sum\limits_{i\in T}X_{i}\right\rVert_{2}^{2}-k^{\star}\left|T\right|\sigma^{2}\geq c\left|T\right|^{2}E^{2}(\epsilon,K,N,\sigma^{2})\right\},$ (B.9)
+$E_{e}:=\left\{\left\lVert P^{\star}\sum\limits_{i\in T}X_{i}\right\rVert_{2}^{2}-k^{\star}\left|T\right|\sigma^{2}\geq c\left|T\right|^{2}E^{2}(\epsilon,K,N,\sigma^{2})\right\},$ (C.9)
 where $T$ is any subset of $\mathbf{X}$, and $c\in\mathbb{R}$ is some universal, properly chosen constant.
@@
 The key to the proof is carefully analyzing the $\ell_{2}$ norm of the sum of the projected original samples
-$\left\lVert P^{\star}\sum\limits_{i\in S^{\prime}}\widetilde{X}_{i}\right\rVert_{2}^{2}=\underbrace{\left\lVert P^{\star}\sum\limits_{i\in S^{\prime}}(\widetilde{X}_{i}-\mu)\right\rVert_{2}^{2}}_{:=\text{I}}+\underbrace{\left|S^{\prime}\right|^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}}_{:=\text{II}}+2\underbrace{\left|S^{\prime}\right|\sum\limits_{i\in S^{\prime}}\left[P^{\star}(\widetilde{X}_{i}-\mu)\right]^{\top}\mu}_{:=\text{III}}.$ (B.10)
+$\left\lVert P^{\star}\sum\limits_{i\in S^{\prime}}\widetilde{X}_{i}\right\rVert_{2}^{2}=\underbrace{\left\lVert P^{\star}\sum\limits_{i\in S^{\prime}}(\widetilde{X}_{i}-\mu)\right\rVert_{2}^{2}}_{:=\text{I}}+\underbrace{\left|S^{\prime}\right|^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}}_{:=\text{II}}+2\underbrace{\left|S^{\prime}\right|\sum\limits_{i\in S^{\prime}}\left[P^{\star}(\widetilde{X}_{i}-\mu)\right]^{\top}\mu}_{:=\text{III}}.$ (C.10)
 We attempt to show that the parts I and III, which are $\ell_{2}$ norm of the centralized projected vectors and the crossing term, are both asymptotically less than the part II under the condition (3.4), which vanishes under the null.
@@
 with probability greater than $1-\alpha$.
-For $\text{I}_{2,S^{\prime}}$, since we allow $S^{\prime}$ to be any subset of $S$, refined analysis is needed.
-Using the Hanson-Wright inequality A.10 and observing that $\left\lVert P^{\star}\right\rVert_{F}=\sqrt{k},\left\lVert P^{\star}\right\rVert_{2}\leq 1$, for each fixed $S^{\prime}\subset S$, we have
+For $\text{I}_{2,S^{\prime}}$, since we allow $S^{\prime}$ to be any subset of $[N]$, refined analysis is needed.
+Using the Hanson–Wright inequality B.10 and observing that $\left\lVert P^{\star}\right\rVert_{F}=\sqrt{k^{\star}},\left\lVert P^{\star}\right\rVert_{2}\leq 1$, for each fixed $S^{\prime}\subset[N]$, we have
 $\mathbb{P}\left(\left|\text{I}_{2}-k^{\star}\left|[N]\backslash S^{\prime}\right|\sigma^{2}\right|\geq t\right)\leq 2\exp\left\{-c\min\limits\left\{\frac{t^{2}}{k^{\star}\left|[N]\backslash S^{\prime}\right|^{2}\sigma^{4}},\frac{t}{\left|[N]\backslash S^{\prime}\right|\sigma^{2}}\right\}\right\}.$
@@
 $\text{III}_{2,S^{\prime}}$ 0 $\sqrt{2\epsilon N^{3}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\sigma^{2}}\cdot\sqrt{2\epsilon N\ln\left(\frac{e}{2\epsilon}\right)}$ $\alpha$
-Table 1: terms from (B.10) and corresponding means and variations.
+Table 1: corresponding means and variations of the terms from (C.10).
 To complete the proof, we finally need to compare the variation of each term above with the part II.
@@
 […]
-B.6 Proof of Theorem 3.8
+C.6 Proof of Theorem 3.11
 Motivated by Canonne et al., (2023), we demonstrate the rationale of Algorithm 8 in this section.
@@
 We depart from the definition of regularity, and arrive at the final complete testing procedure, Algorithm 8.
-The pseudo code is provided in Appendix C.
+The pseudo code is provided in Appendix A.
 Figure 5: roadmap of this section.
 We start with the definition of regularity and $\omega$-regularity, which controls the behaviour of the $\ell_{2}$ norm of the individual and sum, and the inner product.
-Definition B.6 (regularity).
+Definition C.6 (regularity).
 $\mathbf{Y}=\left\{Y_{1},\dots,Y_{N}\right\}$ is said to be $(\epsilon,\beta_{1},\beta_{2})$-regular if for all subsets $S\subset[N]$ with $\left|S\right|\leq\epsilon N$, we have the following properties:
@@
 $\omega$-regularity is a natural generalization of the definition above, which involves the interaction of a weight vector $\omega$ with $\mathbf{Y}$.
-Definition B.7 ($\omega$-regularity).
+Definition C.7 ($\omega$-regularity).
 Given a weight vector $\omega=(\omega_{1},\dots,\omega_{N})^{\top}$, $\mathbf{Y}$ is said to be $(\epsilon,\beta_{1},\beta_{2})$-regular if for all subsets $S\subset[N]$ with $\left|S\right|\leq\epsilon N$, we have the following properties:
@@
 Recall that $\omega_{S}\in\mathbb{R}^{N}$ means the restriction of $\omega$ on the set $S$.
-B.7 immediately recovers B.6 if $\omega=\mathbf{1}_{S}$.
+C.7 immediately recovers C.6 if $\omega=\mathbf{1}_{S}$.
 The next lemma states that any uncontaminated $\mathbf{Y}$ from $\mathcal{N}\left(\nu,\mathbf{I}_{k^{\star}}\right)$ is in fact $(\epsilon,\beta_{1},\beta_{2})$-regular with high probability and specified $\beta_{1}$ and $\beta_{2}$ below depending on $\epsilon,\left\lVert\nu\right\rVert_{2},N,k^{\star}$.
-Lemma B.8.
+Lemma C.8.
 Assume $Y_{1},\dots,Y_{N}$ are i.i.d.
 samples from $\mathcal{N}\left(\nu,\mathbf{I}_{k^{\star}}\right)$, then with probability greater than $1-\alpha$, $\mathbf{Y}=\left\{Y_{1},\dots,Y_{N}\right\}$ is $(\epsilon,\beta_{1},\beta_{2})$-regular with
-$\beta_{1}=\epsilon N\left(\sqrt{k^{\star}}+\sqrt{N}\left\lVert\nu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\sqrt{N}\left\lVert\nu\right\rVert_{2}^{2},$ (B.11)
+$\beta_{1}=\epsilon N\left(\sqrt{k^{\star}}+\sqrt{N}\left\lVert\nu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\sqrt{N}\left\lVert\nu\right\rVert_{2}^{2},$ (C.11)
 $\beta_{2}=\epsilon N\sqrt{\epsilon Nk^{\star}\ln\left(\frac{1}{\epsilon}\right)}+(\epsilon N)^{2}\ln\left(\frac{1}{\epsilon}\right)+\left\lVert\nu\right\rVert_{2}(\epsilon N)^{2}\sqrt{\ln\left(\frac{1}{\epsilon}\right)}+\left\lVert\nu\right\rVert_{2}^{2}(\epsilon N)^{2},$
@@
 with probability greater than $1-\frac{\alpha}{3N}$. Consequently, this holds for all $Y_{i},1\leq i\leq N$ with probability greater than $1-\frac{\alpha}{3}$ by the union bound. Therefore, we have
-$\left|\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\left[\epsilon N\sqrt{(k^{\star}+\left\lVert\nu\right\rVert_{2}^{2})\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\left\lVert\nu\right\rVert_{2}^{2}\right]$ (B.12)
+$\left|\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\left[\epsilon N\sqrt{(k^{\star}+\left\lVert\nu\right\rVert_{2}^{2})\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\left\lVert\nu\right\rVert_{2}^{2}\right]$ (C.12)
 Check of (ii).
@@
 $\mathbb{P}\left(\left|\left\lVert\frac{1}{\sqrt{\left|S\right|}}\sum\limits_{i\in S}(Y_{i}-\nu)\right\rVert_{2}^{2}-k^{\star}\right|\geq t\right)\leq 2\exp\left\{-c\min\limits\left\{\frac{t^{2}}{k^{\star}},t\right\}\right\}.$
-By Lemma A.18, we have at most $\left(\frac{e}{\epsilon}\right)^{\epsilon N}$ different possible $S$ in total.
+By Lemma B.18, we have at most $\left(\frac{e}{\epsilon}\right)^{\epsilon N}$ different possible $S$ in total.
 Setting the probability to be $\frac{\alpha}{6}\left(\frac{\epsilon}{e}\right)^{\epsilon N}$ and solving for $t$, we obtain
@@
 $\mathbb{P}\left(\left|\nu^{\top}\sum\limits_{i\in S}(Y_{i}-\nu)\right|\geq c\left\lVert\nu\right\rVert_{2}\sqrt{\epsilon N}\cdot\sqrt{\epsilon N\ln\left(\frac{1}{\epsilon}\right)}\right)\leq\frac{\alpha}{6}\left(\frac{\epsilon}{e}\right)^{\epsilon N}.$
-Combining the results together, and using the union bound over all possible $T$, we have
+Combining the results together, and using the union bound over all possible $S$, we have
 $\left|\left\lVert\sum\limits_{i\in S}Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\left[\epsilon N\sqrt{\epsilon Nk^{\star}\ln\left(\frac{1}{\epsilon}\right)}+(\epsilon N)^{2}\ln\left(\frac{1}{\epsilon}\right)+\left\lVert\nu\right\rVert_{2}(\epsilon N)^{2}\sqrt{\ln\left(\frac{1}{\epsilon}\right)}+\left\lVert\nu\right\rVert_{2}^{2}(\epsilon N)^{2}\right]$
-with probability greater than $1-\frac{\alpha}{3}$. The RHS is exactly $\beta_{2}$ in (B.11).
+with probability greater than $1-\frac{\alpha}{3}$. The RHS is exactly $\beta_{2}$ in (C.11).
 Check of (iii).
-Finally, we are going to evaluate the term $\left\langle\sum\limits_{i\in S}Y_{i},\sum\limits_{j\in S}Y_{j}\right\rangle$.
-Since this is equal to $\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}+\sum\limits_{i\in S}\left\langle Y_{i},\sum\limits_{j\in S\backslash i}Y_{j}\right\rangle$.
+Finally, we are going to evaluate the term $\left\langle\sum\limits_{i\in S}Y_{i},\sum\limits_{j\in[N]}Y_{j}\right\rangle$.
+Since this is equal to $\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}+\sum\limits_{i\in S}\left\langle Y_{i},\sum\limits_{j\in[N]\backslash i}Y_{j}\right\rangle$.
 We have obtained the bound for $\left\lVert Y_{i}\right\rVert_{2}^{2}$ in (i).
 For the term of inner product, we have
-$\left\langle Y_{i},\sum\limits_{j\in S\backslash i}Y_{j}\right\rangle=\left\langle Y_{i}-\nu,\sum\limits_{j\in S\backslash i}(Y_{j}-\nu)\right\rangle+\left\langle Y_{i}-\nu,(N-1)\nu\right\rangle+\left\langle\nu,\sum\limits_{j\in S\backslash i}(Y_{j}-\nu)\right\rangle+(N-1)\left\lVert\nu\right\rVert_{2}^{2}.$
+$\left\langle Y_{i},\sum\limits_{j\in[N]\backslash i}Y_{j}\right\rangle=\left\langle Y_{i}-\nu,\sum\limits_{j\in[N]\backslash i}(Y_{j}-\nu)\right\rangle+\left\langle Y_{i}-\nu,(N-1)\nu\right\rangle+\left\langle\nu,\sum\limits_{j\in[N]\backslash i}(Y_{j}-\nu)\right\rangle+(N-1)\left\lVert\nu\right\rVert_{2}^{2}.$
 Applying the sub-exponential concentration bound for the first term and sub-Gaussian concentration bound for the second as the previous handling, this is bounded by
@@
 $\leq$ $c\left[\epsilon N\left(\sqrt{Nk^{\star}}+N\left\lVert\nu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right]$
-Comparing with (B.12) and Definition B.6, we find that the required $\beta_{1}$ is
+Comparing with (C.12) and Definition C.6, we find that the required $\beta_{1}$ is
 $\epsilon N\left(\sqrt{k^{\star}}+\sqrt{N}\left\lVert\nu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\sqrt{N}\left\lVert\nu\right\rVert_{2}^{2},$
-which is exactly $\beta_{1}$ in (B.11). The proof is completed. ∎
-By Lemma B.8, we should be able to control the inner product between the sum of samples on two sets $S$ and $S^{\prime}$ by the polarization identity. The next corollary exactly states this benefit.
-Corollary B.9.
+which is exactly $\beta_{1}$ in (C.11). The proof is completed. ∎
+By Lemma C.8, we should be able to control the inner product between the sum of samples on two sets $S$ and $S^{\prime}$ by the polarization identity. The next corollary exactly states this benefit.
+Corollary C.9.
 Assume $\mathbf{Y}$ is $(2\epsilon,\beta_{1},\beta_{2})$-regular.
@@
 $\left\langle\sum\limits_{i\in S}Y_{i},\sum\limits_{j\in S^{\prime}}Y_{j}\right\rangle=\frac{1}{2}\left[\left\lVert\sum\limits_{i\in S\cup S^{\prime}}Y_{i}\right\rVert_{2}^{2}+\left\lVert\sum\limits_{i\in S\cap S^{\prime}}Y_{i}\right\rVert_{2}^{2}-\left\lVert\sum\limits_{i\in S\backslash S^{\prime}}Y_{i}\right\rVert_{2}^{2}-\left\lVert\sum\limits_{i\in S^{\prime}\backslash S}Y_{i}\right\rVert_{2}^{2}\right].$
-Since each of $\S,S\SS\cup S^{\prime},S\cap S^{\prime},S\backslash$ $^{\prime},S^{\prime}\backslash$ S has cardinality no more than $2\epsilon N$ by assumption, by Definition B.6 and Lemma B.8, we know
+Since each of $\S,S\SS\cup S^{\prime},S\cap S^{\prime},S\backslash$ $^{\prime},S^{\prime}\backslash$ S has cardinality no more than $2\epsilon N$ by assumption, by Definition C.6 and Lemma C.8, we know
 $\left|\left\langle\sum\limits_{i\in S}Y_{i},\sum\limits_{j\in S^{\prime}}Y_{j}\right\rangle-\frac{1}{2}\left(\left|S\cup S^{\prime}\right|+\left|S\cap S^{\prime}\right|-\left|S\backslash S^{\prime}\right|-\left|S^{\prime}\backslash S\right|\right)k^{\star}\right|\leq c\beta_{2}.$
 […]
-With Lemma A.19, it is not hard to extend the result in Corollary B.9 to the case dealing with the convex combination of the indicator vectors.
-Corollary B.10.
+With Lemma B.19, it is not hard to extend the result in Corollary C.9 to the case dealing with the convex combination of the indicator vectors.
+Corollary C.10.
 Assume $\mathbf{Y}$ is $(2\epsilon,\beta_{1},\beta_{2})$-regular, then for any vectors $a,b\in[0,1]^{N}$ with $\left\lVert a\right\rVert_{1}\leq\epsilon N,\left\lVert b\right\rVert_{1}\leq\epsilon N$, we have with probability greater than $1-\alpha$:
 (i), $\left|\left\langle\sum\limits_{i=1}^{N}a_{i}Y_{i},\sum\limits_{j=1}^{N}b_{j}Y_{j}\right\rangle-\left\langle a,b\right\rangle k^{\star}\right|\leq c_{1}\beta_{2}$; (ii), $\left|\left\langle\sum\limits_{i=1}^{N}a_{i}Y_{i},\sum\limits_{j=1}^{N}Y_{j}\right\rangle-\left\lVert a\right\rVert_{1}k^{\star}\right|\leq c_{2}\sqrt{N}\beta_{1}$
-In Corollary B.9, there is no corresponding (ii) since that is exactly the third definition of regularity.
+In Corollary C.9, there is no corresponding (ii) since that is exactly the third definition of regularity.
 Proof.
 We proof by direct calculation.
-By Lemma A.19, the vectors $a$ and $b$ can be represent as the convex combination of some indicator vectors, respectively.
+By Lemma B.19, the vectors $a$ and $b$ can be represent as the convex combination of some indicator vectors, respectively.
 Let $a=\sum\limits_{k=1}^{M_{a}}\xi_{k}\mathbf{1}_{S_{k}},b=\sum\limits_{l=1}^{M_{b}}\eta_{l}\mathbf{1}_{S_{l}}$, where $S_{k},S_{l}\subset[N],1\leq k\leq N,1\leq l\leq N$.
@@
 $=k^{\star}\cdot\sum\limits_{i=1}^{N}a_{i}b_{i}+\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}R_{k,l}.$
-Here in the equation, (1) is from Corollary B.9, and $R_{k,l}\leq C\beta_{2}$ for some constant $C$.
-Now that by Lemma A.19, $0\leq\sum\limits_{k=1}^{M_{a}}\xi_{k}\leq 1,0\leq\sum\limits_{l=1}^{M_{b}}\eta_{l}\leq 1$, we know $\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}R_{k,l}\leq C\beta_{2}$.
+Here in the equation, (1) is from Corollary C.9, and $R_{k,l}\leq c\beta_{2}$ for some constant $c$.
+Now that by Lemma B.19, $0\leq\sum\limits_{k=1}^{M_{a}}\xi_{k}\leq 1,0\leq\sum\limits_{l=1}^{M_{b}}\eta_{l}\leq 1$, we know $\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}R_{k,l}\leq c\beta_{2}$.
 Therefore, (i) is checked.
@@
 At this stage, we have all tools needed for the proof.
-The goal of the following Algorithms 3, 4, 5 is to filter the weight and samples such that the resulting weight and samples meet the requirements of $\omega$-regularity.
-We start with the first requirement, which is the simplest one.
+The goal of the following Algorithms 9, 10, 11 is to filter the weight $\omega$ and the samples $\mathbf{Y}$ such that the resulting $\omega$ and $\mathbf{Y}$ meet the requirements of $\omega$-regularity.
+We start with the first requirement.
 1Set $\gamma_{1}=c\left[\sqrt{k^{\star}\ln\left(\frac{N}{\alpha}\right)}+\ln\left(\frac{N}{\alpha}\right)\right]$, $\texttt{count}=0,\texttt{i}=0$
@@
 Algorithm 9 Prefiltering.
-Lemma B.11.
-Upon finishing the algorithm 3, all of samples satisfy
+Lemma C.11.
+Upon finishing Algorithm 9, all of samples satisfy
 $\left|\left\lVert Y_{i}\right\rVert_{2}^{2}-k^{\star}\right|\leq\gamma_{1}.$
-Moreover, under the null hypothesis, the algorithm will terminate without rejecting with probability at least $1-\alpha$.
+Moreover, under the null hypothesis $H_{0}$, the algorithm will terminate without rejecting with probability at least $1-\alpha$.
 Proof.
 The satisfaction of the inequality comes directly from the definition of Algorithm 9.
-According to Lemma B.8 where setting $\left\lVert\nu\right\rVert_{2}=0$, all uncontaminated samples will satisfy the inequality with probability at least $1-\alpha$ and therefore will not be deleted.
+According to Lemma C.8 where setting $\left\lVert\nu\right\rVert_{2}=0$, all uncontaminated samples will satisfy the inequality with probability at least $1-\alpha$ and therefore will not be deleted.
 The remaining follows.
@@
 Let
-$\gamma_{2}:=c\left[\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right].$ (B.13)
+$\gamma_{2}:=c\left[\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right].$ (C.13)
 Denote $D(\omega)=k^{\star}\cdot\text{diag}\{\omega\}$, and $\lambda(\omega)=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$.
@@
 Finally, we check the condition whether $\left\lVert\omega\right\rVert_{1}\leq N(1-2\epsilon)$ and reject the null hypothesis directly if this holds and pass to the following step if not.
-1Set $\gamma_{2}$ as (B.13), and $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
+1Set $\gamma_{2}$ as (C.13), and $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
 2while $\lambda\geq\gamma_{2}$ do
@@
 5  Set II be the smallest index such that $\sum\limits_{i=1}^{I}\omega_{i}\geq 2\epsilon N$
-6  Update $w$ according to (B.14)
+6  Update $w$ according to (C.14)
 7 if $\left\lVert\omega\right\rVert_{1}<N(1-2\epsilon)$ then
@@
 This is formalized as the following lemma.
-Lemma B.12.
+Lemma C.12.
 Given the algorithm 10, we have the following facts:
@@
 Proof of (i) is relatively easy.
-By the update (B.14), the $\omega_{i}$ corresponding to the maximal $\tau_{i}$ will be set to zero in each iteration.
+By the update (C.14), the $\omega_{i}$ corresponding to the maximal $\tau_{i}$ will be set to zero in each iteration.
 Consequently, $\left\lVert\omega\right\rVert_{1}$ will be less than $N(1-2\epsilon)$ in at most $2\epsilon N$ rounds, followed by the rejection.
@@
 Suppose $\omega^{(t)}\in\Lambda_{1}$ and we are still in the iteration.
-By the definition of $\Lambda_{1}$ and the update (B.14), it suffices to prove $\sum\limits_{i\in[I]\cap C}\tau_{i}^{(t)}\omega_{i}^{(t)}\geq\sum\limits_{i\in[I]\cap([N]\backslash C)}\tau_{i}^{(t)}\omega^{(t)}$, i.e.,
+By the definition of $\Lambda_{1}$ and the update (C.14), it suffices to prove $\sum\limits_{i\in[I]\cap C}\tau_{i}^{(t)}\omega_{i}^{(t)}\geq\sum\limits_{i\in[I]\cap([N]\backslash C)}\tau_{i}^{(t)}\omega^{(t)}$, i.e.,
 $v^{\top}\left(\sum\limits_{i\in[I]\cap C}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}\right)v\geq v^{\top}\left(\sum\limits_{i\in[I]\cap([N]\backslash C)}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}\right)v.$
@@
 $v^{\top}\left(\sum\limits_{i=1}^{N}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}-N\mathbf{I}_{k^{\star}}\right)v\geq\gamma_{2}.$
-By Corollary A.14 (set $\left\lVert\nu\right\rVert_{2}=0$) and the definition of $\gamma_{2}$ (B.13), with high probability, we have
+By Corollary B.14 (set $\left\lVert\nu\right\rVert_{2}=0$) and the definition of $\gamma_{2}$ (C.13), with high probability, we have
 […]
@@
 Hence $I-(2\epsilon N+1)\leq\sum\limits_{i\in[I]}(1-\omega_{i}^{(t)})\leq 2\epsilon N$, which leads to $I\leq 4\epsilon N+1$.
-Therefore, according to Lemma A.15, with high probability, we have
-$v^{\top}\left(\sum\limits_{i\in[I]\cap([N]\backslash C)}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}\right)v\leq c\gamma_{2}.$ (B.15)
+Therefore, according to Lemma B.15, with high probability, we have
+$v^{\top}\left(\sum\limits_{i\in[I]\cap([N]\backslash C)}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}\right)v\leq c\gamma_{2}.$ (C.15)
 Now that we sorted $\tau_{i}$ by decreasing order, $\sum\limits_{i\in[I]}\omega_{i}\geq 2\epsilon N$, and $\left|C\right|\leq\epsilon N$, we have
@@
 Therefore $\sum\limits_{i\in[I]}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{2}{3}\sum\limits_{i\in C}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{2(1-c)}{3}\gamma_{2}$.
-Note that by selecting the constant in (B.13) sufficiently large, we can let $c$ sufficiently small.
+Note that by selecting the constant in (C.13) sufficiently large, we can let $c$ sufficiently small.
 Hence, we can assume $c\leq\frac{1}{4}$, by which we have $\sum\limits_{i\in[I]}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{1}{2}\gamma_{2}$.
-Combine with (B.15), we find
+Combine with (C.15), we find
 $\sum\limits_{i\in[I]\cap C}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{1}{4}\gamma_{2}\geq\sum\limits_{i\in[I]\cap([N]\backslash C)}\omega_{i}^{(t)}\tau_{i}^{(t)}.$
@@
 To prove (iii), we need the following lemma, which exploits the property of the norm of the mean from the algorithm.
-Lemma B.13.
+Lemma C.13.
 If $\left\lVert\nu\right\rVert_{2}\geq c$ for some sufficient large but universal constant $c$, then Algorithm 10 will reject $H_{0}$ with high probability.
@@
 $\frac{\sum\limits_{i\in[N]\backslash C}\omega_{i}(v^{\top}Y_{i})^{2}}{N}-1\leq\frac{\sum\limits_{i=1}^{N}\omega_{i}(v^{\top}Y_{i})^{2}}{N}-1<\frac{\gamma_{2}}{N}\leq c,$
-where the last step is from the definition of $\gamma_{2}$ (B.13). If we expand $v^{\top}Y_{i}$, this merely says
-$\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]^{2}+\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}+2\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]\left\lVert\nu\right\rVert_{2}\leq cN.$ (B.16)
+where the last step is from the definition of $\gamma_{2}$ (C.13). If we expand $v^{\top}Y_{i}$, this merely says
+$\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]^{2}+\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}+2\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]\left\lVert\nu\right\rVert_{2}\leq cN.$ (C.16)
 By the fact that $ab\geq-\frac{1}{4}a^{2}-b^{2}$ for $a,b\in\mathbb{R}$, we know
 $2\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]\left\lVert\nu\right\rVert_{2}\geq-\frac{1}{2}\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}-2\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]^{2}.$
-Adding both sides with $\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}$ and plugging (B.16) in, we have
+Adding both sides with $\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}$ and plugging (C.16) in, we have
 $\frac{1}{2}\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}\leq cN+\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]^{2}\leq cN+\sum\limits_{i\in[N]\backslash C}\left[v^{\top}(Y_{i}-\nu)\right]^{2}.$
 Now since $\left\lVert v\right\rVert_{2}=1$, we know that $\left[v^{\top}(Y_{i}-\nu)\right]^{2}$ is subject to a chi-square distribution with degree of freedom 1.
-By the standard argument for the sub-exponential random variables (see A.7), we know that $\sum\limits_{i\in[N]\backslash C}\left[v^{\top}(Y_{i}-\nu)\right]^{2}\leq N+c\sqrt{N}$ with (high) constant probability.
+By the standard argument for the sub-exponential random variables (see B.7), we know that $\sum\limits_{i\in[N]\backslash C}\left[v^{\top}(Y_{i}-\nu)\right]^{2}\leq N+c\sqrt{N}$ with (high) constant probability.
 On the other hand, by the definition of the algorithm we have $\left\lVert\omega_{[N]\backslash C}\right\rVert_{1}\geq\left\lVert\omega\right\rVert_{1}-\epsilon N\geq(1-3\epsilon)N$.
@@
 $\frac{1}{2}(1-3\epsilon)N\left\lVert\nu\right\rVert_{2}^{2}\leq c_{1}\cdot N+c_{2}\cdot\sqrt{N}.$
-Now deviding both sides by $\frac{1}{2}(1-3\epsilon)N$, we see that Lemma B.13 holds if we take $c=\frac{2(c_{1}+c_{2})}{1-3\epsilon}$
+Now deviding both sides by $\frac{1}{2}(1-3\epsilon)N$, we see that Lemma C.13 holds if we take $c=\frac{2(c_{1}+c_{2})}{1-3\epsilon}$
 We are going to prove (iii) by contradiction.
@@
 This indicates that there exists a vector $v\in\mathbb{R}^{k^{\star}}$ with $\left\lVert v\right\rVert_{2}=1$ such that $\sum\limits_{i\in S}\sqrt{\omega_{i}}v^{\top}Y_{i}>\sqrt{2\gamma_{2}\epsilon N}$.
-Therefore, by the Cauchy-Schwarz inequality, we have $\sum\limits_{i\in S\cup C}\omega_{i}(v^{\top}Y_{i})^{2}\geq 2\gamma_{2}$.
+Therefore, by the Cauchy–Schwarz inequality, we have $\sum\limits_{i\in S\cup C}\omega_{i}(v^{\top}Y_{i})^{2}\geq 2\gamma_{2}$.
 This implies that
-$\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}Y_{i}Y_{i}^{\top}+\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}\right\rVert_{2}\geq\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}Y_{i}Y_{i}^{\top}\right\rVert_{2}\geq 2\gamma_{2}.$ (B.17)
+$\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}Y_{i}Y_{i}^{\top}+\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}\right\rVert_{2}\geq\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}Y_{i}Y_{i}^{\top}\right\rVert_{2}\geq 2\gamma_{2}.$ (C.17)
 Now consider the quantity $\left\lVert\sum\limits_{i\in[N]\backslash(S\cup C)}\omega_{i}Y_{i}Y_{i}^{\top}-\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert$, by the triangular inequality, we have
@@
 $+\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}-\left|S\cup C\right|\nu\nu^{\top}\right\rVert_{2}$
-$\leq\underbrace{\left\lVert\sum\limits_{i=1}^{N}\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}-N\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}}_{\text{I}}+\underbrace{\left\lVert\sum\limits_{i=1}^{N}(1-\omega_{i})\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}\right\rVert_{2}}_{\text{II}}$
-$+\underbrace{\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}\widetilde{Y_{i}}\widetilde{Y_{i}}^{\top}\right\rVert}_{\text{III}}+\left|S\cup C\right|\left\lVert\nu\right\rVert_{2}^{2}.$
-By Lemma B.13, without loss of generality we can assume $\left\lVert\nu\right\rVert_{2}\leq c_{0}$ for some constant $c_{0}$.
+$\leq\text{I}+\text{II}+\text{III}+\left|S\cup C\right|\left\lVert\nu\right\rVert_{2}^{2}.$
+where I, II and III are defined as
+$\text{I}=\left\lVert\sum\limits_{i=1}^{N}\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}-N\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2},\text{II}=\left\lVert\sum\limits_{i=1}^{N}(1-\omega_{i})\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}\right\rVert_{2},\text{III}=\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}\widetilde{Y_{i}}\widetilde{Y_{i}}^{\top}\right\rVert$.
+By Lemma C.13, without loss of generality we can assume $\left\lVert\nu\right\rVert_{2}\leq c_{0}$ for some constant $c_{0}$.
 For the last term, since $\left|S\cup C\right|\leq 2\epsilon N$, obviously it can be bounded by $c\gamma_{2}$.
@@
 $\text{I}\leq\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)(\widetilde{Y}_{i}-\nu)^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}+2\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)^{\top}\nu\right\rVert_{2}.$
-Set $\left\lVert\nu\right\rVert_{2}=0$ in Corollary A.14, we have
-$\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)(\widetilde{Y}_{i}-\nu)^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\leq C_{1}\left(\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)\right).$ (B.18)
-By normalization and standard concentration tools (see A.6), we have
-$\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)^{\top}\nu\right\rVert_{2}=\left\lVert\nu\right\rVert_{2}\cdot\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)\right\rVert_{2}=C_{0}\cdot\sqrt{Nk^{\star}}\cdot\left\lVert\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\frac{\widetilde{Y}_{i}-\nu}{\sqrt{k^{\star}}}\right\rVert_{2}\leq c_{2}\sqrt{Nk^{\star}}$ (B.19)
-with (high) constant probability. Now combine (B.18) and (B.19) to get
+Set $\left\lVert\nu\right\rVert_{2}=0$ in Corollary B.14, we have
+$\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)(\widetilde{Y}_{i}-\nu)^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\leq C_{1}\left(\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)\right).$ (C.18)
+By normalization and standard concentration tools (see B.6), we have
+$\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)^{\top}\nu\right\rVert_{2}=\left\lVert\nu\right\rVert_{2}\cdot\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)\right\rVert_{2}=C_{0}\cdot\sqrt{Nk^{\star}}\cdot\left\lVert\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\frac{\widetilde{Y}_{i}-\nu}{\sqrt{k^{\star}}}\right\rVert_{2}\leq c_{2}\sqrt{Nk^{\star}}$ (C.19)
+with (high) constant probability. Now combine (C.18) and (C.19) to get
 $\text{I}\leq(c_{1}+2c_{2})\left(\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)\right)\leq c\gamma_{2}.$
-Bound on II. Since upon the finish the algorithm, we have $\left\lVert\mathbf{1}-\omega\right\rVert_{1}\leq 2\epsilon N$, by Corollary A.16, we know
+Bound on II. Since upon the finish the algorithm, we have $\left\lVert\mathbf{1}-\omega\right\rVert_{1}\leq 2\epsilon N$, by Corollary B.16, we know
 $\text{II}\leq c_{3}\left[\epsilon N\ln\left(\frac{1}{\epsilon}\right)+k^{\star}+\ln\left(\frac{1}{\alpha}\right)\right]\leq c\gamma_{2}.$
-Bound on III. Since $\left|S\cup C\right|\leq 2\epsilon N$, by Corollary A.16 and an argument similar to the one in II, we also have
+Bound on III. Since $\left|S\cup C\right|\leq 2\epsilon N$, by Corollary B.16 and an argument similar to the one in II, we also have
 $\text{III}\leq c\gamma_{2}.$
 Therefore, we conclude that
-$\left\lVert\sum\limits_{i\in[N]\backslash(S\cup C)}\omega_{i}Y_{i}^{\top}Y_{i}-\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\leq 4c\gamma_{2}.$ (B.20)
-Now combine (B.17) and (B.20) and use the triangular inequality, we find that
+$\left\lVert\sum\limits_{i\in[N]\backslash(S\cup C)}\omega_{i}Y_{i}^{\top}Y_{i}-\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\leq 4c\gamma_{2}.$ (C.20)
+Now combine (C.17) and (C.20) and use the triangular inequality, we find that
 $\left\lVert\sum\limits_{i=1}^{N}\omega_{i}Y_{i}^{\top}Y_{i}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\geq(2-4c)\gamma_{2}.$
-Again, by choosing the constant in the definition of $\gamma_{2}$ (B.13) sufficiently large, we can make $c$ above sufficiently small.
+Again, by choosing the constant in the definition of $\gamma_{2}$ (C.13) sufficiently large, we can make $c$ above sufficiently small.
 Therefore, we can assume $c<\frac{1}{4}$, which leads to $\left\lVert\sum\limits_{i=1}^{N}\omega_{i}Y_{i}^{\top}Y_{i}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}>\gamma_{2}$, a contradiction of the control condition of Algorithm 10.
@@
 We now turn to the high-dimensional scenario where $N\leq k^{\star}$. In this case, we let
-$\gamma_{3}=c\left(\sqrt{Nk^{\star}}+\sqrt{k^{\star}\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right)$ (B.21)
+$\gamma_{3}=c\left(\sqrt{Nk^{\star}}+\sqrt{k^{\star}\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right)$ (C.21)
 for some universal constant $c$. The control condition of the iteration, in this case, is
@@
 Next, we update $\omega$ as following
-$\omega_{i}=\left(1-\frac{\tau_{i}}{\max\limits_{i}\tau_{i}}\right)\omega_{i},1\leq i\leq N.$ (B.22)
+$\omega_{i}=\left(1-\frac{\tau_{i}}{\max\limits_{i}\tau_{i}}\right)\omega_{i},1\leq i\leq N.$ (C.22)
 Finally, we check the condition whether $\left\lVert\omega\right\rVert_{1}\leq N(1-6\epsilon)$ and reject the null directly if this holds and pass to the following step if not.
-1Set $\gamma_{3}$ as (B.21), and $\lambda=\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
+1Set $\gamma_{3}$ as (C.21), and $\lambda=\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
 2while $\lambda\geq\gamma_{3}$ do
@@
 4  Compute $\tau_{i}=\frac{v_{i}^{2}}{\omega_{i}}\mathbf{1}_{\left\{w_{i}>0\right\}}$
-5  Update $w$ according to (B.22)
+5  Update $w$ according to (C.22)
 6 if $\left\lVert\omega\right\rVert_{1}<N(1-6\epsilon)$ then
@@
 Similar to the algorithm 10, we will have the following results to guarantee the satisfaction of the second condition of $\omega$-regularity.
-Lemma B.14.
+Lemma C.14.
 Given the algorithm 11, the following statements hold:
@@
 Proof.
-Proof of (i) is same as the one in the proof of Lemma B.12.
+Proof of (i) is same as the one in the proof of Lemma C.12.
 […]
@@
 Suppose this does not hold, then $\left\lVert v_{C}\right\rVert_{2}^{2}<\frac{1}{6}$.
-By Corollary A.14 (set $\left\lVert\nu\right\rVert_{2}=0$), we have $\left\lVert\sqrt{D(\omega_{[N]\backslash C}^{(t)})}Y_{[N]\backslash C}Y_{[N]\backslash C}^{{}^{\top}}\sqrt{D(\omega_{[N]\backslash C}^{(t)})}-D(\omega_{[N]\backslash C}^{(t)})\right\rVert_{2}\leq c\gamma_{3}\leq c\lambda^{(t)}$ with high probability.
+By Corollary B.14 (set $\left\lVert\nu\right\rVert_{2}=0$), we have $\left\lVert\sqrt{D(\omega_{[N]\backslash C}^{(t)})}Y_{[N]\backslash C}Y_{[N]\backslash C}^{{}^{\top}}\sqrt{D(\omega_{[N]\backslash C}^{(t)})}-D(\omega_{[N]\backslash C}^{(t)})\right\rVert_{2}\leq c\gamma_{3}\leq c\lambda^{(t)}$ with high probability.
 However, since $v$ is the eigenvector corresponding to the maximal eigenvalue, we have
@@
 Noting that $\left\lVert v_{C}\right\rVert_{2}^{2}+\left\lVert v_{[N]\backslash C}\right\rVert_{2}^{2}=1$, define $f_{c}(x)=c+(1-c)x+2\sqrt{x(1-x)},0\leq x<\frac{1}{6}$.
-Note that by selecting the constant in (B.21) sufficiently large, we can let $c\rightarrow 0$.
+Note that by selecting the constant in (C.21) sufficiently large, we can let $c\rightarrow 0$.
 It can be proved that in this case $0\leq f_{0}(x)<1$, which is a contradiction.
@@
 Here (i) uses the condition that $\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}\leq\gamma_{3}$. The proof is completed. ∎
-After Algorithms 4 or 5, we have filtered the samples to ensure the second condition of $\omega$-regularity. Now we turn to the check of the final condition.
-Actually, the weight $\omega$ output by Algorithms 4 or 5 already provides us with some useful insight.
-Lemma B.15.
-Set $\beta_{1}$, $\beta_{2}$ according to (B.11) and $\gamma_{i},i\in\left\{2,3\right\}$ according to (B.13) or (B.21) depending on whether $N\geq k^{\star}$ or not.
-For the weight $\omega$ output by the Algorithms 4 or 5 and any S[$\CS\subset[N]\backslash$ C with $\left|S\right|\leq\epsilon N$, we have
+After Algorithms 10 or 11, we have filtered the samples to ensure the second condition of $\omega$-regularity. Now we turn to the check of the final condition.
+Actually, the weight $\omega$ output by Algorithms 10 or 11 already provides us with some useful insight.
+Lemma C.15.
+Set $\beta_{1}$, $\beta_{2}$ according to (C.11) and $\gamma_{i},i\in\left\{2,3\right\}$ according to (C.13) or (C.21) depending on whether $N\geq k^{\star}$ or not.
+For the weight $\omega$ output by the Algorithms 10 or 11 and any S[$\CS\subset[N]\backslash$ C with $\left|S\right|\leq\epsilon N$, we have
 $\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|\leq c(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma)$
@@
 Proof.
-By Lemma B.8, with high probability the original samples $\widetilde{\mathbf{Y}}$ are $(2\epsilon,\beta_{1},\beta_{2})$-regular. By spliting $N]=([N]\backslash C)\cup C$, we have
+By Lemma C.8, with high probability the original samples $\widetilde{\mathbf{Y}}$ are $(2\epsilon,\beta_{1},\beta_{2})$-regular. By spliting $N]=([N]\backslash C)\cup C$, we have
 $\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|$ $=\underbrace{\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},\widetilde{Y}_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|}_{\text{I}}$
@@
 We now bound the terms $\text{I}_{1},\text{I}_{2},\text{II},\text{III}$ respectively.
-Bound on $\text{I}_{1}$. Because $\left|S\right|\leq\epsilon N$, from $\omega$-regularity and Corollary B.10, we know $\text{I}_{1}\leq c_{1}\sqrt{N}\beta_{1}$.
+Bound on $\text{I}_{1}$. Because $\left|S\right|\leq\epsilon N$, from $\omega$-regularity and Corollary C.10, we know $\text{I}_{1}\leq c_{1}\sqrt{N}\beta_{1}$.
 Bound on $\text{I}_{2}$.
-Because $\omega$ is output by the Algorithms 4 or 5, we have $\left\lVert 1-\sqrt{\omega}\right\rVert_{1}\leq\left\lVert 1-\omega\right\rVert_{1}\leq\epsilon N$.
-By Corollary B.10 again, we have $\text{I}_{2}\leq c_{2}\beta_{2}$.
-Bound on II. From Corollary B.10 and the fact $S\cap C=\varnothing$, we know $\text{II}\leq c_{3}\beta_{2}$.
-Bound on III. From Lemma B.12, B.14, and the polarization equality, we have $\text{III}\leq c_{4}\epsilon N\gamma$.
+Because $\omega$ is output by the Algorithms 10 or 11, we have $\left\lVert 1-\sqrt{\omega}\right\rVert_{1}\leq\left\lVert 1-\omega\right\rVert_{1}\leq\epsilon N$.
+By Corollary C.10 again, we have $\text{I}_{2}\leq c_{2}\beta_{2}$.
+Bound on II. From Corollary C.10 and the fact $S\cap C=\varnothing$, we know $\text{II}\leq c_{3}\beta_{2}$.
+Bound on III. From Lemma C.12, C.14, and the polarization equality, we have $\text{III}\leq c_{4}\epsilon N\gamma$.
 Combining them together, we come to the conclusion of the lemma. ∎
 We are now prepared for the final algorithm, which is designed to ensure the third condition of $\omega$-regularity.
-Given samples $S$ and weight $\omega$ by the algorithms 4 or 5, we compute $\tau_{i}=\left|\left\langle\sqrt{\omega_{i}}Y_{i},\sum\limits_{j=1}^{N}\sqrt{\omega_{j}}Y_{j}\right\rangle-\omega_{i}k^{\star}\right|\mathbf{1}_{\left\{\omega_{i}>0\right\}}$.
+Given samples $S$ and weight $\omega$ by the algorithms 10 or 11, we compute $\tau_{i}=\left|\left\langle\sqrt{\omega_{i}}Y_{i},\sum\limits_{j=1}^{N}\sqrt{\omega_{j}}Y_{j}\right\rangle-\omega_{i}k^{\star}\right|\mathbf{1}_{\left\{\omega_{i}>0\right\}}$.
 Then, we set $\omega_{i}$ to zero for the indices $i$ corresponding to the $\epsilon N$ maximal $\tau_{i}$.
@@
 Algorithm 12 Weight filtering.
-Lemma B.16.
+Lemma C.16.
 For the weight output by Algorithm 12 and any $S\subset[N]$ with $\left|S\right|\leq\epsilon N$, we have
 $\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|\leq c(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma).$
-This lemma is very similar to Lemma B.15 but stronger in the sense that we allow any $S\subset[N]$ rather than S[$\CS\subset[N]\backslash$ C.
+This lemma is very similar to Lemma C.15 but stronger in the sense that we allow any $S\subset[N]$ rather than S[$\CS\subset[N]\backslash$ C.
 This advantage is exactly granted by the sorting and pruning processes in Algorithm 12.
@@
 The proof of this lemma is not hard but depends on the following fact.
-Because we have proved in Lemma B.15 that for any S[$\CS\subset[N]\backslash$ C and $\left|S\right|\leq\epsilon N$, we have $\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|\leq c(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma)$, we know that this bound also holds for any $S\subset[N]\backslash\left\{1,2,\dots,\epsilon N\right\}$.
+Because we have proved in Lemma C.15 that for any S[$\CS\subset[N]\backslash$ C and $\left|S\right|\leq\epsilon N$, we have $\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|\leq c(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma)$, we know that this bound also holds for any $S\subset[N]\backslash\left\{1,2,\dots,\epsilon N\right\}$.
 With the proning on $\omega$ to remove the first $\epsilon N$ coordinates, we can assume without loss of generality that $S\subset[N]\backslash\left\{1,2,\dots\epsilon N\right\}$.
@@
 ∎
-We have finished the proofs of Lemma B.11, B.12, B.14, and B.16, which means we have verified the correctness of $\omega$-regularity for our samples with high probability.
+We have finished the proofs of Lemma C.11, C.12, C.14, and C.16, which means we have verified the correctness of $\omega$-regularity for our samples with high probability.
 The final step is to exploit the property of $\omega$-regularity to distinguish between $H_{0}$ and $H_{1}$.
-Lemma B.17.
-Let $\omega$ be the weights after running through the algorithm 3, 4 or 5, and 6, and $\beta_{1},\beta_{2}$ seleced as in Lemma B.8. Then, with high probability we have
-$\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}-N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right|\leq c\left(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma+\underbrace{\left(N^{\frac{3}{2}}\left\lVert\nu\right\rVert_{2}+\sqrt{Nk^{\star}}\right)\sqrt{\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)}_{:=\text{R}}\right).$ (B.23)
+Lemma C.17.
+Let $\omega$ be the weights after running through the algorithm 9, 10 or 11, and 12, and $\beta_{1},\beta_{2}$ seleced as in Lemma C.8. Then, with high probability we have
+$\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}-N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right|\leq c\left(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma+\underbrace{\left(N^{\frac{3}{2}}\left\lVert\nu\right\rVert_{2}+\sqrt{Nk^{\star}}\right)\sqrt{\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)}_{:=\text{R}}\right).$ (C.23)
 Proof.
@@
 Define the vector $r\in\mathbb{R}^{N}$ where $r_{i}=1$ for $i\in C$ and $1-\sqrt{\omega_{i}}$ for i[$\Ci\in[N]\backslash$ C.
-By Algorithm 4 or 5, we have $\left\lVert r\right\rVert_{1}\leq\epsilon N+6\epsilon N=7\epsilon N$.
+By Algorithm 10 or 11, we have $\left\lVert r\right\rVert_{1}\leq\epsilon N+6\epsilon N=7\epsilon N$.
 For the term I, we have
@@
 $\text{I}_{1}=\sum\limits_{i=1}^{N}\left\lVert\widetilde{Y}_{i}-\nu\right\rVert_{2}^{2}+N^{2}\left\lVert\nu\right\rVert_{2}^{2}+2N\nu^{\top}\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu).$
-By the standard concentration bounds (see A.6 and A.7), with high probability we have
+By the standard concentration bounds (see B.6 and B.7), with high probability we have
 $\left|\text{I}_{1}-Nk^{\star}-N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right|\leq c_{1}\left(N^{\frac{3}{2}}\left\lVert\nu\right\rVert_{2}\sqrt{\ln\left(\frac{1}{\alpha}\right)}+\sqrt{Nk^{\star}\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right).\right)$
-For $\text{I}_{2}$, by the argument above, Lemma B.8, and Corollary B.10 with appropriate choice of $\epsilon$, with high probability we have
+For $\text{I}_{2}$, by the argument above, Lemma C.8, and Corollary C.10 with appropriate choice of $\epsilon$, with high probability we have
 $\left|\text{I}_{2}-k^{\star}\left\lVert r\right\rVert_{2}^{2}\right|\leq c_{2}\beta_{2}.$
-Again by Lemma B.8 and Corollary B.10, with high probability we have
+Again by Lemma C.8 and Corollary C.10, with high probability we have
 $\left|\text{I}_{3}-2k^{\star}\left\lVert r\right\rVert_{1}\right|\leq c_{3}\sqrt{N}\beta_{1}.$
@@
 $\left|\text{I}-k^{\star}\left\lVert\omega_{[N]\backslash C}\right\rVert_{1}-N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right|\leq c\left[\sqrt{N}\beta_{1}+\beta_{2}+\left(N^{\frac{3}{2}}\left\lVert\nu\right\rVert_{2}+\sqrt{Nk^{\star}}\right)\sqrt{\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)\right].$
-Bound on II. By Lemma B.12 (recall the definition of $\gamma_{2}$ (B.13) and the fact that $N>k^{\star}$ in this case) or B.14, with high probability we have
+Bound on II. By Lemma C.12 (recall the definition of $\gamma_{2}$ (C.13) and the fact that $N>k^{\star}$ in this case) or C.14, with high probability we have
 $\left|\text{II}-k^{\star}\left\lVert\omega_{C}\right\rVert_{1}\right|\leq c\epsilon N\gamma.$
-Bound on III. By Lemma B.16, with high probability we have
+Bound on III. By Lemma C.16, with high probability we have
 $\left|\text{III}-2k^{\star}\left\lVert\omega_{C}\right\rVert_{1}\right|\leq c\left(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma\right).$
@@
 Proof of Theorem 3.11.
-Substituting $\nu$ with $P^{\star}\mu$, the key is to require each term on the right-hand-side in (B.23) to be asymptotically smaller than the term $N^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}$.
+Substituting $\nu$ with $P^{\star}\mu$, the key is to require each term on the right-hand-side in (C.23) to be asymptotically smaller than the term $N^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}$.
 For convenience, we rewrite the definition of $\beta_{1}$ and $\beta_{2}$ here.
@@
 15 $\ln\left(\frac{1}{\alpha}\right)$ Dominated by the term 3
-Table 2: terms from (B.23) and required conditions on $\left\lVert P^{\star}\mu\right\rVert_{2}$.
+Table 2: terms from (C.23) and corresponding required conditions on $\left\lVert P^{\star}\mu\right\rVert_{2}$.
 Inspecting the column “Required Conditions” and computing the maximal, we conclude the final requirement for $\left\lVert P^{\star}\mu\right\rVert_{2}$:
-$\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\max\limits\left\{\frac{\epsilon\ln\left(\frac{N}{\alpha}\right)}{\sqrt{N}},\epsilon^{2}\ln\left(\frac{N}{\alpha}\right),\sqrt{\frac{\epsilon^{2}k^{\star}\ln\left(\frac{N}{\alpha}\right)}{N}},\frac{\sqrt{k^{\star}}\ln\left(\frac{1}{\alpha}\right)}{N}\right\}.$ (B.24)
-Once this requirement is satisfied, by Lemma B.17 and the inequality (B.23), we know that $\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}P^{\star}X_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}\right|$ will be less than $c_{1}N^{2}P(\epsilon,k^{\star},N,1)$ under $H_{0}$ and greater than $c_{2}N^{2}P(\epsilon,k^{\star},N,1)$ under $H_{1}$ for some $0<c_{1}<c_{2}<1$, and both with high probability.
+$\left\lVert P^{\star}\mu\right\rVert_{2}^{2}\gtrsim\max\limits\left\{\frac{\epsilon\ln\left(\frac{N}{\alpha}\right)}{\sqrt{N}},\epsilon^{2}\ln\left(\frac{N}{\alpha}\right),\sqrt{\frac{\epsilon^{2}k^{\star}\ln\left(\frac{N}{\alpha}\right)}{N}},\frac{\sqrt{k^{\star}}\ln\left(\frac{1}{\alpha}\right)}{N}\right\}.$ (C.24)
+Once this requirement is satisfied, by Lemma C.17 and the inequality (C.23), we know that $\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}P^{\star}X_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}\right|$ will be less than $c_{1}N^{2}P(\epsilon,k^{\star},N,1)$ under $H_{0}$ and greater than $c_{2}N^{2}P(\epsilon,k^{\star},N,1)$ under $H_{1}$ for some $0<c_{1}<c_{2}<1$, and both with high probability.
 $c_{1}$ and $c_{2}$ are universal and their values can be determined following proper choices of $\epsilon,\alpha,N,k^{\star}$.
-From the argument at the end of the proof of Theorem 3.6, we know (B.24) leads to (3.4).
+From the argument at the end of the proof of Theorem 3.8, we know (C.24) leads to (3.6).
 The proof is completed.
 ∎
-B.7 Proof of Corollary 3.12
+C.7 Proof of Corollary 3.15
 Proof.
@@
 ∎
-B.8 Proof of Theorem 3.17
+C.8 Proof of Theorem 4.5
 Proof.
@@
 Again, we impose a uniform prior on the set $\left\{\mu\left\lvert\right.\mu\in K,\mu_{i}=\kappa\gamma_{i}\theta_{i},\gamma_{i}\in\left\{-1,1\right\},i\in I\right\}$ for some positive constant $\kappa$.
-By the same arguments of Appendix B.2, if $\mathbb{E}_{\mu,\mu^{\prime}}\exp\left\{\frac{N\mu^{\top}\mu^{\prime}}{\sigma^{2}}\right\}-1$, or equivalently $\frac{N^{2}\left\lVert\theta\right\rVert_{4}^{4}\kappa^{4}}{2\sigma^{4}}$ is sufficiently small, then there is no test with uniformly small type-I and type-II errors.
+By the same arguments of Appendix C.2, if $\mathbb{E}_{\mu,\mu^{\prime}}\exp\left\{\frac{N\mu^{\top}\mu^{\prime}}{\sigma^{2}}\right\}-1$, or equivalently $\frac{N^{2}\left\lVert\theta\right\rVert_{4}^{4}\kappa^{4}}{2\sigma^{4}}$ is sufficiently small, then there is no test with uniformly small type-I and type-II errors.
 We have
@@
 Consequently, it is not difficult to verify that $\frac{\sqrt{N}\left\lVert v\right\rVert_{4}^{2}}{\sigma^{2}}\leq\epsilon$.
-By Lemma B.1, we establish the second lower bound that $\rho_{\text{critical}}\gtrsim\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{1/p-1/4}}{N^{1/4}}\sigma$.
+By Lemma C.1, we establish the second lower bound that $\rho_{\text{critical}}\gtrsim\frac{\sqrt{\epsilon}(k_{p,2}^{\star})^{1/p-1/4}}{N^{1/4}}\sigma$.
 Suppose now that $\theta$ is the vector from Lemma 4.3 when $k=k_{p,3}^{\star}$.
 Subsequently, we know $\left\lVert\theta\right\rVert_{2}^{2}\leq\left\lVert\theta\right\rVert_{p}^{p}\cdot\left\lVert\theta\right\rVert_{\infty}^{2-p}\leq\epsilon^{2}\sigma^{2}$.
-Following the same reasoning in Appendix B.4, we know there always exists a strategy for $\mathcal{C}$ to discourage the testing.
+Following the same reasoning in Appendix C.4, we know there always exists a strategy for $\mathcal{C}$ to discourage the testing.
 Therefore, we establish the third lower bound.
@@
 ∎
-B.9 Proof of Theorem 3.18
+C.9 Proof of Theorem 4.6
 Before we prove Theorem 4.6, we need the following observation of the p-convex orthosymmetric set.
-Lemma B.18 (Truncation of a $p$-convex orthosymmetric set).
+Lemma C.18 (Truncation of a $p$-convex orthosymmetric set).
 Assume $K\subset\mathbb{R}^{d}$ is a $p$-convex orthosymmetric set.
@@
 Proof.
-Lemma B.18 is straightforward from the definition.
+Lemma C.18 is straightforward from the definition.
 For any $\theta_{J,1},\theta_{J,2}\in K_{J}$, suppose $\theta_{J,1}=\beta_{1,J},\theta_{J,2}=\beta_{2,J},\beta_{1},\beta_{2}\in K$, then we know for any $\lambda\in[0,1]$, we have $\lambda\theta_{J,1}^{p}+(1-\lambda)\theta_{J,2}^{p}=\lambda\beta_{1,J}^{p}+(1-\lambda)\beta_{2,J}^{p}=(\lambda\beta_{1}^{p}+(1-\lambda)\beta_{2}^{p})_{J}\overset{\text{(i)}}{=}\widetilde{\beta}_{J}^{p}\in K_{J}^{p}$, where $\widetilde{\beta}\in K^{p}$ and (i) is from that $K$ is p-convex.
@@
 For $k_{p,1}^{\star}$ and its corresponding $P_{p,1}^{\star}$, without loss of generality we assume $P_{p,1}^{\star}\theta=(\theta_{1},\dots,\theta_{k_{p,1}^{\star}},0,\dots,0)$.
-Regard $P_{p,1}^{\star}\theta$ as a vector in $\mathbb{R}^{k_{p,1}^{\star}}$, by the Hölder’s inequality A.17, we know $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$ for any $\theta\in K$.
+Regard $P_{p,1}^{\star}\theta$ as a vector in $\mathbb{R}^{k_{p,1}^{\star}}$, by the Hölder’s inequality B.17, we know $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$ for any $\theta\in K$.
 Further define $k_{p,1,1}^{\star}:=\max\limits\left\{j\left\lvert\right.0\leq j\leq k_{p,1}^{\star},D_{j-1}(K)>\frac{j^{1/4}}{\sqrt{N}}\sigma\right\}$.
-Since $\epsilon\leq\frac{1}{\sqrt{N}}$, by the proof of Corollary 3.12 and Lemma B.18, we know that if $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, then there exists a desired test.
+Since $\epsilon\leq\frac{1}{\sqrt{N}}$, by the proof of Corollary 3.15 and Lemma C.18, we know that if $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$, then there exists a desired test.
 However, this is already guaranteed since $\left\lVert P_{p,1}^{\star}\theta\right\rVert_{2}\gtrsim\frac{(k_{p,1}^{\star})^{1/4}}{\sqrt{N}}\sigma$ and $k_{p,1}^{\star}\geq k_{p,1,1}^{\star}$.
@@
 ∎
-B.10 Verification of the Minimality in Section 4.2
+C.10 Verification of the Minimality in Section 5.2
 Proof.
@@
 ∎
-B.11 Proof of Lemma 5.3
+C.11 Proof of Lemma 6.3
 Proof.
@@
 The discussion above shows that the minimax rate of estimation is greater than the minimax rate of testing in all possible cases. The proof is completed. ∎
-Appendix C Theoretical and Polynomial Time Algorithms
-C.1 Theoretical Algorithm
-According to Theorem 3.6 and Section B.5, we first attempt to search for a consistent set $T$ with respect to $\phi_{e}$.
-$\phi_{e}(\mathbf{X}_{T})$ then determines whether we are to accept or reject $H_{0}$.
-1
-2import $\mathbf{X},N,K,\sigma$
-3
-4Compute $k_{1}^{\star},P_{1}^{\star},k_{2}^{\star},P_{2}^{\star}$ from $N,K,\sigma,\epsilon$ as 2.3 and 2.4
-5if $k_{1}^{\star}\leq k_{2}^{\star}$ then
-6 $k^{\star}=k_{1}^{\star},P^{\star}=P_{1}^{\star}$
-7
-8else
-9 $k^{\star}=k_{2}^{\star},P^{\star}=P_{2}^{\star}$
-10
-[…]
-12for $T$ in $\left\{\text{all subsets with cardinality greater than }(1-\epsilon)N\right\}$ do
-13
-14 consistent_flag == True
-15 if $\left\lVert\sum\limits_{i\in T}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left|T\right|\sigma^{2}\geq c\left|T\right|^{2}E^{2}(\epsilon,K,N,\sigma)$ then
-16    subset_test == True
-17 else
-18    subset_test == False
-[…]
-20
-21    if $\left\lVert\sum\limits_{i\in T^{\prime}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left|T^{\prime}\right|\sigma^{2}\geq c\left|T^{\prime}\right|^{2}E^{2}(\epsilon,K,N,\sigma)$ then
-22       subsubset_test == True
-23    else
-24       subsubset_test == False
-25    if subset_test XOR subsubset_test then
-26       consistent_flag == False
-27       break
-28
-29 if consistent_flag then
-30    if subset_test then
-31         Reject $H_{0}$
-32       return
-33    else
-34         Accept $H_{0}$
-35       return
-36
-37
-38Accept $H_{0}$
-return
-Algorithm 7 exponential degree algorithm when $\rho\gtrsim E(\epsilon,K,N,\sigma^{2})$
-C.2 Polynomial Time Algorithm
-The polynomial-degree algorithm combines the three sub-algorithms in Section B.6 together to filter the observations $\mathbf{X}$ and through the weight $\omega$.
-The theoretical guarantee is provided in Section B.6.
-1
-2import $\mathbf{X},N,K,\sigma=1$
-3import Algorithm 3, 4, 5, 6
-4
-[…]
-6Compute $k_{1}^{\star},P_{1}^{\star},k_{2}^{\star},P_{2}^{\star}$ from $N,K,\sigma=1,\epsilon$ as 2.3 and 2.4
-7if $k_{1}^{\star}\leq k_{2}^{\star}$ then
-8 $k^{\star}=k_{1}^{\star},P^{\star}=P_{1}^{\star}$
-9
-10else
-11 $k^{\star}=k_{2}^{\star},P^{\star}=P_{2}^{\star}$
-12
-[…]
-14Execute Algorithm 3, and record the returning value $\texttt{R}_{1}$
-15if $\texttt{R}_{1}$ is None then
-16   Reject $H_{0}$
-17 return
-18
-19else
-20 $\mathbf{Y}=\texttt{R}_{1}$
-21
-22Set $\omega=\mathbf{1}_{[k^{\star}]}$
-23if $N>k^{\star}$ then
-24   Execute Algorithm 4, and record the returning value $\texttt{R}_{2}$
-25else
-26   Execute Algorithm 5, and record the returning value $\texttt{R}_{2}$
-27if $\texttt{R}_{2}$ is None then
-28   Reject $H_{0}$
-29 return
-30
-31else
-32 $\omega=\texttt{R}_{2}$
-33
-34Execute Algorithm 6, and record the returning value $\texttt{R}_{3}$
-[…]
-36if $\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}\right|\geq c_{2}N^{2}P(\epsilon,k^{\star},N,1)$ then
-37   Reject $H_{0}$
-38
-39else
-40   Accept $H_{0}$
-41
-return
-Algorithm 8 polynomial degree algorithm when $\rho\gtrsim P(\epsilon,k^{\star},N,1)$.